A Complete Bibliography of Publications in Chymia: Annual Studies in the History of Chemistry: 1948–1967; Historical Studies in the Physical Sciences: 1969–1971;

Historical Studies in the Physical and Biological Sciences: 1986–2007; and Historical Studies in the

Natural Sciences: 2008–date

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 April 2025 Version 1.58

Title word cross-reference

$0 [Mod08]. $105.00 [Lau14]. $116.00 [Mor09]. $130.95 [Zak13]. $14.95 [Oni15]. $15 [Kli13]. $15.46 [Mat15]. $17.00 [Ris10]. $189.00 [Gor14a]. $19.99 [Nov15, Oni15]. $20.75 [Mat15]. $21.00 [Phi10]. $21.95 [Har09, Wol13a, Zak13]. $22 [Joh08]. $22.69 [Nov15]. $22.95 [Bro14, Kli13, Lin09, Oni15]. $23 [Mat15]. $23.99 [Zak13]. $24 [Dea08]. $24.95 [Bra15, Gar09, Jac09, Kli13, Lau14, Rai10].

$25.00 [Coe10, Har09, Jac09, Lin09, Phi10]. $25.44 [Coe10]. $25.95 [Por13, Zak13]. $26 [Ved20]. $26.00 [Rog12]. $26.95 [Zak13]. $26.96 [Nov15]. $27 [Lon08, Phi10, Ved20]. $27.00 [Lau14]. $27.95 [Kev09, Mod08, Rai10]. $28 [Ham10]. $29 [Rai10]. $29.00 [Fin13]. $29.95 [Bra15, Ham10, Har09, Kev09, Rog12, San10, Sta08, Wol13a]. $29.99 [Oni15]. $30 [Bal15, Kli13, Sta08, Ved20]. $32 [Sta08]. $32.50 [Gar09]. $33.00 [Bro14]. $34.00 [Lau14]. $34.95 [Bra15]. $35 [Lon08]. $35.00 [Coe10, Cro20, Gil08, Jac09, Phi10, San10, Sch12a, Wal15a]. $35.96 [Lau14]. $36.00 [Rad13]. $36.50 [Wol13a]. $37.00 [Wol13a]. $37.10 [Str08]. $37.50 [Gar09]. $38.95 [And09]. $39.00 [Fin13]. $39.95 [Mod08, Ris10]. $40 [And09, Joh08, Mod08]. $40.00 [Coe10, Mor09, Wol13a]. $42.00 [Bro14]. $42.70 [Str08]. $43 [Ham10]. $43.00 [Ris10]. $44.95 [Lau14]. $45 [Bal15, Ham10, Joh08, Lon08, Nye09, Rai10]. $45.00 [Fin13, Gil08, Kim14, Phi10, Rad13, Sch12a]. $46 [Str08]. $46.04 [Nov15]. $49 [Mod08, Rai10]. $49.95 [Ham10, San10]. $49.99 [Fle20]. $50.00 [Por13, Wal15a]. $52 [Nye09]. $52.00 [Kev09]. $55 [Bal15, Nye09, SD08]. $55.00 [Cro20, Por13]. $59.50 [Str08]. $60.00 [Fle20, Kim14, Rog12, San10]. $65.00 [Mor09]. $65.33 [Nov15]. $68.95 [Phi10]. $70.00 [Har09, Por13]. $74.50 [SD08]. $74.95 [Har09]. $75.00 [Wal15a]. $76.00 [Por13]. $80 [Lon08]. $85 [Bal15]. $85.00 [Har09, Sch12a]. $87.00 [Wal15a]. $9.95 [Oni15]. $90.00 [Gar09, Har09]. $99.00 [Fle20]. _{2} [Cha12, Gor14a, Wea07]. B = 0 [Dah86]. N [Nye80]. ζ [Rob83].

-Rays [Nye80].

. [Lau14].

//www.gutenberg [San10]. //www.gutenberg-e.org/rentetzi/ [San10].

0-19-927905-5 [Gar09]. 0-226-23348-0 [Nov15]. 0-226-62087-5 [Lon08]. 0-292-70981-1 [Lon08]. 0-520-25750-2 [Nov15]. 0-521-86584-0 [Har09]. 0-674-01495-2 [Gar09]. 0-674-02713-2 [Kev09].

1-4214-0965-8 [Nov15]. 1600 [Bro14]. 1685 [Ris10]. 1730s [Gre83]. 1760 [Kim14].

1763 [Por13]. 1764 [MOBG09, Por13]. 1783 [Coe10]. 1785 [DD59]. 1789 [Mil50]. 17th [Hol04]. 1840s [Nyh17]. 1856 [Rai10]. 1860 [Bal15]. 1870 [And09, Kev09]. 1880s [Sak24]. 1887 [Leg15]. 1890 [Bou12]. 1910s [Sak24]. 1917 [Wol13a]. 1926 [Sta11]. 1930 [Sch12a]. 1930s [Ano05c, BR94, Cas81, Cha20, De 05, Den17, Gas96, Goo91, Hin03, Mas06, Orl98, Rig83, Tur06, Wit04]. 1934 [Phi10]. 1940 [Rad13]. 1940s [Güt24, Rai00]. 1945 [Hop96, Wu20]. 1945/1976 [Har11]. 1949 [Hop96, Rog12]. 1950 [Fin13]. 1950s [Kri08]. 1957 [Sch12a]. 1959 [Wal15a]. 1960s [Den17, Hec14, Wis12b]. 1970 [Joh08]. 1970s [Güt24, Hen19, Men21, Tur24, Wes08a]. 1972 [Rad13]. 1975 [Lau14, Mat15]. 1976 [Gol13]. 1980s [Lam18, Wes08a]. 1986 [HS89]. 1990 [Ham10]. 1990s [Lam18]. 19th [Dar02, Dör91, Gol00, Son53, Wis87]. 19th-century [Dar02, Dör91, Gol00].

2 [DeV87]. 2004 [Joh08]. 2005 [Mod08]. 20th [San10, CK90, Cat98, Dör06, Ore00, Ren09, Shi86, Shi87, Wes20]. 20th-Century [CK90, Wes20]. 2nd [Bro14, Zak13].

43 [SF15].

’50s [Nyh17].

9-780-8018-2509-5 [Rad13]. 9-780-8018-5711-9 [Rad13]. 9-780-8018-9785-6 [Por13]. 9-781-4214-0363-2 [Rad13]. 9-789-5871-6263-9 [Por13]. 978-0-06135-132-7 [Oni15]. 978-0-19-518775-5 [Str08]. 978-0-19-925406-4 [Str08]. 978-0-19-927905-0 [Gar09]. 978-0-19-968696-4 [Bal15]. 978-0-19-974005-5 [Bra15]. 978-0-195306-97-2 [Nye09].

978-0-199-55001-2 [Ris10]. 978-0-22-607968-4 [Bal15]. 978-0-22-620328-7 [Bal15]. 978-0-22-627651-9 [Bal15]. 978-0-226-02939-9 [Bro14]. 978-0-226-06863-3 [Sch12a]. 978-0-226-13949-4 [Ris10]. 978-0-226-23348-2 [Nov15]. 978-0-226-30205-8 [Coe10]. 978-0-226-34623-6 [Ham10]. 978-0-226-39008-6 [Cro20]. 978-0-226-41874-2 [Cro20]. 978-0-226-42972-4 [Har09]. 978-0-226-44046-0 [Cro20]. 978-0-226-44990-6 [Sch12a]. 978-0-226-45412-2 [Sch12a]. 978-0-226-46531-9 [San10]. 978-0-226-47078-8 [Fin13]. 978-0-226-60856-3 [Gar09]. 978-0-226-61063-4 [Wal15a]. 978-0-226-61089-4 [Phi10]. 978-0-226-64076-1 [Mor09]. 978-0-226-70917-8 [Fin13]. 978-0-226-71216-1 [Phi10]. 978-0-226-71287-1 [Rai10]. 978-0-226-72080-7 [Lin09]. 978-0-226-73128-5 [Rai10]. 978-0-226-73149-0 [Phi10]. 978-0-226-73860-4 [Rog12]. 978-0-226-76886-1 [Kim14]. 978-0-226-80347-0 [Fin13]. 978-0-2260-5853-5 [Por13]. 978-0-226111-72-8 [Nye09]. 978-0-2266-1668-1 [Rai10]. 978-0-226737-84-3 [And09]. 978-0-230-23172-6 [Lau14]. 978-0-230-35487-6 [Wal15a]. 978-0-231-13558-0 [San10]. 978-0-231-14628-9 [Wol13a]. 978-0-252-07433-2 [Phi10]. 978-0-253-00634-9 [Wol13a]. 978-0-262-01236-2 [Ham10]. 978-0-262-01658-2 [Kli13]. 978-0-262-01833-3 [Lau14].

978-0-262-07293-9 [Phi10]. 978-0-262-08378-2 [Coe10]. 978-0-262-10116-5 [Mor09]. 978-0-262-12281-8 [Joh08]. 978-0-262-13476-7 [Kev09]. 978-0-262-16242-5 [Jac09]. 978-0-262-19557-7 [Ris10]. 978-0-300-11196-5 [Gar09]. 978-0-300-11796-7 [Lon08]. 978-0-300-21617-2 [Ved20]. 978-0-3001-1075-3 [Joh08]. 978-0-374-25682-1 [Ham10]. 978-0-520-24751-2 [Jac09]. 978-0-520-25750-4 [Nov15]. 978-0-520-25812-9 [Zak13]. 978-0-520-26786-2 [Rog12]. 978-0-520-27144-9 [Zak13]. 978-0-520-27274-3 [Rog12]. 978-0-520-27477-8 [Lau14]. 978-0-520-28930-7 [Wu20]. 978-0-521-74509-3 [Zak13]. 978-0-521-85690-4 [Mor09]. 978-0-521-87087-0 [Lon08]. 978-0-521-87559-2 [Ris10]. 978-0-521-89760-0 [Sch12a]. 978-0-674-02670-4 [Kev09]. 978-0-674-03216-3 [Har09]. 978-0-674-05078-5 [Wol13a]. 978-0-674-06587-1 [Wal15a]. 978-0-674-72581-2 [Wol13a]. 978-0-674-97696-2 [Wu20]. 978-0-691-11345-6 [Rai10]. 978-0-691-12836-8 [Wol13a]. 978-0-691-12854-2 [Ham10]. 978-0-691-14032-2 [Bro14]. 978-0-691-17246-0 [Ved20]. 978-0-7432-9663-2 [Kli13]. 978-0-801-89079-6 [Coe10]. 978-0-801-89101-4 [Jac09]. 978-0-8018-8444-3 [Joh08]. 978-0-8018-8477-1 [Str08]. 978-0-8018-8602-7 [Str08]. 978-0-8047-8411-5 [Wu20]. 978-0-8071-3551-8 [Bra15].

978-0-8090-9355-7 [Oni15]. 978-0-813-54220-1 [San10]. 978-0-813-54331-4 [Ham10]. 978-0-813-54404-5 [San10]. 978-0-8131-2485-8 [Rai10]. 978-0-8135-5163-0 [Kli13]. 978-0-822-33822-2 [Har09]. 978-0-8223-4046-1 [Lin09]. 978-0-87071-609-6 [Bro14]. 978-0-87169-965-7 [Dea08]. 978-0-87328-249-9 [Lau14]. 978-0-905193-07-6 [Nye09]. 978-0262-01231-7 [Mod08]. 978-0262-08355-3 [Mod08]. 978-03001-1078-4 [SD08]. 978-0520-27096-1 [Kli13]. 978-08122-3947-8 [Mod08]. 978-08135-3666-8 [Mod08]. 978-1-107-12697-8 [Fle20]. 978-1-137-02528-9 [Lau14]. 978-1-138-58767-0 [Fle20]. 978-1-4214-0283-3 [Kim14]. 978-1-4214-0788-3 [Lau14]. 978-1-4214-0965-8 [Nov15]. 978-1-4443-3943-7 [Zak13]. 978-1-509-51893-7 [Ved20]. 978-1-55849-975-1 [Bra15]. 978-1-58834-352-9 [Oni15]. 978-1-595-58196-9 [Zak13]. 978-1-59643-827-9 [Oni15]. 978-1-59691-452-0 [Oni15]. 978-1-781-44078-0 [Fle20]. 978-1-845-11459-6 [Har09]. 978-1-86940-727-8 [Bra15]. 978-1-890951-78-8 [And09]. 978-2-200-25563-3 [Kim14]. 978-2-753-50696-1 [Coe10]. 978-3-039-10299-0 [Phi10]. 978-8-400-08692-3 [San10]. 978-84-490-3238-7 [Mat15]. 978-84-9756-311-6 [SD08]. 978-88-222-5496-2 [SD08]. 978-9-5871-6135-9 [Por13]. 978-9-5871-6326-1 [Por13].

978-9-5871-6490-9 [Por13]. 978-94-007-3931-4 [Gor14a]. 978-958-716-498-5 [Mat15]. 978-958-761-142-7 [Mat15]. 978-981-4307-81-9 [Wal15a].

âge [Str06, SD08].

A.D. [God49]. Aaron [Bri15]. Abiko [Dar04]. Abingdon [Fle20]. Aboriginal [And20]. ’Abraded [Wes96a]. Absence [Vaz23]. Absolute [Gal79]. Abstraction [Pye13]. Abstracts [Ful64]. Abuses [Bar18]. Academia [Por13, GUOB11, Por13]. Academic [FHW75, Gru95]. Académie [Hol03]. Academy [Pra49, Sla00, Byc13, NTT10]. accelerate [Hom06]. Accelerating [Sei83]. accelerator [HV06, KH93, Wit04]. Accelerators [Low06, Mun97, de 06]. Acceptance [Fra82, Fra83]. Access [Did18]. Accessibility [FKW24]. Accidents [Bur24]. Acclimatization [Bim21]. Account [Coh50]. Accounts [Löw90c]. Accum [Bro48]. Accuracy [Par09]. accusation [Ric17a]. Acetone [GD59, Gor62]. Acid [Ret65]. Acquired [Bro48]. Across [LGSL^{+}22, Sur22, Tal16]. Act [Han15, Ter14, Hen16]. Action [Lal13, Rad13, Hol62]. Active [Far50]. activities [Tur06]. Activity [Lal13]. Acts [Ben12]. Actual [Den17]. Adam [Rai10, Rud08]. Adaptation [For71]. Adas [Har09]. Addicted [Rog12]. Addiction [Yu23]. address [Abi00]. Adelene [Bal15]. Adhering [Hen19]. Adjuvant [Ras93]. Administration [Mod08]. Adrienne [Ham10, Rai10]. Advanced [Wes08b]. Advantages [Wri14]. Adventure [Mil49]. Adventures [SD08, HS06]. Adversaries [Ful62]. adversary [Wer97]. Adviser [Eck90]. Advisory [Cit15, Ham10, McM22, Wan08]. Aeronaut [Duv67]. Aeronautics [Mod08]. Aerospace [Lau14, Wes12]. Affinities [Hol62]. Affinity [WK67]. Affirmative [Rad13]. Africa [Nov15, CR09, Cro07, Fin13, Har09, Kel07, Til11]. Africans [Wel12, Hec12]. After [Kli13, Tur05, Bro76, Cas01, Csi24, Dah86, Dev00, Kai02, KR18, Pre11, Ric17a, Sch03, Sch12a, Set11, Har09]. Aftereffects [Dör91]. Afterword [Nye17]. Again [Jon18]. Against [Wol13a, Cra00, Cul10, Ian93, Wer97]. Age [Bel18, Gil08, Lin20, Mar21, Mun21, Nee87, Nov15, Nye09, Rai10, Ric12, Rot13, Bar09, Coe07, Cur11, Ham08, McC08, Pri12, Rud08, Sec14, Yi07, Bal15, Bra15, Cah04, Coo07a, Cre15, Ham10, Lon08, Ric17b, San10, Set12]. Agency [Rus11b]. Agent [Bra15, Mar12]. Aging [Mor21a, Par10]. Ago [Eri23]. Agricultural [Mil61, Phi10, Har05]. Agriculture [Cur16, Tur08]. Agronomy [Coo19]. Aguas [Her08, San10]. AI [BV24a, How24]. AIDS [Lau14, Tie13]. Aiken [Rus15]. Aileen [Bal15]. Air [Moo23, Cha01]. Airport [Güt24]. Akademie [Pra49]. Akera [Mod08]. Akkadian [Lev60b]. al [BG10, Dei04, Por13, Lev64b, LAKAK66, Lev62, Lev64b]. al-Kindī [Lev62]. Al-Samarqandi [LAKAK66]. Al-Sumum [Lev64b]. Al-Wahshiya [Lev64b]. Alan [Str10]. Alarm [Rid84]. Albatross [Sla12b]. Albert [Leg15, Hun10, Sch02, Van07a]. Alberto [Por13]. Alberuni [Lev61a]. Alchemical [Hir50, Tay48]. Alchemie [GG50, Gan53]. Alchemist [For49, Gar09, Mil49, Kas05]. Alchemists [DW49]. Alchemy [For53, GG50, Gan53, Gar09, Ihd64, Lev61a, Num07, Rea48, Mor05, Gar09]. Alcobaca [GP17]. Aldo [Koh11]. Alexander [Sch12a, Jan11, Kre11, Lau14, Phi10, Rup08, Bou12]. Alexis [Mat15]. Alfred [SG99, Bor23, For70, Kau67a, Kau67b, Kau67c]. Algeria [Coo19]. Algiers [Pye90a]. Algol [Nof21]. Algorithmic [Mei23]. Algorithmically [Tat19]. Aliphatic [Rei53]. Alix [Lon08].

Alkaloid [Jac14]. Alkaloids [Les81]. Allende [Abb13]. allgemeine [Rhe50]. Allibone [Sei86a]. Allied [Cas94, SR02b]. Allocation [McM22]. allosteric [CG96]. Alloys [Fes62]. Almeling [Kli13]. Almost [Rus13, Wol13a]. Alpinism [FH19]. Alwin [Far66]. Amanda [Ric12]. Ambergris [God49]. America [Bou12, Har09, Nov15, Rad13, Sch12a, Wol13a, Ada06, AM64, Bar16, Coh50, Cro20, Cul10, Ede59, Fab10, Ham10, Hec14, Joh08, Kli53, Koh10, Luc08, Mir21, Mun03, Ndi07, Oli19, Ros82, Ros95, Ros12b, Rus13, Sch09, Shi19, Str08, Wan08, Wat07, Wis12b, Wol13b, Zun12]. American [Bou12, Bri15, Dea08, Fel13, Joh08, Lau14, Lon08, Rai10, Set12, Wol13a, Bac65, BO06, Cah04, Coo12, Dim14, FH11, Gho25, Gol00, Goo03, Hed08, Her12, Jan11, Kai02, Kev90, Koh08, Kri08, Lev00, LK24, Mar15, McM22, Mil50, Mil90, Oli13, Par12, Rai00, Rei49, Swa17, Tur13, Wan95a, Wes07, Wol10]. Americanization [Ass92a]. Americans [GS94, May05, Rai10]. Amherst [Bra15]. among [Lei61, Rob20, Wil82]. Analogues [Jac14]. Analogy [Set16, Whe81]. Analyses [Deb62]. Analysis [Sco50, Sig92, MD03]. Analyst [Gue22]. Analytical [Fri77]. Ananda [Kev94]. Anatomy [Nov15, Cur11]. Ancestry [FR20]. Ancient [Kin11, BJ10, Fes62, Fig66, Lei61, Lev59, LB59, Lev60a]. Andean [Bim21]. Anderson [Kle19]. Andover [HS89]. Andrew [Bou12, Nov15]. Androids [DGH^{+}24]. Anecdote [GM12]. Angel [Gor17, Por13]. Angela [Lin09]. Angeles [Rog12]. Anglo [Goo03]. Animal [Bol22, Güt24, Low23, OD23, Tea23, Ven24, Kli53]. Animals [Lon08, Nyh17, Tyr18, Gol00]. Anja [Wal15a]. Anker [Bra15]. Ann [Cho12, Nov15]. Anna [Bal15, Har09]. Anne [Cro20]. Annie [Kli13]. Anomalies [Jen84, Rue92]. Anomalous [For70]. Answers [Bub19]. Antebellum [Cas09a]. Antecedents [Gue59]. Anthropocene [Wri18]. Anti [Lal13]. Anti-Relativity [Lal13]. Antibody [Dei21]. anticipation [Dar01]. Antiphlogistic [Par59]. Anton [Tai17]. Antonio [Lon08]. Antony [Sta11]. Ants [Por13, Sle25, WG10]. Anxious [Cre01]. Any [GB89]. Apostolos [Oni15]. app [Sch12a]. Appareils [Dau50]. Application [Hod87, RH00]. Applications [JJ15]. Applied [Fel85, Joh08, Kle19, SS18, Sla12b, Ten22, vS67, SG99]. Appreciation [Wit21]. Approach [LVGS^{+}22]. Approaches [Sie78]. Appropriation [Leg15, Sim11]. Apps [Abb13, Sch12a]. Aqrābādhīn [Lev62]. Aquacultural [Eri23]. Arabic [Lev61b, Lev62, Lev67a, Lev67b]. Architectural [Les08]. architecture [Sod94]. Archival [Sem23]. Archive [Mit23]. Archives [Shi23]. Archivo [Por13, BGC08]. Arctic [Sow17]. Argentina [Fes62, HV06, Lar17]. Argonne [Wes10]. Argument [Tay48, Kra07]. Arguments [Kat19]. ARGUS [Mun12]. Aristotelian [Hor66]. Armadillos [Iri23]. Armand [Kim14]. Arming [Ham13, Bra15]. Armstrong [Dei64]. Army [Fle00a]. Arnold [Fle20, Sta11, Set10]. Arnost [Gor17]. Aron [Kat09]. Art [Dör21, Han15, Ric12, For07, Hun13, Ree09]. Arthur [Hag14, Ayr12, Fef89, Ram95]. article [Dar04, Kat05]. Artifact [Kin11, Szc23, BJ10]. Artifacts [Cho12, Kin11]. Artificial [Huf94, Lin09, Liu19, Sme66, Ris07]. Artisan [Lon11, Bro14]. Artisan/practitioners [Lon11, Bro14]. Arts [Pye13]. Ash [Bro14]. Asia [Cul10, Han12, Tin24, Wol13a]. Asian [Wit21, Wan24]. Asif [Sch12a]. Assessments [Shi19]. Assistant [Hil49]. Associating [Pet16]. Association [Bur82, Esp15]. Astroculture [Lau14, Gep11]. Astrologer [Gar09, Kas05]. Astronaut [Bim21, Her12, Lau14].

Astronomer [Cre83]. Astronomers [FM24, Huf81, Dev00]. Astronomy [Ahn22, Bad18, Bal15, Dav86, Het87, Swa17, Dev00, Hen14, Mun97, Mun03, Mun13, Fle20, Lau14]. Asylum [Por18]. Asymmetries [Mat15]. Atlantic [Nov15, CR09, Sur22]. atmosphère [Coe10, Loc08]. Atmospheres [Dun23]. Atmospheric [Wal15b, Dar98a, Hey06]. Atoll [Rai00]. Atom [Cas79, HK69, Kri08, Ser77, Bru03, HA84, Sei86a]. Atomic [Bra15, Fef92, FH11, Ham10, Hof17, Hoo49, Kra79, Lev61c, Lin20, MSD21, Mun21, Oni15, Roc78, Rot13, San10, Sie78, Sla12a, Tho22, vSvS66, vS67, FV13, Gor09, Pri12, Yi07, Cre15, Goo69, Ric17b, Sei86b, Sei01, Set12, Fel13]. atomique [Str06, SD08]. Atomism [Rüg88]. Atoms [Cur16, Fef92, Roc78, Tre75]. Atop [Gil08]. Atsushi [Mod08]. attendants [Tie13, Lau14]. Attitudes [Rue92]. Attractors [Kel09]. Auckland [Bra15]. Auctions [AB23]. Audacious [Pye82]. Audiences [NG12]. Audra [Hec14]. Auerbach [Ric17b]. Auguste [deM53]. Augustine [Ham10]. Augusto [CK90]. Aurora [Kra09b]. Auroral [Ame22]. Austin [Lon08]. Australia [Hom06, Mun97]. Australian [Muk20]. Autarchy [SW10]. Autarkic [NR13]. Autarky [SW10]. Autarky/Autarchy [SW10]. Author [Sch09]. Authority [Gar09, Hec11, Hec14, Num07, Sch02]. Authorship [Csi24]. Automatic [Lav14]. Automation [Moo24]. Automobility [Eis16]. Autònoma [Mat15]. Avery [Ams93, Dei04]. Aviation [Güt24]. Avogadro [Mun67]. Awareness [Rot13]. Axel [Bou12]. Axis [Hei76]. Ayres [Hag14].

B [And20, Far64b, Fra77, Har09, Smo09]. B. [Sol65]. Babcock [Smo09]. Babylonia [Lev60a]. Bacciagaluppi [Sta11]. Bache [Bou12, Jan11]. Back [Ano12a, Ano12b].

Background [Son53]. Backwards [HS18]. Bagory [Nev62]. Balancing [Stu75]. Ballistic [Hop96]. Ballot [AvL16]. Baltimore [Coe10, Jac09, Joh08, Kim14, Lau14, Nov15, Por13, Rad13, Str08, Hen93]. band [Cas07]. Bang [BK21, Zak13, Wil99, Bro07, Kra14, Oli19]. Barcelona [Mat15]. Barnard [Cas09a]. Barometric [Fel85]. Barrera [Lon08]. Barrera-Osorio [Lon08]. Barrow [Ric12]. Based [HS89, Lec92, Nee87, SSD15, Shi19, Sun12, Kri05]. Basic [Kuz66]. Basis [For87, All01]. Baton [Bra15]. Batteries [Eis16]. Battery [Pan90a]. Battle [Wol13a, Cul10]. battlefield [Bru03]. Bauhaus [Bra15, Ank10]. Bay [Eri23, Gil23]. Bayerischen [Pra49]. Bayern [Pra49]. Bayle [Sar50]. Be [How11, Mas22, Par22, Ran22, Aas95, Dar84]. Beam [Cre01, Rig83]. Beauty [Les08]. Became [Jon18, MN18, Bru07]. Become [AL20]. becomes [Col03]. Becoming [Kai10, Pan24, Cho12]. Before [Kat19, Kli13, Rad13, Rai10, Bro76, Csi24, Han79, Hoo49, Low23, Pau11, Rud08, Sch03]. Begin [Kul25]. Beginning [Coh50]. Beginnings [Dim14, Van49, GS94, Hug11]. Behavior [Bri15, Pan14]. Behavioral [Güt24, Sea25]. Behaviorism [Arb19]. Behind [For87]. Being [Hec12, Vel22, Wel12]. Belgian [Bro23]. Belhoste [Kim14]. Belknap [Har09]. Bell [Rus81]. Bellevue [Shi93]. Bellies [Her24]. Benedict [Nes13]. Benioff [Dör21]. Benjamin [Mil53]. Benzer [SD08, HS06]. bequest [Bru03]. Bergman [Sch67]. Bergson [Som20]. Berkeley [Jac09, Kli13, Lau14, Nov15, Rog12, Zak13, CG96, Smo09, Sun12, Wes96a]. Berlin [Cha94, Sch81, Ven24, nye07]. Bernal [Por13]. Bernard [Huf94]. Berthollet [Hol62, Kap65, Par59]. Berzelius [Sol61]. Bethe [Sch12b, Wal15a]. Betrachtung [GG50]. Between [Bri15, Cur16, Dun23, Par09, BJ10, DIR91, Fre81, GG88, God49, Har05, Hop96, Ito02, Jac09, Kim02, Kin11, Lal13, Phi10, Rus86, Sch02, Yi07, Orl04]. Betwixt [Bri15]. Beyond [Cas09b, Ito24, MJ15, Ved19, Wan20, Yu23, Wan99b, Sta11]. bias [Van07a]. bibl [Sch12a, Zak13]. biblio [Lau14]. Bibliography [Ano86b, Dei48, Low86a, Low86b, Low87, Low88b, Low88a, Low89b, Low89a, Low90b, Low90a, Low91, Rid91, Rid92, Rid93, Bey00, Bey01, Bey02, Gro03a, Gro03b, Gro04a, Gro04b, Gro05a, Gro05b, Pan93, Pan94a, Pan94b, Pan95, Wes96b, Wes97a, Wes97b, Wes98a, Wes98b, Wes00, You06b, You06a, You07a, You07b]. Bicycle [Gib12]. Big [ABO10, Bro23, Bro07, Col10, Hal11, Sei86b, Spi11, Zak13, Chr11, Col03, HV06, Kev97, Wan95b, Wil99, BK21, Bro07, Hag14, Hed08, Kev97, Kra14, Oli19, Rai10, Zak13]. Bikini [Rai00]. Binary [Ha15]. Bio [Far50, Iri23]. Bio-Active [Far50]. Bio-Materials [Iri23]. Bioarchaeology [Lan23]. Biobanks [Lin20]. Biochemical [Koh91, Lei61]. Biochemistry [Koh77, Rei17b]. biographical [Sod94]. Biographies [Ng19, Nye09, Wal15a]. Biography [Ber19, Fle20, Sol61, BG13, Oni15, Sta08]. Bioinvasion [Luk24]. Biological [ABO10, Cre20, Gau10, Jia16, NV17, Phi10, Tho17, VD10, All01, Gas96, Nyh09, Vet04]. biologie [Str06, SD08]. Biologies [AL20]. Biologist [Gra25]. Biology [ABO10, Cho12, Hog15, Kev09, Kra09a, Löw90c, Mil10a, Phi10, Rei17b, SD08, Zak13, dC18, AA95, Ben01, Doi09, End07, Mil10b, MWR07, RH00, Rus11a, Kev09, Ric12, Rey21]. Biomedical [Abb13, Nov12, Dim14]. Biomedicine [SSD15, Wu20]. Biophysics [Cul74, Slo14, Wes96a]. Biopolitics [Kli13, Sta19, Tak12]. Biosphere [RO20]. Biot [Fra77]. Biotech [Vet06, Hug11, Dim14, Mod08]. Biotechnology [Hol18, Kev94]. Birds [Cam20]. Birdsell [And20]. Birmingham [Bol22]. Birth [BH08, Bra15, Cre08, Kar77, Kli13, Kra09a, Lee16, Shi93, Ham13, Pau11]. Bizonal [Cas96]. Black [DGH^{+}24, Joh08, Lev64c, Nye74, SS23, Lon08, Wes07, Nur21, Par60]. Blackness [Nov15, Cur11]. Blackwell [Zak13]. Blanc [Gil08]. Bleichmar [Por13]. Blending [Mac24]. Blight [Tur08, Tur05]. Bloch [Kri01]. Blood [Gol00, Sch12a, Kre11]. Bloodworlds [Muk20]. Bloomington [Wol13a]. Bloomsbury [Oni15]. Blue [LU13, Les01, Wes12, Lau14]. Blurry [Csi24]. Bodies [Wu20, Tak12, Kli13]. Body [Bau21, Fel13, Low23]. Bogdanov [Sch12a, Kre11]. Bogotá [Mat15, Por13, BGC08, Por13]. Bohm [Fre05, Koj02]. Bohr [Aas99, Bro71, HK69, Kle70, Kra79, RH00]. Bologna [Hei91]. Bomb [Ber88, Mod08, Cas09b, FV13, Goo03, Hop96, Oni15, Sta11]. Bombs [Joh08, Ndi07]. Bond [Koh71]. Bonds [Sur22]. Bone [Hed08, Kra09a, Rai10]. Bones [Rai10]. Book [Abb13, And09, Bal15, Bou12, Bra15, Bri15, Bro14, Cho12, Coe10, Cro20, Dea08, Dim14, Dör12, Fel13, Fin13, Fle20, Gar09, Gil08, Gol13, Gor14a, Hag14, Ham10, Han15, Har09, Hec14, How12, Jac09, Joh08, Kev09, Kim14, Kin11, Kli13, Lau14, Leg15, Lin09, Lon08, Mat15, Mit21, Mod08, Mor21a, Mor09, Mor21b, Nov15, Nye09, Oni15, Phi10, Pic14, Por13, Rad13, Rad12, Rai10, Rey21, Ric12, Ris10, Rog12, San10, Sch12a, Sei86a, Seo21, Sta11, Sta08, Str08, SD08, Ved20, Wal15a, Wel12, Whe82, Wit21, Wol13a, Wu20, Zak13, Lev64b]. Books [Bal15, Oni15, BG13, Hop96, Sec14]. Booth [Mil66]. Border [Mit21]. Borderland [Bug89, Ser83]. Borderlands [Moo16]. Borders [Cam20, Mac24]. Borealis [Kra09b]. Borrello [Ric12]. Bosio [Por13, BGC08, MOBG09]. Botanical [Nov15, Por13, Ble12, CR09]. Boundaries [Dim14]. Bounding [Pye90b]. Bowl [Cha20]. Bowler [Sch12a]. Bowles [Mod08, Pan90b]. Box [Har24]. Boyden [Str10]. Boyle [Bue53, Deb62, Hoo49, Ihd64, Sar50]. brain [Sma08, Zak13]. Brakes [Cho12, Joh09]. Branch [Gue76]. Brand [Wu20]. Brave [Kli13]. Brazilian [Bur06, RM06, TBV20]. Breakdown [BC91]. Breaking [Sur22]. Breast [Jac09, Par07]. Bred [Tyr15]. Breeding [Cro20, Har10, Tyr15]. Brenngläser [Win49]. Brent [Wol13a]. Brian [Lon08]. Bricolage [LLV^{+}22]. Bridges [GSLV^{+}22b]. Bridging [Hog15]. Brief [Coh50]. Bringing [Les01, Wri14]. Britain [Bow09, Can75, CS78, Dor75, Gar78, Ott08, Rid84, Sch12a, Svi76, WS86]. British [Coe10, Har09, Wal15a, AvL16, AB98, Bur82, CH16, Coe10, Cro07, EG80, Gav90, Gol70, Gol07, Lau12, Mor71, Mul08, Wis82, Wis87]. Broad [Pye11]. Brock [Cho12]. Brodie [Far64b]. Broglie [RF69]. Bronn [Phi10, Gli08]. Bronze [LB59]. Brook [Mod08]. Brookhaven [Cre01, Nee83]. Brown [Zak13]. Browne [Dei48]. Brownian [Ian93]. Bruce [Gar09]. Bruno [SD08, Kim14]. Brunswick [Kli13, Mod08, San10]. Brussels [Aub20]. Bubble [KP86]. Buchwald [Kin11]. Buckland [Bal15]. Bud [Str08]. Buen [NG12]. Build [GB89, Mun03, Par01]. Building [Bou12, Nes13, Par07, Rio01, SW10, Sim19, Tur82, Wu20, Jan11, Jac09]. Built [Sta12]. Bulgaria [Tch67]. Bulletin [Sla12a]. Bullets [Str08]. Bunsen [Rhe50]. Bunsens [Rhe50]. Burden [Fel13]. Burrow [Wei22]. Bursting [Gil08]. Bush [Mir21, Rei87]. Business [Bal15, Rus15, Fyf12]. businessman [Tur06]. Butterflies [Sak24]. Buys [AvL16]. Byrne [Oni15].

C

[Bor23, Cho12, Coe10, Dei48, Dim14, Far64b, Hag14, Har09, Kim14, Lau14, Sta11, SD08]. CA [Wu20, Kra09a]. Calculating [Ake07, Tat16, Mod08]. Calculus [Far64b]. Caldwell [Kli53]. California [Jac09, Kli13, Lau14, Nov15, Rog12, Wu20, Zak13, Goo84, Kar77, Lau14, Wes12, Zul09]. Call [Hui22]. Called [Mar24a]. Calm [Dun23]. Calomel [Urd48]. Caloric [Mel83, Par10]. Caltech [McC04]. Cambridge [And09, Coe10, Fle20, Gar09, Ham10, Har09, Jac09, Joh08, Kev09, Kli13, Lau14, Lon08, Mod08, Mor09, Nye09, Phi10, Ris10, Sch12a, Ved20, Wal15a, Wol13a, Wu20, Zak13, Bec80, Hen84, Hin03, Rob80, Sei86a, ST70, Wil82, Wis87]. Cameralism [ST18]. Camouflage [Gol13, She12]. Camp [Ben01]. Campbell [Cre83]. Campos [Cro20]. Can [Dar84, How24, LVGS^{+}22]. Canada [Wal15b]. Canadian [Cas01, Sow17]. Canal [Bro14, Van07a, Muk09]. Canals [Sim19]. Cancer [Cre15, Fef92, Jac09, San10, San15, Sem08, Kut09, Leo09, Par07]. Candles [Low23]. Cannizzaro [vSvS66]. Capacities [Mil90]. Capital [SD25]. Capture [Lae22]. Capturing [Ame22]. Carbon [Hen19]. Care [Bar15, Col24, Par07, Jac09]. Career [And20, Mil90, Ams93, LKB05]. Careers [Bri15]. Caribbean [Coe10, Mul08]. Carl [Cha11, Hof24, Nyh17, Ros22, dM48]. Carnegie [Wol13a, Par12, Ser83]. Carney [Nov15]. Carson [Hec11, Ric17b]. Cartographies [Bur24]. Cartography [McG18]. Case [Ahn22, Can75, Hac86, Han79, Joh13, LU13, Löw90c, Mel83, Nye12, Sut81, Tab22, Wis79, Dmi04, Wan99b]. Cases [Lin09, CLW07]. Cassidy [Hop96, Sta11]. Catalyse [Lem49]. Catalysis [Kuz66, Lem49]. Catalytic [SS18]. Catastrophic [Bra15, Ham13]. Catching [Ter14, Han15]. Category [Sch12a]. Catherine [Ham10]. Causality [For71]. Causes [Tur71]. Cavendish [Pic14, Sar10b]. CDF [Wes01a]. Celestino [BG10, GUOB11, Por13, WG10]. Cell [Ha15, Kra09a, Lan19, Liu19, SSD15, Vie17]. Cell-Based [SSD15]. Cells [Kli13, Alm11]. Cellular [Rey21]. Censorship [McM22]. Census [Von18]. Center [Bou18, HHH15a, HHH15b, MC13]. Centre [De 02]. Centuries [Dol76]. Century [Bal15, Bar16, Cro20, Dav86, Dör12, Lau14, MJ15, Mor09, Nes13, Nye74, Phi10, Ren09, Rog12, San10, Sch12a, Shi86, Shi87, Str08, Wis87, Wol13a, Bow09, Buc13, Dar02, Gep11, Hol04, Ito24, Jac12, Jan11, Par12, Rhe10a, Sch08a, Ter14, Wes12, Bad18, Bor23, Bou12, Bro69, CK90, Cat98, Dor75, Dör91, Dör06, DW49, Far50, Far64a, Gar78, Gis11, Gol00, Hac86, Han15, HG71, Hol18, Huf71, Jac06, Kat09, Ker13, Lau14, Lev66, Lev67a, NV17, Ols08, Ore00, Pal89, Pan20, Por18, Pye82, Rap60, Rea48, Rob20, Sch09, ST18, Son53, Tea23, Ven24, Von18, Wal15a, Wes20, Wu20, Pic14, Sta16]. CERN [KP86, Kri01]. Certainty [Bug92]. Chakrabarty [Kev94]. Challenges [Wit21]. Chamber [Ber19]. Chambers [Bal15, KP86, Fyf12]. Chandler [Bac65]. Chandra [Bro14]. Change [Cha93, Cho12, Hec14, Huf71, Les81, Mac24, NTT10, Shi79, Doi09, Dör06, Wis12b]. Changes [Kel12]. Changing [DM06, NTT10]. Chapters [Chu67]. Character [Dav86]. Charity [Wol13a, Rus13]. charlatans [Gol00]. Charles [Dea08, Byc13, Kai11, Kli53, Mor09, Pye74, Rus15, SG99, Smi06, Tho17]. Charlotte [Ric17b]. Chatto [Fle20]. Chemical [Bac65, Cof06, Coh50, Ede59, Far61, Far65, Far64b, Gue59, Hir50, Hol62, Hoo49, Jai66, Koh75, Lev61b, Lev64a, Lev66, Luk64, Mau12, Mil50, Mil59, Par48, Pra49, Rei53, Roc78, Sie59, Tho22, Urd48, Mil50]. Chemico [Abi91]. Chemico-Thermal [Abi91]. chemische [Pra49]. chemischen [Win49]. Chemist [Duv67, Mil53, Rap60, Rei49, Rob80, Sar50, Sch59, Siv62].

Chemistry [Bro75, Byk65, Cha93, Coh50, Dei48, Dol76, Don67, Far62, Far66, Fig66, Fre81, Gar09, Gue76, Hal62, HS18, Hof17, Hol03, Hor66, Huf71, Joh13, Kli48, Kli59, LK50, Lev60a, Lev62, Lev64b, LAKAK66, Low23, Par09, Pet16, Rây59, Rhe50, Rob76, Rob80, Sch90, Sem66, Son53, Tch67, Tei10, Tur82, Van49, deM53, Cha94, GS94, Hol04, KW96, Mor05, SG99, PVP50, Rob76, Kli53, Par60, PVP50]. Chemists [DW49, Gav90, Gol60, Joh13, Mil61]. Chenchus [Muk20]. Chesapeake [Gil23]. Chicago [And09, Bal15, Bro14, Coe10, Cro20, Fin13, Gar09, Gil08, Ham10, Har09, Kim14, Lin09, Lon08, Mor09, Nov15, Nye09, Phi10, Por13, Rai10, Ris10, Rog12, San10, Sch12a, Slo14, Wal15a, Fef92, Slo14]. Chichester [Zak13]. Chicken [OCS08]. Chikako [Kli13]. Child [Ham10, Mun11, Rog12, Gre08]. Children [Sch12a, Koh10]. Chile [Abb13]. Chimie [PVP50]. Chimistes [Dau48]. China [Bia22, Gre08, Hu04, Jia16, Mul12, Rog12, Sch08a, Wan99b, Wil99]. Chinese [Rog12, Lam11, LK24, Luk24]. Chlorate [Dav48]. Choh [Zul09]. Choice [Het82, Koz91, KP86, Ric12, Mil10b]. Choosing [Rei95]. Chrestomathy [DGH^{+}24]. Chris [Mor09]. Christian [Zak13, Lev66]. Christiane [Kel96]. Christoph [Wal15a]. Christophe [Cho12, Joh08]. Christophle [Nev65]. Christos [Oni15]. Chromatin [Ha15]. Chrome [Lav14]. Chromosome [dC15]. Chromosomes [San15]. Chronology [Dru20]. Chymie [Nev65]. Ciencias [CA14]. científica [Por13, BGC08]. Científicas [San10, Por13, GUOB11]. cinema [Kir11, Hec14]. circa [FHW75, SF13, SF15]. Circulation [CR16, Leg15, Mat15, Sim11]. Circumstances [Dah86]. Citing [Ala23]. Citrus [Coo19]. City [Tat16]. civic [Fre97]. civil [Hol99, Low99]. Civilianization [Mun12]. Civilizing [Har09, Ada06]. Claiming [Sow17]. claims [Huf07]. Claire [OL50, Sol65]. Clara [McC64]. Clarendon [Gar09, Ris10, Mor92]. Clarke [Sch90]. CLAS [Wes01a]. Classed [Szc23]. Classical [Kuz71, Sie78, Tal16]. Classical-Electromagnetic [Sie78]. Classification [Dea14, Koh08, Rog12, Mul12]. Classroom [Bay24, Ham24]. Clausius [Gar70, Kle69, Pet16, Yag84]. Clément [Lem49]. Clements [Mas06]. Climate [Cha20, Coe10, Dör20, Ger21, Hen16, MN18, NTT10, Wal15b, DM06, Dör06, Fle06, Gol07]. Climatology [Har09, Jen06, Mas06]. Clines [And20]. Clive [Par10]. Closing [Mat15]. cloth [And09, Bra15, Coe10, Cro20, Fin13, Gar09, Gil08, Ham10, Har09, Jac09, Joh08, Kev09, Lin09, Mod08, Mor09, Nov15, Nye09, Phi10, Rai10, San10, Sch12a, Sta08, Ved20, Wu20]. Cloud [Ber19, Ham10, Gor09]. Clouds [Car20, Von01]. CO [Wea07]. Coal [Bou12, Luc08]. Coated [Eri23]. Coats [Hec14, Kir11]. Cockburn [Sei86a]. Cockcroft [HA84, Sei86a]. Coded [Mei23]. Coen [Nye09]. Coins [Lev67b]. Cold [Ake07, Ham10, Hec14, Mod08, San10, Wol13a, AB98, Aro17, Bru03, Cul10, Dör20, Fre05, Hec14, Hou97, Kai02, Kev90, Kri99, Kri05, Lee16, Leo09, Lev13, LK24, Mac24, Mar22, MK20, Mun03, Oli19, SR02a, Sei01, Shi16, Shi19, Sla12a, Sla12b, Tat19, Tur20, VD10, Wal00, Wan95b, Wan99b, Wan08, Wan12, Wan20, Wea97, Wil15, Wol13b]. Colding [Can97]. Colegio [MOBG09, Por13]. Coleman [Hag14]. Colin [Kim14]. Collaborating [Wes01a]. Collaboration [Gue76, Nie89, Nie91, Pye76, Oli09]. collar [Les01]. colleague [BH07]. Collecting [NVHB24]. Collection [Cas09a, Sig92]. Collections [Sun12]. Collections-Based [Sun12]. Collective [Sta19, Can05, Koj02]. collectivism [Koj99b]. Collector [Rus15]. Collectors [Nov15, Fab10]. College [Rob76, Bac65]. Colleges [Phi10, Har05]. Collider [Cre08, Par01, Rio01]. Collins [Oni15]. Colombia [Mat15]. Colonial [Coo19, Don67, Har09, Kel07, LP24, Per25, Wu20, Cro07, Jen06, Cro07, Har09, Pan24]. Colonies [Dim14, McC3 ]. Colonizers [Har09, Jen06]. color [And09, Lau14]. Colorado [Rus15]. Colored [Dav48]. Colorful [Von01]. Columbia [San10, Wol13a]. Columbian [Mil59]. Combinatorics [Dar88, Dar91]. Combining [Rob24]. ComBio [FKW24]. come [Mun03]. Comes [Goo84]. Comfort [Bel18]. Coming [Hop96, Mul12, Yi07, Rog12]. Comment [Lyn12, Kat05]. Commentaries [Lyn12]. Commentary [ST70]. Commerce [Lon08, Coo07a]. Commercial [Kin13, Lev64a]. Commission [Sei86b, Sei01]. Committee [Ham10, McM22, AB98, Cit15, Wan08]. Commodification [Kin13, Sak24]. Common [Ten19]. Commonality [Moo24]. Commons [Ger21]. Communicating [Sim11, Leg15]. Communication [Set12]. Community [Dim14, Koh75, Kra85, Lar17, Lau14, Lek23, Nes13, Shi19, Tei10, Mal00, Mod12, Mun13, Rei95, Shi93]. Community-Building [Nes13]. Comparative [Jac09, Str10, Dar94, Hin03, Par07]. Competing [Hec14, Wol13b, Wis12b, Hec14]. Competition [Lek23, Rob20]. complementarity [Jac07, RH00]. Complex [Lek23]. Complexity [Kel09]. Composition [Sme66]. Compositions [Dav48]. Compton [DR89, Lec92]. Computational [Par09]. Computer [DiM24, Koi07, Mod08, Tat19]. Computers [Mod08, Ake07]. Computing [Abb13, Dim14, HS18, MN18, Sei08, Nov12].

Comteanism [Tho17]. Concept [Bro76, Hei76, Hir69, SF13, SF15, Ten19, Wis79]. Conception [Esp15]. Concepts [Elk70, HG71, Lei61, Par48, FG98, FG99]. Conceptual [Les81]. Concerning [Kus18, Bro48, Far61]. Concrete [Dör12, Rhe10a]. Conference [HS89, Sta11, BV09]. confession [Ric17a]. Conflict [Ung19, Ved19, Kim02]. Conflicting [Vie17]. Confronting [Ric12, Bar09]. Confronts [Cre83]. confusion [KW96]. Congress [Aub20, dM48]. Conjectures [Sie64]. Connecting [Tai17]. Connection [Gre83]. Conquest [Swa17]. Consciousness [Kai11]. Consejo [San10]. Consensus [NTT10, Shi19]. Consensus-Based [Shi19]. Consequences [Pye11]. conservado [BGC08, Por13]. Conservation [Mar16, Wis79, Dar01, Mas06]. Consolidating [Ber19]. Consolidation [Dor75]. Constant [Nep60]. constants [Lel05]. Constructing [Hec11, Wal15b]. Construction [Fin21, Par10, Rog12, Tab22, Wal15a, BV01, Can05, HV06, Lam11, Nye11]. constructivist [Pan97]. Constructs [Kli13, Tak12]. Consultants [Wil15]. Consulting [Bou12, Luc08]. Contact [Hod81]. Contemporaries [Ful67]. Contemporary [Löw90c]. Contentious [Ric12, Bor10]. Contents [Ano05d, Ano07e, Ano07f, Ano08j, Ano08k, Ano08l, Ano08m, Ano09j, Ano09k, Ano09l, Ano09m, Ano10l, Ano10i, Ano10j, Ano10k, Ano11e, Ano11f, Ano12k, Ano12l, Ano12m, Ano12n, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano14i, Ano14j, Ano14k, Ano14l, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano06d, Ano06e]. Contested [Kut09, San10]. Context [Can78, DiM24, Fri89, Rob76, Tur71]. Contextualists [Ste88]. Continental [Ore88]. Continuing [Ung19]. Continuities [Gal23]. Contraception [Kli13, Pre11].

Contraceptive [Kli13, Tak12]. Contributions [Dei48, Hei91, Kli59, Rei53, WK67, Orl98, SG99]. Control [Mir21]. Controlling [Cas94, Cas96]. Controversies [DIR91, Hof24]. Controversy [Bri15, DR89, Ede48, FH19, Kau60, Kin11, Kli53, Ric12, Sch64, SF13, SF15, BJ10, Pan14, Ree09]. Conway [How12]. Cook [Kim14, Lon08, Tak11]. Cool [Moo23]. Cooper [Lon08]. Cooperation [Rob20]. Cooperative [Bug89, Kar77]. Coordination [Kau67b]. Copper [Fes62]. Core [Cas79]. Cork [GP17]. Cornell [MC13]. corners [Wes96a]. coronagraph [Huf94]. Corporate [Cam10, Gol13, Har11]. Corrections [SF15]. correspondiente [MOBG09, Por13]. Corresponding [Gav90]. Corvallis [Bro14]. Cosmic [Cas81, DR89, DIR91, Mon08]. Cosmo [Kra82]. Cosmo-Physics [Kra82]. Cosmology [Kra82, Kra07, Oli19, Rüg88]. Cosmos [Pic14]. Coulson [SG99]. Coulston [Dea08]. Council [Bug89]. Countercultural [Mod08, Vet06]. Country [Hal11]. Courant [Shi16]. Courtier [Mil49]. courtroom [Gol00]. Cover [Ano05a, Ano06a, Ano07a, Ano07b, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13b, Ano13c, Ano13a, Ano13d, Ano13e, Ano14d, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e]. CP [Fra83]. Cradle [Rai10, Hed08]. Crafting [Set10, Sta11]. Craspedacusta [Luk24]. Creager [Lin09]. created [Abi00]. Creation [AvL16, Bro76, Fri77, Ung19, De 02, Kra07, Kri01]. Credit [McM22]. Crepis [Smo09]. crime [All01]. Crisis [Ger21, Tur24, Wis12a]. Critic [Sem08]. Critical [Dry19, Ful64, Mar22, Phi66]. criticism [Hu04]. critique [Ram95, Sur02]. Critiques [Kul25]. Crossings [Mit21]. Crossroads [Kev09, Sta11, BV09, Koi07, MWR07, Rai00]. Crow [Bol22]. Crozier [Har09]. crusade [Aas99]. Crystal [MJ15]. Crystals [Szy84]. CT [SD08, Ved20]. Cuban [Fin13, RH10]. Cullather [Wol13a]. Culotta [Pye74]. cultura [HR12, Mat15, Poh11]. Cultural [Arb19, Har09, Ito24, Set11, Edm06]. Culture [Coo12, For71, Kev09, NR13, Phi10, Por13, Ris10, Tat19, Wal15a, Ble12, DeY11, Gau08, Har06, HR12, Koj99a, Lau12, MWR07, Smi06, Mod08, Mor09, Leg15]. cultures [Rio01]. Curator [Dei64]. Cure [Fin13, Rey22, Yu23, RH10]. Curie [Oni15, Red10]. Curing [Jen06, Har09]. Curious [Bri15, Gue61]. Curran [Nov15]. Current [Bro69, Sie86]. Curry [Cro20]. Curtis [Mil66]. Curve [Phi21]. Cybernetic [Abb13, Lee16]. Cybernetics [Sle25]. cycles [Fle06]. cyclotron [Hin03]. cyclotrons [Kim06]. Cynthia [Zak13]. Cyrus [Dim14]. Cytogenetic [San15]. Cytogenetics [Bar15, Hog15].

D [Fin13, Ham10, Lau14, Mod08, San10]. d’Être [Den17]. Dallas [Bou12, Jan11]. Dalton [Goo69]. Daniel [Zak13]. Daniela [Por13]. Dark [Bra15, Gor17]. Darrigol [Abi04]. Darwin [Kev09, Mor09, Bel18, BG13, Dea14, Oni15, Sch08b, Set16, Smi06]. Darwinism [Phi10, Gli08]. Daston [And09]. Data [ABO10, Bou18, Car20, Dry19, Fra81, Gal23, Igo18, Jon18, Lin20, Por18, Rob83, Ros18, Sep18, dCP18, dC18, Con06, Ste16]. Databases [Did18]. Datafication [Von18]. David [Cho12, Fle20, Hag14, Hop96, Lau14, Sei86a, Sta11, Sta08, Zak13, Fre05, Koj02]. Davie [Mil61]. Davis [Dav49, LK50]. Davy [Ful60, Ful62, Ful64, Ful67, Sie59, Sie64, Str64]. Dawn [Bal15, Ham10, Por13, San10, Gor09, Ham08, McC08, Sec14, WG10]. Days [Fel13, FH11]. DC [Mod08, Oni15]. Dead [Sla12b, Fab10]. Deal [Rei87]. Dear [Ris10]. Death [San10, Kev97]. Debate [Can75, Gho25, Het89, Hof17, Kin11, Pri14, SG25, BJ10, Gas96, Wan95a]. Debated [FH19]. Debates [Dol76, Mar15, Ros22, Cat98]. Debating [Hyu19]. Deborah [Gar09, Nye09]. Debt [Lev60a]. decade [Bru03]. Decades [Buc12]. Decay [Mon08, Coe02]. Decision [GB89, Kai10, Cho12]. deCODE [Jac09, For08]. Decolonization [Yi24]. Decomposition [DD59]. Deconstruction [OCS08]. Decorative [Pye13]. Découvrent [Lem49]. Deep [Gal23, Zak13, Sma08]. Deep-Sea [Gal23]. defeated [De 12]. Defence [AB98]. Defense [Gol70, Jac07]. Definition [Str08, Gre07]. Degradation [Ihd64]. Degree [Mun67]. del [De 12, Mat15, MOBG09, Poh11, Por13]. Delbrück [RH00]. Delicate [Lar17, Lek23]. Deluge [Rus16]. Democracy [Bou12, Lew11, Sch02, Wan99a]. Democratic [Bri15, Tur13]. Deng [Rog12, Gre08]. Denial [Sam23]. Dennis [Joh05]. Density [Eri23]. Departments [Kle19]. Dependency [Pot23]. Depression [FKW24]. derrotado [De 12, Mat15]. Describing [Lon08, Ogi06]. Description [Sig92, Kat03]. Design [Cho12, Gol13, Har09, Ada06, Ank10, Har11, Joh09, Bra15]. Designing [Can21, Cho12]. Désormes [Lem49]. DESY [HHH15a, HHH15b]. Detailed [Stu75]. Detection [Tho12]. Determination [vS67]. determinism [Ian93]. Developed [RF69]. Development [Bri15, Far62, Fin13, Fra82, Gil92, Hen16, Kuz66, Len86, Lev61c, Roc78, Sch90, Sei87, Sta12, Gru95, Hin03, Im95, Lev00, Pan14, Til11]. Deverell [Lau14]. Deville [OL50, Sol65]. d’Expérimentation [Dau50]. Diagnosing [Ha15]. Diagram [VD10]. Diagrams [Wri14]. Dialectical [Gor17]. Dialogue [Kle70, Ved20, Ved20]. Dialogues [Koj99a]. Diane [Kin11]. dictadura [HR12, Mat15]. Dictatorship [GS11, Ham10, HR12, Aug07]. Did [FKW24, Huf71, Wis87]. Different [Les08].

Diffraction [Rus81]. Diffusion [Hir50]. Difuentes [J.23]. Digby [Mil49]. Digital [Abb13, Mod08]. Digitizing [Abb13, Dim14, Nov12]. Dilemma [Phi10, Har05]. Dilemmas [Hop96, Roc17, McC04]. Dimoia [Wu20]. Diodes [LU13]. Dioxide [Hen19]. Diplomacy [Cam20, RO20, Rob20, Shi16, TAR^{+}20, Tur20]. Diplomatic [Aub20, Som20]. Dirac [Kra82]. Disagree [Hof24]. Disappearance [Low23]. Discerning [Nic24]. discharges [Lel05]. Discipline [Sil74, Sim16, BV01]. Disciplines [Ten19]. Discontents [Abb13]. Discontinuous [Sta11]. discourse [Ito02]. Discourses [FR20]. Discover [Lem49]. Discovered [Kau67c]. Discoverer [Bor23]. Discovery [Dah84, Dah86, DR89, DIR91, Fef89, Fra83, Hen93, Hod81, Hod93, Les81, Löw90c, Mel80, Rus81, Dmi04, Neb94, Yi07, Can05]. disease [Gre07, Str08]. Diseases [Str08]. Dispersion [Tal16]. Displacement [Sie86]. Disputations [Mar15]. Dispute [Hol78]. Disregard [Dei21]. Disrupting [Hog15]. Dissemination [Sei08]. Dissent [Dol76, MM75]. Dissipation [Dau70]. Dissociate [Pet16]. distance [Yea03]. distances [Wal00]. Distilling [Mor05, Gar09]. Distinguishing [Gol00]. distrust [Abi03]. Divergent [Mil99]. Divided [TP16]. Divides [Tal16]. Divining [Abb13]. Dmitrii [Dmi04]. DNA [Mod08, Ams93, Dei04, Kev09, Sch08b, Yi07]. Do [Kin11]. Doctors [Hop96]. Doctrine [Pye11]. Documentary [Cho12, Gil92, LB10]. documento [BGC08, Por13]. Documentos [Por13, GUOB11]. Does [Wol10]. Dogma [Hog15]. Dogs [Bol22]. Doing [Cho12, Rob24]. Dolores [Ham10]. Domestication [Lee18]. Dominance [Ada06, Har09]. Donald [Kel96]. Doomed [Eck90]. Door [MSD21, Mod20]. Doors [Mat15]. Dordrecht [Gor14a]. Double [Bou18, Whe81]. Doubler [Hod87]. Doubt [How12, OC10]. Doxiadis [Oni15]. Dr. [Dav48, Dei48, OCS08, Smo12]. dramatic [BG13]. Drawing [J.23]. Dream [Mod08, Kir05]. Dreams [Lee16, HV06]. Driesch [Che24]. Drift [Ore88]. Drops [Fra81]. Drosophila} [Kel96, Koh91, Smo09]. Drowning [Con06]. Drugs [Str08, Gre07, Les07]. duality [Van07a]. DuBridge [McC04]. Duck [Hui22]. Dufay [Byc13]. Duhem [Roc17]. Duke [Har09, Lin09]. Dulong [LO48, vS67]. d’une [Str06, SD08]. DuPont [Joh08, Ndi07]. Durán [Por13]. durée [Wal94a, Rob24]. Durham [Har09, Lin09]. during [Aas99, Ake07, Ame22, Bru03, Cha20, Hen19, Hod93, Huf71, Mod08, Pic14, Rig83, Sar10b, Son53, Wan99b]. Dust [Cha20]. Dutch [Lon08, AvL16, Coo07a, Kle21, Pye11]. Duty [LKB05]. Dynamical [Wis82]. Dynamics [Rei17b, SS18, Tei10, Gas96].

E. [Bas25, Smo09]. e.org/rentetzi/ [San10]. Early [Ass92b, Ber88, Bou12, Bro14, Bro69, Cam10, CK90, Dar88, Dar91, Dav48, Dei04, Dol76, Far61, Fre81, Gav90, Hom06, Ito24, Kai87, Kin13, Lev60a, Lev62, Lev64a, Lev64c, Lev66, Lon08, Mil50, Mil90, Nes13, NTT10, Nof21, Per18, Pye76, Rob76, Rob20, Roc78, San10, Sch12a, Sla12a, Sut81, Tur82, Urd48, Von01, Ahe03, Ano05c, Ash10, BO06, Bow09, Coo07b, De 05, Gru95, Lew11, MD03, Ren09, Sei01, Ran22]. Early-Victorian [Rob76]. Earth [FF17, Goo84, Kre11, Rai10, Sch12a, Bru87, Con06, Con11, O’C07, RO20, Rus11a, Sch67, Zak13]. earthmoving [Kir05, Mod08]. East [Cro07, Ham10, Har09, Pye11, Wit21, Aug07, Bur24, Han12, Tin24]. Eating [Spa12, Kim14]. Eclipse [EG80]. Eclipses [EG80, Huf94]. Eco [Bra15, Stu25]. Eco-House [Bra15]. Eco-Political [Stu25]. ecohouse [Ank10]. Ecological [ABO10, Bra15, Cam20, Ank10]. Ecologies [Hag14, Lek23]. Ecologist [Cit15]. Ecology [Eri23, Hag14, Koh08, Koh11, Ayr12, Bar09, Col10, Ric12]. Economic [Sak24, Cas01]. Economics [Lee16]. Economy [Gau10, Ric12, Ris98, Wan95a]. Ecosystem [Hag14, Wes10, Col10]. Ed [Lau14, SD08, Bro14, Lin09, Sta08]. Edgar [Kle19, Kli59]. Edge [Whe81]. Edición [Por13, MOBG09]. Edition [Bue53, San10]. Editor [CPWW99, Pye07]. Editore [SD08]. Editorial [Mat15, SD08]. Editors [Phi66, Ano09i]. Edmond [Har09]. Edmund [Zak13]. edn [Zak13]. eds [Kev09, Lin09, Mat15]. Educated [Rob80]. Education [AM64, Bou12, Cas83, Les87, Set12, Jan11]. Edward [Por13, Ved20, Lek23, Pan90b, Ros22]. Edwin [Bra15, Het82]. Effect [DR89, SF13, SF15, Whe78, Im95, For70]. Effects [Lin20, Mun12]. Efficiency [Tur24]. efforts [Wal00]. eGalley [McC3 ]. Eggs [Kli13, Alm11]. Egor [Sem66]. Egorovich [Sem66]. Egyptian [Kin11, BJ10]. Ehrenhaft [Hol78]. Eight [Kau67b]. Eighteenth [Han15, Ter14, Bad18, Gis11, HG71, Huf71, Pal89, Pan20, Rap60]. Eighteenth-Century [Bad18, HG71, Pal89, Rap60]. Einstein [Leg15, Abi00, Abi03, Cre83, Gal82, Gol70, HS89, Hu04, Hun10, Kle70, McC70, Nor84, Pye76, Pye13, Sch02, Sta08, Van07a, Van07b]. Einsteins [Pye82]. elaborado [BGC08, Por13]. Elaboration [Gol76]. Elastic [Ian93]. Elborg [Joh08]. Elective [Hol62]. Electric [Bro69, Eis16]. Electrical [Dun23]. Electricity [Fre81, Kat09, Pan90a, Ris98]. électroaimant [Shi93]. Electrochemistry [Cro88]. electrodynamic [Dar96]. Electrodynamics [Can78, Rue92]. Electromagnetic [D’A75, Dor75, Hir69, Hun15, Ker13, Sie78, Abi03, Kat05, Set04, Set05]. Electromagnetism [LP24]. electrometers [Lel05]. Electron [Byk65, Eck87, Fef89, Jen84, Kai87, McC70, Pot23, Rus81, Dar94, Par01]. Electronics [For87, JI08]. Electrons [Dar84, Hir69]. Electrophysiology [Len86]. Electrostatics [Buc77]. Electrotechnology [Pye82]. Element [Par48]. Elemental [Lau12, Wal15a]. Elements [Dru20, Far64a, Par60, PVP50, PVP50]. elephant [Joh05]. Elite [Dim14, McC3 , Cah04]. elitism [Sch02]. Elixir [Str08, Wat07]. Elizabeth [Lin09, Str08]. Elizabethan [Gar09, Har07a, Kas05]. Ellen [San10, LKB05]. Ellyard [Sei86a]. Else [Cro20]. Elster [Dun23]. embryos [Kel96]. Emergence [Cro88, Eck88, GP17, GH17, Hag14, Iri23, Kel09, Koi75, Mel93, Ris10, Sil74, Cas01, Col10, Gau08, Kat04, Lan10]. Emergency [Kli13, Pre11]. Emerging [FM16, Hag14]. Emigration [Rid84]. Émigré [Wal15a]. Emil [Van07a]. Emitting [LU13]. Emma [Kim14]. Emotion [Wan12]. Emotional [Bel18, Lar17]. Emotions [Bay24]. Empathetic [Set20]. Empire [Fin13, Han15, Har09, LP24, Lon08, Por13, BO06, Ble12, Edm06, Til11, Wil13, Gar09, Num07]. Empire-Laden [LP24]. Empirical [Les81, Rob83]. Empiricist [Sch86]. Employment [Kat19]. Encyclopedist [Lev66]. End [Don67, Ham10, Han15, Iri23, Pri14, SG25, Cat98, Gor09, Hol04, Wil13]. Endangered [Ben12]. Endersby [Kev09]. Endorses [Cit15]. Energetic [Kat19]. Energía [Poh11, Mat15]. energies [Cof06]. Energy [Cas81, Che24, Hod87, Huf81, Leg15, Sei86b, Sei01, Sei08, Tur24, Wis79, Dar01, Gol10, Neb94]. engagement [Orl04]. Engineering [Bro14, Cho12, Cro20, Ham10, Mod20, Pan90b, Aug07, Joh09, Muk09]. Engineers [Mod08, Wis12b, Ake07, Hec14].

England [Leg15, Kin13, Qui82, Sim11]. English [Dol76, Tay48]. English-Speaking [Dol76]. Enigma [Kra09b]. Enlightened [Pye11]. Enlightenment [Kim14, Mel93, Gol07, Bel11, Ble12, Coe10, Cur11, Nov15, Por13, Spa12, Tak11]. Enlightenments [Kim14]. Entelechy [Che24]. Enterprise [Pye82]. entomology [Cas01]. entrance [Mal00]. entropic [Kra07]. Entropy [Dau70, KW96]. entry [KW96]. Environment [For71, Nic24, Tab22, Cam16]. Environmental [Rot13, VD10]. Environmentalism [Bra15, Ham13, Lan10, Bra15, Hag14]. Environments [Güt24]. Envisioning [Rey21]. Epic [Oni15, DPPdD11]. Episode [GSLV^{+}22a, Gue61, Lem18, Nye80, Vos20]. Epistemic [McG18, Par22, Sak24]. Epistemology [Dör12, SSD15, Rhe10a, Rhe10b]. Epoch [Don67]. Epsilon [Hop96]. Equally [Mil10a]. Equations [Nor84]. Equilibria [Hol62]. Equilibrium [Lek23]. Equipment [Dau50]. Equivalents [Roc78]. Era [Les90, San10, SSD15, Hop96]. Erasure [Lev61b]. Eric [Bro14, Har09, Mod08]. Erik [How12]. Erika [Ric12]. Ernst [Phi10, Gli08, Mil10a, Ric08, Ten22]. Errington [Koh11]. Error [Hon89]. errors [Coe02]. Erwin [Han79, Rüg88]. escrito [BGC08, Por13]. España [HR12, Mat15, Poh11]. Essay [Dör12, Hop96, Lau14, Rhe10b]. Essays [CR21, Gil07, Lin09, Ris07, Dea08]. Essence [Sta08]. Essential [Gas21]. Establishing [Ber19, JJ15]. Establishment [Hac86, Rob76]. Establishments [FHW75]. Estrogen [Str08, Wat07]. Estudio [Por13, BGC08, MOBG09]. Ether [Gol70, Hir76, Hun86, Nav16]. Ethics [Tin24]. Ethnic [Hyu19, Rog12, Mul12]. Ethnicity [FR20]. Ethnography [Rob24]. Ethnologists [Set16]. Étudier [Coe10, Loc08]. Eugene [Oni15, Sla12a]. Eugenics [DL09]. Euratom [Kri08]. Eurocentric [Wit21]. Europe [Lon08, Bro14, Coo07b, DIR91, FG98, FG99, Lon08, Mil66, Ogi06, Ols08, Phi10, Ran22, Rus11b]. European [Lau14, Ris10, Gep11, Kri99, Rus11b, Sal07]. Eva [Dei64]. Evans [Bim21]. Ever [FKW24]. Every [Ha15, Ham10]. Everyday [Bau21]. Evidence [Cha12, Gor14a]. Evolution [DL09, Fin21, Gro17, Luk24, Rob83, Smo09, Ung19, Wea97, Cro20]. Evolutionary [Bor10, Den17, Phi10, Rad12, Ric12, Rus11a, Sea25, Mil10b, Ric08, Ric12, Zak13]. Example [Dei21, Fel85]. Excavation [Cit15]. Exchange [Lon08, Coo07a, Wan99b]. Exchanges [Kin13]. exciting [BG13]. Exclusion [Hei83]. Exemplary [Lin09, Rad12, CLW07]. exile [Fre05]. Existing [Bia22]. Exner [Dun23, Han79]. Expanded [Nic24]. Expanding [Hog19, Tam18]. Expedition [Sow17]. Expeditions [Por13, Ble12, EG80]. Experience [Rey22]. Experiencing [BO06, Lon08]. Experiment [BH08, Han15, Oli09, Hun13, Bim21]. Experimental [Bec23, Byc13, Coe21, Dau50, Fel85, Fra77, Gal82, Hon89, Hoo49, Neb94, Rhe50, San10, Tho12, Ren09]. Experimentalchemie [Rhe50]. Experimentalists [Wil82]. Experimenting [Hun86, Sch03]. Experiments [Gal82, Jac14, Mon08, Rig83, CG96, Van07a, Van07b]. Expert [Hec11, Kim14, Mas22, Ran22, Tak11]. Expertise [Ash10, Bro14, Tat16, Bro14]. Experts [Aub20, Shi19]. Explain [Lem49]. Explanation [MK20, Sta08, Kat03]. Explanatory [Chu67]. Expliquent [Lem49]. Exploiting [Gif20]. Exploration [Sow17]. Explore [Ser83]. Exploring [Ved20]. Express [Rus11b]. Extinction [Ric12, Szc24, Bar09]. Extraneous [Sch20].

Extraterrestrial [Het89, Von01]. Extreme [Kar13]. Eye [And09, Dör06, Ott08, Sch07, And09, Mor09].

F [Lek23]. F. [Cas09a, Rap60]. Fabian [Nov15]. Fabien [Coe10]. fable [Kri05]. fabrique [Str06, SD08]. Face [Lau14, Oli13]. facs [Por13]. facsimilar [MOBG09, Por13]. Fact [Ten19, Kri05]. Factors [Lei48]. Factory [Sei08]. Facts [Bou12, Byc13, Rog12, Ten19, Gas96, Lam11, Lew11]. Faculty [Shi79]. Fahs [Kli59]. Failed [CK90]. Failing [Wil24]. Failure [Eck90, Shi86, Shi87]. Fairchild [Cho12, LB10]. Faith [Ved20]. Fall [Fox74, Hol18, Lav14, Mir21, Ris10, Slo14, Har07b]. Fallout [Oni15, Red10]. Falsely [Mar24a]. Families [Nye09]. famine [Tur05]. Fancy [Pic14, Sar10b]. Fang [Wil99]. Faraday [Buc77, Goo81]. Farkas [Cha94]. Farm [Hop96, Sei08]. Farrar [Ham10, Sta08]. Fascism [SW10, Sar10a]. Fascist [Sar10a]. Fat [Low23]. Fate [Har10]. Father [Sco50]. Fauna [CA14]. Fé [Por13]. Fear [Arb19, Tho22, Fel13]. February [DD59]. Fechner [Bau21]. FECYT [SD08]. Federation [Set12]. Fee [Sch09]. Felix [Kri01]. Fellow [Bar16]. Fellows [Bar16]. Female [Ric12, Mil10b]. Feminism [Wei25]. Feminist [Kul25]. Fer [Nev62]. Fermi [BR94]. Fermi-Field [BR94]. Fermilab [BH08, Ham10, Hod87, HKW08, Wes02]. Fetter [Oni15]. Fetter-Vorm [Oni15]. Feud [Ros22]. Few [Ric12, Mil10b]. Feynman [Oni15, OM13]. Fiction [Bal15, Buc13]. Fictions [Pic14]. Field [Ber19, BC91, BR94, Cas81, Dor75, Goo81, Hir69, Kuz66, Mil10a, Nor84, Ric12, Ree09, BR94]. Fieldwork [Kar13, Rob24]. Fifteenth [Wu20]. Fifth [Koi07]. Fifty [GM12, Sea25, SG25]. fight [Par01]. Fighting [Mar22, vSvS66]. Figuring [Ric12]. Filosofía [MOBG09, Por13]. Final [Goo81]. financial [HV06]. Finding [Ahn22, Cre20, Lew12, Lin20, Rad13]. Fine [Kra85]. Finite [Wri14]. Fire [Rob59]. First [Cas79, Dru20, Duv67, Hod87, Kle70, Koi75, KP86, Les81, Lin20, Luk64, Neu14, Oni15, Rai10, Str08, Bru03, Cas07, FV13, Les07, May05, Par01, Cro88]. física [Mat15, HR12]. Físicos [Mat15, HR12]. Fission [Hod93]. Fitness [Gro17]. Flame [Dav48]. Flashing [Hen19]. Flats [Fel13]. Fleck [Sta12]. Flight [Lau14, Tie13]. Florence [SD08]. Flory [Cha93]. Flow [Sei87]. fluctuations [Coe02]. Fluid [SS18, Ten19, Dar98a]. Fluids [Lev61b]. Flux [Cre01, Elk70, Mod08]. fly [Cre95]. Food [Kim14, SW10, Spa12]. Football [Cam22]. Force [Hei76, Qui82, Can97]. Forces [Den17, HG71, Wal15a, BR94, Cas94, Sch12b]. Ford [Wol13a, Kri99, Mun21, Par12]. Foreign [Kri08, Nes13, Wal92]. Forest [GP17]. Foreword [Ano07c, Arm48, Edi02, Fle00b, Lau14, Wes01b, CPWW99, Pye07]. Forged [Lau14, Mun13]. Forging [Rad13, Ros12b]. Forgotten [Mir21]. Formalization [Ker13]. Forman [Gar09, Kas05, MK20]. Formation [Cas79, Dei21, Wis79, Im95]. Former [Hil49]. Forms [Mei23]. Formulary [LAKAK66]. Formulation [Kat19]. Forschens [Win49]. Forster [Joh08]. Fortun [Jac09]. Fossil [May05, Rai10]. Fossils [Rai10, O’C07]. Foster [Tho84]. Fought [vSvS66]. Found [Nor84]. Foundation [Lin20, Ser83, Goo91, Kri99, Wan95a]. Foundations [Bec23, Par12, Ris10, Wol13a, Har07b, Par12, Wol13a]. Founder [Tch67, deM53]. Founding [Nee83, Cha94]. Four [Ber88, Buc12]. Fourier [Fri77]. Fraenkel [Den17]. Fragile [Hog15]. Fragmented [Kwa18]. Frames [Mel93]. Francais [Dau48]. France [Leg15, NV17, Sim11, BT10, CS78, Dav86, Fra77, Gil92, Kim14, Nye74, Rap61, Ris98, Sut81, Tak11, Tei10]. Francis [Fre81, Nes13]. Franck [Hon89]. Franco [Cam20, NR13, Tab22]. Francois [Vie17]. Francois-Vincent [Vie17]. Francoist [Cam10]. Frankfurt [Phi10]. Franklinist [Ris98]. Franz [Han79]. Fraud [Zak13, Van07a]. Fraunhofer [Hen93]. Fred [Bri15]. Frederic [SD08, Mas06]. Frederick [Zak13]. Fredrick [Bro48]. Free [Dar84, Cof06]. Freedom [Koj99b]. French [Har09, Bel11, Bro69, Coo19, Dau50, Gue59, Hol04, Jen06, Kel07, Lem49, Nye83, Shi79, Shi86, Shi87, Som20]. Frenkel [Fre97]. Freshwater [Luk24]. Fresnel [Sil74]. Freund [Ras93]. FRG [Eck88, Eck90]. Friedrich [Sco50]. Friendly [Har10]. Frisch [Sta08, Hof24]. Front [Ano86a, Ano05b, Ano06b, Ano06c, Ano07d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09e, Ano09f, Ano09g, Ano09h, Ano10h, Ano10e, Ano10f, Ano10g, Ano11c, Ano11d, Ano12g, Ano12h, Ano12i, Ano12j, Ano13g, Ano13h, Ano13f, Ano13i, Ano13j, Ano14e, Ano14f, Ano14g, Ano14h, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j]. Frontier [Fel13, Ham10, FH11, HKW08]. Frontiers [Gal23]. Fruit [Coo19]. Fuchs [Wal15a, Goo03, Lau12]. Fugue [Lee16]. Full [Fel13]. Fulminans [Dav49]. Fundamental [Mar15, Rus81, Ste03]. Funding [FHW75, Dev00, Wan95b]. Funds [GG88]. Funny [BK21]. Future [Abb13, Dim14, Hag14, McC3 , Stu25, Wes08b, Zak13, Rei95, Spi11, Dim14, Gan53]. Fyfe [Bal15].

G [Gli08, Phi10, Oli09]. G-Stack [Oli09]. G. [Koh71, Rap60, Stu75]. G.-F. [Rap60]. Gabor [Joh05]. Gabrielle [Wel12]. Gaia [Gro23]. Galison [And09]. Gallium [LU13]. Galvanism [Can78, Hei91]. Game [Les87]. Games [Cam22]. Gano [Nes13]. Ganot [Leg15, Sim11, Sim16]. Gas [Kai87, Tho22, Lel05]. Gases [Gar70, Por81]. gates [De 12]. Gateway [CR16]. Gaukroger [Ris10]. Geister [Wol03].

Geitel [Dun23]. Gender [Bel18]. Gendered [San10, Ren09]. Gendering [Abr20]. Gene [Kev09, SD08, HS06, Sch08b]. Genentech [Dim14, Hug11]. General [HS89, Rhe50]. Generation [Koi07, Wal15a, Hou97, Nye11]. Genesis [Coe10, HK69, Kle21, Ris07, Sta12, Har08, Lin09]. Genetic [AL20, BT10, Bur24, Hog15, Jac09, Lin05, Par07]. Genetical [Muk20]. Geneticists [Sar10a]. Genetics [Bar15, Cam10, GP17, Gau10, Hyu19, Jac09, Koh91, Ric17b, SSD15, SW10, Smo09, HS06, Pan14, Bri15, For08, SD08]. genius [Wil83, Sei86a]. Genocide [Hop96]. Genome [GSLV^{+}22a, Kev97, dC20]. Genomic [Did18]. Genomics [FR20, GSLV^{+}22a, GSLV^{+}22b, Jac09, LVGS^{+}22, LGSL^{+}22, LLV^{+}22, For08]. Genre [Wit21]. Genres [Zak12]. Genus [Smo09]. Geochemistry [Bor23]. Geodesy [Gre83]. Geographers [Kwa18]. Geographies [Lan11, San15, Fin13]. Geography [Bur24, Fin13]. Geohistory [Gil08, Rai10, Rud08]. Geological [Set16, Wri18]. geology [Buc13, Bal15]. Geophysical [ABO10, Goo91, Hac00, Ser83]. Geophysics [Pye90a]. Geopolitics [Men21]. George [Ris10]. Geosphere [RO20]. Geppert [Lau14]. Gerald [San10]. Gerhardt [Bro75]. German [Hop96, Phi10, Cah85, Cah88, Can78, Cas83, Cas94, Cas96, Cra00, Cul74, For71, GG50, Gan53, Gli08, Hop96, Joh13, Pra49, Pra53, Rhe50, SR02b, Ten19, Win49]. German-Speaking [Cah88]. Germans [Wal15a, GS94, Lau12]. Germany [Ham10, Phi10, Aug07, Car99, Cas94, Har05, Har10, Hop96, Huf71, Jac06, Joh13, Kat09, Kin13, Mor09, Nyh09, NV17, Phi10, Tur82]. Geschichte [Gan53]. geschichtlichen [GG50]. Ghost [Bou12, Cro20]. Ghosts [Ric12, Bar09]. Gibbs [Kle69]. Gibraltar [CR16]. Gilded [Cah04]. Gillispie [Dea08]. Gills [Gil23]. Gingras [Ved20]. Giroux [Ham10, Sta08]. Giving [Wol13a]. Glacier [FH19]. Glance [FM24]. Gland [Zul09]. Glare [Sch12a, Sid10]. Glaser [Nev65]. Gleb [TBV20]. Gleditsch [LKB05]. Gliboff [Phi10]. Global [AvL16, Cam16, Fin13, Ham24, How12, Kli13, LK24, OCS08, Wea97, Wel12, Hec12, OC10, RH10, Tak12]. Gloeilampenfabrieken [Sch88]. Gloomy [Bra15]. Glow [Sei87]. Glues [Lev61b, Tea23]. Go [Wol10]. God [Lau14, Dar01, MM75, Oli13]. Goethe [Nie89, Nie91]. Gofman [Sem08]. Gold [Ihd64, Lev59, Lev67b, Sme66]. Golden [Coo07a, Lon08]. Golinski [Coe10]. Gómez [Por13]. Gonads [Gil23]. Good [Ric12, Mil10b]. Goose [Mun11]. Gordin [Ham10]. Got [Cro20]. Gottfried [Den17]. Gould [Rad12]. Governed [Mar15]. Governing [Hen16]. Government [Koz91, Cas01, Dev00, Gru95]. government-sponsored [Gru95]. GPT [Wil24]. Grain [Cam10]. Grammars [Zak12]. Granada [Por13, BG10, GUOB11, Por13]. grand [Shi93]. grandfather [Goo03]. Graphic [Oni15, BG13, FV13]. Graphical [VD10]. Grasp [Kwa18]. Grasping [HS17]. grasslands [Kwa93]. Gravitational [FM24, KR18]. Gravity [KR18]. Great [FH19, Pri14, Hun86]. Greater [Coe10, Mul08]. Greatest [Rus16]. Greco [Kin11]. Greeks [Lei61]. Green [Bra15]. Greene [Gar09, Str08]. Greenhalgh [Rog12]. Gregory [Sch12a]. Grieff [Mat15]. Gross [Kra85]. Grotthus [Str64]. Grounding [Abr20]. Grounds [Mod08, Kir05]. Group [Col24, Dim14, Lev60b, McC3 , Ric12, Bor10]. Growing [Fel13, Hal11]. Growth [GH17, Joh08, Tur71, Léc06]. Guido [Sta11]. Guinea [Kev09, End07]. Gulick [Fin21]. Gunn [Ahe03]. Gurr [Oni15]. Gustave [Nye74]. Gutiérrez [Por13]. Guy [Sei86a]. Guyton [Mel83]. Gyromagnetic [Gal82].

H [Gor14a, Lau14, Lin09, Oni15, Cha12]. H. [Gli08, Phi10, Sol65]. Habilitationsschrift [Kau67a]. Habits [Pye90a]. Hackinsack [Wal15a]. Haeckel [Phi10, Gli08, Ric08]. Half [Buc12]. halftones [And09]. Hall [Hop96]. Ham [Har06, Mod08, Mod08]. Hamblin [Bra15, San10]. Hamilton [FKW24]. Handful [How12, OC10]. Hands [Sch12a, Koh10]. Hands-On [Sch12a, Koh10]. Hanford [Fel13, FH11]. Hanover [Hos07, Nye09]. Hans [Dör12, Kev09, Wal15a, Che24, Dör12, Ric17a, Sch12b]. Hans-Jörg [Dör12, Kev09, Dör12]. Hao [Zul09]. hardback [Har09]. hardcover [Bal15, Bro14, Fle20, Gor14a, Joh08, Kim14, Kli13, Lau14, Lon08, Mat15, Mod08, Oni15, Por13, Ris10, Rog12, Str08, SD08, Wal15a, Wol13a, Zak13]. Haring [Mod08]. Harkness [Gar09]. Harmonious [Jac06, Mor09]. Harold [Lon08]. Harper [Coe10, Oni15]. Harrison [Ris10]. Hartcup [Sei86a]. Harvard [Gar09, Har09, Kev09, Wal15a, Wol13a, Wu20, Coh50]. Harwood [Phi10]. Hauksbee [Fre81]. Haunted [Sam23]. Haven [Gar09, Joh08, Lon08, SD08, Ved20]. Hawai’i [FR20]. Hazardous [How11]. Hazards [Hen16]. Hazy [Pot23]. HBT [SF15, SF13]. Health [Bar15, How11, Jac09, Nur21, Wu20, Par07, Hop96]. Heat [Yag84, Fri77]. Heather [Kli13]. Heavy [Cre08, HV06]. Hebrew [Arb19]. Hecht [Wel12]. Hedeen [Rai10]. Heilbron [Hop96, Car24]. Heinrich [DR50, Dör95]. Heisenberg [Sta11, Wal92, Bro71, Car99, Cas79, Cas09b, Eck90, Han79, Hop96, Mac77, Ser77]. Helen [Cro20, Fin13]. helium [Lev00]. Helmholtz [Cah04, Dar98a, Elk70, Jur02]. Hematology [Muk20]. Henchman [Bal15]. Hendry [Sei86a]. Henri [OL50, Som20]. Henry [Fin13, FF17]. Herbert [Lal13]. hereditary [Dei04]. Heredity [Hog19, Kev09, MWR07, Kev09]. Hermann [Kat09]. Herran [San10, Mat15]. Hersch [Lau14]. Herschel [Aga69]. Herschels [Nye09, Hos07]. Hertz [D’A75, Hon89]. Heß [Hof24]. Heuristics [Hen19]. Hidden [LK24, Nic24]. Hide [She12, Gol13]. High [Cas81, Hof17, Joh08, Sei08, Léc06, Neb94, Cre01]. High-Energy [Sei08, Neb94]. Hilfsmittel [Win49]. Hill [HS89, Oni15]. himself [For07]. Hirosige [Nis76]. Hispanic [Por13, Ble12]. historia [Mat15, Poh11]. Historian [Dei64, Kli48, McC64, Sar50]. Historians [Bur82]. Historical [Byk65, Dör12, GG50, MMMSD21, McC15, MK20, Mod20, Rob83, Ten22, Tho22, Vos20, nye07, All01, Sch20]. Historicize [Elm12]. Historicizing [Bur24, Dör12, Rhe10b, Kev09]. Histórico [BGC08, Por13]. Histories [Abb13, DL11, Dör12, Fle20, FKW24, Nye12, Oni15, Sta16, Rhe10a, Gol13]. Historiographical [Bey11]. Historiographies [Mat15]. Historiography [Aro17, Ten19, McE07]. History [And09, Bak64, BV24b, Bar24, Bay24, Ber19, Bou12, Bra15, Bro23, Bru87, Cah88, Cas86, Cho12, Cow17, Dar84, Dea08, Dei48, Dör91, Dry19, Far65, Fig66, FM16, Gal23, Gan53, GSLV^{+}22a, Gif20, Gil92, God49, Gor14a, GD59, Gor62, Ham24, Han15, Het87, Hop96, HS89, Iri23, Ito24, Joh08, Kel08, Kel09, Kev09, Kin13, Kle19, Kli59, Kli13, Kra82, Kuz71, Lan19, Lau14, LK50, Lem18, LVGS^{+}22, Lin09, Lon08, LGSL^{+}22, Mar24b, Mat15, McG18, NVHB24, NG12, Nye80, Nye83, Nye09, Oni15, Per18, Pet16, Por13, Rai10, Rey21, Ric12, Rus15, Sei87, Sep18, ST18, Set11, ST24, Str08, SD08, Szy84, Tam18, Ten22, TAR^{+}20, Urd48, Wan24, Wit21, Wol13a, Yi24, Zak13, vS67, AA95, Ank10, Bor10, Bro07, Chr11, Coo07b, Cre01]. history [Edm06, End07, FV13, Fle06, Gil07, LB10, Lew11, Mar12, Ogi06, Ott08, Pan97, Pre11, Ric07, Ris07, Rus11a, Sch07, Sma08, Spi11, Ter14, Tie13, Van07b, Wat07, WG10, Zun12, Har09, Wol13a, Zak13]. Hitler [Cra00]. Hitting [Joh09, Cho12]. Hoddeson [Ham10]. Hofmann [Jac14]. Hole [SS23]. Holism [Dar91]. Hollywood [Hec14, Kir11]. Holmes [SD08]. Holy [Gar09, Num07]. Home [Sei86b, BG13]. Homogeneity [Hyu19]. Honored [Wu20]. Hopkins [Coe10, Jac09, Joh08, Kim14, Lau14, Nov15, Por13, Rad13, Str08, Fri07]. Horizons [Mil20, Neu14]. Hormone [Str08, Wat07]. Horse [Rey22]. hose [Ahe03]. Hoskin [Nye09]. Hostile [For71]. Hot [Dör20, Kev90]. Hotly [FH19]. House [Bol22, Bra15, Gar09, Hel12, Sta12, Har07a]. Household [Coe21]. hoy [BGC08, Por13]. HSNS [SF15]. HSTM [Ano09i]. http [San10]. Hubbert [Pri14]. Hubble [Het82]. Hughes [Dim14]. Hugo [Dör21]. Human [Bri15, Dry19, Han15, Rog12, SSD15, San15, Sea25, dC15, dC20, Isa12, Sch08a, Wil13, GSLV^{+}22a, Kev97]. Humanism [Hal62]. humanity [Spi11, Zak13]. Humans [AL20, Gol00]. Humboldt [Phi10, Rup08]. Humphrey [Mil61]. Humphry [Ful60, Ful62, Ful64, Sie59, Sie64]. Hundred [vSvS66]. Hungry [Wol13a, Cul10]. Hunting [Ste88]. Huntington [Lau14]. Hurricanes [Mul08, Coe10]. Hurtful [Moo24]. Huxley [Bas25]. Hybridity [Mac24]. Hybridization [Hol18]. Hydride [Far62]. Hydrochloric [Ret65]. hydrodynamics [Dar02]. Hydrogen [Kra85, Goo03]. Hydrogens [Rig83]. Hydrotherapy [Har09, Jen06]. Hygiene [Hop96]. hype [Fle06]. Hypothesis [Fra82, Roc17, Bru07, Pet16]. Hypsometry [Fel85]. Hysteresis [Dör91]. Hysteria [Tho22].

I. [Les08]. IBM [Gol13, Har11, Nof21]. Ibn [Lev64b]. Ice [FH19]. Iceland [Jac09, For08]. Icons [JI08]. Icy [Ros22].

Ideal [Cro20]. Idealist [Pye13]. Ideals [Sim19]. Ideas [Kuz66, RF69, Bru07]. Identities [Bro14]. Identity [Hyu19, Rog12, Shi16, TBV20, Wu20, Ben01, Sch08a]. Ideological [Ros22]. If [Joh08, Mun03]. Ignorance [Zak12]. II [Bri15, Aas99, Cas96, Cas01, Dar91, Dev00, HHH15b, Hod93, Jon18, Kai02, KR18, Kar13, Tei10, Tur13]. Ilich [Fre97]. Illinois [Phi10]. illus [And09, Bal15, Bra15, Bro14, Coe10, Cro20, Fin13, Fle20, Gar09, Gil08, Gor14a, Ham10, Har09, Jac09, Joh08, Kev09, Kim14, Lau14, Lin09, Lon08, Mat15, Mod08, Mor09, Nov15, Nye09, Oni15, Phi10, Por13, Rad13, Rai10, Ris10, San10, Sch12a, Sta08, Str08, SD08, Ved20, Wal15a, Wol13a, Wu20, Zak13]. Illusions [Bia12]. Illustrating [Can75]. Illustration [Elk70]. Illustrative [And20]. Image [And09, HS17, Pye11, ST18, Kim02]. Images [Gol13, Hec11, PA11, San10, Gol13]. Imagination [Mad17, Sch12a, Sid10]. Imaging [SS23]. Imagining [CA14, Gep11, Lau14]. Immunology [Ras93]. Impact [Bro71]. Imperatives [Har09, Par09, Ada06]. Imperfect [Szy84]. Imperial [Gal23]. imponderables [Hei93]. Impossible [Muk09, Ved20, Bro14]. Impressions [Pal89]. Improbable [Kin11, BJ10]. Impulse [Whe81]. impulsos [BG10, Por13]. inaugural [MOBG09, Por13]. includes [San10]. Inclusive [Gro17]. Incommensurabilities [Gar12]. Indebted [Abr20]. Inderjeet [Wol13a]. Indeterminacy [Han79]. index [And09, Bal15, Bra15, Bro14, Coe10, Cro20, Fin13, Fle20, Gar09, Gil08, Gor14a, Ham10, Har09, Kev09, Kim14, Kli13, Lau14, Lin09, Mat15, Mod08, Mor09, Nye09, Phi10, Por13, Rad13, Rai10, Ris10, San10, Sch12a, Sta08, SD08, Ved20, Wal15a, Wol13a, Zak13, San10]. India [Fle20, God49, Sur02]. Indian [Fle20, Lev61a]. Indiana [Wol13a]. Indians [Fle20]. Indicators [Bak64]. Indies [Pye11]. Indifferent [Roc17]. Indigenous [Jia16, Lon08, Coo07b]. Indirect [Mon08]. Indistinguishability [Dar91]. Individualism [Mor71]. Indo [Muk20]. Indoctrination [Set12]. Induced [Far65]. Industrial [Sch59, FG99, Léc95, FG98]. Industrialization [Vos20]. Industrious [Sch88]. Industry [Rus86, Sla00, WS86, Lev00, Vet06, Mod08]. inédito [BGC08, MOBG09, Por13]. inexperience [Orl04]. Infanticide [Ric12, Ree09]. infestans [Tur05]. Infinities [Rue92]. Infinity [Wri14]. Influence [Cam22, Dav86, Van07b]. Information [Bro48, Set12, Tat19, AB98, Con06, Yi07]. Infrastructures [Con11, Pan24]. Initiation [Cha93]. Ink [SS23]. Inks [Lev61b, Lev64c]. Innovation [Cro20, Cur16, Joh08, LU13, Léc06]. Insects [Pan24]. Institute [Kar77, Tur82, Von01, CH16, Goo91, Les08]. Institute-Building [Tur82]. Institutes [Gau10]. Institution [Ser83]. Institutional [Cah85, GSLV^{+}22b, Shi79, Ben01]. Institutions [Cho12]. Instruction [Coh50, Van49]. Instrument [Cas09a, Mor09, Jac06]. Instrumental [Bor23, Mod12, Dim14, Fle20]. Instrumentalists [Jon18]. Instrumentalizing [Yea12]. Instruments [Bro08, Len86, Swa17]. Insufferably [Cas09a]. Intake [Hol87]. Intake-Output [Hol87]. Integration [Bar18, BC91]. integrity [Fri07]. Intellectual [For71]. Intelligence [Han15, Mar24a, Goo03, Hun13, SR02b]. Intelligentsia [Lee16]. Intelligibility [Ris10, Dea07]. Intense [Wes10]. Intensity [Whe81]. Interactions [Cas81]. Intercontinental [Bar16, Wal00]. Interdisciplinary [DK24, Les87, MC13]. Interface [Gol13, Har11]. International [Ame22, Lau14, Nes13, SR02a, Tat16, Mun97, Mun13, ABO10, De 02, RO20]. Internationalism [Sla12a, Wan99b].

Interpretation

[Gor14a, Per18, Fre05, Van07b]. Interpretations [Shi86, Shi87]. Interpreting [Bub19]. Interpretive [Tho12]. Interrupted [McM22]. Intertwined [Nov15]. Interwar [Szc24]. Intimate [Fle20, Pye11]. Intriguing [VD10]. Introduction [BV24a, BV24b, Bar24, CR21, CR22, DM06, FM16, Gue17, HdAG06, MMMSD21, Mil20, MCR23, NV17, Smi01, TAR^{+}20, Wes08a, Zak13, dCP18, Chr11, Hac00]. Intuition [Wri14]. invented [MD03]. Inventing [Coo07b, Her12, Yea12, Lau14, Lon08]. Invention [Bal15, Cam16, Hir50, Kat09, Sim16, Str64, Buc13, CG96]. Investigaciones [San10]. Investigation [Rob76]. Investigations [Bau21]. Investigative [Hol04]. Invincible [Mar21]. Invisible [Ahn22, Tur06]. Inwardness [Dea08]. Ion [Cre08, HV06]. Ionospheric [Yea12]. Ions [Lel05]. Irish [Rei49]. Irreproducibility [Dei21]. Irving [Koh74]. Isaac [Bri15]. Isaacson [Sta08]. ISBN [And09, Bal15, Bra15, Bro14, Coe10, Cro20, Dea08, Fin13, Fle20, Gar09, Gor14a, Ham10, Har09, Jac09, Joh08, Kev09, Kim14, Kli13, Lau14, Lin09, Lon08, Mat15, Mod08, Mor09, Nov15, Nye09, Oni15, Phi10, Por13, Rad13, Rai10, Ris10, Rog12, San10, Sch12a, Str08, SD08, Ved20, Wal15a, Wol13a, Wu20, Zak13]. ISBN-10 [Kev09]. ISBN-13 [Gar09, Kev09]. Islam [Lev64a]. Islamic [Ris10, Sal07]. Isomorphism [Mel80]. Isotopic [Koh77]. Israel [Cha94]. Issue [FM16, MMMSD21, Mil20, NV17, Wes08a]. Issues [How12, OC10]. Italian [Mal00, Mon08, vSvS66]. Italy [GG88, Orl98, Rus86, Sar10a, vSvS66]. Itineraries [Fre81]. Itself [Wu20]. IUD [Kli13, Tak12]. Ives [Lal13]. Ivory [Szc23]. ix [Coe10, Fle20, Mod08, Ris10, Sch12a, SD08].

J [Bou12, Fal88, Fef89, Gil08, Hag14, Hec14, Hun86, Joh08, Lau14, Lon08, Mod08, Nie89, Nie91, Phi10, Por13, Rai10, Sch12a, Sol61, SD08, Ved20]. J. [Ber82, Fal88, Fef89, Fra77, Nev62, Sol61]. Jackson [Mor09]. Jacob [Bra15, San10]. Jacobsen [Wal15a]. Jains [Jai66]. James [Bal15, Kev09, Leg15, Rei49, Cro88, Hun10, Lec92, Mil66]. Jan [Coe10]. Jansen [Bou12]. Japan [FH19, Koi75, Koi07, Low99, Low06, Tea23]. Japanese [Lee18]. Jar [Ris98]. Javeriana [Mat15, Por13]. Javeriano [BGC08, Por13]. Javieriana [Por13]. Jaws [Mar16]. Jay [Hop96, Rad12]. Jean [SS23]. Jean-Pierre [SS23]. Jed [Kin11]. Jefferson [Ric12, Bar09, Wes02]. Jellyfish [Luk24]. Jennings [Har09]. Jeremy [Str08]. Jerusalem [Cha94]. Jessica [Lin09]. Jesús [SD08]. Jewel [Gar09, Har07a]. Jewell [Hen93]. Jewish [Lev64c, Orl98]. Jim [Kev09, Oni15, Bol22]. Jo [Ano17, Wal15a]. Joaquim [GP17]. Joel [Bri15]. John [Aga69, Hop96, Leg15, Rai10, Sei86a, Str08, Car24, DeY11, Fin21, Sch90, Sem08, Tho84, Wu20]. Johns [Coe10, Jac09, Joh08, Kim14, Lau14, Nov15, Por13, Rad13, Str08, Fri07]. Johnson [Cho12]. Joint [Ano09i]. Jonathan [Mor09, Oni15, Phi10]. Jörg [Dör12, Kev09, Dör12]. Jorge [Por13]. Jørgensen [Kau60]. José [Por13, BGC08, BG10, GUOB11, MOBG09, WG10]. Josefowicz [Kin11]. Joseph [Dim14, And20, Fri77, Par60]. Josiah [Sch59]. Joule [Cro88]. Journal [Ano08i]. Journals [Ano09i]. Journey [VD10]. Joydeep [Fle20]. JPL [Joh08, Wes07]. Jr [Ric12]. Judith [Nov15]. Julian [Bas25]. Jun [Wu20]. Jürgen [Sta08]. Just [Gre08, Hof17, TAR^{+}20, Rog12]. justice [Sur02, She16]. justifications [Ste03]. Justus [Tur82]. Jutta [And09].

K. [Fin13, Gor17]. Kahn [Les08]. Kaiser [Gau10, Cho12]. Kaleidoscopic [Bar24]. Kamerlingh [Dah84]. Kapitza [Koz91]. Karl [Hof24, Lec92, Sco50]. Karlsruhe [dM48, vSvS66]. Karyotype [dC15]. Kassell [Gar09]. Katzir [Set05]. Kautilya [Rây59]. Kayser [Dör95]. Keeling [Wea07]. Keep [McC08, Ham10]. Keller [Har09]. Kelvin [WS86]. Kendrick [Lau14]. Kenelm [Mil49]. Kentucky [Rai10]. Key [Wit21]. Khunrath [DR50]. Killer [Abb13, Tab22]. Killian [Lec92]. Killing [Hop96]. Kind [Kel12, Les08, Nie89, Nie91]. Kind-Terms [Kel12]. Kindī [Lev62]. Kinds [Lin20]. Kinetic [Gar70]. Kingdom [WG10, Por13]. Kirby [Hec14]. Kirchhoff [Sch03]. Kirsch [Mod08]. Kitab [Lev64b]. Klaus [Wal15a, Goo03, Lau12]. Knowing [Fin13, Lan11]. Knowledge [Bal15, Bri15, Bro14, Cul74, Fin13, Gar09, Güt24, Hir50, Lon08, Mat15, OCS08, Rey22, Sch09, Tal16, Zak12, Ash10, Bru07, Can05, Coo07b, Fyf12, Hey06, Hou97, Isa12, Joh09, Koj99a, Mor05, Til11, Cho12]. Kohlstedt [Sch12a]. Kolb [Ham10]. Kolbe [Phi66]. Korea [Wu20, Hyu19, Kim02, Wu20, Yi24]. Korean [Wu20, Yi24]. Kraft [Elk70]. Krementsov [Sch12a]. Krieg [Wol03]. Kristen [Mod08]. Kristie [Hop96]. Kristine [Coe10]. Kuhn [Bia12, Buc12, Gib12, Han12, Hel12, Isa12, Lew12, Lyn12, Mau12, Nye12, Ros12a, Smo12, Sta12, Bri15]. Kuhnian [Rad12]. Kutcher [San10]. Kyoto [Abi00].

L [Bro14, Car24, Dav49, Ham10]. Līqs [Lev61b]. Lab [Hag14, Kir11, Lan10, Oti07, Nye09, Hal11, Hec14]. Labographie [Nye09]. Labor [Bol22, Lau14, LK24, Men21, Tie13]. Laboratories [FG98, FG99, Luk64, Str10, Svi76, Tea23, Cas01, Mun03, Sei01, Wes02, Rus81, Wes08a]. Laboratorio [San10, Her08]. Laboratorium [Pra49]. Laboratory [Ben01, Fin13, Kra09b, Mor92, Nes13, Nic24, Sei86b, Van49, Aas95, Lel05, Til11, Ahe03, Cre01, Fle20, Her08, Les08, Nee83, Pra49, Sei83, Ser83, Wes02, Wes08b]. Labs [Sla12b]. Labscapes [Sar10a]. Laden [LP24]. Ladislaus [Cha94]. Laissez [Ore00]. Laissez-tomber [Ore00]. Lam [Rog12]. Lamp [Str64]. Land [MSD21]. Landau [Huf07]. Landé [For70]. Landscape [Kwa18]. Landscapes [BT10, Sar10a]. Lane [Fin13]. Lang [Phi10]. Langevin [Lel05]. Langmuir [Koh74, Koh75]. Language [Bar18, Csi24, Dör12, ST18, Wu20]. Languages [Nof21]. Lannoo [Hag14]. Laplace [Gue76]. Laplacian [Fox74]. Large [Csi24, DD59, Hod87, KP86]. Large-Scale [DD59, Hod87]. Larmor [Dar94]. Larson [Ved20]. Laser [Sei87, MD03]. Late [Dol76, Gar78, Kat09, Kri08, Pye82, Tab22, Tur08, Mal00, Tur05]. Late-Nineteenth-Century [Gar78]. Later [Buc12, Sig92]. Latin [Bar16]. Latitude [DR89]. Laucht [Wal15a]. Launching [Wes08b, Ano08i]. Laura [Nye09]. Lauren [Gar09, Oni15]. Laurent [Bro75, deM53]. Lavoisier [Dau50, DD59, Duv53, Gue61, Gue76, Lei59]. Lavrentiev [Tat16]. Law [AvL16, Cof06, Cro88, Lev64a, Sch66, vS67, Gor02, Kev94, Sch03, Coe02, Dmi04, Gav90, Hol62, Lei48, vSvS66]. Lawrence [SD08, Sei83, Wes08b]. Laws [AvL16, Lin09, CLW07]. Learn [How24, SD08]. Learned [FM24, Kin13, Mod08, Smo12, Bel11]. Learning [Wer97, Her24]. LeBon [Nye74]. L’école [Dau48]. Lectures [Rhe50, Wal92, Par60]. Lécuyer [Cho12, Joh08]. Led [Lei48]. Lee [McC04]. Left [Hec14, Tai17, Wan95a, Lev13]. Legacy [Nov15, Sea25, CR09]. Legends [Rai10, May05]. Legitimating [Gho25, Sem08]. Lehigh [Les01]. Leibnitz [Eck88]. Leipzig [Goo91]. Lenard [Whe78].

Length [Ker13]. Lens [Col24]. Lenses [FM24]. Leo [SD08]. Léon [Jac07, Jac12, Wal15a]. Leopold [San10, Hag14, Koh11, Lan10]. Leprosy [Edm06, Har09]. Lesch [Str08]. Lessons [Kit12, Lon12]. Lethal [Kra09a]. Letters [Sol61, Sol65]. levels [Wea07]. Levin [Hec14]. Lewis [Bou12, For07, Koh71, Koh75, Siv62, Stu75]. Lexicons [Kel12]. Lexington [Rai10]. Leyden [Dah84, Ris98]. Leyland [Oni15]. Li [Zul09]. Liangying [Mil99]. Liberal [Set16, Jur02]. Liberalism [Bri15, Nye09, Coe07, Tur13]. Liberally [Rob80]. Liberally-Educated [Rob80]. Liberals [Wan95a]. Liberation [vSvS66]. Librarian [Dei64]. Library [Dei64, Lau14]. Lick [Rai10, Hed08]. Liebig [Jac14, Kli53, Phi66, Tur82]. Life [Abb13, Ama20, Cro20, Dim14, Dör12, Esp15, Hag14, Het89, Hop96, LO48, Liu19, Pan90a, Per25, Phi10, Pye11, Sta08, Tat16, Zak13, Ayr12, CE81, Coe07, Kev97, Les01, Nov12, Rhe10a, Ric08, Ris07, Rus11a, Cro20, Dör12, Lin09, Nye09, San10, Bue53]. Lifton [Hop96]. Light [Can75, GB89, Hen19, LU13, Leg15, Nie89, Nie91, Nye74, Stu75, Wes08b, Abi03, Bru07, Bur06, Hun10, Ott08]. light-quantum [Bru07]. light-velocity [Abi03]. Lights [Ame22]. LIGO [Col03]. Lillian [Ham10]. limitations [Mun03]. Limitless [Dim14, McC3 ]. Limits [Cha20, Gil08]. Lincoln [Sla12b]. Lindee [Jac09]. Linear [Mun97]. Lines [Hen93]. Linnaeus [Cha11]. Linus [Dei21]. Lisbon [GS11]. Listening [Dör21]. Listing [Sak24]. Literature [Leg15, Lev64c, Pic14, Wis82, Gol10, Hen14, Bal15]. Little [OCS08]. Liverpool [CH16, Hin03]. Lives [Fle20, Pau11, Kli13]. Living [Fin13, Jia16, Til11]. Lizhi [Wil99]. Local [Lon08, Wu20, Coo07b]. Locher [Coe10]. Lockean [HG71]. Lodge [Hun86]. Loeb [Cro20]. Logicomix [DPPdD11, Oni15].

Logics [LU13]. Logistical [NVHB24]. Logistics [Sak24, Ven24]. Lombard [LK50]. London [Fle20, Gar09, Gil08, Har09, Kev09, Wal15a, Zak13, Har07a, Hun13, Kas05, Han15]. Long [Bro14, Pan20, Yea03, ABO10]. long-distance [Yea03]. Longevity [Par10]. longue [Wal94a, Rob24]. Looking [And09, Mil10b, Ric12]. Lord [Zak13, WS86]. Lords [Par01]. Lorentz [Dar94, Hir69, McC70]. Lorraine [And09, Ric12, DW49]. Loss [Les90]. Lot [SD08]. Louis [LO48, Les08]. Louisiana [Bra15]. Love [FF17, Mod08, Oni15, Red10]. Loving [Smo12]. LTER [ABO10]. Lucier [Bou12]. Lucky [Oli19]. Luis [Cro20]. lumières [Bel11, Kim14]. Luminet [SS23]. Lunbeck [Lin09]. Lynn [Phi10]. Lyon [Gar09]. Lyot [Huf94]. Lysenkoism [DL09, Wol10].

M [Dim14, Fin13, How12, Jac09, Les08, Lin09, Phi10]. MA [And09, Coe10, Gar09, Ham10, Har09, Jac09, Joh08, Kev09, Kli13, Lau14, Mod08, Mor09, Ris10, Wal15a, Wol13a, Wu20, Zak13]. Mach [Ten22]. Machine [Hun86]. Machinery [Moo24]. Machines [Kel08, Kel09]. Macmillan [Lau14, Wal15a]. MacNeven [Rei49]. Macquer [Nev62, Sme66]. Macrakis [Hop96]. Mad [Pri12, Bra15]. Madame [Duv53]. Made [Abr20, Cro20, Sle25, Cro20]. Madison [Hec14, Lev13]. Madness [Har09, Wu20, Kel07]. Madrid [Her08, Por13, San10, SD08, BGC08, CA14, Her08, San10]. Mads [Kau60]. Magic [Gar09, Kas05, Str08]. Magnetic [CK90, Goo81]. Magneton [Qué88]. Maier [Eck88]. Main [Pye90b]. Maine [Par10]. Makers [LB10, Mor09, Jac06, Cho12]. Makes [Ris10, Dea07]. Making [Bri15, Bro23, Did18, Fin13, Lec92, Léc06, MW18, NTT10, OD23, Ris10, Shi19, Sla12b, Sta11, Str08, Sur22, Wal15a, Zak12, Isa12, Lau12, Sal07, Sch12b, Joh08]. Malden [Zak13]. Male [Lau14, Tie13]. Males [Ric12, Mil10b]. Mammal [Ben12]. Man [Bim21, Ham10, Hop96, Mil49, Nep60, Ris10, Rog12, Rus16, BG13, Har07b, Sch08a, RO20]. Manageability [Par09]. managerial [Orl04]. Managing [Bug89]. Manhattan [Kra09a]. Mania [Aub20]. manpower [Kai02]. Manufacture [Lev61b]. Manufactures [Pye13]. Manuscript [Kau67c, Nev62]. manuscrito [MOBG09, Por13]. Many [Kit12, Mil90, Ved20]. Mao [Mil99]. Maoist [Bia22]. Maps [Chr11, San10, Zak13]. March [Joh08, Ndi07]. Margaret [Pic14, Rad13, Sar10b]. Maria [Fin13, San10, SD08]. Mariano [AM64]. Marie [Oni15, Red10]. Marine [Ben01, Ben12, Esp15]. Marino [Lau14]. Mark [CE81, Hop96, Mod08, Ric12]. Market [Kli13, Alm11, Mar22]. Markets [Men21]. Marrow [Kra09a]. Mars [Fin13, Lan11, Rus11b, Sim19]. Martian [Sch12a, Kre11]. Martin [Gil08, Rai10]. Martina [Mun11]. Martini [Bra15]. Marxist [Jac07]. Mary [Wal15a, Ano17]. Mask [Pea21]. Masks [Den23]. Mass [Hol62, Pot23]. Massachusetts [Bra15, HS89]. Masses [Sie78]. Master [Zul09]. Material [Kim14, Dei04]. Materialism [Gor17]. Materials [Bec23, BV01, Eis16, GH17, Iri23, LU13, Mac24, MC13, OD23, San10, Sol61, Ren09]. Math [Kle19]. Mathematical [Bar18, Gas96, Lee16, LK24, TP16, Wis79, Yag84]. Mathematician [Abr20]. Mathematicians [For71, Rid84, Wil82]. Mathematics [Bar16, Bec80, Fel85, Jun79, Kle19, LK24, Sch81, Shi16, Shi79, SS18, Wis79, SG99]. Mathematization [Buc77, Fra77]. Matrix [Mac77]. Matter [Ano86a, Ano05b, Ano06b, Ano06c, Ano07d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09e, Ano09f, Ano09g, Ano09h, Ano10h, Ano10e, Ano10f, Ano10g, Ano11c, Ano11d, Ano12a, Ano12b, Ano12g, Ano12h, Ano12i, Ano12j, Ano13g, Ano13h, Ano13f, Ano13i, Ano13j, Ano14e, Ano14f, Ano14g, Ano14h, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ber82, Dar86, HG71, Rei17a, Kuk95]. Matters [Coo07a, Lon08]. Matthew [Coe10, Hec14, Lau14]. Matthews [Fra82]. Max [Hop96, Hal11, Gol76, RH00, Von01]. MAX-lab [Hal11]. Maxwell [BE69, Gar70, Hun15, LP24, Por81, Rei53, Wis82]. May [HS89, How11]. Mayer [Hei76]. Mayor [MOBG09, Por13, Rai10]. Mayr [Mil10a]. Maze [MJ15]. McCay [Par10]. McCormmach [Aas07, BH07, Set20]. McCray [Ham10, Dim14]. McGill [Tho84]. MD [Kim14, Lau14, Por13]. Me [Igo18]. Meadow [Dun23]. Mean [Wol10]. Meaning [Bro08, Löw90c]. Meanings [CG96, Ved20, Can97]. Measure [Lam18, Wal00]. Measurement [Hun15, Pot23, WS86]. Mechanical [Dor75, Ser77, Yag84, Yea12]. Mechanics [Bro76, Bro08, JJ15, Mac77, Mel93, Rüg88, Sta11, Dar98a, Fre05, Van07b]. Mechanistic [Hir76]. Media [Gol13, Nav16, Tab22, She12]. Mediaeval [Lev64c]. Medical [Bar15, Cro07, Har09, Hop96, Kli13, LAKAK66, Wes96a, Alm11, Edm06, Ede59]. Medicina [BGC08, Por13, Por13]. Medicine [Ama20, BV24b, Bay24, Har09, Hop96, Jac09, Kas05, Kli13, Lon08, Mit23, Mor21a, Por18, San10, Wu20, Coo07a, Cro07, Kut09, Les07, Lin05, Par07, Gar09, Jac09, Str08]. médico [BGC08, Por13]. Medieval [Lev67b]. Medium [Nav16]. meetings [Bel11]. Meets [Gro17]. Megascience [BH08, HKW08, Ham10]. Meghnad [Sur02]. Meissen [Pra53]. Meissner [Pra53]. Membrane [Liu19]. Memorial [Dei48, Mar21, MSD21]. Memoriam [Car24, Dei64, Nis76, Pye74]. Memory [Sta11]. men [Low99]. Mendeleev [Dmi04, Dru20, Gor02, Lei48]. Mental [Wu20]. Merchants [OC10, How12]. Merle [Wan12]. Merlin [Tay48]. Merton [Wan99a]. Mesa [Les08]. Mesopotamia [Lev59]. Mesopotamian [LB59]. Mesotrons [Mon08]. Mess [Abb13]. metabiography [Rup08, Phi10]. Metabolism [Nes13]. Metal [Far62]. Metals [Eck87, Sme66]. Metamorphosis [Tre75]. Metaphorical [Wei22]. metaphors [Koj99b]. meteorology [Har08, Coe10]. Meters [Kat09]. Method [Fri89, HS18, Hol87, Rei17b, Yag84]. methodological [All01]. Methodologist [Bor23]. Methodology [Can75]. Methods [LVGS^{+}22, Ten22, Tho12, VD10]. Metropolis [Bey11, Lau14, Wes12]. Mexican [MSD21]. Mexico [Bar15, de 06]. Meyer [Dav48]. Michael [Hag14, Ham10, Har09, Hop96, Jac09, Nye09, Ved20, Nye11, Ram95, Ric17a, Rob60, Wal15a, nye07]. Michigan [Mar21]. Microbial [Lee18]. Microchip [Cho12, LB10]. Microscope [And09, Iri23, Sch07]. Microscopy [Dim14, Mod12]. Mid [Bar16, Dav86, Hol18, Ore00, Bor23]. mid-20th-Century [Ore00]. Mid-Nineteenth [Dav86]. Mid-Twentieth [Bar16]. Mid-Twentieth-Century [Hol18, Bor23]. Middle [Bur24]. Midi [Bro14, Muk09]. Migration [Zul09]. Mikado [Ito24]. Mikhail [Tat16]. Milam [Ric12]. Militarized [Arb19]. Military [Hac00, Les87, Les90, Ore00, Sei87, Kut09, San10]. Milky [Tai17]. Millikan [DR89, Fra81, Hol78]. Milt [McC64]. Mind [Bal15, Bau21, Kuk95, Pye90a, Hen14]. Mine [Nev62]. Mineral [Gal23]. Mineralogy [Lev67a]. Mining [AM64]. Minkowski [Gal79]. Minting [Lev67b]. Miracle [Str08, Les07]. mirage [KH93]. Misbehaving [Pan14, Bri15]. Miserably [Cas09a]. Mises [SS18]. Missing [Joh13]. Mission [Neu14, Ada06, Har09, Lam18, Rus11b]. Missions [Som20]. Mistake [Qué88]. mitad [Mat15, Poh11]. Mitscherlich [Mel80]. Mittasch [Far66]. Mixed [CH16, LVGS^{+}22]. Mixed-Methods [LVGS^{+}22]. Mobilization [Joh13]. Model [Cas79, FM24, Het82, Lin09, Stu86, CLW07]. Modeling [DM06, Gro17, Hey06, Kwa93]. Models [Csi24, Joh13, Len86, Mac77, MN18, Car99]. Moderation [Hen19]. Modern [Bar16, Bar18, Bas25, Bro14, Coe10, Fle20, Joh08, Kin11, Kin13, Leg15, Lon08, Nyh09, Ran22, Rog12, Swa17, Ten19, Wan20, deM53, Ash10, BJ10, Coo07b, DeY11, Har08, Mul12, Ndi07, FH19, Phi10]. Modernism [McE07]. Modernity [Ric17a, Ris10, Wit21, Gau08]. Modernization [BT10, Gau10]. Modernizing [Joh13]. Modes [Hun15]. modification [Fle06]. Mody [Dim14]. Mohr [Sco50]. moléculaire [Str06, SD08]. Molecular [Ass92a, Ass92b, Cas07, Gar78, Hog15, Hoo49, Joh13, Mac24, Rei17b, Rey21, Rig83, SD08, AA95, Jac09, Kat03]. Molecularization [Tur08]. Molecularizing [Slo14]. Molecules [Den17, Mun67, Pet16, Dar01]. Moments [Cho12, Lin05, Kai10, Jac09]. Money [Wea07]. Monika [Hop96]. Monitoring [Wea07]. monopoly [Gor09, Ham10]. Monotheism [Ris10]. Mont [Gil08]. Months [Kli13, Pau11]. Montreal [Tho84]. Moonwatch [Ham10, McC08]. Moral [Koz91, Sch02]. Moran [Gar09]. Morien [Tay48]. Moritz [Dav48]. Morley [Fra82]. Morning [Kli13, Pre11]. Morris [Han15, Wil13]. Moseley [FF17]. Most [Lek23, Rei17a]. mostly [BG13]. Mother [Bra15, FF17, Ham13]. motion [Ian93]. motions [Dar98a]. Mount [Ahn22, Bim21]. Mountain [Dun23]. Mountains [Rei17a]. Move [Bar18]. Movement [Lae22, Koj02].

Mukerji [Bro14]. Mulcahy [Coe10].

Mullaney [Rog12]. Müller

[Kev09, Nye09, Oti07]. Müller-Wille [Kev09]. Mulliken [But94]. Multidisciplinary [Par10]. Multimedia [Bri15, Tur13]. Multispecies [Coe21]. Mumford [For07]. München [Pra49]. Munich [Pra49, Pye82]. Munns [Lau14]. Munro [Kli13]. Murphy [Kli13]. muscles [San05]. músculos [San05, SD08]. Museo [CA14]. Museum [Lar17, Sun12]. Museums [ST24, Str10]. Music [Dör21]. Musicians [Mor09, Jac06]. Muslim [Lev60a]. Musschenbroek [Bec23]. Mutation [Cre15]. Mutis [BGC08, BG10, Por13, GUOB11, WG10]. Mutisiana [Por13, BGC08, MOBG09, GUOB11]. My [Igo18, Mun11]. Myles [Mor09]. Myrick [Oni15]. MYRRHA [Bro23]. Myself [Lew12]. Mythology [Abb13].

N [Hop96, Koh71, Lin09, Stu75, Tch67]. N. [Sch88]. Nachwirkung [Dör91, Ian93]. Nacional [CA14, Mat15, Por13]. Naming [Ito24, Kra14, Wri18, Wu20]. Nanotechnologies [Dim14, McC3 ]. Nanotechnology [MC13, Mod12, Dim14]. Naomi [How12]. Napoleonic [Fra77, Sut81]. Narrative [Por18]. Narratives [Per18, CLW07, Lin09]. NASA [Mod08]. nascendi [Dmi04]. Nation [Bou12, Cam10, Rog12, SD25, Wu20, Jan11, Lam11, Mul12]. Nation-Building [Wu20]. Nation-State [Rog12]. National [Bou18, CA14, Cre01, For87, Gor20, Hop96, Hyu19, Mod08, Nee83, Aas95, Mun03, Sei01, Bug89, Goo91, Hen16, NTT10, Wal92, Wal94a, Wan95a, Wes08a]. nationalism [Ito02]. Nations [Rob20]. Natividade [GP17]. Natural [Bou12, Han15, Iri23, Ito24, Kin13, Kle19, Lar17, Lon08, Mar24b, Mod08, NVHB24, NG12, Pic14, Rai10, Rob80, Rus15, Sch20, Sep18, ST18, Tab22, Tam18, Ten22, Wil82, Wit21, Ake07, Coo07b, Lew11, MOBG09, Ogi06, Por13, Ric07, Sar10b, Ter14, WG10]. Naturales [CA14]. Naturalists [Dea14]. Naturalized [Cow17]. Nature [Bar09, Bra15, CA14, Cha11, Dea14, Dor75, Gol76, Han15, Kin13, Lon08, MM75, Phi10, Por13, Pye11, Ris10, Sch12a, Sla12a, Stu75, BO06, Bru03, Coe02, Dea07, Koh10, Nyh09, Ter14, Vet04, Ham13, Ric12, Ric12, Wit21]. Natures [NVHB24]. Naval [Hac86, Ahe03]. Navy [Rai00]. Nazi [Hop96, Gau10, Gif20, Har10, Joh13]. Nazis [Hop96]. Nazism [Hop96, Hop96]. NC [Har09, Lin09]. Ndiaye [Joh08]. Necessity [Bel18]. Needham [TAR^{+}20]. Neffe [Sta08]. Neither [Ric17a]. Neo [Pye13]. Neo-Idealist [Pye13]. Neocolonialism [Swa17]. Néstor [Mat15, San10]. Netherlands [Kin13]. Network [ABO10, GSLV^{+}22b]. Networks [LGSL^{+}22, AB98]. Neufeld [Hop96]. Neurospora [Koh91]. Neutron [Ano05c, Bro71, De 05, Eck88, Wes10]. Neutrons [Eck88]. Never [Liu19]. Newlands [vSvS66]. Newton [For49, Ris10]. Newtonian [HG71]. Newtonianism [Ste88]. Nicholas [Nov15]. Nick [Wol13a]. Nicolaas [Phi10]. Niels [Aas99, Kra79, RH00]. Nier [Bor23]. Nierenberg [OCS08]. Nikolai [Sch12a]. Nine [Kli13, Pau11]. Nineteenth [Bal15, Bou12, Mor09, Phi10, Buc13, Jan11, Bro69, Dav86, Dol76, Dor75, Far50, Far64a, Gar78, Jac06, Kat09, NV17, Ols08, Por18, Pye82, ST18, Von18]. Nineteenth-Century [Bal15, Mor09, Phi10, Buc13, Bro69, Dor75, Jac06, Kat09, NV17, Ols08, Por18, Pye82, Von18]. Ninth [Lev66]. Ninth-Century [Lev66]. Nishina [Ito02, Kim06]. Nitride [LU13]. Nixon [TAR^{+}20]. NJ [Bro14, Ham10, Kli13, San10, Ved20, Wol13a]. NMR [Rei17b]. No [Abr20, SG25, SF15]. Nobel [Cra00, Fri89, Joh05]. Noble [Lar17].

Noise [Nic24, Oli19]. Nomenclature

[Luk24]. Nonbinary [Vel22]. Nonclassical [Kuz71, Kuz74]. Nonintegral [Sie78]. Nonlinear [Wil15]. nor [Ric17a]. Normal [Sla12b]. Normalization [dC20]. North [Har09, HS89, Kel07, Sch12a, Koh10, Rob20]. Northern [Ame22]. Norton [Lin09]. Norwegian [Kle21]. Notes [Chu67, Lau14]. Notions [Lev66]. nouvelle [Str06, SD08]. Novel [Buc13, Bal15]. November [Dim14]. Nuclear [Cit15, Eck90, Fel13, Hop96, Kle21, Mar21, Mod08, Mun12, Mun21, Nee83, Ros22, Rot13, San10, Sch12b, Sei86a, Sem08, Tho84, Tho22, Tur06, Wal15a, Wel12, Ahe03, BR94, Bru03, Ham08, Hec12, Hom06, Kir05, Lau12, Wal15a]. Nucleosynthesis [Far64a]. Nucleus [Stu86]. Nuestra [MOBG09, Por13]. Nueva [Por13, GUOB11]. Nuevo [BG10, Por13]. Number [Kau67b]. Numbers [Coe10, Str08, Gre07, Har08]. Nummedal [Gar09]. Nüsslein [Kel96]. Nüsslein-Volhard [Kel96]. Nye [Wal15a, Ano17]. Nyhart [Phi10]. Nylon [Ndi07, Joh08]. NZ [Bra15].

O [Bas25, Bor23, Bro14, Gor14a, Por13, SD08, Cha12]. Oak [GP17]. Oakland [Wu20]. Objective [Cul74]. Objectivity [DG07, SB25, Str10, Yea12, And09]. Objects [Mei23, Wri14, Kel96]. Obscured [How12, OC10]. Observation [Het82, Mon08, DL11, Gol13]. Observational [Cre83]. Observations [Por18, Nev62]. Observatory [Ahn22]. Obsolete [Sch20]. occupation [Cas96]. Ocean [CR16, Gal23, Mar16, Rob20]. Oceanography [CH16, Lam18, Mil90, Con06, Ore00, Rai00]. Oceans [Gal23, Gol60, San10, Ham08]. Ochoa [SD08, San05]. O’Connor [Rai10]. Octaves [vSvS66]. Octet [Koh74]. Octubre [Por13, MOBG09]. Odóe [NG12]. Office [Fle00a]. Ogilvie [Lon08]. Oil [Bou12, Fra81, Pri14, Luc08]. Old [Bub19, Jen84, Nav16, Rei87, Ret65]. Oliphant [CE81, Sei86a]. Oliver [Lau14, Hun86]. Olivier [Wol13a, Abi04]. Olschki [SD08]. Olson [Phi10]. Omission [Per18]. Once [Tre75]. One [Abr20, Cam10, Elm12, Jen84, Kel08, Rog12, Sch09, vSvS66, Gre08, Ric17a]. One-Electron [Jen84]. Online [San10]. Onnes [Dah84]. Ooze [Gal23]. Open [Did18]. Open-Access [Did18]. Opening [Mat15]. Operation [Ham10, Hop96, McC08, Mun12]. Operations [Far64b]. Oppenheimer [Ber82, Cas07]. Opportunity [Rid84]. Optical [Tal16]. Optics [Wil15]. Optimism [Kuz74, Mar21]. oración [Por13, MOBG09]. Oral [FKW24]. Orange [Bra15, Mar12]. Order [Cro20, Rei87, MMMSD21]. Orders [Tho17]. Oregon [Bro14]. Oreskes [How12]. Orfila [PVP50]. organ [Dar98a]. Organic [Byk65, HS18, Jac14, Tch67, deM53, Gor02]. Organicist [Sle25]. Organism [Che24, Nyh17]. Organismal [Esp15]. Organismic [Mil10a]. Organisms [Kel08, Kel09]. Organization [Bug92, Kel08, Kel09, NV17]. Organized [Hu04]. Organizing [DeV87]. Origin [Cah82, Dar86, DR50, For53, Hoo49, Koh71, Sie86, Abi03]. Originality [Csi24]. Origins [Abi91, BK21, Bar15, Dor75, Hei83, Hir69, Hir76, Hyu19, Jac14, Ker13, Mod08, Neu14, NTT10, Pau11, Phi10, Ten19, Tur20, Wal15a, Dar96, Gli08, Huf07, KH93, Nye11, Vet06, Gru95, Kli13]. Ornithology [Rus15]. Orogeny [Rei17a]. Ørsted [Can97]. Ortiz [Por13]. Osborne [BE69]. Osgood [HS89]. Osiris [Bro14]. Osorio [Lon08]. Ostwald [Hil49]. Oswald [Ams93]. Other [Rei17b, Hei93]. Otis [Nye09]. Ottaviani [Oni15]. Otter [Mor09]. Our [Kli13, Kre21, Pau11]. Ourselves [Kre21]. Outer [Lau14, Gep11]. Output [Hol87]. overhead [Gru95]. overload [Con06]. owe [Wis87]. Owl [DR50]. Oxford [Bal15, Bra15, Gar09, Nye09, Ris10, Str08, Mor92]. Oxidation [Far65]. Oxygen [Mel83]. Oysters [Gil23].

P [Cas09a, Lau14, Wu20, Coe10]. P. [Nev62, Tch67]. P.-J. [Nev62]. Pacific [AL20]. Paduan [Gre83]. Page [Sch09]. Pages [Ede59]. Pair [Koh71]. Paleoecology [Cha20]. Paleontological [Mad17]. paleontology [Hed08, Rai10]. Palgrave [Lau14, Wal15a]. Pamela [Bro14]. Panama [Sim19]. Panel [DeV87]. Pangloss [OCS08]. Pankaj [Fle20]. Pannekoek [Tai17]. Panofsky [Bri15]. Pap [Joh08]. Papadimitriou [Oni15]. paper [Bra15, Bro14, Cas07, Coe10, Cro20, Dea08, Dei04, Fin13, Fle20, Gil08, Har09, Joh08, Kim14, Kli13, Lau14, Lon08, Mat15, Nov15, Oni15, Phi10, Por13, Rad13, Rai10, Ris10, Rog12, Ved20, Wol13a, Zak13]. paperback [Jac09, SD08]. Paradox [Far65, Fel13]. Paradoxes [Joh13]. Parasites [Hal15]. Parcours [Kim14, Bel11]. Paris [Kim14, Kin11, Bel11, BJ10, CG96, Gre83, Kim14, Spa12]. Park [Cho12]. Parmar [Wol13a]. Parsimony [Dei21]. Parsons [Bri15, Isa12]. Part [Kel08, Kel09, Nie91, HHH15a, HHH15b, Nie89]. Parthasarathy [Jac09]. Particle [Ber19, BH08, Bro76, Tho12, de 06, Neb94, Van07a]. partir [BGC08, Por13]. Parts [Bar18, Esp15]. Pascal [Rai10]. Passion [Rog12, Lam11]. Past [Hag14, Jia16]. patent [Kev94]. patenting [Tur06]. patents [MD03]. Path [Dim14, Pan90a, Rus11b, Mod12]. pathological [Fle06]. Pathology [Tur08, Tur05]. Paths [Nov15]. Pathway [Wei22, Sod94]. pathways [Bel11]. Patrick [Dim14, Ham10]. Patronage [CH16, KR18, Hac00, Ore00]. Paul [Bou12, Hop96, Kli13, Cha93, Cit15, Koh11, Lel05]. Pauli [Ser77]. Pauling [Dei21]. Paupers [Wol13a, Rus13]. Paying [Sch09]. Peace [MSD21, Shi16, KH93]. Peaceful [Kri08].

Peak [Pri14]. Pearl [Eri23]. Peasant [Har10]. Peasant-Friendly [Har10]. Peculiarity [MMMSD21]. pedagogical [Hol04]. Pedagogy [BV24b]. Peder [Bra15]. Pedro [Por13]. Pei [Les08]. Peierls [Wal15a, Lau12]. Peking [Rog12, Sch08a]. Penicillin [Bud07, Lee18, SR02b, Str08]. Pennemunde [Hop96]. Pennsylvania [Mod08]. Penrose [Wri14]. People [Rog12, Sch08a]. Peoples [Lin20]. Performativity [Ito24]. Perfumery [Lev60b]. period [Mil99]. Periodic [Dmi04, Dru20, Gor02, Lei48]. peripheries [De 02]. Periphery [GS11, HHH15a, HHH15b]. Perpetual [Hec14]. Persistence [Gil23, Tea23]. Personal [Hil49]. Personnel [FHW75]. Perspective [Gra25, Het89, Joh08, Phi10, Str10, Wes10, Nyh09]. Perspectives [Cas83, Hog19, Wan99a]. Pertinent [Sch20]. Peter [And09, Hag14, Joh08, Lau14, Phi10, Ris10, Sch12a]. Petit [vS67]. Petrus [Bec23]. Phage [SD08, HS06]. Pharmaceutical [Urd48]. Pharmacy [Son53]. Phase [Kle70]. Pheasant [Ito24]. Phenomena [Goo81]. phenomenological [Kat03]. Phil [Lau14]. Philadelphia [Dea08, Mod08, Mil50]. Philanthropy [Wol13a, Zun12, Wol13a]. Philip [Sch88]. Philipp [Whe78]. philologist [Dör95]. Philosopher [Som20]. Philosophical [Dea08, Het82, Lan19, Mar15, Ric17a]. Philosophy [Aga69, Bro75, Che24, Far66, Gol76, Kuz74, Lin09, Nye74, Pic14, Sie59, Wal15a, Jac12, Ris07, Sar10b, Sch20]. Phlogistic [Sie64, Rap61]. Phlogiston [Mel83]. Phoenix [Mar21, MMMSD21, MSD21]. Phosphate [Sur22]. phosphorus [Kri05]. phosphorus-32 [Kri05]. Photoelectric [Whe78]. Photographic [Pot23]. Photography [Ame22, Gol13, dC15, She12]. Photon [SF13, SF15]. Physical [Bad18, Dol76, FM16, For87, Jun79, McC15, Mil90, Sch20, Shi79, ST70, VD10, Cha94, Lel05]. physician [Kas05, Gar09]. Physicist [Wal15a, Cat98, Orl98, Sch12b, Wil99]. Physicists [Ber88, For71, GG88, Gav90, Koi75, Mor09, Rid84, Tur24, Wol03, HR12, Jac06, Kai02, Low99]. Physics [Asp80, Ass92a, Ber19, BH08, Bro69, Cah85, Cah88, Can78, CK90, Cas81, Cho12, CS78, Dav86, Dör91, Fel85, FHW75, Fox74, Fra77, Fre81, GG88, Gal82, GS11, Gue76, Ham10, Hei76, Hol99, KR18, Kat09, Kev90, Kra82, Kra85, Kuz71, Leg15, MJ15, Mar15, Mor92, NR13, Nye74, Orl98, Pet16, Por81, Sch09, Sch86, Sei86a, Sei08, Set11, Shi86, Shi87, Sil74, Sim16, Sta11, Svi76, Szy84, Tal16, TBV20, Ten19, Ten22, Tho84, Tho12, Wal92, Wal15a, Wan12, Wan20, Wes20, Wil82, Wis87, Ano05c, Cah04, Cas09b, Cat98, De 05, Doi09, Dör95, Hen84, HKW08, Hom06, Hun10, Ito02, Jac12, Kat04, Kim02, Koj99b, Kri99, Mil99, Neb94, SR02a, Sim11, Ste03, Tur06, Wes96a, De 02, HR12, Von01]. Physikalisch [Cah82]. Physikalisch-Technische [Cah82]. Physiology [Hol87]. Phytophthora [Tur05]. Picasso [Pye13]. Pick [Rai10]. Pickering [Fle20]. Pierre [Oni15, Red10, Fre81, LO48, Roc17, SS23]. piezoelectricity [Kat03]. Pig [Kev09, LLV^{+}22, End07]. Pigeonholes [Koh08]. Pigs [Kre21]. Pilatre [Duv67]. Pile [Eck88, Sut81]. Pills [Str08]. Pilot [Per18]. Pioneer [AM64, Kli48]. Piotr [Koz91]. Pipeline [Wei22]. pipes [Dar98a]. Piscataway [Ham10]. Place [Bay24, Cas09a, Wri14]. Placing [DiM24]. Plain [Nic24]. Planck [Hop96, Von01, Gol76, Kuh84, Sch03]. Plane [Tie13, Lau14]. Planet [Fin13, Lan11]. Planetarium [SD25]. Planetary [Ros22]. Plans [Sch81, Wea97]. Plant [Cro20, Den17, Har10, Kwa18, Smo09, Tur05, Tur08]. Plants [Koh08, Por13, Lon08]. plata [Don67]. Plates [Pot23, Lau14]. Players [Cam22]. Playing [Les87]. Plenty [MSD21]. Plowshare [Mod08, Cit15, Kir05]. Plum [Mod08]. Pluralism [Cha12, Gor14a]. Pluto [Neu14]. Plutonium [Hod93]. pneumonia [Ams93]. poder [HR12, Mat15]. Poem [Tay48]. Poetics [Rai10, O’C07]. Poetry [Ful60]. Pohl [Mat15]. Point [Hod81, MJ15]. Point-Contact [Hod81]. Poison [Fle20, Ham08, San10]. Poisons [Lev64b]. Poland [Szc24]. Polanyi [Wal15a, Nye11, Ric17a, nye07]. Polar [Ame22]. Polarized [Dar84]. Policy [AB98, Eck90, Hal11, Kri08, Mir21, Neu14, Rog12, Shi19, Ved19, Wal15b, Cas96, Gre08]. policymakers [Low99]. Polinière [Fre81]. Political [Gau10, HV06, Kri08, MN18, McG18, Nyh17, Stu25, Tyr18, Aas99, Koj99a, Ott08, Wan95a]. politicization [Lev00]. Politics [Abb13, Arb19, Bra15, Eck88, Eri23, Fef92, Hen19, Hop96, Hyu19, Jac09, Kev09, Kin11, Neu14, Nof21, NV17, SB25, Sut81, Tat16, TP16, Wal15a, Wu20, But94, Car99, Jac12, Kev97, Kri05, Low06, Mar12, Mil99, MWR07, Par07, Wan95b, Sta08]. Polity [Ved20]. Pollution [Fle20]. Polymer [Cha93]. Polytechnic [Sch81]. Pontificia [Mat15, Por13]. Pool [Mun21]. Poor [Wol13a, Rus13, Ris98]. Popular [Rai10, Rog12, Sch12a, O’C07, Sch08a]. Popularization [Sch12a, Bow09]. Population [Gro17, Hyu19, Gas96]. Populations [MW18]. Porcelain [Pra53]. portrait [Fre97]. Portugal [GP17, Sar10a]. Porzellans [Pra53]. POSEIDON [Lam18]. Positive [Fal88]. Positivism [Jon18]. positron [Par01]. post [Mil99, Pan97]. post-constructivist [Pan97]. post-Mao [Mil99]. postmodernism [McE07]. Postmodernist [Pye11]. postulate [Abi03]. Postwar [Dim14, Les90, Mir21, Ras93, Sei83, Wal15b, Low06, Wan95a]. Potassium [Dav48]. Potato [Tur08]. potatoes [Tur05]. Potential [Shi79]. Poulson [Kel96].

Poverty [Wol13a, Cul10]. Power

[Aub20, Hop96, Ito24, Leg15, NR13, Per18, Sur22, HR12, Hun10, Koj99a, Par12, Wol13a]. powered [Fyf12, Bal15]. Powers [HG71, Hop96]. pp [And09, Bal15, Bra15, Bro14, Coe10, Cro20, Dea08, Fin13, Fle20, Gar09, Gil08, Gor14a, Ham10, Har09, Jac09, Joh08, Kev09, Kim14, Kli13, Lau14, Lin09, Lon08, Mat15, Mod08, Mor09, Nov15, Nye09, Oni15, Phi10, Por13, Rad13, Rai10, Ris10, Rog12, San10, Sch12a, Sta08, Str08, SD08, Ved20, Wal15a, Wol13a, Wu20, Zak13]. Practical [Bro14, Ash10]. Practice [Ame22, CH16, Han15, Koh08, Phi10, Sta11, VD10, Dar98b, Har05, Set10, Ter14, nye07, Hey06]. Practices [Aub20, Bro14, Dea14, FR20, San10, San15, FG98, FG99, Ren09, Tur06]. Practising [Cro07, Har09]. practitioners [Lon11, Bro14]. Precision [Lav14]. PreCrime [Lem18]. Precursor [Bac65]. Predecessors [EG80]. Predicament [Sta19]. Prediction [Cha20]. Predispositions [Gal82]. Preface [Gor14b]. Prehistory [Pra53, vS67]. Prejudice [Cah88]. Preliminary [Lev61b]. preneogranadinos [GUOB11, Por13]. Preparing [Nee87]. Prescott [Kli13]. Prescribing [Gre07, Str08]. Present [ABO10, Hag14, Bro07, Zak13]. President [Ham10, Wan08]. Press [And09, Bal15, Bra15, Bro14, Coe10, Cro20, Fin13, Fle20, Gar09, Gil08, Ham10, Har09, How24, Jac09, Joh08, Kev09, Kim14, Kli13, Lau14, Lin09, Lon08, Mod08, Mor09, Nov15, Nye09, Phi10, Por13, Rad13, Rai10, Ris10, Rog12, San10, Sch12a, Sta08, Str08, SD08, Ved20, Wal15a, Wol13a, Wu20, Zak13]. Presses [Coe10]. prestige [Mun97]. Presuppositions [Hol78]. prévoir [Coe10, Loc08]. Price [KR18]. Pride [Cah88, Qué88]. Priestley [Bra15, Ede48, MM75]. Primacy [Eck90]. Primatology [Ric12, Ree09]. Prince [Mil61]. Princeton [Bro14, Ham10, Rai10, Sta08, Ved20, Wol13a, Aas95]. Principle [Hei83, Kat04]. Printing [Hir50, How24]. Prior [Coh50, Deb62, Dör91]. Prison [Ham24]. Private [KR18, Nye09, Coe07]. Prize [Joh05]. prizes [Cra00, Fri89]. Probe [Dim14, Mod12]. Problem [Fin13, Hir76, Huf81, McM22, Sie78, Til11, Cre83]. Problems [Nye83]. Proceedings [HS89]. Process [Wil24]. Processes [Far65]. Proclamation [Wol03]. Proctor [Hop96]. Produced [Kev09, MWR07]. Production [Cho12, Koh91, Lee18, Leg15, Mat15, SW10, Sme66, Joh09, Kai02, SR02b, Sim11]. Productive [Bia12, Het89, Sch12a]. Productivity [FHW75]. Profane [Lau14, Oli13]. Professional [Cas79, Rus15, Cas07]. Professionally [Vel22]. Professorial [Tur71]. Profit [Did18, Les90]. Profit-Making [Did18]. Program [CK90, Gau10, Hen16, Joh08, Lau14, LK24, Mod08, Par10, Smo09, Oli13, Wes07, ABO10, Slo14]. Programmatic [Cas81]. Programme [RO20]. Programming [Nof21]. Programs [MC13, Von01]. Progress [Sim19]. progressive [Léc95, Wan95a]. Project [Mar21, Mod08, GSLV^{+}22a, Kev97, Kir05, Koi07, MSD21, McC04, Nof21]. Projected [Bue53]. Projecting [Dim14]. Proletarian [Sch12a, Kre11]. Prometheus [Ham10, Aug07]. promise [Fle06]. Promised [MSD21]. Promising [For08, Jac09]. promoción [GUOB11, Por13]. pronunciada [MOBG09, Por13]. proofs [Sch03]. Propaganda [Eck87, Wal92]. propagation [Yea03]. Proportions [Kap65]. propulsion [Ahe03]. Prospects [FM16]. protean [Vet04]. Protection [Ben12]. proteínas [San05, SD08]. proteins [San05]. Proto [Kat19, Mei23]. Proto-Algorithmic [Mei23]. Proto-Energetic [Kat19]. Proton [KP86, Wit04]. Proust [Kap65]. Proving [Kir05, Mod08]. Provoked [Kin11, BJ10].

Prussia [Tur71]. Prussian [Tur82]. PSAC [Wan95b]. Psychedelic [Bri15, Tur13]. Psychiatry [Har09, Kel07]. Psychological [Kar13]. Psychology [Hop96, Nye80]. Public [Hec11, Nur21, Ric17b, Set12, Ved19, Wal94b, Wol10, Dör06, Low99]. Publication [Dru20, Moo24]. Publications [Ano17, Nye09, McM22]. Published [Fra81]. Publishing [Bal15, BV24a, Fyf12]. puertas [De 12, Mat15]. Pulsed [Wes10]. Pulvis [Dav49]. Puppis [Rob83]. pure [Ito02]. Purifying [BT10]. Purpose [Ved20, Ved20]. Pursued [Dim14, McC3 ]. Pursuing [Hun10, Leg15]. Pursuit [Kev09, Sch08b]. Puzzle [Eri23]. Pyrotechny [Dav48].

qua [SG99]. Qualities [Hil49]. Quantification [Fel85, Hol87]. quantitative [Hei93]. Quantized [Dar86]. Quantum [Ass92b, Bro76, Bro08, BC91, Bub19, Cas81, Dar88, Dar91, For71, For87, JJ15, Jen84, Joh13, Kuz71, Par09, Per18, Rue92, Sch90, Set04, SG99, Sta11, Tal16, Bru07, Cas09b, Fre05, GS94, Koj99b, Set10, Van07b, BV09, Sta11]. Quantum-Relativistic [Kuz71]. quasiparticles [Koj99b]. Queer [Lau14, Tie13]. Quest [Bug92, Hec14, Hop96, Str10, RM06]. Question [Dar84, Kus18, Pye13, SF13, SF15]. Questions [Bub19, Ras93]. Quicksand [Fri89]. quicksilver [Cha01]. Química [PVP50].

R [Nye09]. Rabies [Bol22]. Rabinowitch [Sla12a]. Race [Bim21, FR20, Gis11, Hop96, Nov15, Tyr15, Zul09, Fab10]. Rachel [Hec11, Ric17b]. Racial [And20, Tyr15, Hop96]. Racialism [Set16]. Radar [San10, Leo09]. radiaciones [Her08, San10]. Radiactividad [San10, Her08]. Radiant [Whe81]. Radiation [Cre15, Hal15, HHH15a, HHH15b, Kra09a, Lin20, Mon08, Sei83, Sem08, Her08, Sch03]. Radicalism [Tai17]. Radio [Lau14, Mod08, Oli19, Pan90b, Yea12, Har06, Mun97, Mun13, Yea03, Mun03]. radio-wave [Yea03]. Radioactive [Coe02, Oni15, Red10, San10, Ham08]. Radioactivity [Her08]. Radioecologists [Rot13]. Radiometer [BE69]. Radium [Bra15, San10, Pri12, Ren09, Cro20]. Raikov [Tch67]. Raison [Den17]. Ralph [Rai10]. Ramsauer [Im95]. RAND [Hou97]. Rapid [Rob24]. rara [BG10, Por13]. Rare [FF17]. Raspail [Vie17]. Rating [FKW24]. Rational [MM75]. Rationalism [Kuz71]. Ray [Cas81, Coo19, DR89, Het87, Van07a]. Rays [CK90, DIR91, Fal88, Nye80, Whe81]. Re [Ano08i, FKW24, MK20, Rey21]. Re- [FKW24]. Re-Envisioning [Rey21]. Re-launching [Ano08i]. Re-reading [MK20]. Reach [Bal15, Hen14]. Reaction [Gav90]. Reactions [Far61]. Reactor [Bro23, Kle21, Mod08, Mun21, Cre01]. Reactors [Nee83]. Readers [Bal15, Sec14]. Reading [Szc23, MK20]. Ready [Ger21]. Readying [Wes02]. real [Cha01]. Realism [Cha12, Gor14a]. Realistic [SS23]. reality [Hey06]. Realization [Ras93, Orl98]. Really [Bia22, BG13]. Reason [Fin13, Pic14, RH10, Sar10b]. Réaumur [Han15, Ter14]. Rebecca [Bra15]. Rebirth [Slo14]. Recalled [Hil49]. Recently [Bro48, Kau67c]. Reception [Can75, Fra82]. receptions [Huf07]. Recoil [Tre75]. Recombinant [Cre08, Mod08]. Reconceiving [HS06, SD08]. reconnaissance [She12, Gol13]. Reconsidered [Wol10]. Reconsidering [Che24, Sta11, BV09, Phi10]. Reconstructing [Cha20, Wu20]. Reconstruction [Gil08, Rai10, Joh05, Rud08]. Recruitment [Bia12]. Red [Fin13, Gor09, Sch12a, Aug07, Ham10, Lan11, Sid10]. Redefining [Jac09, Tur24]. Redesigning [Cho12]. Rediscovery [BC91]. Redniss [Oni15]. Redshift [Hen93]. Reduction [Far65]. Redux [Lin09, Roc17, Ris07]. Rees [Ric12]. Reference [Mel93]. Refining [Lev59]. Reflection [Yi24]. Reflections [And09, Tur82, Sch07, FKW24, Kev97]. Reform [Lec92, Léc95, Rud08, Rai10]. Reframing [Pan20]. Regime [BT10]. Regnant [Sch86]. Regulation [Ben12, CG96]. Reich [Hop96]. Reichenbach [Ric17a]. Reichsanstalt [Cah82]. Reid [Fin13]. Reid-Henry [Fin13]. Reimagining [Col24, Wes20]. Reino [BG10, Por13]. Rejection [Ore88, Ser77]. Relations [Sol61, SR02a]. Relationship [Bau21, Jac09]. Relatively [DD59]. Relativistic [Kuz71, Sie78, Cre08]. Relativity [HS89, Lal13, Abi91, Abi00, Cas86, Cre83, Dar96, EG80, Gol70, Gol76, Hir76, Hu04, Mal00]. Religion [Kin11, She16, Ved19, Ved20, BJ10]. Remaking [Mor21a]. Remarkable [Gil23]. Renaissance [Lon08, Sal07, Tho84, Ogi06, Ris10]. rencontres [Bel11, Kim14]. Rene [Kli13]. Renneberg [Hop96]. Rennes [Coe10]. Renormalization [Men21]. Rentetzi [San10]. repair [Yi07]. Replacement [Str08, Wat07]. Reply [Abi04, ST70]. reports [SR02b]. Repository [GM12, Ede59]. Representation [Cha11, Pan97]. Representative [Son53]. Reprint [Sig92]. Reprints [Sig92]. Reproductive [Kli13, Kus18]. republic [Lew11, Bou12, Tho17]. Republican [NG12, Jia16]. Repulsive [Qui82]. requisitions [Kai02]. Research [AB98, Ahe03, ABO10, Bec23, Ben12, Ber19, Bro23, Bug89, Byc13, CK90, DeV87, Dei21, Eck88, Eis16, For87, FF17, GG88, GS11, Gau10, Hac86, Hal11, Hog15, Hop96, Iri23, Jia16, Jun79, Kar77, Lan19, Les87, Lin20, Mor92, Nee87, Nes13, Par10, Rei87, Rus81, San10, Sch88, Sei87, Shi79, Sla12b, Smo09, Sun12, Tur71, Wal15b, Ake07, Ams93, Cas01, FG98, FG99, Hom06, Kut09, LKB05, Rei95, Ren09, SR02b, Shi93, Von01, Wea97, Wes02, Mod08]. research-technology [Shi93]. Researches [D’A75]. Reshaping [Ng19]. Residential [Koh11]. Resistance [Fin13, RH10]. resolución [BG10, Por13]. Resource [Gal23]. responded [Dev00]. Response [Ano09i, Gol70, Set05, Wol10]. Responses [Rue92, Dei04, Mil99]. Responsibility [Kli13, LKB05]. Rest [Kli13, Pau11]. Restoration [Han15, Hun13]. Restraints [Ric12, Bor10]. Restriction [Par10]. Rethinking [JJ15]. Retooling [Wes08b]. Retouching [Jia16]. Retrospection [Kuz71]. Reveal [LVGS^{+}22]. Reverse [Ihd64]. reversibility [Yi07]. Review [Bia22, Bou12, Bro14, Bru87, Bug92, Bur82, Cah88, Cam16, Cas83, Cas86, Cho12, Dea08, Dim14, Dör12, Fel13, Gil08, Gol13, Gor14a, Hag14, Ham10, Hec14, Het87, Het89, How12, Jac09, Kel11, Kim14, Kin11, Lan19, Mat15, Mit21, Moo16, Mor21a, Mor09, Mor21b, Nye83, Pal89, Pic14, Pye90b, Rad13, Rad12, Rai10, Rey21, Ric12, Rus16, Sei86a, Seo21, She16, Shi86, Shi87, Sta11, Sta16, Ste16, Ste88, Ved20, Wal15a, Wel12, Wil24, Wis87, Wit21, Wu20, Zak13, Hop96]. Reviews [Abb13, And09, Bal15, Bra15, Bri15, Coe10, CR21, Cro20, Dea08, Fin13, Fle20, Gar09, Han15, Har09, Joh08, Kev09, Kli13, Lau14, Leg15, Lin09, Lon08, Mod08, Nov15, Nye09, Oni15, Phi10, Por13, Ris10, Rog12, San10, Sch12a, Sta08, Str08, SD08, Whe82, Wol13a, Gil07]. Revisited [Rad12, Wan95a]. Revisiting [Kuh84]. Revolution [Bor23, Cah85, Cho12, Ede59, Gue59, Mau12, Pic14, BO06, Har07a, Mor05, Sar10b, Doi09, Gar09, Gil08, Lon08, Rap61]. Revolutionaries [Abb13]. Revolutionary [NV17, Gil92]. Revolutions [GM12, Lon12, Nye12, Ros12a]. Revolving [Mod20]. Reworking [Wei22]. Reynolds [BE69]. Rheinberger [Dör12, Kev09, Dör12, Dör12]. Rhetorical [Per18]. RHIC [Cre08]. Rice [Cam10]. Richard [Har09, Nov15, Phi10, Ris98, Shi16, SS18]. Richards [Phi10]. Richet [Rai10]. Ricketts [Hag14, Lan10, Lek23]. Riding [Gib12]. Righi [CK90]. ring [Par01]. Rio [Don67]. Rise [Bou18, Bro14, Fel13, Fox74, Hol18, Lav14, Mac77, Mac24, Men21, Mod08, Phi10, Shi19, Slo14, ST70, Svi76, Wol13a, Ake07, Lon11, Nyh09, Par12]. Risk [Kli13]. Riskin [Lin09]. Risks [Sem08]. Ritter [Esp15]. Rivalries [DIR91]. Rivero [AM64]. Robert [Ber82, Hop96, Phi10, Sta08, Str08, Bue53, But94, Deb62, Ihd64]. Robin [Tho17]. Robotland [Rob24]. Rockefeller [Wol13a, Goo91, Par12]. Rocket [DeV87, Hop96, Sch12a, Sid10]. rockets [Von01]. Rocky [Fel13]. Rod [Har09]. Role [Can75, Eck90, JJ15, Leg15, Mar16, Rus15, Sem66, SS18, Wis79, DeY11, Kim02]. Roman [Gar09, Num07]. Romancing [DR89]. Romanticism [Cul74]. room [Cre95]. Rooster [Som20]. Roosth [Cro20]. Roots [LP24, Ros22, Gor02]. Roqué [Mat15]. Rosario [Mat15, MOBG09, Por13]. Rosenfeld [Wal15a, Jac07, Jac12]. Rosomoff [Nov15]. Ross [Ahe03]. Rossiter [Rad13]. Rouelle [Rap60, Rap61]. Rouge [Bra15]. Route [Wis79]. Routes [LK24]. Routledge [Fle20]. Rowe [Sta08]. Rowland [Hen93]. Royal [Rob76]. Royale [Hol03]. Rozier [Duv67]. Rudolf [Wal15a, Koh77, Lau12]. Rudwick [Gil08, Rai10]. Rule [Coo19, Dei21]. Rupke [Phi10]. Rupp [Van07a, Van07b]. Ruse [Ved20]. rush [Hom06, Mil53]. Russell [Zak13, Aas07, BH07, Set20]. Russia [Fig66, Lei59, Luk64]. Russian [Sol61]. Ruswick [Wol13a]. Rutgers [Ham10, Kli13, Mod08, San10]. Rutherford [Sei86a, Stu86, Tre75, Wil83, Sei86a].

Rydberg [Nep60].

S [But94, Den17, Fin13, Gil08, Mau12, Rai10, Sol65, Wal15a]. Sacred [Lau14, Oli13]. Safety [Sem08, Str64]. Sagan [Ros22]. Saha [Sur02]. Said [Kul25]. Saint [Dar01]. Sainte [OL50, Sol65]. Sainte-Claire [OL50, Sol65]. Salazar [GS11]. Saliba [Ris10]. Salk [Les08]. sallies [Huf07]. Sally [Dim14, Sch12a]. Salvage [Bol22]. Salvation [Fin21]. Samarqandi [LAKAK66]. Samuel [Fre81]. San [Lau14]. Sander [Phi10]. Santa [Por13]. Santafé [MOBG09]. Santesmases [SD08]. Satellite [Con06, Stu86, Orl04]. savant [Bel11, Coe10, Kim14, Loc08]. saw [For07]. Saxony [Jun79]. Scale [DD59, FKW24, Hod87, Cha01]. Sceptical [Sar50]. Schickore [And09]. Schmalzer [Rog12]. Schoenheimer [Koh77]. Scholar [FH19, Mil49]. Scholarly [BV24a]. Scholars [Sol61, Wil24]. Scholarship [Wit21]. School [GS11, Hof17]. Schools [Son53]. Schrödinger [Han79, RF69, Rüg88]. Schulmann [Sta08]. Schuster [Kli13, Sta08, Fef89]. Schwann [Vie17]. Schwartz [Kev09]. Schweber [Wal15a]. Science [And09, Arb19, ABO10, Aro17, Bal15, BV24b, Bar24, Bay24, Ben12, Bia22, Bou12, Bow09, Bra15, Bri15, Bru03, Bug89, CH16, Che24, Coo12, CLW07, Cre08, Csi24, Dea08, Dim14, Dör12, Eck87, Fef92, Fin13, Fle20, Fre05, Fri77, Fri89, Gar78, Gho25, Gif20, Gil92, GH17, Hac86, Hal11, Ham10, Ham24, Han15, Hec14, Hei76, How24, Jac09, Joh08, Kar77, Kev90, Kli13, Koh11, Kra09b, Kuz71, Kuz74, Lau14, Lav14, Lec92, Leg15, Les81, Lin09, Lon08, Low99, Lyn12, Mac24, Mar16, Mar21, Mat15, McC64, McG18, McM22, Mir21, MC13, Moo16, Mor21a, Mor71, Mor09, NR13, Neu14, Ng19, Nov15, Nye80, Nye83, Nye09, Oli19, Ols08, Pal89, Phi10, Pic14, Pra49, Rai00, Rai10, Ris10, RO20, Rob20, Rog12, Ros22, Rus86, Sch12a, Sch88]. Science [Sei83, Sei86b, ST18, She16, ST24, Shi79, Sim16, Sow17, Sut81, ST70, Swa17, Tat16, Tat19, Ten22, TAR^{+}20, Tur20, Tur08, Tyr15, Ved19, Ved20, Wal94a, Wal15b, Wan24, Wan08, Wei22, Wil15, Wol13a, Yi24, Zul09, BV01, BJ10, Buc13, But94, Car99, Cas94, Cas96, Coe07, Col10, Col03, Coo07a, Cur11, DeY11, Dea07, Doi09, Fab10, For07, Gau08, Gil07, Gol10, Gre08, Gru95, Hac00, Har07b, Har05, Hei93, Hop96, Hun13, Ito02, Jan11, Jur02, Kir11, Koh10, Kre11, Les01, Mar12, McE07, Mun13, Nye11, O’C07, Ogi06, Pan97, Pan14, Ree09, RH10, Rus13, Sal07, Sch08a, Sec14, Str06, SD08, Sur02, Tak12, Wal94b, Wan95a, Wan95b, Wan99a, Wol13b, nye07, Bro23, Cho12, Dea08, Fin13, Hag14, Ham10, Hop96, Kev97, Kin11]. Science [Leg15, Lin09, Mod08, Phi10, Ric12, Ris10, Sch12a, Ved20, Wal15a, HV06, Sch20]. science-as-social-practice [nye07]. Science-City [Tat16]. Sciences [Bri15, Bro14, FM16, Jun79, Kim14, Mar24b, McC15, Mod20, Pan20, Rob80, Sea25, Shi79, Sta16, Tab22, VD10, Wil82, Con06, Isa12, Lon11, Spa12, Vet04, Hol03, Sch20]. Scientific [Can75, Cas09a, Cha93, Cur16, Dmi04, Eck90, Fin13, Gar09, Gol13, GM12, Hop96, Lal13, Les08, Lon08, Lon12, McM22, Moo24, OCS08, Pic14, Pye76, Ris10, Ros12a, Shi16, Sim19, Sla12a, Sol61, Son53, Sow17, Tai17, TBV20, Tei10, Van07a, VD10, Wal15a, Wan12, BO06, Can05, Dar98b, DL11, Gau08, Har07a, Kai02, Kri01, Mal00, Mor05, Orl04, Sar10b, Til11, Wan99b, Nye09]. Scientism [Phi10, Sur02, Ols08]. Scientist [Coe02, Elm12, Hec11, Leg15, Sch67, Aas99, DeY11, Sch02]. Scientists [Dim14, GB89, Hec14, Hof24, How12, Lar17, Luc08, McC3 , Mod08, Rad13, Ric17b, AB98, Ake07, But94, Cra00, Gol00, Kir11, OC10, Ros82, Ros95, Ros12b, Bou12, Set12, Sla12a, Mor21b]. Scott [Mod08]. Scottish [Rea48]. Scripps [Esp15]. Scrutinizing [dCP18]. Sea [Gal23, Rob20]. Sealab [Kar13]. Sean [Kim14]. Search [Hec14, Oni15, dC20, Cha01, CG96, DPPdD11, Mun97]. Searching [Har09, Rad13]. Sears [Cit15]. Seaside [Ben01]. Second [Kra79, Oni15, Ame22]. Secord [Bal15]. Secrecy [Kle21]. Secret [Cro20, Gil92, Hop96]. Secretary [Dei64]. Security [For87, Kev90, Aas95]. Sedimentary [Gal23]. See [FM24]. Seebeck [Nie89, Nie91]. seeds [Her08]. Seeing [Fin13, Mar16, Lan11, Ste16]. Seek [Gol13, She12]. segunda [Mat15, Poh11]. Seismology [Dör21, Goo84]. Seiya [Dar04]. Sekhsaria [Fle20]. Selected [Ano86b, Bey00, Bey01, Bey02, Gro03a, Gro03b, Gro04a, Gro04b, Gro05a, Gro05b, Low86a, Low86b, Low87, Low88b, Low88a, Low89b, Low89a, Low90b, Low90a, Low91, Pan93, Pan94a, Pan94b, Pan95, Rid91, Rid92, Rid93, Wes96b, Wes97a, Wes97b, Wes98a, Wes98b, Wes00, You06b, You06a, You07a, You07b, Lau14]. selection [Bor10, Ric12]. Self [Abr20, Kel08, Kel09, Lan23, Wan12, Yi24]. Self-Made [Abr20]. Self-Organization [Kel08, Kel09]. Self-Reflection [Yi24]. Semantics [Wri18]. Semiconductor [Sch88, Cho12, LB10]. semillas [Her08, San10]. Semipermeable [Liu19]. Sen [Fle20]. Señora [MOBG09, Por13]. Sensational [Mor09]. Sense [Phi10, Ris10, Dea07, Ric08]. Sentiments [Lar17]. Sequencers [LVGS^{+}22]. Sequences [LVGS^{+}22]. Sequencing [GSLV^{+}22b]. Serendipity [Oli19]. Series [Bro14]. Serological [Str10]. Service [Har09, Léc95, Cro07]. Seth [Kat05, Sta11]. Settles [Ede48]. Seven [Chu67]. Seventeenth [Pic14, DW49, Rea48]. several [Wes96a]. Severo [San05, SD08]. Sex [Alm11, Bel18, Ha15, Kli13]. Sexuality [Lau14, Tie13]. Seymour [SD08, HS06].

Shack [Hag14, Lan10]. Shadow [Fel13, Ham10, Nov15, CR09, Wan99a, Wan08]. Shanghai [Pye90a]. Shape [Kli13, Pau11]. Shaped [Ten19]. Shaping [Ayr12, Bar16, Ng19, Ris10, Cah04, Gau08, Hag14]. Shared [Koh71]. Shark [Mar16]. Shaul [Set05]. She-Wolf [Tab22]. Shelley [Sta08]. Shells [Gil23]. Shifting [Ama20, Dim14]. Shobita [Jac09]. Shovel [Ger21]. Shovel-Ready [Ger21]. Show [Rai10, O’C07]. Showcase [Tat16]. Showers [Cas81]. Siberian [Tat16]. Siddiqi [Sch12a]. Side [Bra15]. Sidney [Mac24]. siècle [Loc08, Coe10]. Siegel [Str08]. Siemens [Cah82]. Sight [Nic24, SG25]. siglo [Mat15, Poh11]. Signal [Fle00a, Nic24]. Significance [DR50]. Sigrid [Rog12]. silence [KW96]. Silicon [Joh08, Léc06]. Silvan [Wal15a]. Silver [Lev67b, Sme66]. Simon [Gar09, Kli13, Oni15, Sta08, Kas05]. Simple [Sei86a, Wil83]. Simpson [Rei53]. Simulation [Abb13, Hey06]. Simulations [Mac24]. Simultaneous [Fre81]. Sin [Qué88]. Since [Rus16, Tei10, Jon18, Rad13, Ros12b, Wan99a, Wu20]. Single [Lau14, Mun13]. Singular [GSLV^{+}22a]. Síntesis [SD08]. Síntesis/FECYT [SD08]. Sir [Aga69, CE81, Far64b, Ful60, Mil49, Mil61]. SIRIO [Orl04]. site [Can05]. Sites [Cha93, FG98, FG99]. sixties [Lev13, Tur13, Bri15, Hec14]. Skaar [Wal15a]. Sketch [Byk65]. Sketches [Fre97, Ful67]. Skies [AvL16, Ham10, McC08]. Skowron [DL09]. Skull [Nov15, Fab10]. Sky [Bal15, Con11, Lau14, Hen14, Mun13, Wes12]. Skylab [Kar13]. SLAC [Hal15, Wan95b]. Slater [Sch90]. Slavery [Mit23, Nov15, CR09, Cur11]. Sleepwalkers [Dea14]. Smail [Zak13]. Small [Hal11]. Smith [Dim14, Mor09, Kli59]. Smithsonian [Oni15]. Smoke [How12, OC10]. Snails [Fin21]. So-Called [Mar24a]. Soap [Low23].

Social [Bro14, Can78, Gro17, Huf71, OCS08, Por81, Rob76, Rog12, She16, Szy84, Tho17, Wal15a, Wil15, Lam11, Nye11, Sur02, nye07, Koj99b, Mor21b]. Socialism [Bia22, Hop96, Goo91, Hop96, Wal92, Wal94a]. Socialist [Tai17, Tat19]. sociedades [GUOB11, Por13]. Societies [Mil50]. Society [Ama20, Coe10, Dea08, Jac09, Mil50, Hol99, Kev94, Low99, Mul08, Bac65, LK24, Mil59]. Sociobiology [Gho25, Gra25, Kul25, Per25, Sea25, SB25, Sle25, Stu25, SG25, Wei25]. Sociologists [Kwa18]. Soft [Aub20]. Solar [Bad18, Hen93]. Solid [MJ15, Tei10]. Solid-State [Tei10]. Solution [Deb62, Dol76]. Solutions [Tea23]. Solvay [BV09, Sta11]. Somatic [Hol18]. Some [Bey11, DW49, Gue59, Hil49, Lev64c, Pal89, BG13]. Something [Cro20]. Sommerfeld [Sta11, Eck87, Set10]. Sonar [Hac86]. Sonderweg [Phi10]. Sooner [KW96]. Sophia [Cro20]. Sophomore [Ros12a]. Sophus [Kau60]. Sort [Lek23]. Sounders [Yea12]. Source [Wes08b, Wes10]. Sources [Nye83, nye07]. South [Wu20, AM64, Cas09a, Hyu19, Kim02, Yi24]. Southern [Lau14, Goo84, Wes12]. Sovereignty [CR16]. Soviet [Sch12a, Hol99, Lee16, Sid10]. Soviets [Hec14, Wol13b]. sowerbii [Luk24]. Soyoung [Wu20]. SP [Mod08]. SP-2006-4317 [Mod08]. Space [AB23, DeV87, Dim14, Gal79, Ham10, Joh08, Lam18, Lau14, McC3 , Mod08, Nee87, Rüg88, Rus11b, Gep11, McC08, Oli13, Wes07]. Spaceflight [Sch12a, Sid10]. Spaces [Kar13, Sow17]. Spain [Cam10, Cam20, HR12, NR13, NG12, Tab22]. Spanish [Lon08, BO06, CA14, De 12, Her08, HR12, Mat15, PVP50, SR02a, San05]. Sparking [Mac24]. Spary [Kim14]. spas [Jen06, Har09]. Spatial [Von18]. Speaking [Cah88, Dol76]. speaks [Dev00]. Special [FM16, MMMSD21, Mil20, NV17, Wes08a, Abi91, Cas86, Gol70, Gol76]. specialties [Cas01]. Species [Ben12, Jia16]. Specimen [Kin13]. Spectacular [Mor09]. spectra [Cas07]. Spectro [Lav14]. Spectro-Chrome [Lav14]. Spectrum [Kra09b, Rob83]. Speculation [Mit23, For08, Jac09]. Spend [Bey11]. sperm [Alm11, Kli13]. Spier [Zak13]. Spirit [Cha01, Kri01]. Spokesman [Hop96]. sponsored [Gru95, Wan99b]. Spontaneous [BC91, Hod93]. Spots [Pot23]. Spread [Eck87, Lei59]. Springer [Gor14a]. Sputnik [Ham10, Aas95, Wan08]. Squeeze [Wes08a]. SSC [Kev97, KH93]. Stability [Bad18]. Stable [Kel09]. Stack [Oli09]. Staffan [Kev09]. stages [Hom06]. Stahl [Rap61]. Stalin [Ham10, Gor09, Koz91]. Standardizing [Ame22]. Standards [Ker13]. Stands [Bad18]. Stanford [Wu20, Les87, Vet04]. Stanisław [DL09]. Stanley [Rai10]. Starry [Bal15, Hen14]. Start [Oli19]. Started [Smo12]. State [Bra15, Bro14, Cam10, Dah86, Hec14, MJ15, Rog12, Tei10, Ash10, Hol99, Lam11, Sur02, Wan99b, Wil99, Wol13b, Bro14, Gau10, Kev90]. state-sponsored [Wan99b]. Statecraft [Szc24]. States [Abb13, McG18, Nyh17, BV01, DIR91, Dim14, For87, Gav90, Hac00, Kev97, Kli13, Nov12, Pre11, Rid84, Sch86, Son53, Tyr15]. statesman [Aas99]. Station [Ben01, Mod08, GP17]. Statistical [Por81]. Statistics [Dar88, Dar91, Kle19, Sep18, ST18, Von18]. statu [Dmi04]. stayed [BG13]. Steal [Kel11]. Steam [Fyf12, Bal15]. Steam-powered [Fyf12, Bal15]. Stefan [Mat15]. Stellar [Huf81, Huf07]. Stellar-Energy [Huf81]. Stem [Kra09a, Lan19]. Stephen [Rad12, Ris10]. Steps [Goo81]. stereochemistry [Ram95]. steroids [Sla00]. Stevenson [Han15, Wil13]. Stewart [Sei86a]. Sticky [Tea23]. Still [Ric12, Sch64, SD08, Zak13]. Stillborn [Tre75]. Stimuli [Den17]. Stokes [Zak13]. Stolen [Den23]. Stommel [VD10]. Stop [Mod08, Smo12]. Stories [How12, Wes01a]. Storm [CH16, Dun23]. Storms [Fle00a, HV06]. Story [Ham10, Shi23, BG13, McC08]. Stranded [Sch12a, Kre11]. Strange [Byc13]. Strasser [SD08]. Strategies [Nes13, Rad13, Set12, Tho12, Ros82]. Straus [Ham10, Sta08]. Strength [Bec23]. strikes [Fle00a]. String [BH08]. Strong [Bec23]. Structure [Bia12, Buc12, Das12, Gib12, GM12, GD59, Gro17, Han12, Kit12, Kra85, Lon12, Lyn12, Mor71, Rad12, Ros12a, Rüg88, Smo12, Sta12]. Structures [Pye11, Tei10]. Struggle [Phi10, Cas96, Ric08]. Struggles [Rad13, Ros82]. STS [Rad13]. Stuart [Tho84]. Students [MSD21]. Studies [And09, Far61, Lev61c, Lyn12, McC15, Mod20, Sch20, Dör06]. Study [Can75, Cur16, Dol76, Goo81, Hac86, Hof24, Hon89, Koz91, LB59, Lon08, Mar24b, Nye74, Phi10, Sch12a, She16, Wis79, All01, Dar94, Gli08, Hin03, Koh10, Wan99b, Yea03]. Studying [Fri89, Ved19]. Stupid [Cas09a]. Style [Cas79, Les08, Cas07, Neb94]. Styles [Gor20, Hol04]. Stylish [Gor20]. Subelectrons [Hol78]. Submolecularity [Mun67]. Subsequent [JJ15]. Subsidiary [JJ15]. Substance [Par48]. Substances [Far50]. Subversive [Cit15]. Success [Aga69, Shi86, Shi87, Tre75]. Suh [Wu20]. Sulfa [Str08, Les07]. Suman [Sta11, Kat05]. Summer [Ben01]. Summers [SD08]. Sumum [Lev64b]. Sunshine [Coo19]. Super [Eis16, Rio01]. Superbomb [GB89]. Superconducting [Rio01]. Superconductivity [Dah84, Dah86, Hod87]. Superior [San10]. Superpowers [TP16]. Support [Huf71]. Supporting [Fef92]. Surge [CH16]. Surmise [Mar21]. Surround [Bri15, Tur13]. Surrounding [Dah86, SF13, SF15]. Surveillance [CR16, Dry19, Fle00a]. Survey [Lev61b, Muk20, Por81]. Surveys [Rog12, Lam11]. Surviving [Wes08a, Wes10, Hop96]. Susan [Jac09, Rog12]. Swansong [Hei07]. Swastika [Hop96]. Swedish [GH17, Hal11]. Swimming [Mun21]. Swindlers [Bou12, Luc08]. Symmetry [BC91, Stu75, Kat04]. Symptoma [Dar91]. synchrocyclotron [Bur06, RM06]. Synchrotron [Cho12, Hal15, HHH15a, HHH15b, KP86, Doi09]. Synthesis [Bas25, DD59, Har09, Jac14, Rei53, Tho22, Zak12, Sla00]. Synthetic [HS18, Cro20]. Synthetical [Jac14]. System [Bad18, Bar15, Dru20, Sei86b, Gru95, Hal11, Shi79]. Systems [Cur16, Koh91, Lin09, CLW07].

T [Ams93, Har09, Lau14, Sei86a]. T. [Nie89, Nie91]. Table [Ano05d, Ano06d, Ano06e, Ano07e, Ano07f, Ano08j, Ano08k, Ano08l, Ano08m, Ano09j, Ano09k, Ano09l, Ano09m, Ano10l, Ano10i, Ano10j, Ano10k, Ano11e, Ano11f, Ano12k, Ano12l, Ano12m, Ano12n, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano14i, Ano14j, Ano14k, Ano14l, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o]. Tables [Coo19, Sch12a]. Tackles [Ros12a]. Taiwan [Ito24, Pan24]. Takats [Kim14]. Take [Huf81]. Takeshita [Kli13]. Tale [Gar09, Oni15, Wes10, De 02, Red10, Rio01, Wes02]. Talking [Shi23]. Tallahassee [Wal00]. Tallow [Low23]. Tansley [Ayr12, Hag14]. Tara [Gar09]. task [Orl98]. Tasks [Gan53]. Tauris [Har09]. Taxonomy [Str10]. Teacher [Rap60, Sch81, Aas07]. Teachers [Son53]. Teaching [Coh50, DK24, Ham24, Jun79, Koh10, Lon12, Mun21, Sta12, Yi24, LKB05, Sch12a]. Team [DK24]. Tech [Joh08, Léc06]. Technical [Bro14, Cas83, Gol13, Güt24, HS17, Mod08, PA11, Har06]. Technische [Cah82]. Technological [Cro20, Cur16, Har09, LP24, Lec92, Sim19, Ada06].

Technologies [Mar22]. Technology [Abb13, Bro14, DiM24, Gif20, Ham10, Har05, Hec14, Hop96, Jac09, Koi07, Kus18, Leg15, Lev61b, Lev64a, Mod20, Mor21a, Neu14, Rot13, Hun10, Muk09, Par07, Shi93, Wis12b, Wol13b, Kar77, Phi10, Moo16]. Technomyths [Abb13]. Technopolitics [Eis16]. Technoscience [Lee16, Tin24]. Teeth [Szc23]. Tell [How12, Wes10]. Teller [Ros22]. Telling [Sta16]. tempête [Coe10, Loc08]. Temper [Sch86]. temperature [Cha01]. Template [Dei21]. Temple [Kar77]. Temporal [Mil20, Sep18]. temps [Bel11, Coe10, Kim14, Loc08]. Tenney [Dav49, LK50]. Tenth [Lev67a]. tercera [MOBG09, Por13]. Term [ABO10]. termodinámica [Mat15, Poh11]. Terms [Kel12, Rog12, Mul12]. Terpene [Sem66]. Terrain [Wei22]. Territoriality [Bro14, Muk09]. Test [Mun12]. Testing [Fel13]. Tetu [Nis76]. Texas [Lon08]. Text [Fri89]. Textbook [Sim16]. Textbooks [Leg15, Sim11]. Texts [Lev60b, Ste88]. Their [EG80, FKW24, Her24, Huf07, Van07b]. Theodor [Vie17]. Theodore [Wu20, Str64]. Theological [Fin21]. Theophrastos [Rob59]. Theoretical [Bel18, De 02, Gal82, Sch86, Van07a]. Theories [Byk65, Cas81, Dei21, Jai66, Kai87, LP24, Vie17, BR94, Dar94, Gas96, Kat03]. Theory [Ass92b, Bar18, BC91, Can75, Dar88, Dar91, Dol76, Dor75, Eck87, Far64a, Fin21, For71, Fri77, Gar70, Gol70, Gol76, Goo81, Goo69, Hof17, Hoo49, Hun15, Jen84, Koh71, Koh74, Koh75, Kra79, LP24, Lei59, Lev61c, McC70, Men21, Oli19, Per18, Roc78, Sta11, Tre75, WK67, Wis82, Yag84, Yea12, Abi03, BV09, BR94, Dar96, Huf07, Sch03, Set04, Set10, Abi00, BK21, Cre83, Hir69, Hir76, Par59, Rad12]. Therapy [Lav14, Str08, Wat07]. There [SD08]. Thermal [Abi91]. Thermochemistry [Sch66].

Thermodynamics [Cof06, Cro88, Sch66]. ThermoPoetics [Gol10, Leg15]. Thesis [Ved19]. Thickening [LGSL^{+}22]. Things [Jia16, Lam18, McG18, dC18]. Thinking [Gal79, Har24, HS18, Mei23, Tat19]. Thinly [Eri23]. Third [Cof06, Sch66, Tho17]. Thirties [Kra82, Hen84, Sei86a]. Thomas [Fin21, Hop96, Rog12, Kli48, Mau12, Nye12, Ros12a]. Thomson [Buc77, Fal88, Fef89, Kli48, Wis79]. Thoroughbred [Tyr15]. Thought [HG71, Ric08, Phi10]. Threat [Ano09i]. Three [Bey11, Fle06, Aas95]. Thunderstorms [Kel08, Kel09]. Tidal [CH16]. tide [Ian93]. Tidepool [Lek23]. Tiemeyer [Lau14]. TikTok [Wan24]. Tilley [Fin13]. Timbuktu [Wal00]. Time [Bey11, Far61, Gal79, Gil08, OD23, Rüg88, Sch09, Sep18, Stu75, Wu20, Zak13, Bel11, Chr11, Ric07, Rai10, Seo21]. Time-Honored [Wu20]. Time/Travel [Seo21]. Times [Ama20, CE81]. Tissue [Gro23]. Tobacco [How12, OC10]. Tocqueville [Bou12]. Today [Oli19]. Together [Bri15, Wes01a]. Token [Den17]. Tolerance [Fri07]. Tomás [Por13]. tomber [Ore00]. Tong [Rog12]. Took [Elm12]. Tool [Bia12, Did18, Ten22]. TOPEX [Lam18]. TOPEX/POSEIDON [Lam18]. topology [Dar98b]. Torbern [Sch67]. Torque [Vel22]. Tortoise [Ven24]. tortuous [KW96]. Tory [Gor14a]. totalitarian [Koj99a]. Touch [Lau14, Oli13]. Townsend [Hon89]. Toxic [Fle20]. TPC [Wes01a]. Tracers [Koh77]. Tracing [Hyu19, Vaz23]. Trade [NVHB24, Hec12, Wel12]. trading [Tur06]. Tradition [Ass92b, Jia16, Sle25]. Traffic [Ven24]. Trafficking [Ren09, San10]. Traffics [NVHB24]. tragedy [Bud07, Str08]. Tragic [Phi10, Ric08]. Train [Gue22]. Training [Mun21, Sch81]. Traité [Nev65]. Trans [Joh08, Rai10, Sta08]. Transcripts [Hop96]. Transfer [Gif20, Kra09a]. Transformation [Can78, Gol13, KR18, Mel93, Sei83, Gli08, Har11, Phi10]. Transformations [Lee18, NVHB24, GG50]. Transformed [Str08, Wu20, Les07]. Transfusions [Sch12a, Kre11]. Transistor [Hod81, Les01]. Transition [Oli09]. transitional [Kel96]. Translation [Arb19, LK24, Phi10, Gli08]. Transmission [CS78]. Transmitting [Tal16]. Transnational [Aub20]. Transnationalism [TBV20]. Transplantation [Kra09a]. Transport [Güt24, Hey06]. Transubstantiation [Lin09]. Travel [Seo21]. Travelers [Bar16]. Traveling [Bar16]. Trends [Aro17, AA95]. Trespasser [Par22]. Triads [Mor09, Jac06]. Trials [Gor17]. Tribute [Dei48]. Trieste [De 02]. Trinity [FV13, Oni15]. Tripos [Rob80, Wil82]. Triumph [Han15, Rei87, Str08, Bud07, Wil13]. Trondheim [Wit04]. Trouble [AB23]. Truman [Ham10, Gor09]. Truth [How12, Jac09, DPPdD11, Lin05, OC10, Oni15]. Tswett [Rob60]. Tumors [San15]. Turbulence [Dar02]. Turn [Nye74, Set11]. turned [Ahe03]. Turner [Bri15]. Tuve [Wan12]. Twentieth [Bar16, Bor23, Cro20, Dör12, Ito24, Lau14, MJ15, Nes13, Rog12, Sch12a, Sta16, Str08, Wal15a, Bow09, Gep11, Jac12, Rhe10a, Sch08a, Dol76, Hac86, Hol18, Ker13, Rob20, Sch09, Tea23, Ven24]. Twentieth-Century [Dör12, MJ15, Nes13, Rog12, Sch12a, Ito24, Bow09, Rhe10a, Sch08a, Hac86, Ker13, Rob20, Sch09, Tea23, Ven24]. Two [Hon89, Jen84, Kel09, Rus86, Wit21, De 02, Kim06, Rio01, Wes02]. Tyndall [Leg15, DeY11]. Types [And20, Hon89].

U.K [Zak13]. U.S [Mod08]. U.S. [Ake07, Ben12, Cas94, Fle00a, Par01, Rei95, Wan99b]. U.S.A. [Van49]. UA1 [Wes01a]. Ubiquitous [Abb13]. Unbound [Luk24]. unburied [Fab10, Nov15]. uncertainties [HV06]. Uncertainty [Nye09, Sta11, Cas09b, Coe07, Hey06, Mar12, Bra15, Hop96]. Understand [Zak13, Rus11a]. Understanding [Cas86, Fri89]. Unification [Hop96, Cat98, Jur02]. unified [Asp80]. Union [Bac65, Hol99]. United [Abb13, Dim14, Kli13, BV01, DIR91, For87, Hac00, Kev97, Nov12, Pre11, Rid84, Sch86, Son53, Tyr15]. Uniting [Zak13, Rus11a]. Universality [TP16]. Universe [Gol13, Het82, PA11, De 12, Sta08]. Universidad [Mat15, Por13, San10, Her08]. Universitaires [Coe10]. Universitat [Mat15]. Universities [GG88, GH17]. University [And09, Bal15, Bra15, Bro14, Coe10, Cro20, Fin13, Fle20, Gar09, Gil08, Ham10, Har09, Hec14, Jac09, Joh08, Kev09, Kim14, Kli13, Lau14, Lec92, Lin09, Lon08, Mod08, Mor09, Nee87, Nov15, Nye09, Phi10, Por13, Rad13, Rai10, Ris10, Rog12, San10, Sch12a, Sla12b, Sta08, Str08, SD08, Ved20, Wal15a, Wol13a, Wu20, Zak13, Lev13, Fef92, GS11, Her08, MC13, Slo14, Tho84, Vet04, Zul09]. University-Based [Nee87]. universo [De 12, Mat15]. Unknown [DD59]. Unmaking [MW18]. Unmechanischer [Ser77]. Unpublished [Bue53, Fra81, Nev62, Sol61, Sol65]. Unrealized [Mod08, Kir05]. unruly [Von01]. Unsaid [Ala23]. Untwisting [Gar09]. Upright [Hop96]. upsilon [Neb94]. Uranium [Wel12, Hec12]. Urbana [Phi10]. Uribe [Por13]. Use [Dav48, FKW24, Ker13, FKW24]. Useful [Mil90, Sch12a]. Users [Kli13, Tak12]. Uses [Kat19].

V [Ric12, Sch88, DeV87]. V-2 [DeV87]. Vaccine [Var21]. Vagner [Sem66]. Valdivieso [Por13]. Valence [Koh75, WK67, Koh74]. Valentini [Sta11]. Valero [Mat15]. Valley [Joh08, Léc06, Les01]. Value [Byc13].

Values [Het82, Ito02]. Vannevar [Mir21, Rei87]. Vantage [MJ15]. Variable [Kau67b]. Variances [Löw90c]. Variation [dC20]. Velocity [Pot23, Abi03]. Velocity-Dependency [Pot23]. Velvet [Cho12, Doi09]. Venant [Dar01]. vendetta [Cra00]. Venerella [Rai10]. Venial [Qué88]. Ventilator [Wal21]. Verne [Han15, Wil13]. vers [Dau48]. versa [CG96]. versus [Gas96, Hon89, SS18, Tur82]. Vertebrate [Sun12]. Verticality [Rei17a]. Vettel [Mod08]. via [FKW24]. viaje [BG10, Por13]. vice [CG96]. vice-versa [CG96]. Vichy [BT10]. Victorian [Bal15, Leg15, Mor09, DeY11, Gol10, Hen14, Ott08, Rob76, Sec14, Smi06, ST70]. Vieira [GP17]. Vienna [Coe07, Nye09, Ren09, San10]. View [Dor75, GG50, Kat05, Set04, Set05]. Viewpoint [Coo12]. Vignettes [CA14]. vii [Kli13, Lau14]. viii [Bra15, Wal15a]. Vincent [Vie17]. Vine [Fra82]. Violation [Fra83]. Virtues [Tai17]. Virus [Can21]. Visible [Ble12, Pan24, Por13]. Vision [Leg15, DeY11, Ott08]. Visioneers [Dim14, McC3 ]. Visions [Hec14, Sec14, Wis12b, Bal15]. Visited [Kle19]. Visits [Nes13]. Vista [McC04]. Visual [And09, Cha11, Gal79, Han15, Kwa18, Mor09, Pan97, Por13, ST18, Ble12, Hun13, Smi06]. Visualization [Sep18, Von18]. Visualizing [Gis11, Por13]. Vitalism [Che24]. Vito [Sig92]. Vivian [Dei64]. Vogt [Nyh17]. vol [Bro14]. Volcanology [Dör06]. Volhard [Kel96]. Volta [Pan90a]. Voltaic [Sut81]. Volterra [Sig92]. Volumetric [Sco50]. Vorgeschichte [Pra53]. Vorlesung [Rhe50]. Vorm [Oni15]. vs [Kin13, Kwa18].

W [Dim14, Ham10, Lon08, Mor09]. W. [Esp15]. Waals [Gav90]. Wahshiya [Lev64b]. Walk [Bau21]. Walker [Hop96]. Wall [Fre81]. Wallace [Cre83]. Walter [Sta08]. Wandlungen [GG50]. Wang [Ham10, Oni15]. War [Aas99, Bri15, Cas01, Dah86, Dev00, Ham10, Hec14, Hod93, Hop96, Jon18, Kai02, KR18, Mir21, Mod08, Mun03, Oli19, Ric17a, Shi19, Szc24, Tei10, Tur13, Wan08, Wol13a, Wol13b, AB98, Ake07, Aro17, Bru03, Cul10, Dör20, Fre05, Hou97, Kai02, Kev90, Kri99, Kri05, Lee16, Leo09, Lev13, LK24, Mac24, Mar22, MK20, SR02a, San10, Sei01, Shi16, Sla12a, Sla12b, Tat19, Tur20, VD10, Wal00, Wan95b, Wan99b, Wan12, Wan20, Wea97, Wil15]. Warburg [Wer97]. Warfare [Hac86]. Warming [OCS08, OC10, Wea97, How12]. Wars [Rus86]. wartime [Ito02, SR02b]. Was [For49, Liu19, RF69, vSvS66, Aas95]. Washington [Mod08, Oni15, Ser83]. Waste [San10, Ham08]. Wataghin [TBV20]. Watching [Ham10, McC08]. Water [Cha12, DD59, Ede48, Gor14a, Hel12, Her08, Sch64]. waterwheels [Dar01]. Watkins [Str08]. Watt [Leg15, Hun10]. Wave [Can75, Per18, Rüg88, Van07a, Yea03]. wave-particle [Van07a]. wavefront [Joh05]. Waves [D’A75, Dar86, Goo84]. Way [Far66, Gol13, HS18, MM75, Tai17]. Ways [Bey11]. Weapon [Kri08, Mun12]. Weapons [Gil92, Tho22]. Weather [Coe10, Har08, Fle06, Gol07, Coe10]. Weatherwiser [Coe10]. Webs [Col24]. Website [FKW24]. Wedgwood [Sch59]. Weighing [Hei93]. Weights [vSvS66, vS67]. Weimar [For71, MK20]. Weindling [Hop96]. Weiss [Qué88]. Well [San10, Ham08]. Weltzein [dM48]. Were [Gor20]. Werner [Hop96, Cah82, Kau60, Kau67a, Kau67b, Kau67c, Wal92]. West [Fel13, Car99, Cas96, FH11]. Westfall [Ham10]. Westwick [Joh08, Lau14]. Wheat [Sar10a]. Whewell [Bec80]. Which [Lei48, Mod20]. Whirling [Hun86]. white [Joh05, Lon08]. Whiteness [Szc23]. Who [Dev00, MD03, RF69, BG13, Mun03]. Whole [Bru87, Zak12]. Wicked [Hun13, Han15]. Wieland [Wer97]. Wien [Wol03]. Wild [Moo16]. Wildlife [Koh11, Szc24]. Wiley [Zak13]. Wiley-Blackwell [Zak13]. Wilhelm [Gau10, Hil49, Wol03]. will [Gol00, Mun03]. Wille [Kev09]. William [Bal15, Lau14, SD08, Bec80, Buc77, Cre83, Fyf12, OCS08, Rei49, Siv62, Wis79]. Williamson [Pet16]. Wilson [Ahn22, Por13, Sei86a, Bas25, Ber19, Sea25, Wei25]. Wilsonian [Men21]. Win [Cam22, Les87]. wins [Kev94]. Winter [Ros22]. Wireless [Nav16]. Wise [Lin09]. Wisnioski [Hec14]. Wissenschaften [Pra49]. Within [Bal15, Hen14, LGSL^{+}22]. Without [Cam20, Lin09, CLW07]. Wolf [Tab22]. Wolfe [Hec14]. Wollaston [Goo69]. Woman [Ham10]. Women [Kli13, Rad13, Ric17b, Ros82, Ros95, Ros12b, Ore00, Tak12, Rad13]. Wonder [Str08]. Wonderful [Mil10a]. Wonders [Byc13]. Word [Ros18]. Wordplay [Bar18]. Work [Col24, Cro88, Fal88, Gav90, LO48, Nie89, Nie91, Pal89, Rüg88, Str64, WS86, Kel96, Lel05, Ore00]. Worker [Gas21]. Workers [Ahn22]. Working [Bro14, Isa12, Sta16, Bri15]. Works [Bia22, GP17, Bue53]. Workshops [Tea23, FG98, FG99]. World [Bri15, Bub19, Cah88, Dol76, Gal79, Jac09, Kel12, Kli13, Mod08, Rad13, TP16, Wal15a, Wil15, Wol13a, dCP18, Ake07, CR09, Cul10, Dea07, For08, Kat05, KH93, Ros12b, Set04, Set05, Wil13, Aas99, Cas01, Dah86, Dev00, Hod93, Jon18, Kai02, KR18, Nov15, Por13, Ric17a, Ris10, Rus86, Tei10, Tur13, WG10, Han15]. world-view [Kat05, Set04, Set05]. Worlds [Rud08, Sta16, Vie17, Rai10]. Worldview [Hir76, Set20]. Worrying [Mod08, Smo12]. Worthy [Wol13a, Rus13]. Wounds [Sem23]. Wrenching [Vel22]. Writing [Sim16, TAR^{+}20]. Wrote [Joh08, BG13]. WWF [Cam20]. www.gutenberg-e.org [San10]. x [Bra15, Har09, Jac09, Kev09, Kim14, Lau14, Mor09, Nye09, San10, Sta08, Het87, Hog15, Whe81]. X-Ray [Het87]. X-Rays [Whe81]. Xavier [Mat15]. xi [Ham10, Har09, Kli13, Lon08, Mod08, Nye09, Ris10, Wal15a]. xii [Bra15, Bro14, Fin13, Jac09, Kev09, Lau14, Phi10, Por13, Ris10, San10, Wal15a, Wol13a]. xiii [Bal15, Fin13, Fle20, Ham10, Kev09, Rai10]. xiv [Fin13, Har09, Lon08, Phi10, Rai10, Sch12a, Wal15a, Wol13a, Zak13]. xix [Bro14, Dea08, Nye09, Poh11]. XIXe [Coe10, Loc08]. Xu [Mil99]. xv [Coe10, Lin09, Mod08, Wu20, Zak13]. xvi [Bal15, Bra15, Lon08, Sch12a, Zak13]. xvii [Gar09, Mod08]. xviii [Gar09, Rad13, Rai10]. xx [Phi10, Rad13, Mat15]. xxi [Gor14a, Zak13]. xxii [Fle20, Rai10, Sta08]. xxiii [Mor09]. xxiv [Gil08]. xxix [Lau14, Mod08]. xxvii [Ham10, Zak13]. xxxiii [SD08]. xxxiv [Sta08]. xxxix [Rai10].

Yakobi [Sol65]. Yakov [Fre97]. Yale [Gar09, Joh08, Lon08, SD08, Ved20]. Year [Ame22, ABO10]. Years [Ber88, Dah84, GM12, Sea25, vSvS66]. Yeast [GSLV^{+}22b]. Yellow [Hen19]. Yes [SD08]. Yip [Mac24]. Yoo [Wu20]. York [Bra15, Ham10, Kli13, Lau14, Oni15, San10, Sch12a, Sta08, Wol13a, Zak13]. Yoshio [Ito02, Kim06]. Young [Sch90]. youthful [Huf07]. Yves [Ved20].

Z [Kin11]. Zealand [Bra15, Pri12]. Zeeman [For70]. Zodiac [Kin11, BJ10]. Zoo [Güt24, Ven24]. Zoology [Sun12]. Zukunftsaufgaben [Gan53]. Zunz [Wol13a]. Zuoxiu [Mil99]. Zuoyue [Ham10]. Zur [Pra53]. Zwang [Ser77].

References

Abir-Am:1995:NTH

[AA95] Pnina Abir-Am. ’new’ trends in the history of molecular biology. Historical Studies in the Physical and Biological Sciences, 26 (1):167–196, ???? 1995. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Aaserud:1995:SPT

[Aas95] Finn Aaserud. Sputnik and the ‘Princeton three:’ the national security laboratory that was not to be. Historical Studies in the Physical and Biological Sciences, 25(2):185–239, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Aaserud:1999:SSN

[Aas99] Finn Aaserud. The scientist and the statesman: Niels Bohr’s political crusade during World War II. Historical Studies in the Physical and Biological Sciences, 30(1):1–47, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Aaserud:2007:RMT

[Aas07] Finn Aaserud. Russell Mc-Cormmach as a teacher. Historical Studies in the Physical and Biological Sciences, 37(2): 453–462, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Agar:1998:BSC

[AB98] Jon Agar and Brian Balmer. British scientists and the Cold War: The Defence Research Policy Committee and information networks, 1947–1963. Historical Studies in the Physical and Biological Sciences, 28(2): 209–252, ???? 1998. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Armstrong:2023:TSA

[AB23] Eleanor S. Armstrong and Jordan Bimm. The trouble with space auctions. Historical Studies in the Natural Sciences, 53(4):425–??, September 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/4/425/197390/The-
Trouble-with-Space-
Auctions.

Abbate:2013:BRK

[Abb13] Janet Abbate. Book reviews: Killer apps and technomyths: Divining a Digital Future: Mess and Mythology in Ubiquitous Computing Histories of Computing Cybernetic Revolutionaries: Technology and Politics in Allende’s Chile Biomedical Computing: Digitizing Life in the United States Simulation and Its Discontents. Historical Studies in the Natural Sciences, 43(1):105–112, February 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.1.105.pdf.

Abiko:1991:CTO

[Abi91] Seiya Abiko. On the chemicothermal origins of Special Relativity. Historical Studies in the Physical and Biological Sciences, 22(1):1–24, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757671.

Abiko:2000:EKA

[Abi00] Seiya Abiko. Einstein’s Kyoto address: ‘how I created the Theory of Relativity’. Historical Studies in the Physical and Biological Sciences, 31(1):1–35, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Abiko:2003:EDE

[Abi03] Seiya Abiko. On Einstein’s distrust of the electromagnetic theory: The origin of the lightvelocity postulate. Historical Studies in the Physical and Biological Sciences, 33(2):193– 215, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See comments [Dar04] and reply [Abi04].

Abiko:2004:ROD

[Abi04] Seiya Abiko. Reply to Olivier Darrigol. Historical Studies in the Physical and Biological Sciences, 35(1):157–160, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See [Abi03, Dar04].

Aronova:2010:BSB

[ABO10] Elena Aronova, Karen S. Baker, and Naomi Oreskes. Big science and big data in biology: From the International Geophysical Year through the International Biological Program to the Long Term Ecological Research (LTER) Network, 1957– present. Historical Studies in the Natural Sciences, 40(2):183– 224, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Abrams:2020:INO

[Abr20] Ellen Abrams. “Indebted to no one”: Grounding and gendering the self-made mathematician. Historical Studies in the Natural Sciences, 50(3):217–247, June 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/3/217/110324/Indebted-
to-No-One-Grounding-and-
Gendering-the.

Adas:2006:DDT

[Ada06] Michael Adas. Dominance by design: technological imperatives and America’s civilizing mission. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 2006. ISBN 0-674-01867-2, 0-674-03216-0. 542 pp. LCCN E744 .A243 2006eb. URL http:

//site.ebrary.com/id/
10312789.

Agassi:1969:SJH

[Aga69] Joseph Agassi. Sir John Herschel’s philosophy of success. Historical Studies in the Physical Sciences, 1(??):1–36, ???? 1969. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757293.

Ahern:2003:WHT

[Ahe03] Joseph-James Ahern. ’we had the hose turned on us!’: Ross Gunn and the Naval Research Laboratory’s early research into nuclear propulsion, 1939–1946. Historical Studies in the Physical and Biological Sciences, 33 (2):217–236, March 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Ahn:2022:FIW

[Ahn22] Eun-Joo Ahn. Finding the invisible workers in astronomy: The case of Mount Wilson Observatory, 1900–1930. Historical Studies in the Natural Sciences, 52(5):555–??, November 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/5/555/194655/Finding-
the-Invisible-Workers-in-
AstronomyThe-Case.

Akera:2007:CNW

[Ake07] Atsushi Akera. Calculating a natural world: scientists, engineers, and computers during the rise of U.S. Cold War research. Inside technology. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-01231-6. ix + 427 pp. LCCN QA76.17 .A42 2007.

Anderson:2020:PBH

[AL20] Warwick Anderson and M. Susan Lindee. Pacific biologies: How humans become genetic. Historical Studies in the Natural Sciences, 50(5):483–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

                                        online.ucpress.edu/hsns/
                                        article/50/5/483/114252/
                                        Pacific-Biologies-How-
                                        Humans-Become-Genetic.

Alam:2023:CU

[Ala23] Eram Alam. Citing the unsaid. Historical Studies in the Natural Sciences, 53(1):73–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        53/1/73/195535/Citing-the-
                                        Unsaid.

Allen:2001:BBC

[All01] Garland E. Allen. The biological basis of crime: an historical and methodological study. Historical Studies in the Physical and Biological Sciences, 31 (2):183–222, March 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Almeling:2011:SCM

[Alm11] Rene Almeling. Sex cells: the medical market for eggs and sperm. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-27095-9 (hardcover), 0-520-27096-7 (paperback). x + 228 pp. LCCN HQ761 .A46 2011eb.

Alcalde-Mongrut:1964:MRP

[AM64] Arturo Alcalde-Mongrut. Mariano de Rivero, pioneer of mining education in South America. Chymia: Annual Studies in the History of Chemistry, 9 (??):77–95, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757234.

Amado:2020:STM

[Ama20] Eduardo Díaz Amado. Shifting times: Medicine, society, and life. Historical Studies in the Natural Sciences, 50(1–2):183– 198, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Amery:2022:CNL

[Ame22] Fiona Amery. Capturing the Northern Lights: Standardizing the practice of auroral photography during the Second International Polar Year, 1932–1933. Historical Studies in the Natural Sciences, 52(2):147–??, April 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/2/147/144555/Capturing-
the-Northern-
LightsStandardizing-the.

Amsterdamska:1993:PDR

[Ams93] Olga Amsterdamska. From pneumonia to DNA: The research career of Oswald T. Avery. Historical Studies in the Physical and Biological Sciences, 24(1):1–40, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://

profiles.nlm.nih.gov/CC/A/
A/P/S/_/ccaaps.pdf.

Anderson:2009:BRE

[And09] Nancy Anderson. Book reviews: Eye and Image: Looking at a Visual Studies of Science: Lorraine Daston and Peter Galison. Objectivity. Cambridge, MA: MIT Press, 2007. 500 pp., 140 illus., 32 color, index. ISBN 978-1-890951-78-8. $38.95 (cloth). Jutta Schickore. The Microscope and the Eye: A History of Reflections, 1740– 1870. Chicago: University of Chicago Press, 2007. 320 pp., 16 halftones, index. ISBN 978-0-226737-84-3. $40 (cloth). Historical Studies in the Natural Sciences, 39(1):115–125, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anderson:2020:RTA

[And20] Warwick Anderson. From racial types to Aboriginal Clines: The illustrative career of Joseph B. Birdsell. Historical Studies in the Natural Sciences, 50(5):498– ??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/50/5/498/
114251/From-Racial-Types-
to-Aboriginal-ClinesThe.

Anker:2010:BEH

[Ank10] Peder Anker. From Bauhaus to ecohouse: a history of ecological design. Louisiana State University Press, Baton Rouge, LA, USA, 2010. ISBN 0-8071-3551-8 (hardcover). viii + 188 pp. LCCN NA2542.35 .A54 2010.

Anonymous:1986:FM

[Ano86a] Anonymous. Front matter. Historical Studies in the Physical and Biological Sciences, 17 (1):??, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757573.

Anonymous:1986:SB

[Ano86b] Anonymous. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 16(1):189–195, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757562.

Anonymous:2005:C

[Ano05a] Anonymous. Cover. Historical Studies in the Physical and Biological Sciences, 36 (1):C1, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2005:FM

[Ano05b] Anonymous. Front matter. Historical Studies in the Physical and Biological Sciences, 36(1): ii–v, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2005:NPE

[Ano05c] Anonymous. Neutron physics in the early 1930s. Historical Studies in the Physical and Biological Sciences, 36(1):207, September 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2005:TC

[Ano05d] Anonymous. Table of contents. Historical Studies in the Physical and Biological Sciences, 36 (1):i, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2006:C

[Ano06a] Anonymous. Cover. Historical Studies in the Physical and Biological Sciences, 36 (2):C1, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2006:FMa

[Ano06b] Anonymous. Front matter. Historical Studies in the Physical and Biological Sciences, 36 (2):C2, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2006:FMb

[Ano06c] Anonymous. Front matter. Historical Studies in the Physical and Biological Sciences, 37(1): ii–iv, September 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2006:TCa

[Ano06d] Anonymous. Table of contents. Historical Studies in the Physical and Biological Sciences, 36(2):i, March 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2006:TCb

[Ano06e] Anonymous. Table of Contents.

Historical Studies in the Physical and Biological Sciences, 37 (1):i, September 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2007:Ca

[Ano07a] Anonymous. Cover. Historical Studies in the Physical and Biological Sciences, 37 (2):C1–C4, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2007:Cb

[Ano07b] Anonymous. Cover. Historical Studies in the Physical and Biological Sciences, 37(S1): C1, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.cover.pdf.

Anonymous:2007:F

[Ano07c] Anonymous. Foreword. Historical Studies in the Physical and Biological Sciences, 37(S1): 1, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.1.pdf.

Anonymous:2007:FM

[Ano07d] Anonymous. Front matter. Historical Studies in the Physical and Biological Sciences, 37(S1): C2–ii, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.fm.pdf.

Anonymous:2007:TCa

[Ano07e] Anonymous. Table of contents. Historical Studies in the Physical and Biological Sciences, 37 (2):1–5, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Anonymous:2007:TCb

[Ano07f] Anonymous. Table of contents.

Historical Studies in the Physical and Biological Sciences, 37 (S1):i, September 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.toc.pdf.

Anonymous:2008:Ca

[Ano08a] Anonymous. Cover. Historical Studies in the Natural Sciences, 38(1):C1–C4, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:Cb

[Ano08b] Anonymous. Cover. Historical Studies in the Natural Sciences, 38(2):C1–C4, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:Cc

[Ano08c] Anonymous. Cover. Historical Studies in the Natural Sciences, 38(3):C1–C4, August 2008. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:Cd

[Ano08d] Anonymous. Cover. Historical Studies in the Natural Sciences, 38(4):C1–C4, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:FMa

[Ano08e] Anonymous. Front matter. Historical Studies in the Natural Sciences, 38(1):C2–ii, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:FMb

[Ano08f] Anonymous. Front matter. Historical Studies in the Natural Sciences, 38(2):C2–ii, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:FMc

[Ano08g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 38(3):C2–ii, August 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:FMd

[Ano08h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 38(4):C2–ii, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:RLJ

[Ano08i] Anonymous. On re-launching a journal. Historical Studies in the Natural Sciences, 38(1):1–3, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:TCa

[Ano08j] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 38(1):i, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:TCb

[Ano08k] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 38(2):i, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:TCc

[Ano08l] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 38(3):i, August 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2008:TCd

[Ano08m] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 38(4):i, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:Ca

[Ano09a] Anonymous. Cover. Historical Studies in the Natural Sciences, 39(1):C1–C4, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:Cb

[Ano09b] Anonymous. Cover. Historical Studies in the Natural Sciences, 39(2):C1–C4, May 2009.

CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:Cc

[Ano09c] Anonymous. Cover. Historical Studies in the Natural Sciences, 39(3):C1–C4, August 2009. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:Cd

[Ano09d] Anonymous. Cover. Historical Studies in the Natural Sciences, 39(4):C1–C4, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:FMa

[Ano09e] Anonymous. Front matter. Historical Studies in the Natural Sciences, 39(1):C2–ii, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:FMb

[Ano09f] Anonymous. Front matter. Historical Studies in the Natural Sciences, 39(2):C2–ii, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:FMc

[Ano09g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 39(3):C2–ii, August 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:FMd

[Ano09h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 39(4):C2–ii, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:JUT

[Ano09i] Anonymous. Journals under threat: a joint response from HSTM Editors. Historical Studies in the Natural Sciences, 39(1):126–128, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:TCa

[Ano09j] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 39(1):i, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:TCb

[Ano09k] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 39(2):i, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:TCc

[Ano09l] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 39(3):i, August 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2009:TCd

[Ano09m] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 39(4):i, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:Ca

[Ano10a] Anonymous. Cover. Historical Studies in the Natural Sciences, 40(1):C1–C4, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:Cb

[Ano10b] Anonymous. Cover. Historical Studies in the Natural Sciences, 40(2):C1–C4, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:Cc

[Ano10c] Anonymous. Cover. Historical Studies in the Natural Sciences, 40(3):C1–C4, August 2010. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:Cd

[Ano10d] Anonymous. Cover. Historical Studies in the Natural Sciences, 40(4):C1–C4, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2010.
40.4.cover.

Anonymous:2010:FMa

[Ano10e] Anonymous. Front matter. Historical Studies in the Natural Sciences, 40(1):C2–ii, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:FMb

[Ano10f] Anonymous. Front matter.

Historical Studies in the Natural Sciences, 40(2):C2–ii, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:FMc

[Ano10g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 40(3):C2–ii, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:FMF

[Ano10h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 40(4):C2–ii, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2010.
40.4.fm.

Anonymous:2010:TCa

[Ano10i] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 40(1):i, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:TCb

[Ano10j] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 40(2):i, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:TCc

[Ano10k] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 40(3):i, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Anonymous:2010:TCT

[Ano10l] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 40(4):i, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2010.
40.4.toc.

Anonymous:2011:Ca

[Ano11a] Anonymous. Cover. Historical Studies in the Natural Sciences, 41(1):C1–C4, Winter 2011. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.1.cover.

Anonymous:2011:Cb

[Ano11b] Anonymous. Cover. Historical Studies in the Natural Sciences, 41(2):C1–C4, Spring 2011. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.2.cover.

Anonymous:2011:FMa

[Ano11c] Anonymous. Front matter. Historical Studies in the Natural Sciences, 41(1):C2–ii, Winter 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.1.fm.

Anonymous:2011:FMb

[Ano11d] Anonymous. Front matter.

Historical Studies in the Natural Sciences, 41(2):c2–ii, Spring 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.2.fm.

Anonymous:2011:TCa

[Ano11e] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 41(1):i, Winter 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.1.toc.

Anonymous:2011:TCb

[Ano11f] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 41(2):i, Spring 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.2.toc.

Anonymous:2012:BMa

[Ano12a] Anonymous. Back matter. Historical Studies in the Natural Sciences, 42(4):??, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.4.BM.pdf.

Anonymous:2012:BMb

[Ano12b] Anonymous. Back matter. Historical Studies in the Natural Sciences, 42(5):??, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.BM.pdf.

Anonymous:2012:Ca

[Ano12c] Anonymous. Cover. Historical Studies in the Natural Sciences, 42(2):??, April 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.2.COVER.pdf.

Anonymous:2012:Cb

[Ano12d] Anonymous. Cover. Historical Studies in the Natural Sciences, 42(3):??, June 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.3.COVER.pdf.

Anonymous:2012:Cc

[Ano12e] Anonymous. Cover. Historical Studies in the Natural Sciences, 42(4):??, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.4.COVER.pdf.

Anonymous:2012:Cd

[Ano12f] Anonymous. Cover. Historical Studies in the Natural Sciences, 42(5):??, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.COVER.pdf.

Anonymous:2012:FMa

[Ano12g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 42(2):??, April 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.2.FM.pdf.

Anonymous:2012:FMb

[Ano12h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 42(3):??, June 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.3.FM.pdf.

Anonymous:2012:FMc

[Ano12i] Anonymous. Front matter. Historical Studies in the Natural Sciences, 42(4):??, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.4.FM.pdf.

Anonymous:2012:FMd

[Ano12j] Anonymous. Front matter. Historical Studies in the Natural Sciences, 42(5):??, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.FM.pdf.

Anonymous:2012:TCa

[Ano12k] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 42(2):??, April 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.2.TOC.pdf.

Anonymous:2012:TCb

[Ano12l] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 42(3):??, June 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.3.TOC.pdf.

Anonymous:2012:TCc

[Ano12m] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 42(4): ??, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.4.TOC.pdf.

Anonymous:2012:TCd

[Ano12n] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 42(5): ??, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.TOC.pdf.

Anonymous:2013:Ca

[Ano13a] Anonymous. Cover. Historical Studies in the Natural Sciences, 43(1):??, February 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.1.cover.pdf.

Anonymous:2013:CIb

[Ano13b] Anonymous. Cover. Historical Studies in the Natural Sciences, 43(2):??, April 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.2.cover.pdf.

Anonymous:2013:CIc

[Ano13c] Anonymous. Cover. Historical Studies in the Natural Sciences, 43(3):c1–c2, June 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.3.cover.pdf.

Anonymous:2013:Cb

[Ano13d] Anonymous. Cover. Historical Studies in the Natural Sciences, 43(4):c1–c2, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2013.43.4.cover.

Anonymous:2013:Cc

[Ano13e] Anonymous. Cover. Historical Studies in the Natural Sciences, 43(5):c1–c2, November 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2013.43.5.cover.

Anonymous:2013:FMa

[Ano13f] Anonymous. Front matter. Historical Studies in the Natural Sciences, 43(1):??, February 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.1.fm.pdf.

Anonymous:2013:FMIb

[Ano13g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 43(2):??, April 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.2.fm.pdf.

Anonymous:2013:FMIc

[Ano13h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 43(3):i–ii, June 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.3.fm.pdf.

Anonymous:2013:FMb

[Ano13i] Anonymous. Front matter. Historical Studies in the Natural Sciences, 43(4):i–ii, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2013.43.4.fm.

Anonymous:2013:FMc

[Ano13j] Anonymous. Front matter. Historical Studies in the Natural Sciences, 43(5):i–ii, November 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2013.43.5.fm.

Anonymous:2013:TCa

[Ano13k] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 43(1): ??, February 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.1.toc.pdf.

Anonymous:2013:TCb

[Ano13l] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 43(2):??, April 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.2.toc.pdf.

Anonymous:2013:TCc

[Ano13m] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 43(3):iii, June 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.3.toc.pdf.

Anonymous:2013:TCd

[Ano13n] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 43(4):iii, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2013.43.4.toc.

Anonymous:2013:TCe

[Ano13o] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 43(5):iii, November 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2013.43.5.toc.

Anonymous:2014:Ca

[Ano14a] Anonymous. Cover. Historical Studies in the Natural Sciences, 44(1):c1–c2, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2014.44.1.cover.

Anonymous:2014:Cb

[Ano14b] Anonymous. Cover. Historical Studies in the Natural Sciences, 44(2):c1–c2, April 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.2.cover.

Anonymous:2014:Cc

[Ano14c] Anonymous. Cover. Historical Studies in the Natural Sciences, 44(3):c1–c2, June 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.3.cover.

Anonymous:2014:CI

[Ano14d] Anonymous. Cover. Historical Studies in the Natural Sciences, 44(4):c1–c2, September 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2014.44.4.cover.

Anonymous:2014:FMa

[Ano14e] Anonymous. Front matter. Historical Studies in the Natural Sciences, 44(1):i–iii, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2014.44.1.fm.

Anonymous:2014:FMb

[Ano14f] Anonymous. Front matter. Historical Studies in the Natural Sciences, 44(2):i–iii, April 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.2.fm.

Anonymous:2014:FMc

[Ano14g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 44(3):i–iii, June 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.3.fm.

Anonymous:2014:FMd

[Ano14h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 44(4):i–ii, September 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2014.44.4.fm.

Anonymous:2014:TCa

[Ano14i] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 44(1):ii, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2014.44.1.toc.

Anonymous:2014:TCb

[Ano14j] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 44(2):ii, April 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.2.toc.

Anonymous:2014:TCc

[Ano14k] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 44(3):ii, June 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.3.toc.

Anonymous:2014:TCd

[Ano14l] Anonymous. Table of contents.

Historical Studies in the Natural Sciences, 44(4):iii, September 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2014.44.4.toc.

Anonymous:2015:Ca

[Ano15a] Anonymous. Cover. Historical Studies in the Natural Sciences, 45(1):c1–c2, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.1.cover.

Anonymous:2015:Cb

[Ano15b] Anonymous. Cover. Historical Studies in the Natural Sciences, 45(2):c1–c2, April 2015.

CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.2.cover.

Anonymous:2015:Cc

[Ano15c] Anonymous. Cover. Historical Studies in the Natural Sciences, 45(3):c1–c2, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.3.cover.

Anonymous:2015:Cd

[Ano15d] Anonymous. Cover. Historical Studies in the Natural Sciences, 45(4):C1–C2, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.4.cover.

Anonymous:2015:Ce

[Ano15e] Anonymous. Cover. Historical Studies in the Natural Sciences, 45(5):C1–C2, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.5.cover.

Anonymous:2015:FMa

[Ano15f] Anonymous. Front matter. Historical Studies in the Natural Sciences, 45(1):i–ii, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.1.fm.

Anonymous:2015:FMb

[Ano15g] Anonymous. Front matter. Historical Studies in the Natural Sciences, 45(2):i–ii, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.2.fm.

Anonymous:2015:FMc

[Ano15h] Anonymous. Front matter. Historical Studies in the Natural Sciences, 45(3):i–iii, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.3.fm.

Anonymous:2015:FMd

[Ano15i] Anonymous. Front matter. Historical Studies in the Natural Sciences, 45(4):i–iii, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.4.fm.

Anonymous:2015:FMe

[Ano15j] Anonymous. Front matter. Historical Studies in the Natural Sciences, 45(5):i–iii, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.5.fm.

Anonymous:2015:TCa

[Ano15k] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 45(1):iii, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2015.45.1.toc.

Anonymous:2015:TCb

[Ano15l] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 45(2):iii, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2015.
                                        45.2.toc.

Anonymous:2015:TCc

[Ano15m] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 45(3):ii, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2015.
                                        45.3.toc.

Anonymous:2015:TCd

[Ano15n] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 45(4):ii, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2015.45.4.toc.

Anonymous:2015:TCe

[Ano15o] Anonymous. Table of contents. Historical Studies in the Natural Sciences, 45(5):ii, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.5.toc.

Anonymous:2017:MJN

[Ano17] Anonymous. Mary Jo Nye publications (1972–2016). Historical Studies in the Natural Sciences, 47(3):450–457, June 2017. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Arbel:2019:FHM

[Arb19] Tal Arbel. Fear in Hebrew: Militarized behaviorism and the cultural politics of science in translation. Historical Studies in the Natural Sciences, 49(5):471– 503, November 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Armstrong:1948:F

[Arm48] Eva V. Armstrong. Foreword.

Chymia: Annual Studies in the History of Chemistry, 1(??):ii, ix–xiv, ???? 1948. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757109.

Aronova:2017:RTH

[Aro17] Elena Aronova. Recent trends in the historiography of science in the Cold War. Historical Studies in the Natural Sciences, 47(4):568–577, September 2017. ISSN 1939-1811 (print), 1939-182X (electronic).

Ash:2010:EPK

[Ash10]  Eric H. Ash, editor.  Ex-

pertise: practical knowledge and the early modern state, volume 25 (second series) of Osiris, 0369-7827. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-02939-5 (paperback). 259 pp. LCCN Q125 .E89 2010. URL http:/

/www.journals.uchicago.
edu/loi/osiris/.

Aspden:1980:PUH

[Asp80] Harold Aspden. Physics unified. Sabberton Publications, Southampton, UK, 1980. ISBN 0-85056-009-8 (hardcover), 0-85056-010-1 (paperback). xiii + 206 pp. LCCN QC173.7 .A85.

Assmus:1992:AMP

[Ass92a] Alexi Assmus. The Americanization of molecular physics. Historical Studies in the Physical and Biological Sciences, 23 (1):1–34, ???? 1992. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757691.

Assmus:1992:MTE

[Ass92b] Alexi Assmus. The molecular tradition in early quantum theory. Historical Studies in the Physical and Biological Sciences, 22(2):209–231, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757681.

Aubin:2020:CMB

[Aub20] David Aubin. Congress mania in Brussels, 1846–1856: Soft power, transnational experts, and diplomatic practices. Historical Studies in the Natural Sciences, 50(4):340– ??, September 2020. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/4/340/111893/Congress-
Mania-in-Brussels-1846-
1856Soft-Power.

Augustine:2007:RPE

[Aug07] Dolores L. Augustine. Red Prometheus: engineering and dictatorship in East Germany, 1945–1990. Transformations. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-01236-7 (hardcover). xxx + 381 + 4 pp. LCCN TA73.1 .A94

2007. URL http://www.loc.
gov/catdir/toc/ecip078/
2007000656.html.

Achbari:2016:DSG

[AvL16] A. Achbari and F. van Lunteren.

Dutch skies, global laws: The British creation of “Buys Ballot’s Law”. Historical Studies in the Natural Sciences, 46(1): 1–43, February 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Ayres:2012:SEL

[Ayr12] P. G. (Peter G.) Ayres. Shaping ecology: the life of Arthur Tansley. Wiley-Blackwell, Chich-

ester,  West  Sussex,

UK, 2012. ISBN 0-470-67156-4 (hardcover), 0-470-67154-8 (paperback). xii + 213 pp. LCCN QH31.T37 A87 2012.

Bacon:1965:PAC

[Bac65] Egbert K. Bacon. A precursor of the American Chemical Society — Chandler and the Society of Union College. Chymia: Annual Studies in the History of Chemistry, 10(??): 183–197, ???? 1965. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757250.

Badino:2018:YIS

[Bad18]  Massimiliano Badino.  And

yet it stands: The stability of the solar system in Eighteenth-Century physical astronomy. Historical Studies in the Natural Sciences, 48(2):123– 179, April 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Baker:1964:HI

[Bak64] A. Albert Baker, Jr. A history of indicators. Chymia: Annual Studies in the History of Chemistry, 9(??):147–167, ???? 1964. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757238.

Baldwin:2015:BRA

[Bal15] Melinda Baldwin. Book reviews: Adelene Buckland. Novel Science: Fiction and the Invention of Nineteenth-Century Geology.

Chicago: University of Chicago Press, 2013. 400 pp., illus., index. ISBN 978-0-22-607968-4. $45 (hardcover). Aileen Fyfe. Steam-Powered Knowledge: William Chambers and the Business of Publishing, 1820–1860. Chicago: University of Chicago Press, 2012. xvi + 336 pp., illus., index. ISBN 978-0-22-627651-9. $55 (hardcover). Anna Henchman. The Starry Sky Within: Astronomy and the Reach of the Mind in Victorian Literature. Oxford: Oxford University Press, 2014. xvi + 320 pp., illus., index. ISBN 978-0-19-968696-4. $85 (hardcover). James Secord. Visions of Science: Books and Readers at the Dawn of the Victorian Age. Chicago: University of Chicago Press, 2014. xiii + 256 pp., illus., index. ISBN 978-0-22-620328-7. $30 (hardcover). Historical Studies in the Natural Sciences, 45(4):610– 620, September 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.4.610.

Barrow:2009:NGC

[Bar09] Mark V. Barrow. Nature’s ghosts: confronting extinction from the age of Jefferson to the age of ecology. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-03815-7 (e-book), 0-226-03814-9. xi + 497 pp. LCCN QL84.2 .B377

Barahona:2015:MGM

[Bar15] Ana Barahona. Medical genetics in Mexico: The origins of cytogenetics and the health care system. Historical Studies in the Natural Sciences, 45(1): 147–173, February 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2015.
                                        45.1.147.

Barany:2016:FTT

[Bar16] Michael J. Barany. Fellow travelers and traveling fellows: The intercontinental shaping of modern mathematics in midtwentieth century Latin America. Historical Studies in the Natural Sciences, 46(5):669– 709, November 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Barany:2018:IPW

[Bar18] Michael J. Barany. Integration by parts: Wordplay, abuses of language, and modern mathematical theory on the move. Historical Studies in the Natural Sciences, 48(3):259–299, June 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Barany:2024:KIH

[Bar24] Michael J. Barany. A kaleidoscopic introduction to the history of science. Historical Studies in the Natural Sciences, 54 (1):86–??, February 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/86/200192/A-
Kaleidoscopic-
Introduction-to-the-
History-of.

Bashford:2025:NMS

[Bas25] Alison Bashford. The new modern synthesis: E. O. Wilson and Julian Huxley. Historical Studies in the Natural Sciences, 55(1):61–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/61/207543/The-New-
Modern-SynthesisE-O-
Wilson-and-Julian.

Bauchner:2021:FWE

[Bau21] Joshua Bauchner. Fechner on a walk: Everyday investigations of the mind–body relationship. Historical Studies in the Natural Sciences, 51(1):1–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/1/1/116282/Fechner-on-
a-WalkEveryday-
Investigations-of-the.

Bayoumi:2024:WPE

[Bay24] Soha Bayoumi. What place for emotions in the history of science and medicine classroom? Historical Studies in the Natural Sciences, 54(1):89–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/1/89/200203/What-Place-
                                        for-Emotions-in-the-
                                        History-of-Science.

Brown:1991:SBS

[BC91] Laurie M. Brown and Tian Yu Cao. Spontaneous breakdown of symmetry: Its rediscovery and integration into quantum field theory. Historical Studies in the Physical and Biological Sciences, 21(2):211–235, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757663.

Brush:1969:MOR

[BE69] S. G. Brush and C. W. F. Everitt. Maxwell, Osborne Reynolds, and the radiometer. Historical Studies in the Physical Sciences, 1(??): 105–125, ???? 1969. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757296.

Becher:1980:WWC

[Bec80] Harvey W. Becher. William Whewell and Cambridge mathematics. Historical Studies in the Physical Sciences, 11 (1):1–48, ???? 1980. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757470.

Beck:2023:SFP

[Bec23] Pieter T. L. Beck. Strong foundations: Petrus van Musschenbroek’s experimental research on the strength of materials. Historical Studies in the Natural Sciences, 53(2):109–??, April 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/2/109/196001/Strong-
FoundationsPetrus-van-
Musschenbroek-s.

Belhoste:2011:PSP

[Bel11] Bruno Belhoste. Paris savant:

parcours et rencontres au temps des lumières. (French) [Paris learned: pathways and meetings at the time of the Enlightenment]. Colin, Paris, France, 2011. ISBN 2-200-25563-2. 310 pp. LCCN Q127.F8 B44

2011. URL http://www.loc.
gov/catdir/toc/fy12pdf03/
2012428139.html.

Bellon:2018:ECT

[Bel18] Richard Bellon. Emotional comfort and theoretical necessity: Sex and gender in the age of Darwin. Historical Studies in the Natural Sciences, 48(2):246– 257, April 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Benson:2001:SCS

[Ben01] Keith R. Benson. Summer camp, seaside station, and marine laboratory: Marine biology and its institutional identity.

Historical Studies in the Physical and Biological Sciences, 32 (1):11–18, September 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Benson:2012:ESR

[Ben12] Etienne Benson. Endangered science: The regulation of research by the U.S. Marine Mammal Protection and Endangered Species Acts. Historical Studies in the Natural Sciences, 42 (1):30–61, February 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2012.
42.1.30.

Bernstein:1982:MJR

[Ber82] Barton J. Bernstein. In the matter of J. Robert Oppenheimer. Historical Studies in the Physical Sciences, 12(2): 195–252, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757496.

Bernstein:1988:FPB

[Ber88] Barton J. Bernstein. Four physicists and the bomb: The early years, 1945–1950. Historical Studies in the Physical and Biological Sciences, 18(2): 231–263, ???? 1988. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757603.

Bertozzi:2019:ECR

[Ber19] Eugenio Bertozzi. Establishing and consolidating a research field: The biography of the Wilson cloud chamber in the history of particle physics. Historical Studies in the Natural Sciences, 49(2):117–150, April 2019. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Beyer:2000:SB

[Bey00] Kurt Beyer. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 31(1):175–??, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Beyer:2001:SB

[Bey01] Kurt W. Beyer. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 31(2):381–387, March 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Beyer:2002:SB

[Bey02]  Kurt Beyer.  Selected bib-

liography. Historical Studies in the Physical and Biological Sciences, 32(2):409–416, March 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Beyler:2011:TWS

[Bey11] Richard H. Beyler. Three ways to spend some time in the historiographical metropolis. Historical Studies in the Natural Sciences, 41(3):354–364, Summer 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.3.354.

BernalVillegas:2010:IRR

[BG10] J. E. (Jaime E.) Bernal Villegas and Alberto Gómez Gutiérrez. A impulsos de una rara resolución: el viaje de José Celestino Mutis al Nuevo Reino de Granada, 1760–1763. Pontificia Universidad Javeriana, Bogotá, Colombia, 2010. ISBN 958-716-326-5. 352 pp. LCCN F2261 .B47 2010.

Byrne:2013:DGB

[BG13] Eugene Byrne and Simon Gurr. Darwin: a graphic biography: the really exciting and dramatic story of a man who mostly stayed at home and wrote some books. Smithsonian Books, Washington, DC, 2013. ISBN 1-58834-352-9 (paperback). 96 pp. LCCN QH31.D2 B97 2013.

BernalVillegas:2008:MCM

[BGC08] J. E. (Jaime E.) Bernal Villegas, Alberto Gómez Gutiérrez, and José Celestino. Medicina científica mutisiana: estudio elaborado a partir de un documento médico inédito de José Celestino Mutis y Bosio, escrito en Madrid en 1759 y conservado hoy en el Archivo Histórico Javeriano de Bogotá. Editorial Pontificia Universidad Javeriana, Bogotá, Colombia, 2008.

ISBN 958-716-135-1. 144 pp. LCCN QP21 .B45 2008.

Brooke:2007:RMC

[BH07] John Hedley Brooke and Peter Harman. Russell McCormmach as a colleague. Historical Studies in the Physical and Biological Sciences, 37(2): 475–478, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Bodnarczuk:2008:MPP

[BH08] Mark Bodnarczuk and Lillian Hoddeson. Megascience in particle physics: The birth of an experiment string at Fermilab. Historical Studies in the Natural Sciences, 38(4):508– 534, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Biagioli:2012:PIK

[Bia12] Mario Biagioli. Productive illusions: Kuhn’s structure as a recruitment tool. Historical Studies in the Natural Sciences, 42(5):479–484, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.479.pdf.

Bian:2022:SRE

[Bia22] He Bian. Science and really existing socialism in Maoist China: a review of recent works. Historical Studies in the Natural Sciences, 52(2):265–??, April 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/2/265/144552/Science-
and-Really-Existing-
Socialism-in-Maoist.

Bimm:2021:AMA

[Bim21] Jordan Bimm. Andean man & the astronaut: Race and the 1958 Mount Evans Acclimatization Experiment. Historical Studies in the Natural Sciences, 51(3):285–??, June 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/3/285/117359/Andean-
Man-amp-the-AstronautRace-
and-the-1958.

Buchwald:2010:ZPH

[BJ10]  Jed  Z.  Buch-

wald and Diane Greco Josefowicz. The zodiac of Paris: how an improbable controversy over an ancient Egyptian artifact provoked a modern debate between religion and science. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14576-8. vi + 428 + 8 pp. LCCN Q127.E8 B83 2010.

Bagdonas:2021:FOB

[BK21] Alexandre Bagdonas and Alexei Kojevnikov. Funny origins of the Big Bang Theory. Historical Studies in the Natural Sciences, 51(1):87–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/1/87/116291/Funny-
Origins-of-the-Big-Bang-
Theory.

Bleichmar:2012:VEB

[Ble12] Daniela Bleichmar. Visible empire: botanical expeditions and visual culture in the Hispanic Enlightenment. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-05853-0 (hardcover). xii + 286 pp. LCCN QK21.S7 B54 2012.

Barrera-Osorio:2006:ENS

[BO06] Antonio Barrera-Osorio. Experiencing nature: the Spanish American empire and the early scientific revolution. University of Texas Press, Austin, TX, USA, 2006. ISBN 0-292-70981-1. xi + 211 pp. LCCN Q127.L38

B37 2006. URL http://www.
loc.gov/catdir/
enhancements/fy0622/
2005011166-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0512/2005011166.html.

Bolman:2022:AHR

[Bol22] Brad Bolman. In the animal house: Rabies, labor, and salvage dogs in Jim Crow Birmingham. Historical Studies in the Natural Sciences, 52(5):589–??, November 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/52/5/589/194656/

In-the-Animal-HouseRabies-
Labor-and-Salvage-Dogs.

Borrello:2010:ERC

[Bor10] Mark E. Borrello. Evolutionary restraints: the contentious history of group selection. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-06701-7 (hardcover). xi + 215 pp. LCCN QH376 .B67 2010.

Borg:2023:DMA

[Bor23] George Borg. Discoverer and methodologist: Alfred O. C. Nier and the Mid-Twentieth-Century instrumental revolution in geochemistry. Historical Studies in the Natural Sciences, 53(5):445–??, November 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/5/445/198193/
Discoverer-and-
MethodologistAlfred-O-C-
Nier-and.

Bouk:2012:BRT

[Bou12] Dan Bouk. Book review: Tocqueville’s ghost: Alexander Dallas Bache: Building the American Nation through Science and Education in the Nineteenth Century, by Axel Jansen; A Democracy of Facts: Natural History in the Early Republic, by Andrew J. Lewis; Scientists and Swindlers: Consulting on Coal and Oil in America, 1820– 1890, by Paul Lucier. Historical Studies in the Natural Sciences, 42(4):329–339, September 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2012.
42.4.329.

Bouk:2018:NDC

[Bou18] Dan Bouk. The national data center and the rise of the data double. Historical Studies in the Natural Sciences, 48(5):627– 636, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Bowler:2009:SAP

[Bow09] Peter J. Bowler. Science for all: the popularization of science in early twentieth-century Britain. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-06866-8 (e-book), 0-226-06863-3 (hardcover). x + 339 pp. LCCN Q225.2.G7 B69

2009. URL http://proxy2.
hec.ca/login?url=http://
site.ebrary.com/lib/hecm/
Doc?id=10354019.

Brown:1994:FTN

[BR94] Laurie M. Brown and Helmut Rechenberg. Field theories of nuclear forces in the 1930s: The Fermi-Field theory. Historical Studies in the Physical and Biological Sciences, 25(1):1–24, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Brain:2015:BRD

[Bra15] Stephen Brain. Book reviews:

Dark green: The gloomy side of environmentalism: Peder Anker. From Bauhaus to Eco-House: A History of Ecological Design. Baton Rouge: Louisiana State University Press, 2010. viii + 188 pp., illus., index. ISBN 978-0-8071-3551-8. $34.95 (cloth).; Jacob Hamblin. Arming Mother Nature: The Birth of Catastrophic Environmentalism. New York: Oxford University Press, 2013. x + 298 pp., index. ISBN 978-0-19-974005-5. $29.95 (cloth).; Edwin A. Martini. Agent Orange: History, Science, and the Politics of Uncertainty. Amherst: University of Massachusetts Press, 2012. xvi + 302 pp., illus., index. ISBN 978-1-55849-975-1. $24.95 (paper).; Rebecca Priestley. Mad on Radium: New Zealand in the Atomic Age. Auckland, NZ: University of Auckland Press, 2012. xii + 284 pp., illus., index. ISBN 978-1-86940-727-8. $34.95 (paper). Historical Studies in the Natural Sciences, 45(2):340– 347, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2015.45.2.340.

Brick:2015:BRR

[Bri15] Howard Brick. Book reviews: Betwixt, between, and all together: The curious careers of new sciences: Working Knowledge: Making the Human Sciences from Parsons to Kuhn by Joel Isaac; Misbehaving Science: Controversy and the Development of Behavior Genetics by Aaron Panofsky; The Democratic Surround: Multimedia & American Liberalism from World War II to the Psychedelic Sixties by Fred Turner. Historical Studies in the Natural Sciences, 45(5):805–815, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.5.805.

Browne:1948:RAI

[Bro48]  C. A. Browne.  Recently

acquired information concerning Fredrick Accum, 1769–1838. Chymia: Annual Studies in the History of Chemistry, 1(??): 1–9, ???? 1948. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757110.

Brown:1969:ECE

[Bro69] Theodore M. Brown. The electric current in early Nineteenth-Century French physics. Historical Studies in the Physical Sciences, 1(??):61–103, ???? 1969. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (elec-

tronic). URL http://www.
jstor.org/stable/27757295.

Bromberg:1971:INB

[Bro71] Joan Bromberg. The impact of the neutron: Bohr and Heisenberg. Historical Studies in the Physical Sciences, 3(??): 307–341, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757321.

Brooke:1975:LGP

[Bro75] John Hedley Brooke. Laurent, Gerhardt, and the philosophy of chemistry. Historical Studies in the Physical Sciences, 6 (??):405–429, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757345.

Bromberg:1976:CPC

[Bro76] Joan Bromberg. The concept of particle creation before and after quantum mechanics. Historical Studies in the Physical Sciences, 7(??):161–191, ???? 1976. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757356.

Brown:2007:BHB

[Bro07] Cynthia Stokes Brown. Big history: from the Big Bang to the present. New Press, New York, NY, USA, 2007. ISBN 1-59558-196-0, 1-59558-414-5 (paperback), 1-59558-558-3 (e-book). xvi + 288 pp. LCCN D20 .B77 2007. URL http://

                                        www.loc.gov/catdir/toc/
                                        ecip0711/2007006741.html.

Bromberg:2008:NIM

[Bro08] Joan Lisa Bromberg. New instruments and the meaning of quantum mechanics. Historical Studies in the Natural Sciences, 38(3):325–352, August 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Broman:2014:BRW

[Bro14] Thomas Broman. Book review: Working knowledge technical practices, social identities, and expertise in early modern Europe Pamela O. Long. Artisan/Practitioners and the Rise of the New Sciences, 1400–1600. Corvallis: Oregon State University Press, 2011. xii + 196 pp., illus., index. ISBN 978-0-87071-609-6. $22.95 (paper).; Eric L. Ash, ed. Expertise: Practical Knowledge and the Early Modern State. Osiris 2nd Series, vol. 25. Chicago: University of Chicago Press, 2010. 259 pp., illus., index. ISBN 978-0-226-02939-9. $33.00 (paper).; Chandra Mukerji. Impossible Engineering: Technology and Territoriality on the Canal du Midi. Princeton, NJ: Princeton University Press, 2009. xix + 304 pp., illus., index. ISBN 978-0-691-14032-2. $42.00 (hardcover). Historical Studies in the Natural Sciences, 44 (1):80–89, February 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.1.80.

Brookhuis:2023:MBB

[Bro23] Hein Brookhuis. Making Belgian Big Science: A history of the MYRRHA research reactor (1994–2010). Historical Studies in the Natural Sciences, 53 (1):35–??, February 2023. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        53/1/35/195533/Making-
                                        Belgian-Big-ScienceA-
                                        History-of-the-MYRRHA.

Brush:1987:RWE

[Bru87] Stephen G. Brush. Review: Whole Earth history. Historical Studies in the Physical and Biological Sciences, 17(2):345– 355, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757586.

Bruno:2003:BNB

[Bru03] Laura A. Bruno. The bequest of the nuclear battlefield: Science, nature, and the atom during the first decade of the Cold War. Historical Studies in the Physical and Biological Sciences, 33 (2):237–260, March 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Brush:2007:HIB

[Bru07] Stephen G. Brush. How ideas became knowledge: The lightquantum hypothesis 1905–1935. Historical Studies in the Physical and Biological Sciences, 37 (2):205–246, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Bonneui:2010:PLV

[BT10] Christophe Bonneui and Frederic Thomas. Purifying landscapes: The Vichy Regime and the genetic modernization of France. Historical Studies in the Natural Sciences, 40(4):532– 568, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2010.40.4.532.

Bub:2019:IQW

[Bub19] Jeffrey Bub. Interpreting the quantum world: Old questions and new answers. Historical Studies in the Natural Sciences, 49(2):226–239, April 2019. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Buchwald:1977:WTM

[Buc77] Jed Z. Buchwald. William Thomson and the mathematization of Faraday’s electrostatics. Historical Studies in the Physical Sciences, 8(??): 101–136, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757369.

Buchwald:2012:KSF

[Buc12] Jed Z. Buchwald. Kuhn’s structure four and a half decades later. Historical Studies in the Natural Sciences, 42(5):485– 490, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.485.pdf.

Buckland:2013:NSF

[Buc13] Adelene Buckland. Novel science: fiction and the invention of nineteenth-century geology. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07968-6 (hardcover), 0-226-92363-0 (e-book). 377 pp. LCCN PR468.G38 B83 2013.

Bud:2007:PTT

[Bud07] Robert Bud. Penicillin: triumph and tragedy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-925406-0. ix + 330 pp. LCCN RM666.P35 B83

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0725/2006102287-d.html;
                                         http://www.loc.gov/
                                        catdir/toc/ecip077/
                                        2006102287.html.

Buehler:1953:PUE

[Bue53] Curt F. Buehler. A projected but unpublished edition of the “Life and Works” of Robert Boyle. Chymia: Annual Studies in the History of Chemistry, 4(??):79–83, ???? 1953. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757163.

Bugos:1989:MCR

[Bug89] Glenn E. Bugos. Managing cooperative research and borderland science in the National Research Council, 1922–1942. Historical Studies in the Physical and Biological Sciences, 20(1): 1–32, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757633.

Bugos:1992:ROQ

[Bug92] Glenn E. Bugos. Review: The organization of the quest for certainty. Historical Studies in the Physical and Biological Sciences, 23(1):181–191, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757694.

Burchfield:1982:RBA

[Bur82] Joe D. Burchfield. Review:

The British Association and its historians. Historical Studies in the Physical Sciences, 13(1): 165–174, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757510.

Burgos:2006:BSL

[Bur06]  Marcelo  Bau-

mann Burgos. Brazilian synchrocyclotron light. Historical Studies in the Physical and Biological Sciences, 36(2):329– 341, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Burton:2024:AGH

[Bur24] Elise K. Burton. Accidents of geography: Historicizing genetic cartographies of the Middle East. Historical Studies in the Natural Sciences, 54(1): 3–??, February 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/3/200205/Accidents-
of-GeographyHistoricizing-
Genetic.

Butler:1994:RMP

[But94] Loren Butler. Robert S. Mulliken and the politics of science and scientists, 1939–1946. Historical Studies in the Physical and Biological Sciences, 25 (1):25–45, ???? 1994. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Bensaude-Vincent:2001:CDM

[BV01] Bernadette Bensaude-Vincent. The construction of a discipline: Materials science in the United States. Historical Studies in the Physical and Biological Sciences, 31(2):223– 248, March 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Bacciagaluppi:2009:QTC

[BV09]  Guido  Bacciagaluppi

and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay conference. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9 (hardcover). LCCN QC173.96 .B33 2009. URL http://www.

loc.gov/catdir/toc/
ecip0817/2008019585.html.

Baldwin:2024:IAS

[BV24a] Melinda Baldwin and Brigid Vance. Introduction: AI and scholarly publishing. Historical Studies in the Natural Sciences, 54(5):609–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/5/609/203883/
IntroductionAI-and-
Scholarly-Publishing.

Baldwin:2024:IPH

[BV24b] Melinda Baldwin and Brigid Vance. Introduction: Pedagogy in the history of science and medicine. Historical Studies in the Natural Sciences, 54 (1):84–??, February 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/84/200190/
IntroductionPedagogy-in-
the-History-of-Science-
and.

Bycroft:2013:WAV

[Byc13]  Michael Bycroft.  Wonders

in the Academy: The value of strange facts in the experimental research of Charles Dufay. Historical Studies in the Natural Sciences, 43(3): 334–370, June 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.3.334.pdf.

Bykov:1965:HSE

[Byk65] G. V. Bykov. Historical sketch of the electron theories of organic chemistry. Chymia: Annual Studies in the History of Chemistry, 10(??):199– 253, ???? 1965. CODEN CHYMAP. ISSN 0095-9367.

                                        URL http://www.jstor.org/
                                        stable/27757251.

Casado:2014:VSN

[CA14] Santos Casado and Santiago Aragón. Vignettes of Spanish nature: Imagining a national fauna at the Museo Nacional de Ciencias Naturales, Madrid (1910–1936). Historical Studies in the Natural Sciences, 44 (3):197–233, June 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.3.197.

Cahan:1982:WSO

[Cah82] David Cahan. Werner Siemens and the origin of the Physikalisch-Technische Reichsanstalt, 1872–1887. Historical Studies in the Physical Sciences, 12(2):253–283, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757497.

Cahan:1985:IRG

[Cah85] David Cahan. The institutional revolution in German physics, 1865–1914. Historical Studies in the Physical Sciences, 15 (2):1–65, ???? 1985. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757549.

Cahan:1988:RPP

[Cah88] David Cahan. Review: Pride and prejudice in the history of physics: The German-speaking world, 1740–1945. Historical Studies in the Physical and Biological Sciences, 19(1):173– 191, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757620.

Cahan:2004:HSA

[Cah04] David Cahan. Helmholtz and the shaping of the American physics elite in the Gilded Age. Historical Studies in the Physical and Biological Sciences, 35 (1):1–34, September 2004. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Camprubi:2010:OGO

[Cam10] Lino Camprubí. One grain, one nation: Rice genetics and the corporate state in early Francoist Spain (1939–1952). Historical Studies in the Natural Sciences, 40(4):499–531, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2010.
40.4.499.

Camprubi:2016:RIG

[Cam16] Lino Camprubí. Review: The Invention of the Global Environment. Historical Studies in the Natural Sciences, 46(2):243– 251, April 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Camprubi:2020:BBE

[Cam20] Lino Camprubí. Birds without borders: Ecological diplomacy and the WWF in Franco’s Spain. Historical Studies in the Natural Sciences, 50(4):433– ??, September 2020. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/4/433/111897/Birds-
Without-BordersEcological-
Diplomacy-and-the.

Campagnolo:2022:HWG

[Cam22] Gian Marco Campagnolo. How to win games and influence football players. Historical Studies in the Natural Sciences, 52 (1):120–??, February 2022. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/1/120/120497/How-to-
Win-Games-and-Influence-
Football-Players.

Cantor:1975:RWT

[Can75] Geoffrey Cantor. The reception of the wave theory of light in Britain: a case study illustrating the role of methodology in scientific debate. Historical Studies in the Physical Sciences, 6(??):109–132, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757341.

Caneva:1978:GET

[Can78] Kenneth L. Caneva. From Galvanism to electrodynamics: The transformation of German physics and its social context. Historical Studies in the Physical Sciences, 9(??):63–159, ???? 1978. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757377.

Caneva:1997:CRM

[Can97] Kenneth L. Caneva. Colding, Ørsted, and the meanings of force. Historical Studies in the Physical and Biological Sciences, 28(1):1–138, ???? 1997. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Caneva:2005:DSC

[Can05] Kenneth L. Caneva. ‘Discovery’ as a site for the collective construction of scientific knowledge. Historical Studies in the Physical and Biological Sciences, 35(2):175–291, March 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Candela:2021:DV

[Can21] Emily Candela. Designing the virus. Historical Studies in the Natural Sciences, 51(1):140–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/51/1/140/116284/
                                        Designing-the-Virus.

Carson:1999:NMS

[Car99] Cathryn Carson. New models for science in politics: Heisenberg in West Germany. Historical Studies in the Physical and Biological Sciences, 30(1): 115–171, ???? 1999. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Carson:2020:CD

[Car20] Cathryn Carson. Clouds of data. Historical Studies in the Natural Sciences, 50(1–2):81– 89, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Carson:2024:JHM

[Car24] Cathryn Carson. John L. Heilbron (1934–2023): In memoriam. Historical Studies in the Natural Sciences, 54(4):521–??, September 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/54/4/521/203067/

John-L-Heilbron-1934-2023-
In-Memoriam.

Cassidy:1979:HFC

[Cas79] David C. Cassidy. Heisenberg’s first core model of the atom: The formation of a professional style. Historical Studies in the Physical Sciences, 10 (??):187–224, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757390.

Cassidy:1981:CRS

[Cas81] David C. Cassidy. Cosmic ray showers, high energy physics, and quantum field theories: Programmatic interactions in the 1930s. Historical Studies in the Physical Sciences, 12 (1):1–39, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757488.

Cassidy:1983:RRG

[Cas83] David C. Cassidy. Review:

Recent German perspectives on German technical education. Historical Studies in the Physical Sciences, 14(1): 187–200, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757529.

Cassidy:1986:RUH

[Cas86] David Cassidy. Review: Understanding the history of Special Relativity. Historical Studies in the Physical and Biological Sciences, 16(1):177–188, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757561.

Cassidy:1994:CGS

[Cas94] David Cassidy. Controlling German science, I: U.S. and Allied forces in Germany, 1945–1947. Historical Studies in the Physical and Biological Sciences, 24 (2):197–235, ???? 1994. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Cassidy:1996:CGS

[Cas96] David Cassidy. Controlling German science, II: Bizonal occupation and the struggle over West German science policy, 1946–1949. Historical Studies in the Physical and Biological Sciences, 26(2):197–239, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Castonguay:2001:ERS

[Cas01] Stephane Castonguay. The emergence of research specialties in economic entomology in Canadian government laboratories after World War II. Historical Studies in the Physical and Biological Sciences, 32(1): 19–40, September 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Cassidy:2007:OFP

[Cas07] David C. Cassidy. Oppenheimer’s first paper: Molecular band spectra and a professional style. Historical Studies in the Physical and Biological Sciences, 37(2):247–270, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Case:2009:ISM

[Cas09a] Stephen Case. ‘Insufferably stupid or miserably out of place’: F. A. P. Barnard and his scientific instrument collection in the Antebellum South. Historical Studies in the Natural Sciences, 39(4):418–443, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Cassidy:2009:BUH

[Cas09b] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Cat:1998:PDU

[Cat98] Jordi Cat. The physicists’ debates on unification in physics at the end of the 20th Century. Historical Studies in the Physical and Biological Sciences, 28 (2):253–300, ???? 1998. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Cockburn:1981:OLT

[CE81] Stewart Cockburn and David Ellyard. Oliphant, the life and times of Sir Mark Oliphant. Axiom Books, Adelaide, SA, Australia, 1981. ISBN 0-9594164-0-4. xviii + 369 pp. LCCN QC16.O46 C6 1981.

Creager:1996:MSE

[CG96] Angela N. H. Creager and Jean-Paul Gaudillière. Meanings in search of experiments and viceversa: The invention of allosteric regulation in Paris and Berkeley (1959–1968). Historical Studies in the Physical and Biological Sciences, 27(1):1–89, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Carlsson-Hyslop:2016:PPB

[CH16] Anna Carlsson-Hyslop. Patronage and practice in British oceanography: The mixed patronage of storm surge science at the Liverpool Tidal Institute, 1919–1959. Historical Studies in the Natural Sciences, 46(3):270– 312, June 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Chayut:1993:NSS

[Cha93] Michael Chayut. New sites for scientific change: Paul Flory’s initiation into polymer chemistry. Historical Studies in the Physical and Biological Sciences, 23(2):193–218, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757698.

Chayut:1994:BJL

[Cha94] Michael Chayut. From Berlin to Jerusalem: Ladislaus Farkas and the founding of physical chemistry in Israel. Historical Studies in the Physical and Biological Sciences, 24(2):237–263, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Chang:2001:SAQ

[Cha01] Hasok Chang. Spirit, air, and quicksilver: The search for the ‘real’ scale of temperature. Historical Studies in the Physical and Biological Sciences, 31(2): 249–284, March 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Charmantier:2011:CLV

[Cha11] Isabelle Charmantier. Carl Linnaeus and the visual representation of nature. Historical Studies in the Natural Sciences, 41(4):365–404, November 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.4.405.

Chang:2012:WHE

[Cha12] Hasok Chang. Is Water H_{2}O?:

Evidence, Realism and Pluralism, volume 293 of Boston Studies in the Philosophy and History of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-3931-1, 94-007-3932-X (ebook). ISSN 0068-0346. xxi + 316 pp. LCCN QD169.W3

C43 2012. URL https://link.
springer.com/book/10.1007/
978-94-007-3932-1.

Charenko:2020:RCP

[Cha20] Melissa Charenko. Reconstructing climate: Paleoecology and the limits of prediction during the 1930s Dust Bowl. Historical Studies in the Natural Sciences, 50(1–2):90–128, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Chen:2024:EER

[Che24] Bohang Chen. Entelechy and energy: Reconsidering Hans Driesch’s vitalism in the science and philosophy of the organism. Historical Studies in the Natural Sciences, 54(4):423– ??, September 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/4/423/203066/Entelechy-
and-EnergyReconsidering-
Hans-Driesch-s.

Choi:2012:BRD

[Cho12] Hyungsub Choi. Book review: Designing and redesigning artifacts and institutions: Hitting the Brakes: Engineering Design and the Production of Knowledge, by Ann Johnson; Makers of the Microchip: A Documentary History of Fairchild Semiconductor, by Christophe Lécuyer and David C. Brock; Velvet Revolution at the Synchrotron: Biology, Physics, and Change in Science, by Park Doing; Becoming MIT: Moments of Decision, by David Kaiser. Historical Studies in the Natural Sciences, 42(4):322– 328, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2012.
42.4.322;
https://www.math.utah.edu/
pub/tex/bib/hsns.bib.

Christian:2011:MTI

[Chr11] David Christian. Maps of time:

an introduction to big history, volume 2 of The California world history library. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-24476-1, 0-520-27144-0, 0-520-95067-4 (e-book). xxvii + 642 pp. LCCN CB19 .C47

2011eb. URL http://site.
ebrary.com/lib/princeton/
Doc?id=10496862.

Churchill:1967:SCE

[Chu67]  Mary S. Churchill.  The

seven chapters, with explanatory notes. Chymia: Annual Studies in the History of Chemistry, 12(??):27, 29–57, ???? 1967. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757273.

Cittadino:2015:PSP

[Cit15] Eugene Cittadino. Paul Sears and the Plowshare Advisory Committee: “subversive” ecologist endorses nuclear excavation? Historical Studies in the Natural Sciences, 45(3):397– 446, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

                                        //www.jstor.org/stable/10.
                                        1525/hsns.2015.45.3.397.

Carazza:1990:ARM

[CK90]  Bruno  Carazza  and

Helge Kragh. Augusto Righi’s magnetic rays: a failed research program in early 20th-Century physics. Historical Studies in the Physical and Biological Sciences, 21(1):1–28, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757653.

Creager:2007:SLM

[CLW07] Angela N. H. Creager, Elizabeth Lunbeck, and M. Norton Wise, editors. Science without laws: model systems, cases, exemplary narratives. Science and cultural theory. Duke University Press, Durham, NC, USA, 2007. ISBN 0-8223-4046-1 (cloth), 0-8223-4068-2 (paperback). 287 pp. LCCN QH324.8 .S25 2007. URL http://www.

                                        loc.gov/catdir/toc/
                                        ecip0710/2007006303.html.

Coen:2002:SEN

[Coe02] Deborah R. Coen. Scientists’ errors, nature’s fluctuations, and the Law of Radioactive Decay, 1899–1926. Historical Studies in the Physical and Biological Sciences, 32(2):179–205, March 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Coen:2007:VAU

[Coe07] Deborah R. Coen. Vienna in the age of uncertainty: science, liberalism, and private life. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-11172-5 (cloth). xi + 380 pp. LCCN DB844.E98 C64

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0707/2007003723-b.html;
http://www.loc.gov/
catdir/enhancements/
fy0707/2007003723-d.html;
http://www.loc.gov/
catdir/toc/ecip0710/
2007003723.html.

Coen:2010:BRW

[Coe10] Deborah R. Coen. Book reviews: Weatherwiser? Matthew Mulcahy. Hurricanes and Society in the British Greater Caribbean, 1624–1783. Baltimore: Johns Hopkins University Press, 2006. ix + 257 pp., illus., index. ISBN 978-0-801-89079-6. $25.00 (paper). Jan Golinski. British Weather and the Climate of Enlightenment. Chicago: University of Chicago Press, 2007. xv + 284 p., illus., index. ISBN: 978-0-226-30205-8. $35.00 (cloth). Fabien Locher. Le savant et la tempête: Étudier l’atmosphère et prévoir le temps au XIXe siècle. Presses Universitaires Rennes, 2008. 221 pp., illus. ISBN 978-2-753-50696-1. $25.44 (paper). Kristine C. Harper. Weather by the Numbers: The Genesis of Modern Meteorology. Cambridge, MA: MIT Press, 2008. ix + 308 pp., illus., index. ISBN 978-0-262-08378-2. $40.00 (cloth). Historical Studies in the Natural Sciences, 40(1):125–135, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Coen:2021:EMH

[Coe21] Deborah R. Coen. The experimental multispecies household. Historical Studies in the Natural Sciences, 51(3):330–??, June 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/3/330/117356/The-
Experimental-Multispecies-
Household.

Coffey:2006:CFE

[Cof06] Patrick Coffey. Chemical free energies and the Third Law of Thermodynamics. Historical Studies in the Physical and Biological Sciences, 36(2):365– 396, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Cohen:1950:BCI

[Coh50] I. Bernard Cohen. The beginning of chemical instruction in America: a brief account of the teaching of chemistry at Harvard prior to 1800. Chymia: Annual Studies in the History of Chemistry, 3(??):17–44, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757144.

Collins:2003:LBB

[Col03] H. M. Collins. LIGO becomes big science. Historical Studies in the Physical and Biological Sciences, 33(2):261–297, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Coleman:2010:BEE

[Col10] David C. Coleman. Big ecology: the emergence of ecosystem science. University of California Press, Berkeley, CA, USA, 2010. ISBN 0-520-26475-4 (hardcover). xii + 236 pp. LCCN QH541.2 .C63 2010.

Colon:2024:RGW

[Col24]  Sigma Colón.  Reimagining

group work through the lens of care Webs. Historical Studies in the Natural Sciences, 54 (1):92–??, February 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/92/200195/
Reimagining-Group-Work-
through-the-Lens-of-Care.

Conway:2006:DDS

[Con06] Erik M. Conway. Drowning in data: Satellite oceanography and information overload in the Earth sciences. Historical Studies in the Physical and Biological Sciences, 37(1): 127–151, September 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Conway:2011:IES

[Con11] Erik M. Conway. Infrastructures of Earth and sky. Historical Studies in the Natural Sciences, 41(3):344–353, Summer 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2011.41.3.344.

Cook:2007:MEC

[Coo07a] Harold John Cook. Matters of exchange: commerce, medicine, and science in the Dutch Golden Age. Yale University Press, New Haven, CT, USA, 2007. ISBN 0-300-11796-5. xiv + 562 pp. LCCN Q127.N2 C66

                                        2007. URL http://www.loc.
                                        gov/catdir/toc/ecip0619/
                                        2006026973.html.

Cooper:2007:IIL

[Coo07b] Alix Cooper. Inventing the indigenous: local knowledge and natural history in early modern Europe. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87087-9 (hardcover). xiii + 218 pp. LCCN GR880 .C67

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0703/2006037259-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0703/2006037259-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0729/2006037259-b.html.

Cook:2012:SCA

[Coo12] Harold J. Cook. Science as culture: An American viewpoint. Historical Studies in the Natural Sciences, 42(5):491– 495, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.491.pdf.

Cooper:2019:RSF

[Coo19] Austin R. Cooper. “A ray of sunshine on French tables”: Citrus fruit, colonial agronomy, and French rule in Algeria (1930–1962). Historical Studies in the Natural Sciences, 49(3): 241–272, June 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Cowles:2017:HN

[Cow17] Henry M. Cowles. History naturalized. Historical Studies in the Natural Sciences, 47(1):107– 116, February 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Carson:1999:EF

[CPWW99] Cathryn Carson, Ethan Pollock, Peter Westwick, and James H. Williams. Editors’ foreword. Historical Studies in the Physical and Biological Sciences, 30 (1):i–??, ???? 1999. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Carney:2009:SSA

[CR09]  Judith  Ann  Car-

ney and Richard Nicholas Rosomoff. In the shadow of slavery: Africa’s botanical legacy in the Atlantic world. University of California Press, Berkeley, CA, USA, 2009. ISBN 0-520-25750-2 (hardcover). xiv + 280 + 7 pp. LCCN E29.N3 C375 2009.

Camprubi:2016:GOC

[CR16] Lino Camprubí and Sam Robinson. A gateway to ocean circulation: Surveillance and sovereignty at Gibraltar. Historical Studies in the Natural Sciences, 46(4):429–459, September 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Cowles:2021:ERI

[CR21] Henry M. Cowles and Chitra Ramalingam. Essays & reviews: Introduction. Historical Studies in the Natural Sciences, 51(1):138–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/

51/1/138/116283/Essays-
amp-ReviewsIntroduction.

Cowles:2022:I

[CR22] Henry M. Cowles and Chitra Ramalingam. Introduction. Historical Studies in the Natural Sciences, 52(1):118–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/52/1/118/120495/
Introduction.

Crawford:2000:GSH

[Cra00] Elisabeth Crawford. German scientists and Hitler’s vendetta against the Nobel prizes. Historical Studies in the Physical and Biological Sciences, 31 (1):37–53, ???? 2000. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Crelinsten:1983:WWC

[Cre83] Jeffrey Crelinsten. William Wallace Campbell and the “Einstein Problem”: An observational astronomer confronts the Theory of Relativity. Historical Studies in the Physical Sciences, 14(1):1–91, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757525.

Creager:1995:FR

[Cre95] Angela N. Creager. In the fly room. Historical Studies in the Physical and Biological Sciences, 25(2):357–360, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Crease:2001:AHH

[Cre01] Robert P. Crease. Anxious history: The High Flux Beam Reactor and Brookhaven National Laboratory. Historical Studies in the Physical and Biological Sciences, 32(1):41–56, September 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Crease:2008:RSB

[Cre08] Robert P. Crease. Recombinant science: The birth of the Relativistic Heavy Ion Collider (RHIC). Historical Studies in the Natural Sciences, 38(4):535– 568, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Creager:2015:RCM

[Cre15] Angela N. H. Creager. Radiation, cancer, and mutation in the Atomic Age. Historical Studies in the Natural Sciences, 45(1):14–48, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2015.45.1.14.

Creager:2020:FB

[Cre20] Angela N. H. Creager. Finding the biological. Historical Studies in the Natural Sciences, 50(1– 2):74–80, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Cropper:1988:JJW

[Cro88] William H. Cropper. James Joule’s work in electrochemistry and the emergence of the First Law of Thermodynamics. Historical Studies in the Physical and Biological Sciences, 19(1): 1–15, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757615.

Crozier:2007:PCM

[Cro07] Anna Crozier. Practising colonial medicine: the Colonial Medical Service in British East Africa. I. B. Tauris, London, UK, 2007. ISBN 1-84511-459-0. xiv + 225 pp. LCCN RA395.A353 C76 2007.

Crowe:2020:BRE

[Cro20] Nathan Crowe. Book reviews:

The engineering ideal: Loeb’s ghost, or something else?, Luis A. Campos, Radium and the Secret of Life. Chicago: University of Chicago Press, 2015. 378 pp., illus., index. ISBN 978-0-226-41874-2. $35.00 (paper). Helen Anne Curry, Evolution Made to Order: Plant Breeding and Technological Innovation in Twentieth Century America. Chicago: University of Chicago Press, 2016. 285 pp., illus., index. ISBN 978-0-226-39008-6. $55.00 (cloth). Sophia Roosth, Synthetic: How Life Got Made. Chicago: University of Chicago Press, 2017. 251 pp., illus., index. ISBN 978-0-226-44046-0. $35.00 (paper). Historical Studies in the Natural Sciences, 50(5):624–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        50/5/624/114257/The-
                                        Engineering-IdealLoeb-s-
                                        Ghost-or-Something.

Crosland:1978:TPF

[CS78] Maurice Crosland and Crosbie Smith. The transmission of physics from France to Britain: 1800–1840. Historical Studies in the Physical Sciences, 9 (??):1–61, ???? 1978. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757376.

Csiszar:2024:BAO

[Csi24] Alex Csiszar. Blurry authorship: Originality in science before and after large language models. Historical Studies in the Natural Sciences, 54(5):611– ??, November 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/5/611/203887/Blurry-
                                        AuthorshipOriginality-in-
                                        Science-before-and.

Culotta:1974:GBO

[Cul74] Charles A. Culotta. German biophysics, objective knowledge, and romanticism. Historical Studies in the Physical Sciences, 4(??):3–38, ???? 1974.

CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (elec-

tronic). URL http://www.
jstor.org/stable/27757326.

Cullather:2010:HWA

[Cul10] Nick Cullather. The hungry world: America’s Cold War battle against poverty in Asia. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-05078-9 (hardcover), 0-674-72581-6 (paperback). xi + 348 + 8 pp. LCCN HD2056.Z8 C85 2010.

Curran:2011:ABS

[Cur11]  Andrew S. Curran.  The

anatomy of blackness: science and slavery in an age of Enlightenment. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0150-9 (hardcover). xiv + 310 pp. LCCN PQ265 .C87 2011.

Curry:2016:AAS

[Cur16] Helen Anne Curry. Atoms in agriculture: A study of scientific innovation between technological systems. Historical Studies in the Natural Sciences, 46(2): 119–153, April 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

DAgostino:1975:HRE

[D’A75]  Salvo D’Agostino.  Hertz’s

researches on electromagnetic waves. Historical Studies in the Physical Sciences, 6(??): 261–323, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757343.

Dahl:1984:KOD

[Dah84] Per F. Dahl. Kamerlingh Onnes and the discovery of superconductivity: The Leyden years, 1911–1914. Historical Studies in the Physical Sciences, 15 (1):1–37, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757541.

Dahl:1986:SAW

[Dah86] Per F. Dahl. Superconductivity after World War I and circumstances surrounding the discovery of a state B = 0. Historical Studies in the Physical and Biological Sciences, 16(1): 1–58, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757556.

Darrigol:1984:HQC

[Dar84] Olivier Darrigol. A history of the question: Can free electrons be polarized? Historical Studies in the Physical Sciences, 15 (1):39–79, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757542.

Darrigol:1986:OQM

[Dar86] Olivier Darrigol. The origin of quantized matter waves. Historical Studies in the Physical and Biological Sciences, 16(2): 197–253, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757565.

Darrigol:1988:SCE

[Dar88] Olivier Darrigol. Statistics and combinatorics in early quantum theory. Historical Studies in the Physical and Biological Sciences, 19(1):17–80, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757616.

Darrigol:1991:SCE

[Dar91] Olivier Darrigol. Statistics and combinatorics in early quantum theory, II: Early symptoma of indistinguishability and holism. Historical Studies in the Physical and Biological Sciences, 21 (2):237–298, ???? 1991. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757664.

Darrigol:1994:ETL

[Dar94] Olivier Darrigol. The electron theories of Larmor and Lorentz: a comparative study. Historical Studies in the Physical and Biological Sciences, 24(2):265–336, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:1996:EOR

[Dar96] Olivier Darrigol. The electrodynamic origins of Relativity theory. Historical Studies in the Physical and Biological Sciences, 26(2):241–312, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:1998:OPA

[Dar98a] Olivier Darrigol. From organ pipes to atmospheric motions: Helmholtz on fluid mechanics. Historical Studies in the Physical and Biological Sciences, 29 (1):1–54, ???? 1998. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:1998:TNT

[Dar98b] Olivier Darrigol. Toward a new topology of scientific practice. Historical Studies in the Physical and Biological Sciences, 28 (2):337–351, ???? 1998. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:2001:GWM

[Dar01] Olivier Darrigol. God, waterwheels, and molecules: Saint– Venant’s anticipation of energy conservation. Historical Studies in the Physical and Biological Sciences, 31(2):285–353, March 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:2002:TCH

[Dar02] Olivier Darrigol. Turbulence in 19th-century hydrodynamics. Historical Studies in the Physical and Biological Sciences, 32 (2):207–262, March 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Darrigol:2004:RAS

[Dar04] Olivier Darrigol. On a recent article by Seiya Abiko. Historical Studies in the Physical and Biological Sciences, 35(1): 153–155, September 2004. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See [Abi03] and reply [Abi04].

Daston:2012:S

[Das12] Lorraine Daston. Structure. Historical Studies in the Natural Sciences, 42(5):496–499, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.496.pdf.

Daumas:1948:LCF

[Dau48] Maurice Daumas. L’école des Chimistes Français vers 1840. Chymia: Annual Studies in the History of Chemistry, 1(??):55– 65, ???? 1948. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757114.

Daumas:1950:AEL

[Dau50] Maurice Daumas. Les Appareils d’Expérimentation de Lavoisier. (French) [Lavoisier’s experimental equipment]. Chymia: Annual Studies in the History of Chemistry, 3(??):45–62, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757145.

Daub:1970:ED

[Dau70] Edward E. Daub. Entropy and dissipation. Historical Studies in the Physical Sciences, 2(??): 321–354, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757310.

Davis:1948:EUP

[Dav48] Tenney L. Davis. The early use of potassium chlorate in pyrotechny: Dr. Moritz Meyer’s colored flame compositions. Chymia: Annual Studies in the History of Chemistry, 1(??):75–92, ???? 1948. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757116.

Davis:1949:PFT

[Dav49] Tenney L. Davis. Pulvis Fulminans Tenney L. Davis. Chymia: Annual Studies in the History of Chemistry, 2(??):ii + 99– 110, ???? 1949. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757126.

Davis:1986:IAC

[Dav86] John L. Davis. The influence of astronomy on the character of physics in mid-Nineteenth Century France. Historical Studies in the Physical and Biological Sciences, 16(1):59–82, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757557.

deChadarevian:2015:CPH

[dC15] Soraya de Chadarevian. Chromosome photography and the human karyotype. Historical Studies in the Natural Sciences, 45(1):115–146, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.1.115.

deChadarevian:2018:TDR

[dC18] Soraya de Chadarevian. Things and data in recent biology. Historical Studies in the Natural Sciences, 48(5):648–658, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

deChadarevian:2020:NSV

[dC20] Soraya de Chadarevian. Normalization and the search for variation in the human genome. Historical Studies in the Natural Sciences, 50(5):578–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/

                                        article/50/5/578/114255/
                                        Normalization-and-the-
                                        Search-for-Variation-in-
                                        the.

deChadarevian:2018:ISD

[dCP18] Soraya de Chadarevian and Theodore M. Porter. Introduction: Scrutinizing the data world. Historical Studies in the Natural Sciences, 48(5):549– 556, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Daumas:1959:LRU

[DD59] Maurice Daumas and Denis Duveen. Lavoisier’s relatively unknown large-scale decomposition and synthesis of water, February 27 and 28, 1785. Chymia: Annual Studies in the History of Chemistry, 5(??): 113–129, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757179.

DeGreiff:2002:TTP

[De 02] Alexis De Greiff. The tale of two peripheries: The creation of the International Centre for Theoretical Physics in Trieste. Historical Studies in the Physical and Biological Sciences, 33 (1):33–59, September 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

DeGregorio:2005:NPE

[De 05] Alberto G. De Gregorio. Neutron physics in the early 1930s.

Historical Studies in the Physical and Biological Sciences, 35 (2):293–340, March 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

delaPazRamosLara:2006:PAM

[de 06] María de la Paz Ramos Lara.

Particle accelerators in Mexico. Historical Studies in the Physical and Biological Sciences, 36 (2):297–309, March 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

DeGreiffA:2012:PDU

[De 12] Alexis De Greiff A. A las puertas del universo derrotado. (Spanish) [At the gates of the defeated universe]. Colección obra selecta. Universidad Nacional de Colombia, Vicerrectoría Académica, Bogotá, D.C., Colombia, 2012. ISBN 958-761-141-1 (hardcover), 958-761-142-X (paperback), 958-761-143-8 (e-book). 180 pp. LCCN QC7.5 .D425 2012.

Dear:2007:INH

[Dea07] Peter Robert Dear. The intelligibility of nature: how science makes sense of the world. Science.culture. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-13948-4, 0-226-13949-2. xii + 242 pp. LCCN Q175.32.R45. URL

http://www.gbv.de/dms/
bowker/toc/9780226139487;
 http://www.loc.gov/

                                        catdir/enhancements/
                                        fy0702/2005026032-b.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0702/2005026032-d.html;
                                         http://www.loc.gov/
                                        catdir/toc/ecip0518/
                                        2005026032.html.

Dear:2008:BRI

[Dea08] Peter Dear. Book review: The Inwardness of Science: Charles Coulston Gillispie. Essays and Reviews in History and History of Science. Philadelphia: American Philosophical Society, 2007. xix + 424 pp. ISBN 978-0-87169-965-7. $24 (paper). Historical Studies in the Natural Sciences, 38(4):621–625, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Dear:2014:DSN

[Dea14] Peter Dear. Darwin’s sleepwalkers: Naturalists, nature, and the practices of classification. Historical Studies in the Natural Sciences, 44(4):297– 318, September 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.4.297.

Debus:1962:SAP

[Deb62] Allen G. Debus. Solution analyses prior to Robert Boyle. Chymia: Annual Studies in the History of Chemistry, 8(??):41– 61, ???? 1962. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757217.

Deischer:1948:MTD

[Dei48] Claude K. Deischer. A memorial tribute to Dr. C. A. Browne: With a bibliography of his contributions to the history of chemistry. Chymia: Annual Studies in the History of Chemistry, 1(??):11–22, ???? 1948. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757111.

Deischer:1964:MSL

[Dei64] Claude K. Deischer. In memoriam to a secretary, librarian, curator of a library, and historian: Eva Vivian Armstrong — 1877–1962. Chymia: Annual Studies in the History of Chemistry, 9(??):13–17, ???? 1964. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757227.

Deichmann:2004:ERA

[Dei04] Ute Deichmann. Early responses to Avery et al.’s paper on DNA as hereditary material. Historical Studies in the Physical and Biological Sciences, 34 (2):207–232, March 2004. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Deichmann:2021:TTR

[Dei21] Ute Deichmann. Template theories, the rule of parsimony, and disregard for irreproducibility — the example of Linus Pauling’s research on antibody formation. Historical Studies in the Natural Sciences, 51(4):427– ??, September 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/51/4/427/
118595/Template-Theories-
the-Rule-of-Parsimony-and.

deMilt:1953:ALF

[deM53] Clara deMilt. Auguste Laurent, founder of modern organic chemistry. Chymia: Annual Studies in the History of Chemistry, 4(??):85–114, ???? 1953. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757164.

Dentinger:2017:AFT

[Den17] Rachel Mason Dentinger. From “actual forces” to “token stimuli”: Gottfried S. Fraenkel and the evolutionary ‘raison d’Être” of plant molecules, 1930s–1960s. Historical Studies in the Natural Sciences, 47(2):127–163, April 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Dent:2023:SM

[Den23] Rosanna Dent. Stolen masks. Historical Studies in the Natural Sciences, 53(1):76–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/1/76/195534/Stolen-
Masks.

DeVorkin:1987:OSR

[DeV87] David DeVorkin. Organizing for space research: The V-2 Rocket Panel. Historical Studies in the Physical and Biological Sciences, 18(1):1–24, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757594.

Devorkin:2000:WSA

[Dev00]  David H. Devorkin.  Who

speaks for astronomy? How astronomers responded to government funding after World War II. Historical Studies in the Physical and Biological Sciences, 31(1):55–92, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

DeYoung:2011:VMS

[DeY11] Ursula DeYoung. A vision of modern science: John Tyndall and the role of the scientist in Victorian culture. Palgrave studies in the history of science and technology. Palgrave Macmillan, New York, NY, 2011. ISBN 0-230-11053-3 (hardcover). 270 pp. LCCN Q143.T96

D43 2011. URL http://www.
palgraveconnect.com/
doifinder/10.1057/
9780230118058.

Daston:2007:O

[DG07] Lorraine Daston and Peter Galison. Objectivity. Zone Books, New York, NY, USA, 2007.

ISBN 1-890951-78-1. 501 + 38 pp. LCCN BD220 .D37

2007. URL http://www.loc.
gov/catdir/toc/ecip0720/
2007023997.html.

Diagne:2024:BAC

[DGH^{+}24] Awa  Hanane  Diagne,

Emily Grenon, Syeda Hasan, Holly K. M. Johnstone, Edward Jones-Imhotep, Alexander Offord, Neve Ostry Young, and Sarai Rudder. The black androids: a chrestomathy. Historical Studies in the Natural Sciences, 54(3):407–??, June 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/3/407/201970/The-Black-
AndroidsA-Chrestomathy.

Didier:2018:OAG

[Did18] Emmanuel Didier. Open-access genomic databases: A profitmaking tool? Historical Studies in the Natural Sciences, 48(5): 659–672, November 2018. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Dimoia:2014:BRP

[Dim14] John P. Dimoia. Book review: Projecting the future? the shifting boundaries of postwar American science. Biomedical Computing: Digitizing Life in the United States, by Joseph A. November; Instrumental Community: Probe Microscopy and the Path to Nanotechnology, by Cyrus C. M. Mody. Genentech:

Beginnings of Biotech, by Sally Smith Hughes. The Visioneers: How a Group of Elite Scientists Pursued Space Colonies, Nanotechnologies, and a Limitless Future, by W. Patrick Mc-Cray. Historical Studies in the Natural Sciences, 44(1):90–98, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2014.44.1.90.

DiMoia:2024:PCT

[DiM24] John P. DiMoia. Placing computer technology in context. Historical Studies in the Natural Sciences, 54(1):95–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/95/200199/Placing-
Computer-Technology-in-
Context.

DeMaria:1991:DCR

[DIR91] M. De Maria, M. G. Ianniello, and A. Russo. The discovery of cosmic rays: Rivalries and controversies between Europe and the United States. Historical Studies in the Physical and Biological Sciences, 22(1): 165–192, ???? 1991. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757676.

Dunning:2024:ITT

[DK24] David E. Dunning and Judith R. H. Kaplan. Interdisciplinary team teaching. Historical Studies in the Natural Sciences, 54(1):99–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/1/99/200197/
                                        Interdisciplinary-Team-
                                        Teaching.

Dejong-Lambert:2009:ELE

[DL09] William Dejong-Lambert. From eugenics to Lysenkoism: The evolution of Stanisław Skowron. Historical Studies in the Natural Sciences, 39(3):269–299, August 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Daston:2011:HSO

[DL11] Lorraine Daston and Elizabeth Lunbeck, editors. Histories of scientific observation. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-13677-9 (hardcover), 0-226-13678-7 (paperback). 460 + 1 pp. LCCN Q174.8 .H57 2011.

deMilt:1948:CWC

[dM48] Clara de Milt. Carl Weltzein and the Congress at Karlsruhe. Chymia: Annual Studies in the History of Chemistry, 1(??): 153–169, ???? 1948. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757121.

Dorries:2006:ICC

[DM06]  Matthias  Dörries  and

Christophe Masutti. Introduction: Changing climate — modeling climate. Historical Studies in the Physical and Biological Sciences, 37(1):1–2, September 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Dmitriev:2004:SDS

[Dmi04] Igor S. Dmitriev. Scientific discovery in statu nascendi: The case of Dmitrii Mendeleev’s Periodic Law. Historical Studies in the Physical and Biological Sciences, 34(2):233–275, March 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Doing:2009:VRS

[Doi09] Park Doing. Velvet Revolution at the synchrotron: biology, physics, and change in science. Inside technology. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-04255-X (hardcover). viii + 152 pp. LCCN QC51.N7 D65 2009.

Dolby:1976:DTS

[Dol76] R. G. A. Dolby. Debates over the theory of solution: a study of dissent in physical chemistry in the English-speaking world in the late Nineteenth and early Twentieth Centuries. Historical Studies in the Physical Sciences, 7(??):297–404, ???? 1976. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757359.

Donghi:1967:CRP

[Don67] Letícá Halperín Donghí. Chemistry in the rio de la plata at the end of the colonial epoch. Chymia: Annual Studies in the History of Chemistry, 12(??): 119–126, ???? 1967. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757276.

Doran:1975:OCF

[Dor75] Barbara Giusti Doran. Origins and consolidation of field theory in Nineteenth-Century Britain: From the mechanical to the electromagnetic view of nature. Historical Studies in the Physical Sciences, 6(??): 133–260, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757342.

Dorries:1991:PHA

[Dör91] Matthias Dörries. Prior history and aftereffects: Hysteresis and “Nachwirkung” in 19th-Century physics. Historical Studies in the Physical and Biological Sciences, 22(1):25–55, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757672.

Dorries:1995:HKP

[Dör95] Matthias Dörries. Heinrich Kayser as philologist of physics.

Historical Studies in the Physical and Biological Sciences, 26 (1):1–33, ???? 1995. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Dorries:2006:PEV

[Dör06] Matthias Dörries. In the public eye: Volcanology and climate change studies in the 20th Century. Historical Studies in the Physical and Biological Sciences, 37(1):87–125, September 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Dorries:2012:BRL

[Dör12] Matthias Dörries. Book review: Life, language, and science: Hans-Jörg Rheinberger’s historical epistemology: An Epistemology of the Concrete: Twentieth-Century Histories of Life by Hans-Jörg Rheinberger; On Historicizing Epistemology: An Essay by Hans-Jörg Rheinberger. Historical Studies in the Natural Sciences, 42(1):71–82, February 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2012.42.1.71. See

[Rhe10a, Rhe10b].

Dorries:2020:HCC

[Dör20] Matthias Dörries. Hot climate, Cold War. Historical Studies in the Natural Sciences, 50(1– 2):67–73, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Dorries:2021:ALH

[Dör21] Matthias Dörries. The art of listening: Hugo Benioff, seismology, and music. Historical Studies in the Natural Sciences, 51(4):468–??, September 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        51/4/468/118596/The-Art-
                                        of-ListeningHugo-Benioff-
                                        Seismology-and.

Doxiadis:2011:LES

[DPPdD11] Apostolos Doxiadis, Christos H. Papadimitriou, Alecos Papadatos, and Annie di Donna. Logicomix: an epic search for truth. Bloomsbury, New York, NY, USA, 2011. ISBN 1-59691-452-1. 347 pp. LCCN PN6790.D69 L64 2009.

Deischer:1950:OHK

[DR50]  Claude  K.

Deischer and Joseph L. Rabinowitz. The owl of Heinrich Khunrath: Its origin and significance. Chymia: Annual Studies in the History of Chemistry, 3(??):ii + 243–250, ???? 1950. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757142.

DeMaria:1989:CRR

[DR89] M. De Maria and A. Russo. Cosmic ray romancing: The discovery of the latitude effect and the Compton–Millikan controversy. Historical Studies in the Physical and Biological Sciences, 19 (2):211–266, ???? 1989. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757626.

Druzhinin:2020:FPM

[Dru20] Petr A. Druzhinin. The first publication of Mendeleev’s periodic system of elements: A new chronology. Historical Studies in the Natural Sciences, 50(1– 2):129–182, April 2020. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Dryer:2019:NCH

[Dry19] Theodora Dryer. The new critical history of surveillance and human data. Historical Studies in the Natural Sciences, 49(5): 556–565, November 2019. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Duncan:2023:BMM

[Dun23] Katy Duncan. Between the mountain, the meadow, the calm, and the storm: Exner’s and Elster and Geitel’s electrical atmospheres. Historical Studies in the Natural Sciences, 53(4):349–??, September 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/4/349/197391/Between-
the-Mountain-the-Meadow-
the-Calm-and-the.

Duveen:1953:ML

[Duv53] Denis I. Duveen. Madame Lavoisier 1758–1836. Chymia: Annual Studies in the History of Chemistry, 4(??):13–29, ???? 1953. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757161.

Duval:1967:PRC

[Duv67] Clément Duval. Pilatre de Rozier (1754–1785), chemist and first aeronaut. Chymia: Annual Studies in the History of Chemistry, 12(??):99–117, ???? 1967. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757275.

Duveen:1949:SSC

[DW49] Denis I. Duveen and Antoine Willemart. Some Seventeenth Century chemists and alchemists of Lorraine. Chymia: Annual Studies in the History of Chemistry, 2(??):111–117, ???? 1949. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757137.

Eckert:1987:PSS

[Eck87] Michael Eckert. Propaganda in science: Sommerfeld and the spread of the electron theory of metals. Historical Studies in the Physical and Biological Sciences, 17(2):191–233, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757582.

Eckert:1988:NPM

[Eck88] Michael Eckert. Neutrons and politics: Maier–Leibnitz and the emergence of pile neutron research in the FRG. Historical Studies in the Physical and Biological Sciences, 19(1): 81–113, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757617.

Eckert:1990:PDF

[Eck90] Michael Eckert. Primacy doomed to failure: Heisenberg’s role as scientific adviser for nuclear policy in the FRG. Historical Studies in the Physical and Biological Sciences, 21(1):29–58, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757654.

Edelstein:1948:PSW

[Ede48] Sidney M. Edelstein. Priestley settles the water controversy. Chymia: Annual Studies in the History of Chemistry, 1 (??):123–137, ???? 1948. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757119.

Edelstein:1959:CRA

[Ede59] Sidney M. Edelstein. The chemical revolution in America from the pages of the “Medical Repository”. Chymia: Annual Studies in the History of Chemistry, 5 (??):155–179, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757183.

Editors:2002:F

[Edi02] Guest Editors. Foreword. Historical Studies in the Physical and Biological Sciences, 33(1): 1–2, September 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Edmond:2006:LEM

[Edm06] Rod Edmond. Leprosy and empire: a medical and cultural history. Cambridge social and cultural histories. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-86584-0. x + 255 pp. LCCN RC154.4 .E36

                                        2006. URL http://www.h-net.
                                        org/review/hrev-a0f5c6-aa;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0733/2007277435-b.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0733/2007277435-d.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0733/2007277435-t.html.

Earman:1980:REB

[EG80] John Earman and Clark Glymour. Relativity and eclipses: The British Eclipse Expeditions of 1919 and their predecessors. Historical Studies in the Physical Sciences, 11(1):49–85, ???? 1980. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757471.

Eisler:2016:MRS

[Eis16] Matthew N. Eisler. Materials research, super batteries, and the technopolitics of electric automobility. Historical Studies in the Natural Sciences, 46(1): 44–66, February 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Elkana:1970:HKI

[Elk70] Yehuda Elkana. Helmholtz’ “Kraft”: An illustration of concepts in flux. Historical Studies in the Physical Sciences, 2 (??):263–298, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757308.

Elman:2012:ITS

[Elm12] Benjamin A. Elman. It took a scientist to historicize one! Historical Studies in the Natural Sciences, 42(5):500–503, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.500.pdf.

Endersby:2007:GPH

[End07] Jim Endersby. A guinea pig’s history of biology. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02713-2. xii + 499 pp. LCCN QH431 .E615 2007. URL http://www.

loc.gov/catdir/toc/
ecip0718/2007020824.html.

Ericson:2023:PTC

[Eri23] Kjell David Ericson. The puzzle of the thinly coated pearl: Aquacultural ecology and the politics of density in Ago Bay. Historical Studies in the Natural Sciences, 53(3):256–??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/3/256/196658/The-
Puzzle-of-the-Thinly-
Coated-PearlAquacultural.

Esposito:2015:MTP

[Esp15] Maurizio Esposito. More than the parts: W. E. Ritter, the Scripps Marine Association, and the organismal conception of life. Historical Studies in the Natural Sciences, 45(2):273– 302, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2015.45.2.273.

Fabian:2010:SCR

[Fab10] Ann Fabian. The skull collectors: race, science, and America’s unburied dead. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-23348-0 (hardcover). xi + 270 pp. LCCN GN90 .F33 2010.

Falconer:1988:JJT

[Fal88] Isobel Falconer. J. J. Thomson’s work on positive rays, 1906–1914. Historical Studies in the Physical and Biological Sciences, 18(2):265–310, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757604.

Farber:1950:BAS

[Far50] Eduard Farber. Bio-active substances in the Nineteenth Century. Chymia: Annual Studies in the History of Chemistry, 3 (??):63–76, ???? 1950. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757146.

Farber:1961:ESC

[Far61] Eduard Farber. Early studies concerning time in chemical reactions. Chymia: Annual Studies in the History of Chemistry, 7(??):135–148, ???? 1961. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757210.

Farber:1962:DMH

[Far62] Eduard Farber. The development of metal hydride chemistry. Chymia: Annual Studies in the History of Chemistry, 8 (??):165–180, ???? 1962. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757223.

Farber:1964:TEN

[Far64a] Eduard Farber. The theory of the elements and nucleosynthesis in the Nineteenth Century. Chymia: Annual Studies in the History of Chemistry, 9 (??):181–200, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757240.

Farrar:1964:SBC

[Far64b] W. V. Farrar. Sir B. C. Brodie and his calculus of chemical operations. Chymia: Annual Studies in the History of Chemistry, 9(??):169–179, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757239.

Farber:1965:IOR

[Far65]  Eduard Farber.  In-

duced oxidation–reduction processes, the history of a chemical paradox. Chymia: Annual Studies in the History of Chemistry, 10(??):121–181, ???? 1965. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757249.

Farber:1966:CPW

[Far66] Eduard Farber. From chemistry to philosophy: The way of Alwin Mittasch (1869–1953). Chymia: Annual Studies in the History of Chemistry, 11(??): 157–178, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757266.

Feffer:1989:ASJ

[Fef89] Stuart M. Feffer. Arthur Schuster, J. J. Thomson, and the discovery of the electron. Historical Studies in the Physical and Biological Sciences, 20(1): 33–61, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757634.

Feffer:1992:ACP

[Fef92] Stuart M. Feffer. Atoms, cancer, and politics: Supporting atomic science at the University of Chicago, 1944–1950. Historical Studies in the Physical and Biological Sciences, 22(2): 233–261, ???? 1992. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757682.

Feldman:1985:AMQ

[Fel85] Theodore S. Feldman. Applied mathematics and the quantification of experimental physics: The example of barometric hypsometry. Historical Studies in the Physical Sciences, 15(2): 127–195, ???? 1985. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757551.

Feldman:2013:BRT

[Fel13] James W. Feldman. Book review: Testing the nuclear paradox: Atomic Frontier Days: Hanford and the American West Full Body Burden: Growing Up in the Nuclear Shadow of Rocky Flats The Rise of Nuclear Fear. Historical Studies in the Natural Sciences, 43(4):

531–538, September 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2013.
                                        43.4.531.

Fester:1962:CCA

[Fes62] G. A. Fester. Copper and copper alloys in Ancient Argentina. Chymia: Annual Studies in the History of Chemistry, 8(??):21–31, ???? 1962. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757215.

Frederick-Frost:2017:LMH

[FF17] K. M. Frederick-Frost. For the love of a mother — Henry Moseley’s rare earth research. Historical Studies in the Natural Sciences, 47(4):529–567, September 2017. ISSN 1939-1811 (print), 1939-182X (electronic).

Fox:1998:LWS

[FG98] Robert Fox and Anna Guagnini. Laboratories, workshops, and sites. Concepts and practices of research in Industrial Europe, 1800–1914. Historical Studies in the Physical and Biological Sciences, 29(1):55–140, ???? 1998. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Fox:1999:LWS

[FG99] Robert Fox and Anna Guagnini. Laboratories, workshops, and sites. Concepts and practices of research in industrial Europe, 1800–1914. Historical Studies in the Physical and Biological Sciences, 29(2):191–??, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Findlay:2011:AFD

[FH11] John M. Findlay and Bruce William Hevly. Atomic frontier days: Hanford and the American West. Emil and Kathleen Sick series in Western history and biography. Center for the Study of the Pacific Northwest in association with University of Washington Press, Seattle, WA, USA, 2011. ISBN 0-295-99097-X (paperback). xv + 368 pp. LCCN TD898.12.W2 F56 2011.

Fedman:2019:HDI

[FH19]  David  Fedman

and Martin Hood. Hotly debated ice: Scholar alpinism and the great glacier controversy in Modern Japan. Historical Studies in the Natural Sciences, 49 (3):273–299, June 2019. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Forman:1975:PCP

[FHW75] Paul Forman, John L. Heilbron, and Spencer Weart. Physics circa 1900: Personnel, funding, and productivity of the academic establishments. Historical Studies in the Physical Sciences, 5(??):1–185, ???? 1975. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (elec-

tronic). URL http://www.
jstor.org/stable/27757335.

Figurovskii:1966:HCA

[Fig66] N. A. Figurovskii. The history of chemistry in Ancient Russia. Chymia: Annual Studies in the History of Chemistry, 11 (??):45–79, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757259.

Finnegan:2013:BRM

[Fin13] Diarmid A. Finnegan. Book reviews: Making science, making geography: K. Maria D. Lane, Geographies of Mars: Seeing and Knowing the Red Planet. Chicago: University of Chicago Press, 2010. xiii + 266 pp., illus., index. ISBN 978-0-226-47078-8. $45.00 (cloth).; S. M. Reid-Henry, Cuban Cure: Reason and Resistance in Global Science. Chicago: University of Chicago Press, 2010. xii + 200 pp., illus., index. ISBN 978-0-226-70917-8. $39.00 (cloth).; Helen Tilley, Africa as a Living Laboratory: Empire, Development and the Problem of Scientific Knowledge, 1870–1950. Chicago: University of Chicago Press, 2011. xiv + 496 pp., illus., index. ISBN 978-0-226-80347-0. $29.00 (paper). Historical Studies in the Natural Sciences, 43(1):113–119, February 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/

stable/pdfplus/10.1525/
hsns.2013.43.1.113.pdf.

Finnegan:2021:SST

[Fin21] Diarmid A. Finnegan. Of snails and salvation: The theological construction of John Thomas Gulick’s theory of evolution. Historical Studies in the Natural Sciences, 51(5):567–??, November 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/51/5/567/119068/
Of-Snails-and-
SalvationThe-Theological.

Freeborn:2024:DHE

[FKW24] Alfred Freeborn, Lara Keuck, and Hanna Lucia Worliczek. Did Hamilton ever use the Hamilton rating scale for depression? Reflections on the (re-)Use of oral histories and their accessibility via the Com-Bio Website. Historical Studies in the Natural Sciences, 54(4):527–??, September 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/4/527/203065/Did-
Hamilton-Ever-Use-the-
Hamilton-Rating-Scale.

Fleming:2000:SSS

[Fle00a] James Rodger Fleming. Storms, strikes, and surveillance: The U.S. Army Signal Office, 1861– 1891. Historical Studies in the

Physical  and  Bio-

logical Sciences, 30(2):315–332, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Fleming:2000:F

[Fle00b] James Roger Fleming. Foreword. Historical Studies in the Physical and Biological Sciences, 30(2):307–??, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Fleming:2006:PHW

[Fle06] James Rodger Fleming. The pathological history of weather and climate modification: Three cycles of promise and hype. Historical Studies in the Physical and Biological Sciences, 37(1): 3–25, September 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Fleetwood:2020:BRS

[Fle20] Lachlan Fleetwood. Book reviews: Science in India and Indians in science: [David Arnold. Toxic Histories: Poison and Pollution in Modern India. Cambridge: Cambridge University Press, 2016. ix + 241 pp., illus., index. ISBN 978-1-107-12697-8. $49.99 (paper). Pankaj Sekhsaria. Instrumental Lives: An Intimate Biography of an Indian Laboratory. Abingdon: Routledge, 2018. xxii + 126 pp., illus., index. ISBN 978-1-138-58767-0. $60.00 (hardcover). Joydeep Sen. Astronomy in India, 1784 1876. London: Pickering & Chatto, 2014. xiii + 268 pp., illus., index. ISBN 978-1-781-44078-0. $99.00 (hardcover)]. Historical Studies in the Natural Sciences, 50(3):302–310, June 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/3/302/110323/Science-
in-India-and-Indians-in-
Science.

Fisher:2016:SII

[FM16] Amy A. Fisher and Joseph D.

Martin. Special issue introduction: Emerging prospects for history of the physical sciences. Historical Studies in the Natural Sciences, 46(3):261–269, June 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Fernandez-Mulligan:2024:MGH

[FM24] Sebastian Fernandez-Mulligan.

From the model to the glance: How astronomers learned to see gravitational lenses, 1960–2020. Historical Studies in the Natural Sciences, 54(4):461–??, September 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/54/4/461/203068/
From-the-Model-to-the-
GlanceHow-Astronomers.

Forbes:1949:WNA

[For49] R. J. Forbes. Was Newton an alchemist? Chymia: Annual Studies in the History of Chemistry, 2(??):27–36, ???? 1949. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757130.

Forbes:1953:OA

[For53] R. J. Forbes. On the origin of alchemy. Chymia: Annual Studies in the History of Chemistry, 4(??):1–11, ???? 1953. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757160.

Forman:1970:ALA

[For70] Paul Forman. Alfred Landé and the Anomalous Zeeman Effect, 1919–1921. Historical Studies in the Physical Sciences, 2(??): 153–261, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757307.

Forman:1971:WCC

[For71] Paul Forman. Weimar culture, causality, and quantum theory, 1918–1927: Adaptation by German physicists and mathematicians to a hostile intellectual environment. Historical Studies in the Physical Sciences, 3 (??):1–115, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757315.

Forman:1987:BQE

[For87] Paul Forman. Behind quantum electronics: National security as basis for physical research in the United States, 1940–1960. Historical Studies in the Physical and Biological Sciences, 18 (1):149–229, ???? 1987. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757599.

Forman:2007:HLM

[For07]  Paul Forman.  How Lewis

Mumford saw science, and art, and himself. Historical Studies in the Physical and Biological Sciences, 37(2):271–336, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Fortun:2008:PGI

[For08] Michael Fortun. Promising genomics: Iceland and deCODE Genetics in a world of speculation. University of California Press, Berkeley, CA, USA, 2008. ISBN 0-520-24750-7 (cloth), 0-520-24751-5 (paperback). x + 330 pp. LCCN QH438.7 .F67

2008. URL http://www.loc.
gov/catdir/toc/ecip083/
2007044597.html.

Fox:1974:RFL

[Fox74] Robert Fox. The rise and fall of Laplacian physics. Historical Studies in the Physical Sciences, 4(??):89–136, ???? 1974. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757328.

Fujimura:2020:REA

[FR20]  Joan  H.

Fujimura and Ramya M. Rajagopalan. Race, ethnicity, ancestry, and genomics in Hawai’i: Discourses and practices. Historical Studies in the Natural Sciences, 50(5):596–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

                                        online.ucpress.edu/hsns/
                                        article/50/5/596/114256/
                                        Race-Ethnicity-Ancestry-
                                        and-Genomics-in-Hawai.

Frankel:1977:JBB

[Fra77] Eugene Frankel. J. B. Biot and the mathematization of experimental physics in Napoleonic France. Historical Studies in the Physical Sciences, 8(??): 33–72, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757367.

Franklin:1981:MPU

[Fra81] Allan D. Franklin. Millikan’s published and unpublished data on oil drops. Historical Studies in the Physical Sciences, 11 (2):185–201, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757478.

Frankel:1982:DRA

[Fra82] Henry Frankel. The development, reception, and acceptance of the Vine–Matthews–Morley hypothesis. Historical Studies in the Physical Sciences, 13 (1):1–39, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757504.

Franklin:1983:DAC

[Fra83] Allan Franklin. The discovery and acceptance of CP violation. Historical Studies in the Physical Sciences, 13(2): 207–238, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757515.

Freudenthal:1981:EEB

[Fre81] Gad Freudenthal. Early electricity between chemistry and physics: The simultaneous itineraries of Francis Hauksbee, Samuel Wall, and Pierre Polinière. Historical Studies in the Physical Sciences, 11(2):203–229, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757479.

Frenkel:1997:YIF

[Fre97] Victor Ya Frenkel. Yakov Ilich Frenkel: Sketches toward a civic portrait. Historical Studies in the Physical and Biological Sciences, 27(2):197–236, ???? 1997. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Freire:2005:SED

[Fre05] Olival Freire Jr. Science and exile: David Bohm, the Cold War, and a new interpretation of quantum mechanics. Historical Studies in the Physical and Biological Sciences, 36(1): 1–34, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Friedman:1977:CNS

[Fri77] Robert Marc Friedman. The creation of a new science: Joseph Fourier’s Analytical Theory of Heat. Historical Studies in the Physical Sciences, 8 (??):73–99, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757368.

Friedman:1989:TCQ

[Fri89] Robert Marc Friedman. Text, context, and quicksand: Method and understanding in studying the Nobel Science Prizes. Historical Studies in the Physical and Biological Sciences, 20(1): 63–77, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757635.

Friedman:2007:TIJ

[Fri07] Robert Marc Friedman. Tolerance and integrity at Johns Hopkins. Historical Studies in the Physical and Biological Sciences, 37(2):463–474, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Fullmer:1960:PSH

[Ful60] J. Z. Fullmer. The poetry of Sir Humphry Davy. Chymia: Annual Studies in the History of Chemistry, 6(??):102–126, ???? 1960. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757194.

Fullmer:1962:HDA

[Ful62] J. Z. Fullmer. Humphry Davy’s adversaries. Chymia: Annual Studies in the History of Chemistry, 8(??):147–164, ???? 1962. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757222.

Fullmer:1964:HDC

[Ful64] J. Z. Fullmer. Humphry Davy’s critical abstracts. Chymia: Annual Studies in the History of Chemistry, 9(??):97–115, ???? 1964. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757235.

Fullmer:1967:DSH

[Ful67] J. Z. Fullmer. Davy’s sketches of his contemporaries. Chymia: Annual Studies in the History of Chemistry, 12(??):127– 150, ???? 1967. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757277.

Fetter-Vorm:2013:TGH

[FV13] Jonathan Fetter-Vorm. Trinity: a graphic history of the first atomic bomb. Novel graphics. Hill and Wang: a division of Farrar, Straus and Giroux, New York, NY, USA, 2013. ISBN 0-8090-9355-3. 154 pp. LCCN QC773.3.U5 F47 2013.

Fyfe:2012:SPK

[Fyf12] Aileen Fyfe. Steam-powered knowledge: William Chambers and the business of publishing, 1820–1860. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-27651-1 (hardcover). xvi + 313 pp. LCCN Z325.C44 F94 2012.

Galison:1979:MST

[Gal79]  Peter  Louis  Gal-

ison. Minkowski’s space–time: From visual thinking to the absolute world. Historical Studies in the Physical Sciences, 10 (??):85–121, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757388.

Galison:1982:TPE

[Gal82]  Peter Galison.  Theoretical

predispositions in experimental physics: Einstein and the gyromagnetic experiments, 1915– 1925. Historical Studies in the Physical Sciences, 12(2): 285–323, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757498.

Galka:2023:OOD

[Gal23] Jonathan Galka. Oceans of ooze: Deep-sea sedimentary data, mineral resource frontiers, and Imperial continuities in ocean history. Historical Studies in the Natural Sciences, 53(5):481–??, November 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/5/481/198192/Oceans-of-
OozeDeep-Sea-Sedimentary-
Data-Mineral.

Ganzenmueller:1953:ZGA

[Gan53] W. Ganzenmueller. Zukunftsaufgaben der Geschichte der Alchemie. (German) [Future tasks in the history of alchemy]. Chymia: Annual Studies in the History of Chemistry, 4(??):31–36, ???? 1953. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757162.

Garber:1970:CMK

[Gar70] Elizabeth Wolfe Garber. Clausius and Maxwell’s kinetic theory of gases. Historical Studies in the Physical Sciences, 2 (??):299–319, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757309.

Garber:1978:MSL

[Gar78]  Elizabeth Garber.  Molecu-

lar science in late-Nineteenth-Century Britain. Historical Studies in the Physical Sciences, 9(??):265–297, ???? 1978. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757380.

Garber:2009:BRU

[Gar09] Margaret D. Garber. Book reviews: Untwisting the Greene Lyon’s Tale: Deborah E. Harkness. The Jewel House: Elizabethan London and the Scientific Revolution. New Haven: Yale University Press, 2007. xviii + 349 pp., illus., index. ISBN: 978-0-300-11196-5. $32.50 (cloth). Lauren Kassell. Medicine and Magic in Elizabethan London: Simon Forman: Astrologer, Alchemist & Physician. Oxford: Clarendon Press, 2005. xviii + 281 pp., illus., index. ISBN: 0-19-927905-5, 978-0-19-927905-0. $90.00 (cloth). Bruce Moran. Distilling Knowledge: Alchemy, Chemistry, and the Scientific Revolution. Cambridge, MA: Harvard University Press, 2005. 210 pp., illus., index. ISBN: 0-674-01495-2. $24.95 (cloth). Tara Nummedal. Alchemy and Authority in the Holy Roman Empire. Chicago: University of Chicago Press, 2007. xvii + 260 pp., illus., index. ISBN-13: 978-0-226-60856-3. $37.50 (cloth). Historical Studies in the Natural Sciences, 39(4):491– 500, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Garber:2012:I

[Gar12] Daniel Garber. Incommensurabilities. Historical Studies in the Natural Sciences, 42(5): 504–509, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.504.pdf.

Gasca:1996:MTV

[Gas96] Ana Millán Gasca. Mathematical theories versus biological facts: a debate on mathematical population dynamics in the 1930s. Historical Studies in the Physical and Biological Sciences, 26(2):347–403, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Gaston:2021:EW

[Gas21] Pablo Gastón. The essential worker. Historical Studies in the Natural Sciences, 51(1):146–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/51/1/146/116285/
The-Essential-Worker.

Gaukroger:2008:ESC

[Gau08] Stephen Gaukroger. The emergence of a scientific culture: science and the shaping of modernity 1210–1685. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-929644-8 (hardcover), 0-19-955001-8 (paperback). ix + 563 pp. LCCN BL240.3 .I58 v.

199. US$19.99. URL http://

www.gbv.de/dms/bowker/toc/
9780199550012.

Gausemeier:2010:GMP

[Gau10] Bernd Gausemeier. Genetics as a modernization program: Biological research at the Kaiser Wilhelm Institutes and the political economy of the Nazi State. Historical Studies in the Natural Sciences, 40(4):429– 456, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2010.40.4.429.

Gavroglu:1990:RBP

[Gav90] Kostas Gavroglu. The reaction of the British physicists and chemists to van der Waals’ early work and to the Law of Corresponding States. Historical Studies in the Physical and Biological Sciences, 20(2):199– 237, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757643.

Galison:1989:LSD

[GB89] Peter Galison and Barton Bernstein. In any light: Scientists and the decision to build the superbomb, 1952–1954. Historical Studies in the Physical and Biological Sciences, 19(2): 267–347, ???? 1989. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757627.

Gorman:1959:HSA

[GD59] Mel Gorman and Charles Doering. History of the structure of acetone. Chymia: Annual Studies in the History of Chemistry, 5(??):202–208, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757186.

Geppert:2011:IOS

[Gep11] Alexander C. T. Geppert, editor. Imagining outer space: European astroculture in the twentieth century. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2011. ISBN 0-230-23172-1 (hardcover). xvii + 393 pp. LCCN TL789.6.E9

I43 2012. URL http://www.
netread.com/jcusers2/
bk1388/726/9780230231726/
image/lgcover.
9780230231726.jpg.

Gere:2021:SRC

[Ger21] Cathy Gere. “Shovel-Ready”:

The commons and the climate crisis. Historical Studies in the Natural Sciences, 51(4):542–??, September 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/51/4/542/118592/

                                        Shovel-Ready-The-Commons-
                                        and-the-Climate-Crisis.

Ganzenmueller:1950:WGB

[GG50]  W.  Ganzenmueller  and

W. Ganzenmüller. Wandlungen in der geschichtlichen Betrachtung der Alchemie. (German) [Transformations in the historical view of alchemy]. Chymia: Annual Studies in the History of Chemistry, 3(??):143–154, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757150.

Galdabini:1988:PIB

[GG88] Silvana Galdabini and Giuseppe Giuliani. Physics in Italy between 1900 and 1940: The universities, physicists, funds, and research. Historical Studies in the Physical and Biological Sciences, 19(1):115–136, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757618.

Gribbe:2017:EGM

[GH17] Johan Gribbe and Olof Hallonsten. The emergence and growth of materials science in Swedish universities. Historical Studies in the Natural Sciences, 47(4):459–493, September 2017. ISSN 1939-1811 (print), 1939-182X (electronic).

Ghosh:2025:LSD

[Gho25] Nayanika Ghosh. Legitimating the sociobiology debate in American science. Historical Studies in the Natural Sciences, 55(1):48–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/48/207539/
Legitimating-the-
Sociobiology-Debate-in-
American.

Gibbs:2012:RBK

[Gib12] Frederick W. Gibbs. Riding the bicycle of Kuhn’s structure. Historical Studies in the Natural Sciences, 42(5):510– 513, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.510.pdf.

Giffard:2020:ENS

[Gif20] Hermione Giffard. Exploiting Nazi science and technology and the history of technology transfer. Historical Studies in the Natural Sciences, 50(1–2):199– 208, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Gillispie:1992:SSW

[Gil92] Charles Coulston Gillispie. Science and secret weapons development in Revolutionary France, 1792– 1804: a documentary history. Historical Studies in the Physical and Biological Sciences, 23 (1):35–152, ???? 1992. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757692.

Gillispie:2007:ERH

[Gil07] Charles Coulston Gillispie. Essays and reviews in history and history of science, volume 96, part 5 of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 2007. ISBN 0-87169-965-6 (paperback). xix + 424 pp. LCCN Q126.8 .G55 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy0710/2006052602.html.

Gillispie:2008:BRA

[Gil08] Charles C. Gillispie. Book review: Atop Mont Blanc: Bursting the Limits of Time: The Reconstruction of Geohistory in the Age of Revolution, by Martin J. S. Rudwick. Chicago and London: University of Chicago Press, 2005. xxiv + 708 pp., illus., index. $45.00 (cloth), $35.00 (paper). Historical Studies in the Natural Sciences, 38(1):163–171, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Gilbert:2023:SGG

[Gil23] Scott F. Gilbert. Shells, gills, and gonads: On the remarkable persistence of oysters in the Chesapeake Bay. Historical Studies in the Natural Sciences, 53(3):221–??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/3/221/196655/Shells-
Gills-and-GonadsOn-the-
Remarkable.

Gissis:2011:VRE

[Gis11] Snait B. Gissis. Visualizing “race” in the Eighteenth Century. Historical Studies in the Natural Sciences, 41(1):41–103, Winter 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2011.41.1.41.

Gliboff:2008:HGB

[Gli08] Sander Gliboff. H. G. Bronn, Ernst Haeckel, and the origins of German Darwinism: a study in translation and transformation. Transformations. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-07293-9 (hardcover). xii + 259 pp. LCCN QH366.2 .G55

2008. URL http://www.loc.
gov/catdir/toc/ecip081/
2007039859.html.

Gordin:2012:RMT

[GM12] Michael D. Gordin and Erika Lorraine Milam. A repository for more than anecdote: Fifty years of the structure of scientific revolutions. Historical Studies in the Natural Sciences, 42(5):476–478, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.476.pdf.

Gode:1949:HAI

[God49] P. K. Gode. History of ambergris in India between about A.D. 700 and 1900. Chymia: Annual Studies in the History of Chemistry, 2(??):51–56, ???? 1949. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757133.

Goldberg:1960:CO

[Gol60] Edward D. Goldberg. Chemists and the oceans. Chymia: Annual Studies in the History of Chemistry, 6(??):162–179, ???? 1960. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757197.

Goldberg:1970:DEB

[Gol70] Stanley Goldberg. In defense of ether: The British response to Einstein’s Special Theory of Relativity, 1905–1911. Historical Studies in the Physical Sciences, 2(??):89–125, ???? 1970. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757305.

Goldberg:1976:MPP

[Gol76] Stanley Goldberg. Max Planck’s philosophy of nature and his elaboration of the Special Theory of Relativity. Historical Studies in the Physical Sciences, 7(??):125–160, ???? 1976. CODEN HSPSAS. ISBN 0-691-08169-7. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757355.

Golan:2000:BWD

[Gol00] Tal Golan. Blood will out: Distinguishing humans from animals and scientists from charlatans in the 19th-Century American courtroom. Historical Studies in the Physical and Biological Sciences, 31(1):93–124, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Golinski:2007:BWC

[Gol07] Jan Golinski. British weather and the climate of enlightenment. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-30205-9 (cloth). xv + 284 pp. LCCN QC989.G69 G66

2007. URL http://www.h-net.
org/review/hrev-a0f8a2-aa;
http://www.loc.gov/
catdir/enhancements/
fy0702/2006031722-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0707/2006031722-b.html;
http://www.loc.gov/
catdir/enhancements/
fy0707/2006031722-d.html.

Gold:2010:TEV

[Gol10] Barri J. Gold. ThermoPoetics:

energy in Victorian literature and science. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01372-X (hardcover). x + 343 pp. LCCN PR468.S34 G65

Golec:2013:BRW

[Gol13] Michael J. Golec. Book review: In the way of technical images: Into the Universe of Technical Images; Hide and Seek: Camouflage, Photography, and the Media of Reconnaissance; Histories of Scientific Observation; The Interface: IBM and the Transformation of Corporate Design, 1945–1976. Historical Studies in the Natural Sciences, 43(4): 539–549, September 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2013.
43.4.539.

Goodman:1969:WAT

[Goo69] D. C. Goodman. Wollaston and the Atomic Theory of Dalton. Historical Studies in the Physical Sciences, 1(??):37–59, ???? 1969. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757294.

Gooding:1981:FSF

[Goo81] David Gooding. Final steps to the field theory: Faraday’s study of magnetic phenomena, 1845–1850. Historical Studies in the Physical Sciences, 11(2): 231–275, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757480.

Goodstein:1984:WES

[Goo84] Judith R. Goodstein. Waves in the earth: Seismology comes to Southern California. Historical Studies in the Physical Sciences, 14(2):201–230, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757533.

Good:1991:RFL

[Goo91] Gregory A. Good. The Rockefeller Foundation, the Leipzig Geophysical Institute, and National Socialism in the 1930s. Historical Studies in the Physical and Biological Sciences, 21 (2):299–316, ???? 1991. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757665.

Goodman:2003:GHB

[Goo03] Michael S. Goodman. The grandfather of the hydrogen bomb?: Anglo–American intelligence and Klaus Fuchs. Historical Studies in the Physical and Biological Sciences, 34(1): 1–22, September 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Gorman:1962:HA

[Gor62] Mel Gorman. The history of acetone, 1600–1850. Chymia: Annual Studies in the History of Chemistry, 8(??):97–104, ???? 1962. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757220.

Gordin:2002:ORM

[Gor02] Michael D. Gordin. The organic roots of Mendeleev’s periodic law. Historical Studies in the Physical and Biological Sciences, 32(2):263–290, March 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Gordin:2009:RCD

[Gor09] Michael D. Gordin. Red cloud at dawn: Truman, Stalin, and the end of the atomic monopoly. Farrar, Straus and Giroux, New York, NY, USA, 2009. ISBN 0-374-25682-9. xii + 402 pp. LCCN U264 .G67 2009.

Gordin:2014:BRT

[Gor14a]  Michael D. Gordin.  Book

review: The Tory interpretation of history: Is Water H_{2}O?: Evidence, Realism and Pluralism, Dordrecht: Springer, 2012. xxi + 316 pp., illus., index. ISBN 978-94-007-3931-4. $189.00 (hardcover). Historical Studies in the Natural Sciences, 44(4):413–423, September 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.4.413.

Gordin:2014:P

[Gor14b] Michael D. Gordin. Preface. Historical Studies in the Natural Sciences, 44(1):1–2, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2014.44.1.1.

Gordin:2017:TAK

[Gor17] Michael D. Gordin. The trials of Arnošt K.: The dark angel of dialectical materialism. Historical Studies in the Natural Sciences, 47(3):320–348, June 2017. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Gordin:2020:WNS

[Gor20] Michael D. Gordin. When national styles were stylish. Historical Studies in the Natural Sciences, 50(1–2):11–16, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Garcia-Pereda:2017:EFG

[GP17] Ignacio García-Pereda. The emergence of forest genetics in Portugal: The works of Joaquim Vieira Natividade (1899–1968) in the Alcobaça Cork Oak Station. Historical Studies in the Natural Sciences, 47(1):76–106, February 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Graves:2025:STN

[Gra25] Joseph L. Graves, Jr. Sociobiology then and now: a biologist’s perspective. Historical Studies in the Natural Sciences, 55(1): 82–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/55/1/82/
207536/Sociobiology-Then-
and-NowA-Biologist-s-
Perspective.

Greenberg:1983:GPP

[Gre83] John Greenberg. Geodesy in Paris in the 1730s and the Paduan Connection. Historical Studies in the Physical Sciences, 13(2):239–260, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757516.

Greene:2007:PND

[Gre07] Jeremy A. Greene. Prescribing by numbers: drugs and the definition of disease. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8477-2 (hardcover). xv + 318 pp. LCCN RM262.G74 2007.

Greenhalgh:2008:JOC

[Gre08] Susan Greenhalgh. Just one child: science and policy in Deng’s China. University of California Press, Berkeley, CA, USA, 2008. ISBN 0-520-25338-8 (hardcover), 0-520-25339-6 (paperback). xxii + 403 + 6 pp. LCCN HQ766.5.C6 G74

2008. URL http://www.loc.
gov/catdir/toc/ecip0727/
2007038377.html.

Groppi:2003:SBa

[Gro03a] Susan M. Groppi. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 33(2):405–413, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Groppi:2003:SBb

[Gro03b] Susan M. Groppi. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 34(1):199–206, September 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Groppi:2004:SBa

[Gro04a] Susan M. Groppi. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 34(2):399–406, March 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Groppi:2004:SBb

[Gro04b] Susan M. Groppi. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 35(1):161–174, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Groppi:2005:SBa

[Gro05a] Susan M. Groppi. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 35(2):371–382, March 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Groppi:2005:SBb

[Gro05b] Susan M. Groppi. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 36(1):197–205, September 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Grodwohl:2017:MSE

[Gro17] Jean-Baptiste Grodwohl. Modeling social evolution, 1964– 1973: Inclusive fitness meets population structure. Historical Studies in the Natural Sciences, 47(1):1–41, February 2017. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Grote:2023:GT

[Gro23] Mathias Grote. Gaia’s tissue. Historical Studies in the Natural Sciences, 53(4):434–??, September 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/53/4/434/197389/
Gaia-s-Tissue.

Gruber:1995:OSG

[Gru95] Carol Gruber. The overhead system in government-sponsored academic science: Origins and early development. Historical Studies in the Physical and Biological Sciences, 25(2):241–268, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Gavroglu:1994:AGB

[GS94] Kostas Gavroglu and Ana Simoes. The Americans, the Germans and the beginnings of quantum chemistry. Historical Studies in the Physical and Biological Sciences, 25(1):47–110, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Gaspar:2011:PPR

[GS11] Júlia Gaspar and Ana Simões.

Physics on the periphery: a research school at the University of Lisbon under Salazar’s dictatorship. Historical Studies in the Natural Sciences, 41(3): 303–343, Summer 2011. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.3.303.

Garcia-Sancho:2022:HGP

[GSLV^{+}22a] Miguel García-Sancho, Rhodri Leng, Gil Viry, Mark Wong, Niki Vermeulen, and James Lowe. The Human Genome Project as a singular episode in the history of genomics. Historical Studies in the Natural Sciences, 52(3):320–??, June 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/3/320/185531/The-Human-
Genome-Project-as-a-
Singular-Episode-in.

Garcia-Sancho:2022:YSN

[GSLV^{+}22b] Miguel García-Sancho, James Lowe, Gil Viry, Rhodri Leng, Mark Wong, and Niki Vermeulen. Yeast sequencing: “Network” genomics and institutional bridges. Historical Studies in the Natural Sciences, 52(3):361–??, June 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        52/3/361/185528/Yeast-
                                        Sequencing-Network-
                                        Genomics-and.

Guerlac:1959:SFA

[Gue59] Henry Guerlac. Some French antecedents of the chemical revolution. Chymia: Annual Studies in the History of Chemistry, 5(??):73–112, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757178.

Guerlac:1961:CLE

[Gue61] Henry Guerlac. A curious Lavoisier episode. Chymia: Annual Studies in the History of Chemistry, 7(??):103–108, ???? 1961. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757207.

Guerlac:1976:CBP

[Gue76] Henry Guerlac. Chemistry as a branch of physics: Laplace’s collaboration with Lavoisier. Historical Studies in the Physical Sciences, 7(??):193–276, ???? 1976. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757357.

Guerrini:2017:I

[Gue17] Anita Guerrini. Introduction.

Historical Studies in the Natural Sciences, 47(3):263–267, June 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Guenther:2022:HTY

[Gue22] Katja Guenther. How to train your analyst. Historical Studies in the Natural Sciences, 52 (1):123–??, February 2022. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/1/123/120498/How-to-
Train-Your-Analyst.

GomezGutierrez:2011:AMD

[GUOB11] Alberto Gómez Gutiérrez, Jorge Tomás Uribe Angel, Pedro Ortiz Valdivieso, and J. E. (Jaime E.) Bernal Villegas. Academia Mutisiana: documentos preneogranadinos de José Celestino Mutis y la promoción de sociedades científicas en la Nueva Granada. Pontificia Universidad Javieriana, Academia Nacional de Medicina, Bogotá, Colombia, 2011. ISBN 958-716-490-3. 325 pp. LCCN QK31.M8 G655 2011.

Guttler:2024:AZB

[Güt24] Nils Güttler. The airport and the zoo: Behavioral knowledge, animal transport, and the technical environments of aviation, 1940s–1970s. Historical Studies in the Natural Sciences, 54(2): 135–??, April 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/54/2/135/
200664/The-Airport-and-
the-ZooBehavioral-
Knowledge-Animal.

Hartcup:1984:CA

[HA84] Guy Hartcup and T. E. Allibone. Cockcroft and the atom. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-759-4. xii + 320 pp. LCCN QC774.C6 H37 1984. US$18.95.

Ha:2015:DSC

[Ha15] Nathan Q. Ha. Diagnosing sex chromatin: A binary for every cell. Historical Studies in the Natural Sciences, 45(1):49–84, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2015.45.1.49.

Hackmann:1986:SRN

[Hac86] Willem D. Hackmann. Sonar research and naval warfare 1914– 1954: a case study of a Twentieth-Century establishment science. Historical Studies in the Physical and Biological Sciences, 16 (1):83–110, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757558.

Hacker:2000:MPG

[Hac00] Barton Hacker. Military patronage and the geophysical science in the United States: An introduction. Historical Studies in the Physical and Biological Sciences, 30(2):309–??, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hagen:2014:BRE

[Hag14]  Joel B. Hagen.  Book re-

view: Emerging ecologies: Past, present, and future: Shaping Ecology: The Life of Arthur Tansley, by Peter Ayres; Big Ecology: The Emergence of Ecosystem Science by David C. Coleman; Leopold’s Shack and Ricketts’s Lab: The Emergence of Environmentalism by Michael J. Lannoo. Historical Studies in the Natural Sciences, 44(3):285–295, June 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.3.285.

Hall:1962:HC

[Hal62] Marie Boas Hall. Humanism in chemistry. Chymia: Annual Studies in the History of Chemistry, 8(??):33–39, ???? 1962. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757216.

Hallonsten:2011:GBS

[Hal11] Olof Hallonsten. Growing big science in a small country: MAX-lab and the Swedish Research Policy System. Historical Studies in the Natural Sciences, 41(2):179–215, Spring 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2011.41.2.179.

Hallonsten:2015:PSR

[Hal15] Olof Hallonsten. The parasites: Synchrotron radiation at SLAC, 1972–1992. Historical Studies in the Natural Sciences, 45(2):217– 272, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

                                        //www.jstor.org/stable/10.
                                        1525/hsns.2015.45.2.217.

Hamblin:2008:PWR

[Ham08] Jacob Darwin Hamblin. Poison in the well: radioactive waste in the oceans at the dawn of the nuclear age. Rutgers University Press, New Brunswick, NJ, USA, 2008. ISBN 0-8135-4220-0 (hardcover). x + 311 pp. LCCN TD898 .H35 2008. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        ecip0715/2007015498.html.

Hamblin:2010:BRS

[Ham10] Jacob Darwin Hamblin. Book review: Science and Technology for Every Man, Woman, and Child: Zuoyue Wang. In Sputnik’s Shadow: The President’s Science Advisory Committee and Cold War America. Piscataway, NJ: Rutgers University Press, 2008. xiii + 454 pp., illus., index. ISBN 978-0-813-54331-4. $49.95 (cloth). Lillian Hoddeson, Adrienne W. Kolb, and Catherine Westfall. Fermilab: Physics, the Frontier, and Megascience. Chicago: University of Chicago Press, 2008. xi + 497 pp., illus., index. ISBN 978-0-226-34623-6. $45 (cloth). W. Patrick McCray. Keep Watching the Skies! The Story of Operation Moonwatch and the Dawn of the Space Age. Princeton: Princeton University Press, 2008. xiii + 308 pp., illus., index. ISBN 978-0-691-12854-2. $29.95 (cloth). Michael D. Gordin. Red Cloud at Dawn: Truman, Stalin, and the End of the Atomic Monopoly. New York: Farrar, Straus and Giroux, 2009. 416 pp., illus., index. ISBN 978-0-374-25682-1. $28 (cloth). Dolores L. Augustine. Red Prometheus: Engineering and Dictatorship in East Germany, 1945–1990. Cambridge, MA: MIT Press, 2007. xxvii + 381 pp., illus., index. ISBN 978-0-262-01236-2. $43 (cloth). Historical Studies in the Natural Sciences, 40(2): 259–267, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Hamblin:2013:AMN

[Ham13] Jacob Darwin Hamblin. Arming Mother Nature: the birth of catastrophic environmentalism.

Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-974005-4. x + 298 pp. LCCN GE180 .H35 2013.

Hamza:2024:TGH

[Ham24] Shireen Hamza. Teaching the global history of science in a prison classroom. Historical Studies in the Natural Sciences, 54(1):102–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/1/102/200196/Teaching-
                                        the-Global-History-of-
                                        Science-in-a-Prison.

Hanle:1979:IBH

[Han79] Paul A. Hanle. Indeterminacy before Heisenberg: The case of Franz Exner and Erwin Schrödinger. Historical Studies in the Physical Sciences, 10 (??):225–269, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757391.

Hanson:2012:KSE

[Han12] Marta Hanson. Kuhn’s structure in East Asia. Historical Studies in the Natural Sciences, 42(5):514–520, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.514.pdf.

Hanson:2015:BRW

[Han15] Craig Ashley Hanson. Book reviews: Wicked Intelligence: Visual Art and the Science of Experiment in Restoration London; Catching Nature in the Act: Réaumur and the Practice of Natural History in the Eighteenth Century; The Triumph of Human Empire: Verne, Morris, and Stevenson at the End of the World. Historical Studies in the Natural Sciences, 45(3):493– 500, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2015.45.3.493.

Harwood:2005:TDA

[Har05] Jonathan Harwood. Technology’s dilemma: agricultural colleges between science and practice in Germany, 1860–1934. P. Lang, Oxford, UK, 2005. ISBN 3-03910-299-0 (paperback), 0-8204-6987-4 (US: paperback). 288 pp. LCCN S535.G3 H37 2005.

Haring:2006:HRT

[Har06] Kristen Haring. Ham radio’s technical culture. Inside technology. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-08355-8 (hardcover). xvii + 218 pp. LCCN TK9956 .H3626 2007.

Harkness:2007:JHE

[Har07a] Deborah E. Harkness. The Jewel house: Elizabethan London and the scientific revolution. Yale University Press, New Haven, CT, USA, 2007. ISBN 0-300-11196-7. xviii + 349 pp. LCCN Q127.G4 H37 2007.

URL http://www.h-net.org/
review/hrev-a0f5v3-aa;
http://www.loc.gov/catdir/
enhancements/fy0833/
2007002683-b.html;
http://www.loc.gov/catdir/
enhancements/fy0833/
2007002683-d.html;   http:
//www.loc.gov/catdir/toc/
ecip079/2007002683.html.

Harrison:2007:FMF

[Har07b] Peter Harrison. The fall of man and the foundations of science. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87559-5. xi + 300 pp. LCCN BL245 .H355

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2008272575-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0827/2008272575-t.html.

Harper:2008:WNG

[Har08] Kristine Harper. Weather by the numbers: the genesis of modern meteorology. Transformations. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-08378-7. ix + 308 pp. LCCN QC855 .H374 2008. URL http:

//www.loc.gov/catdir/toc/
ecip081/2007039860.html.

Harrison:2009:BRA

[Har09] Mark Harrison. Book reviews: After Empire: Searching for A New Synthesis: Michael Adas. Dominance by Design: Technological Imperatives and America’s Civilizing Mission. Cambridge, MA: Harvard Belknap Press, 2006. 542 pp., illus., index. ISBN 978-0-674-03216-3. $29.95 (cloth). Anna Crozier. Practising Colonial Medicine: The Colonial Medical Service in British East Africa. London: I. B. Tauris, 2007. xiv + 225 pp., illus., index. ISBN 978-1-845-11459-6. $85.00 (hardback). Rod Edmond. Leprosy and Empire: A Medical and Cultural History. Cambridge: Cambridge University Press, 2006. x + 255 pp., illus., index. ISBN 0-521-86584-0. $90.00 (hardback). Eric T. Jennings. Curing the Colonizers: Hydrotherapy, Climatology, and French Colonial Spas. Durham, NC: Duke University Press, 2006. xi + 271 pp., illus., index. ISBN 978-0-822-33822-2. $74.95 (cloth), $21.95 (paper). Richard C. Keller. Colonial Madness: Psychiatry in French North Africa. Chicago: University of Chicago Press, 2007. xi + 294 pp., illus., index. ISBN 978-0-226-42972-4. $70.00 (cloth), $25.00 (paper). Historical Studies in the Natural Sciences, 39(2):258–268, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Harwood:2010:FPF

[Har10] Jonathan Harwood. The fate of peasant-friendly plant breeding in Nazi Germany. Historical Studies in the Natural Sciences, 40(4):569–603, Fall 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2010.
                                        40.4.569.

Harwood:2011:IIT

[Har11] John Harwood. The interface: IBM and the transformation of corporate design, 1945/ 1976. A Quadrant book. University of Minnesota Press, Minneapolis, MN, USA, 2011. ISBN 0-8166-7849-9, 0-8166-7039-0. 278 + 8 pp. LCCN HD9696.2.U64. URL http://

                                        site.ebrary.com/lib/
                                        alltitles/Doc?id=10519770.

Hardy:2024:TIB

[Har24] Penelope K. Hardy. Thinking inside the box. Historical Studies in the Natural Sciences, 54(1):105–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/1/105/200204/Thinking-
                                        Inside-the-Box.

Home:2006:I

[HdAG06] R. W. Home, Ana M. Ribeiro de Andrade, and Carlos D. Galles. Introduction. Historical Studies in the Physical and Biological Sciences, 36(2): 209–211, March 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hecht:2011:CSE

[Hec11] David K. Hecht. Constructing a scientist: Expert authority and public images of Rachel Carson. Historical Studies in the Natural Sciences, 41(3):277–302, Summer 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.3.277.

Hecht:2012:BNA

[Hec12] Gabrielle Hecht. Being nuclear:

Africans and the global uranium trade. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01726-1 (hardcover). xx + 451 pp. LCCN HD9539.U72 A43 2012.

Hecht:2014:BRP

[Hec14] David K. Hecht. Book review: The perpetual quest: Science and the search for authority: Lab Coats in Hollywood: Science, Scientists, and Cinema, by David A. Kirby; Cold War University: Madison and the New Left in the Sixties, by Matthew Levin; Engineers for Change: Competing Visions of Technology in 1960s America, by Matthew Wisnioski; Competing with the Soviets: Science, Technology, and the State in Cold War America, by Audra J. Wolfe. Historical Studies in the Natural Sciences, 44(3):277– 284, June 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

                                        //www.jstor.org/stable/10.
                                        1525/hsns.2014.44.3.277.

Hedeen:2008:BBL

[Hed08] Stanley Hedeen. Big Bone Lick: the cradle of American paleontology. University Press of Kentucky, Lexington, KY, USA, 2008. ISBN 0-8131-2485-9 (hardcover). xviii + 182 pp. LCCN QE705.U6 H43

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0812/2007040474-b.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0812/2007040474-d.html;
                                         http://www.loc.gov/
                                        catdir/toc/fy0803/
                                        2007040474.html.

Heimann:1976:MCF

[Hei76] P. M. Heimann. Mayer’s concept of “force”: The “axis” of a new science of physics. Historical Studies in the Physical Sciences, 7(??):277–296, ???? 1976. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757358.

Heilbron:1983:OEP

[Hei83] J. L. Heilbron. The origins of the exclusion principle. Historical Studies in the Physical Sciences, 13(2):261–310, ???? 1983. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757517.

Heilbron:1991:CBG

[Hei91] J. L. Heilbron. The contributions of Bologna to Galvanism. Historical Studies in the Physical and Biological Sciences, 22 (1):57–85, ???? 1991. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757673.

Heilbron:1993:WIO

[Hei93] J. L. Heilbron. Weighing imponderables and other quantitative science around 1800. Historical Studies in the Physical and Biological Sciences, 24(1 (Supplement)):1–337, ???? 1993. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Heilbron:2007:S

[Hei07] J. L. Heilbron. Swansong. Historical Studies in the Physical and Biological Sciences, 37(2): 185–187, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Helmreich:2012:HKW

[Hel12] Stefan Helmreich. The house of Kuhn, by the water. Historical Studies in the Natural Sciences, 42(5):521–526, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.521.pdf.

Hendry:1984:CPT

[Hen84] John Hendry, editor. Cambridge physics in the thirties. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-761-6. 209 pp. LCCN QC51.G72 C352 1984.

Hentschel:1993:DRS

[Hen93] Klaus Hentschel. The discovery of the redshift of solar Fraunhofer lines by Rowland and Jewell in Baltimore around 1890. Historical Studies in the Physical and Biological Sciences, 23 (2):219–277, ???? 1993. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757699.

Henchman:2014:SSW

[Hen14] Anna Henchman. The starry sky within: astronomy and the reach of the mind in Victorian literature. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-968696-3. xvi + 294 pp. LCCN PR468.A79 H46

                                        2014. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1410/2013942193-b.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1410/2013942193-d.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1410/2013942193-t.html.

Henderson:2016:GHC

[Hen16] Gabriel Henderson. Governing the hazards of climate: The development of the National Climate Program Act, 1977–1981. Historical Studies in the Natural Sciences, 46(2):207–242, April 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Henderson:2019:ASF

[Hen19] Gabriel Henderson. Adhering to the “Flashing Yellow Light”: Heuristics of moderation and carbon dioxide politics during the 1970s. Historical Studies in the Natural Sciences, 49(4):384– 419, September 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Herran:2008:ASR

[Her08] Nestor Herran. Aguas, semillas y radiaciones: el laboratorio de radiactividad de la universidad de Madrid 1904–1929. (Spanish) [Water, seeds and radiation: the University of Madrid Radioactivity Laboratory 1904– 1929], volume 52 of Estudios sobre la ciencia. Consejo Superior de Investigaciones Cinetificas, Madrid, Spain, 2008. ISBN 84-00-08692-9. 271 pp. LCCN MLC.

Hersch:2012:IAA

[Her12] Matthew H. Hersch. Inventing the American astronaut. Palgrave studies in the history of science and technology. Palgrave Macmillan, New York, NY, USA, 2012. ISBN 1-137-02527-1 (hardcover), 1-137-02528-X. xiii + 219 pp. LCCN TL789.8.U5 H53

                                        2012. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1311/2012275134-b.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1311/2012275134-d.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1311/2012275134-t.html.

Hersch:2024:LTB

[Her24] J. D. PhD Matthew H. Hersch. “Learning on their bellies”. Historical Studies in the Natural Sciences, 54(1):109–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/54/1/109/200198/
                                        Learning-on-Their-Bellies.

Hetherington:1982:PVO

[Het82] Norriss S. Hetherington. Philosophical values and observation in Edwin Hubble’s choice of a model of the universe. Historical Studies in the Physical Sciences, 13(1):41–67, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757505.

Hetherington:1987:RTH

[Het87] Norriss S. Hetherington. Review: Toward the history of X-ray astronomy. Historical Studies in the Physical and Biological Sciences, 17(2):357– 360, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757587.

Hetherington:1989:REL

[Het89] Norriss Hetherington. Review:

The extraterrestrial life debate: a productive perspective. Historical Studies in the Physical and Biological Sciences, 20(1): 179–182, ???? 1989. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757638.

Heymann:2006:MRP

[Hey06] Matthias Heymann. Modeling reality. Practice, knowledge, and uncertainty in atmospheric transport simulation. Historical Studies in the Physical and Biological Sciences, 37(1):49– 85, September 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Heimann:1971:NFL

[HG71] P. M. Heimann and J. E. Mc Guire. Newtonian forces and Lockean powers: Concepts of matter in Eighteenth-Century thought. Historical Studies in the Physical Sciences, 3(??): 233–306, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757320.

Heinze:2015:PCSa

[HHH15a] Thomas Heinze, Olof Hallonsten, and Steffi Heinecke. From periphery to center: Synchrotron radiation at DESY, Part I: 1962–1977. Historical Studies in the Natural Sciences, 45(3):447–492, June 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.3.447.

Heinze:2015:PCSb

[HHH15b] Thomas Heinze, Olof Hallonsten, and Steffi Heinecke. From periphery to center: Synchrotron radiation at DESY, Part II: 1977–1993. Historical Studies in the Natural Sciences, 45(4):513–548, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.4.513.

Hillpern:1949:SPQ

[Hil49] Edmund P. Hillpern. Some personal qualities of Wilhelm Ostwald Recalled by a former assistant. Chymia: Annual Studies in the History of Chemistry, 2(??):57–64, ???? 1949. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757134.

Hinokawa:2003:CSC

[Hin03] Shizue Hinokawa. A comparative study of cyclotron development at Cambridge and Liverpool in the 1930s. Historical Studies in the Physical and Biological Sciences, 34(1):

23–39, September 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hirsch:1950:IPD

[Hir50] Rudolf Hirsch. The invention of printing and the diffusion of alchemical and chemical knowledge. Chymia: Annual Studies in the History of Chemistry, 3 (??):115–141, ???? 1950. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757149.

Hirosige:1969:OLT

[Hir69]  Tetu Hirosige.  Origins of

Lorentz’ Theory of Electrons and the concept of the electromagnetic field. Historical Studies in the Physical Sciences, 1 (??):151–209, ???? 1969. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757298.

Hirosige:1976:EPM

[Hir76] Tetu Hirosige. The ether problem, the mechanistic worldview, and the origins of the Theory of Relativity. Historical Studies in the Physical Sciences, 7 (??):3–82, ???? 1976. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757354.

Heilbron:1969:GBA

[HK69] John L. Heilbron and Thomas S.

Kuhn. The genesis of the Bohr atom. Historical Studies in the Physical Sciences, 1(??):vi, 211–290, ???? 1969. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757291.

Hoddeson:2008:FPF

[HKW08] Lillian Hoddeson, Adrienne W. Kolb, and Catherine Westfall. Fermilab: physics, the frontier, and megascience. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-34623-4 (cloth). xi + 497 pp. LCCN QC789.2.U62 F474

2008. URL http://www.loc.
gov/catdir/toc/ecip0810/
2008006254.html.

Hoddeson:1981:DPC

[Hod81] Lillian Hoddeson. The discovery of the point-contact transistor. Historical Studies in the Physical Sciences, 12(1): 41–76, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757489.

Hoddeson:1987:FLS

[Hod87] Lillian Hoddeson. The first large-scale application of superconductivity: The Fermilab Energy Doubler, 1972–1983. Historical Studies in the Physical and Biological Sciences, 18(1): 25–54, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757595.

Hoddeson:1993:DSF

[Hod93] Lillian Hoddeson. The discovery of spontaneous fission in plutonium during World War II. Historical Studies in the Physical and Biological Sciences, 23 (2):279–300, ???? 1993. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757700.

Hoffman:2017:JTA

[Hof17] Michelle D. Hoffman. Just a theory: The atomic theory debate and high school chemistry, 1905–1917. Historical Studies in the Natural Sciences, 47(4): 494–528, September 2017. ISSN 1939-1811 (print), 1939-182X (electronic).

Hoffmann:2024:WSD

[Hof24] Christoph Hoffmann. When scientists disagree: Carl von Heß, Karl von Frisch, and the study of controversies. Historical Studies in the Natural Sciences, 54(5):535–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/5/535/203884/When-
Scientists-DisagreeCarl-
von-He-Karl-von.

Hogan:2015:DGD

[Hog15] Andrew J. Hogan. Disrupting genetic dogma: Bridging cytogenetics and molecular biology in fragile X research. Historical Studies in the Natural Sciences, 45(1):174–197, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2015.
                                        45.1.174.

Hogan:2019:EPH

[Hog19] Andrew J. Hogan. Expanding perspectives on heredity. Historical Studies in the Natural Sciences, 49(1):104–115, February 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Holmes:1962:EAC

[Hol62] Frederic L. Holmes. From elective affinities to chemical equilibria: Berthollet’s Law of Mass Action. Chymia: Annual Studies in the History of Chemistry, 8(??):105–145, ???? 1962. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757221.

Holton:1978:SPM

[Hol78] Gerald Holton. Subelectrons, presuppositions, and the Millikan– Ehrenhaft dispute. Historical Studies in the Physical Sciences, 9(??):161–224, ???? 1978. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757378.

Holmes:1987:IOM

[Hol87] Frederic L. Holmes. The intakeoutput method of quantification in physiology. Historical Studies in the Physical and Biological Sciences, 17(2):235– 270, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757583.

Holloway:1999:PSC

[Hol99] David Holloway. Physics, the state, and civil society in the Soviet Union. Historical Studies in the Physical and Biological Sciences, 30(1):173–193, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Holmes:2003:CAR

[Hol03]  Frederic  Lawrence

Holmes. Chemistry in the Académie Royale des Sciences. Historical Studies in the Physical and Biological Sciences, 34 (1):41–68, September 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Holmes:2004:IPS

[Hol04] Fredric Lawrence Holmes. Investigative and pedagogical styles in French chemistry at the end of the 17th century. Historical Studies in the Physical and Biological Sciences, 34 (2):277–309, March 2004. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Holmes:2018:SHR

[Hol18]  Matthew Holmes.  Somatic

hybridization: The rise and fall of a mid-Twentieth-Century biotechnology. Historical Studies in the Natural Sciences, 48 (1):1–23, February 2018. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Home:2006:RAE

[Hom06] R. W. Home. The rush to accelerate: Early stages of nuclear physics research in Australia. Historical Studies in the Physical and Biological Sciences, 36 (2):213–241, March 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hon:1989:FHV

[Hon89] Giora Hon. Franck and Hertz versus Townsend: a study of two types of experimental error. Historical Studies in the Physical and Biological Sciences, 20 (1):79–106, ???? 1989. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757636.

Hooykaas:1949:EOC

[Hoo49] R. Hooykaas. The experimental origin of chemical atomic and molecular theory before Boyle. Chymia: Annual Studies in the History of Chemistry, 2(??):65– 80, ???? 1949. CODEN CHYMAP. ISSN 0095-9367.

                                        URL http://www.jstor.org/
                                        stable/27757135.

Hopper:1996:ERR

[Hop96] Teresa Hopper. Essay review:

Recent books on Nazism and science: Uncertainty: The Life and Science of Werner Heisenberg by David Cassidy; The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science by John Heilbron; The Nazi Doctors: Medical Killing and the Psychology of Genocide by Robert Jay Lifton; Surviving the Swastika: Scientific Research in Nazi Germany by Kristie Macrakis; The Rocket and the Reich: Pennemunde and the Coming of the Ballistic Era by Michael Neufeld; Operation Epsilon: The Farm Hall Transcripts; Heisenberg’s War: The Secret History of the German Bomb by Thomas Powers; Racial Hygiene: Medicine under the Nazis by Robert N. Proctor; Science, Technology and National Socialism by Monika Renneberg and Mark Walker; German National Socialism and the Quest for Nuclear Power, 1939–1949 by Mark Walker; Health, Race and German Politics between National Unification and Nazism, 1870–1945 by Paul Weindling. Historical Studies in the Physical and Biological Sciences, 27(1):163– 176, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757772.

Horne:1966:AC

[Hor66] R. A. Horne. Aristotelian chemistry. Chymia: Annual Studies in the History of Chemistry, 11(??):21–27, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757256.

Hoskin:2007:HH

[Hos07] Michael A. Hoskin. The Herschels of Hanover. Science History Publications, Cambridge, UK, 2007. ISBN 0-905193-07-5 (hardcover). x + 182 pp. LCCN QB35 .H67 2007.

Hounshell:1997:CWR

[Hou97] David Hounshell. The Cold War, RAND, and the generation of knowledge, 1946–1962. Historical Studies in the Physical and Biological Sciences, 27 (2):237–267, ???? 1997. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Howell:2011:MHY

[How11] David L. Howell. May be hazardous to your health. Historical Studies in the Natural Sciences, 41(1):104–111, Winter 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2011.41.1.104.

Howe:2012:BRS

[How12] Joshua P. Howe. Book review: The stories we tell: Merchants of Doubt: How a Handful of Scientists Obscured the Truth on Issues from Tobacco Smoke to Global Warming, by Naomi Oreskes and Erik M. Conway. Historical Studies in the Natural Sciences, 42(3):244– 254, June 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2012.42.3.244.

Howard:2024:ASW

[How24] Nicole Howard. AI, science, and what we can learn from the printing press. Historical Studies in the Natural Sciences, 54(5):617–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/5/617/203880/AI-
Science-and-What-We-Can-
Learn-from-the-Printing.

Herran:2012:FDF

[HR12]  Néstor  Herrán

and Xavier Roqué. La física en la dictadura: físicos, cultura y poder en España, 1939– 1975. (Spanish) [Physics under the Dictatorship: physicists, culture, and power in Spain], volume 10 of Congressos. Universitat Autónoma de Barcelona, Servei de Publicacions, Bellaterra, España, 2012. ISBN 84-490-3238-5. 310 pp. LCCN QC9.S7 F58

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1503/2013408785-b.html;

                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1503/2013408785-d.html.

Howard:1989:EHG

[HS89] Don Howard and John J. Stachel, editors. Einstein and the History of General Relativity: Based on the Proceedings of the 1986 Osgood Hill Conference, North Andover, Massachusetts, 8–11 May 1986, volume 1 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1989. ISBN 0-8176-3392-8. LCCN QC173.6 .E38 1989.

Holmes:2006:RGS

[HS06] Frederic Lawrence Holmes and William C. Summers. Reconceiving the gene: Seymour Benzer’s adventures in phage genetics. Yale University Press, New Haven, CT, USA, 2006. ISBN 0-300-11078-2 (hardcover). ix + 334 pp. LCCN QH429.2.B46 H65 2006.

Hepler-Smith:2017:GTI

[HS17] Evan Hepler-Smith. Grasping the technical image. Historical Studies in the Natural Sciences, 47(1):117–125, February 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Hepler-Smith:2018:WTB

[HS18] Evan Hepler-Smith. A way of thinking backwards: Computing and method in synthetic organic chemistry. Historical Studies in the Natural Sciences, 48(3):300–337, June 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Hu:2004:OCE

[Hu04] Danian Hu. Organized criticism of Einstein and Relativity in China, 1949–1989. Historical Studies in the Physical and Biological Sciences, 34(2): 311–338, March 2004. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hufbauer:1971:SSC

[Huf71] Karl Hufbauer. Social support for chemistry in Germany during the Eighteenth Century: How and why did it change? Historical Studies in the Physical Sciences, 3(??): 205–231, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757319.

Hufbauer:1981:ATS

[Huf81] Karl Hufbauer. Astronomers take up the stellar-energy problem, 1917–1920. Historical Studies in the Physical Sciences, 11(2):277–303, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757481.

Hufbauer:1994:AEB

[Huf94] Karl Hufbauer. Artificial eclipses:

Bernard Lyot and the coronagraph, 1929–1939. Historical Studies in the Physical and Biological Sciences, 24(2):337–394, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hufbauer:2007:LYS

[Huf07] Karl Hufbauer. Landau’s youthful sallies into stellar theory: Their origins, claims, and receptions. Historical Studies in the Physical and Biological Sciences, 37(2):337–354, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hughes:2011:GBB

[Hug11] Sally Smith Hughes. Genentech: the beginnings of biotech. Synthesis. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-35918-2 (hardcover). xv + 213 pp. LCCN HD9999.B444 H85 2011.

Hui:2022:HCD

[Hui22] Alexandra Hui. How to call a duck. Historical Studies in the Natural Sciences, 52(1):128–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/52/1/128/120499/
                                        How-to-Call-a-Duck.

Hunt:1986:EEO

[Hun86] Bruce Hunt. Experimenting on the ether: Oliver J. Lodge and the Great Whirling Machine. Historical Studies in the Physical and Biological Sciences, 16 (1):111–134, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757559.

Hunt:2010:PPL

[Hun10]  Bruce J. Hunt.  Pursuing

power and light: technology and physics from James Watt to Albert Einstein. Johns Hopkins introductory studies in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9358-5 (hardcover), 0-8018-9359-3 (paperback). 182 pp. LCCN T173.8 .H92 2010.

Hunter:2013:WIV

[Hun13] Matthew C. Hunter. Wicked intelligence: visual art and the science of experiment in Restoration London. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-01729-X (hardcover), 0-226-01732-X (e-book). xv + 329 + 10 pp. LCCN Q222 .H86 2013.

Hunt:2015:MMM

[Hun15] Bruce J. Hunt. Maxwell, measurement, and the modes of electromagnetic theory. Historical Studies in the Natural Sciences, 45(2):303–339, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2015.45.2.303.

HurtadoDeMendoza:2006:PSF

[HV06] Diego Hurtado De Mendoza and Ana Maria Vara. Political storms, financial uncertainties, and dreams of “big science:” The construction of a heavy ion accelerator in Argentina. Historical Studies in the Physical and Biological Sciences, 36(2): 343–364, March 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Hyun:2019:TNO

[Hyu19] Jaehwan Hyun. Tracing national origins, debating ethnic homogeneity: Population genetics and the politics of national identity in South Korea. Historical Studies in the Natural Sciences, 49(4):351–383, September 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Ianniello:1993:ENB

[Ian93] Maria Grazia Ianniello. Elastic Nachwirkung, Brownian motion and the tide against determinism: 1835–1920. Historical Studies in the Physical and Biological Sciences, 24(1):41–100, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Igo:2018:MMD

[Igo18] Sarah E. Igo. Me and my data. Historical Studies in the Natural Sciences, 48(5):616– 626, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Ihde:1964:ARR

[Ihd64] Aaron J. Ihde. Alchemy in reverse: Robert Boyle on the degradation of gold. Chymia: Annual Studies in the History of Chemistry, 9(??):47–57, ???? 1964. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757231.

Im:1995:FDR

[Im95] Gyeong Soon Im. The formation and development of the Ramsauer effect. Historical Studies in the Physical and Biological Sciences, 25(2):269–300, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

IrinaPodgorny:2023:AUM

[Iri23] Susana V. García Irina Podgorny . Armadillos under the microscope: The end of natural history and the emergence of bio-materials research. Historical Studies in the Natural Sciences, 53(3):332–??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/3/332/196656/
Armadillos-under-the-
MicroscopeThe-End-of-
Natural.

Isaac:2012:WKM

[Isa12] Joel Isaac. Working knowledge: making the human sciences from Parsons to Kuhn. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06574-3 (hardcover), 0-674-06522-0 (ebook). vi + 314 pp. LCCN H62 .I77 2012.

Ito:2002:VPS

[Ito02] Kenji Ito. Values of ‘pure science’: Nishina Yoshio’s wartime discourse between nationalism and physics, 1940–1945. Historical Studies in the Physical and Biological Sciences, 33(1): 61–86, September 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Ito:2024:PPN

[Ito24] Takashi Ito. The power and performativity of naming: a natural and cultural history of the mikado pheasant in early twentieth-century Taiwan and beyond. Historical Studies in the Natural Sciences, 54 (2):157–??, April 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/2/157/200670/The-Power-
                                        and-Performativity-of-
                                        NamingA-Natural.

J:2023:DD

[J.23] Amrah Salomón J. Drawing on the difuentes. Historical Studies in the Natural Sciences, 53(1):91–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/1/91/195528/Drawing-on-
the-Difuentes.

Jackson:2006:HTP

[Jac06] Myles W. Jackson. Harmonious triads: physicists, musicians, and instrument makers in Nineteenth-Century Germany. Transformations. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-10116-5. x + 395 pp. LCCN ML3805 .J33

2006. URL http://www.loc.
gov/catdir/toc/fy0710/
2006044991.html.

Jacobsen:2007:LRM

[Jac07] Anja Skaar Jacobsen. Léon Rosenfeld’s Marxist defense of complementarity. Historical Studies in the Physical and Biological Sciences, 37(S1):3– 34, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.3.pdf.

Jackson:2009:BRM

[Jac09] Myles W. Jackson. Book review:

Molecular Genetics: Redefining the Relationship between Science and Society: M. Susan Lindee. Moments of Truth in Genetic Medicine. Baltimore: Johns Hopkins University Press, 2005. xii + 270 pp., illus. ISBN 978-0-801-89101-4. $25.00 (paperback). Shobita Parthasarathy. Building Genetic Medicine: Breast Cancer, Technology, and the Comparative Politics of Health Care. Cambridge, MA: MIT Press, 2007. x + 271 pp., illus. ISBN 978-0-262-16242-5. $35.00 (cloth). Michael A. Fortun. Promising Genomics: Iceland and deCODE Genetics in a World of Speculation. Berkeley: University of California Press, 2008. x + 330 pp., illus. ISBN 978-0-520-24751-2. $24.95 (paperback). Historical Studies in the Natural Sciences, 39(3):367– 376, August 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Jacobsen:2012:LRP

[Jac12] Anja Skaar Jacobsen. Léon Rosenfeld: physics, philosophy, and politics in the twentieth century. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2012. ISBN 981-4307-81-5 (hardcover), 981-4307-82-3 (e-book), 1-280-66920-9. xii + 354 pp. LCCN QC16.R493 J326 2012. URL http://www.

worldscientific.com/
worldscibooks/10.1142/
7776.

Jackson:2014:SEA

[Jac14] Catherine M. Jackson. Synthetical experiments and alkaloid analogues: Liebig, Hofmann, and the origins of organic synthesis. Historical Studies in the Natural Sciences, 44(4): 319–363, September 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.4.319.

Jain:1966:CTJ

[Jai66] N. L. Jain. Chemical theories of the Jains. Chymia: Annual Studies in the History of Chemistry, 11(??):11–19, ???? 1966. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757255.

Jansen:2011:ADB

[Jan11] Axel Jansen. Alexander Dallas Bache: building the American nation through science and education in the nineteenth century. Campus Verlag, Frankfurt, Germany, 2011. ISBN 3-593-39355-7. 352 pp. LCCN Q143.B2 J36 2011; Q143.B11 J36 2011.

Jensen:1984:TOE

[Jen84] Carsten Jensen. Two oneelectron anomalies in the old quantum theory. Historical Studies in the Physical Sciences, 15(1):81–106, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757543.

Jennings:2006:CCH

[Jen06] Eric Thomas Jennings. Curing the colonizers: hydrotherapy, climatology, and French colonial spas. Duke University Press, Durham, NC, USA, 2006. ISBN 0-8223-3808-4 (cloth), 0-8223-3822-X (paperback). xi + 271 pp. LCCN RM813 .J46

2006. URL http://www.h-net.
org/review/hrev-a0f3m5-aa;
 http://www.loc.gov/
catdir/toc/ecip0611/
2006010438.html.

Jones-Imhotep:2008:IE

[JI08] Edward Jones-Imhotep. Icons and electronics. Historical Studies in the Natural Sciences, 38 (3):405–450, August 2008. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Jiang:2016:RPL

[Jia16] Lijing Jiang. Retouching the past with living things: Indigenous species, tradition, and biological research in Republican China, 1918–1937. Historical Studies in the Natural Sciences, 46(2):154–206, April 2016. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

James:2015:SSR

[JJ15] Jeremiah James and Christian Joas. Subsequent and subsidiary? Rethinking the role of applications in establishing quantum mechanics. Historical Studies in the Natural Sciences, 45(5):641–702, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.5.641.

Johnston:2005:WEN

[Joh05] Sean F. Johnston. From white elephant to Nobel Prize: Dennis Gabor’s wavefront reconstruction. Historical Studies in the Physical and Biological Sciences, 36(1):35–70, September 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Johnson:2008:BRW

[Joh08] Ann Johnson. Book reviews:

What If We Wrote the History of Science from the Perspective of Applied Science? Christophe Lécuyer. Making Silicon Valley: Innovation and the Growth of High Tech, 1930–1970. Cambridge, MA: MIT Press, 2005. 405 pp., illus. ISBN 978-0-262-12281-8. $40 (cloth), $22 (paper). Pap A. Ndiaye. Nylon and Bombs: DuPont and the March of Modern America. Trans. Elborg Forster. Baltimore: Johns Hopkins University Press, 2006. 289 pp. ISBN 978-0-8018-8444-3. $45 (hardcover). Peter J. Westwick. Into the Black: JPL and the American Space Program 1976–2004. New Haven: Yale University Press, 2006. 416 pp., illus. ISBN 978-0-3001-1075-3. $40 (cloth). Historical Studies in the Natural Sciences, 38(4):610–620, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Johnson:2009:HBE

[Joh09] Ann Johnson. Hitting the brakes: engineering design and the production of knowledge. Duke University Press, Durham, NC, USA, 2009. ISBN 0-8223-4526-9 (hardcover), 0-8223-4541-2 (paperback). xviii + 207 pp. LCCN TL269.3 .J64 2009.

Johnson:2013:CMG

[Joh13] Jeffrey Allan Johnson. The case of the missing German quantum chemists: On molecular models, mobilization, and the paradoxes of modernizing chemistry in Nazi Germany. Historical Studies in the Natural Sciences, 43(4):391–452, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2013.
                                        43.4.391.

Jones:2018:HWB

[Jon18] Matthew L. Jones. How we became instrumentalists (again): Data positivism since World War II. Historical Studies in the Natural Sciences, 48(5):673– 684, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Jungnickel:1979:TRP

[Jun79] Christa Jungnickel. Teaching and research in the physical sciences and mathematics in Saxony, 1820–1850. Historical Studies in the Physical Sciences, 10 (??):iv + 3–47, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757386.

Jurkowitz:2002:HLU

[Jur02] Edward Jurkowitz. Helmholtz and the liberal unification of science. Historical Studies in the Physical and Biological Sciences, 32(2):291–317, March 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kaiser:1987:ETE

[Kai87] Walter Kaiser. Early theories of the electron gas. Historical Studies in the Physical and Biological Sciences, 17(2):271– 297, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757584.

Kaiser:2002:CWR

[Kai02] David Kaiser. Cold War requisitions, scientific manpower, and the production of American physicists after World War II. Historical Studies in the Physical and Biological Sciences, 33 (1):131–159, September 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kaiser:2010:BMD

[Kai10] David Kaiser, editor. Becoming MIT: moments of decision. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-11323-6 (hardcover). vi + 207 pp. LCCN T171.M428 B43 2010.

Kaiser:2011:CC

[Kai11] David Kaiser. Consciousness on the Charles. Historical Studies in the Natural Sciences, 41 (2):265–275, Spring 2011. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.2.265.

Kapoor:1965:BPP

[Kap65] Satish C. Kapoor. Berthollet, Proust, and proportions. Chymia: Annual Studies in the History of Chemistry, 10(??): 53–110, ???? 1965. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757247.

Kargon:1977:TSC

[Kar77] Robert H. Kargon. Temple to science: Cooperative research and the birth of the California Institute of Technology. Historical Studies in the Physical Sciences, 8(??):iv + 3–31, ???? 1977. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757365.

Karafantis:2013:SIS

[Kar13] Layne Karafantis. Sealab II and Skylab: Psychological fieldwork in extreme spaces. Historical Studies in the Natural Sciences, 43(5):551–588, November 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/

stable/10.1525/hsns.2013.
43.5.551.

Kassell:2005:MME

[Kas05] Lauren Kassell. Medicine and magic in Elizabethan London: Simon Forman, astrologer, alchemist, and physician. Oxford historical monographs. Clarendon Press, New York, NY, USA, 2005. ISBN 0-19-927905-5. xviii + 281 pp. LCCN R489.F585 K37

2005. URL http://www.h-net.
org/review/hrev-a0e9l9-aa;
 http://www.loc.gov/
catdir/enhancements/
fy0637/2005298872-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0637/2005298872-t.html.

Katzir:2003:EDM

[Kat03] Shaul Katzir. From explanation to description: Molecular and phenomenological theories of piezoelectricity. Historical Studies in the Physical and Biological Sciences, 34(1): 69–94, September 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Katzir:2004:EPS

[Kat04] Shaul Katzir. The emergence of the principle of symmetry in physics. Historical Studies in the Physical and Biological Sciences, 35(1):35–65, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Katzir:2005:EWV

[Kat05] Shaul Katzir. On “The electromagnetic world-view”: a comment on an article by Suman Seth. Historical Studies in the Physical and Biological Sciences, 36(1):189–192, September 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See [Set04] and response [Set05].

Katzir:2009:HAE

[Kat09] Shaul Katzir. Hermann Aron’s electricity meters: Physics and invention in late Nineteenth-Century Germany. Historical Studies in the Natural Sciences, 39(4):444–481, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Katzir:2019:EBF

[Kat19] Shaul Katzir. Employment before formulation: Uses of protoenergetic arguments. Historical Studies in the Natural Sciences, 49(1):1–40, February 2019. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kauffman:1960:SMJ

[Kau60]  George B. Kauffman.  So-

phus Mads Jørgensen and the Werner–Jørgensen controversy. Chymia: Annual Studies in the History of Chemistry, 6(??): 180–204, ???? 1960. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757198.

Kauffman:1967:AWH

[Kau67a] George B. Kauffman. Alfred Werner’s “Habilitationsschrift”. Chymia: Annual Studies in the History of Chemistry, 12(??): 183–187, ???? 1967. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757281.

Kauffman:1967:CNE

[Kau67b] George B. Kauffman. On coordination number “eight” and the variable coordination number, by Alfred Werner. Chymia: Annual Studies in the History of Chemistry, 12(??):221– 232, ???? 1967. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757284.

Kauffman:1967:RDM

[Kau67c] George B. Kauffman. A recently discovered manuscript by Alfred Werner. Chymia: Annual Studies in the History of Chemistry, 12(??):217–219, ???? 1967. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757283.

Keller:1996:DET

[Kel96] Evelyn Fox Keller. Drosophila embryos as transitional objects: The work of Donald Poulson and Christiane Nüsslein-Volhard. Historical Studies in the Physical and Biological Sciences, 26(2):313–346, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Keller:2007:CMP

[Kel07] Richard C. (Richard Charles) Keller. Colonial madness: psychiatry in French North Africa. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-42972-5 (cloth), 0-226-42973-3 (paperback). xi + 294 pp. LCCN RC451.A42 K45 2007.

URL http://www.h-net.org/
review/hrev-a0f7o5-aa;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006026923-b.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006026923-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0619/2006026923.html.

Keller:2008:OMT

[Kel08] Evelyn Fox Keller. Organisms, machines, and thunderstorms: a history of self-organization, part one. Historical Studies in the Natural Sciences, 38(1):45–75, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Keller:2009:OMT

[Kel09] Evelyn Fox Keller. Organisms, machines, and thunderstorms: a history of self-organization, part two. Complexity, emergence, and stable attractors. Historical Studies in the Natural Sciences, 39(1):1–31, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kelty:2011:SR

[Kel11] Christopher Kelty. Steal this review! Historical Studies in the Natural Sciences, 41(2):255– 264, Spring 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2011.41.2.255.

Keller:2012:LKT

[Kel12]  Evelyn Fox Keller.  Lexi-

cons, kind-terms, and world changes. Historical Studies in the Natural Sciences, 42(5):527– 531, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.527.pdf.

Kershaw:2013:TCL

[Ker13] Michael Kershaw. Twentieth-Century length: The origins, use, and formalization of electromagnetic standards. Historical Studies in the Natural Sciences, 43(2):162–201, April 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.2.162.pdf.

Kevles:1990:CWH

[Kev90] Dan Kevles. Cold War and hot physics: Science, security, and the American State, 1945–56. Historical Studies in the Physical and Biological Sciences, 20 (2):239–264, ???? 1990. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757644.

Kevles:1994:ACW

[Kev94] Daniel Kevles. Ananda Chakrabarty wins a patent: Biotechnology, law, and society, 1972–1980. Historical Studies in the Physical and Biological Sciences, 25 (1):111–135, ???? 1994. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kevles:1997:BSB

[Kev97] Daniel J. Kevles. Big Science and big politics in the United States: Reflections on the death of the SSC and the life of the Human Genome Project. Historical Studies in the Physical and Biological Sciences, 27(2): 269–297, ???? 1997. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kevles:2009:BRH

[Kev09] Daniel J. Kevles. Book reviews: Historicizing Heredity: Jim Endersby. A Guinea Pig’s History of Biology. Cambridge, MA: Harvard University Press, 2007. xii + 499 pp., illus., index. ISBN-10: 0-674-02713-2. $27.95 (cloth). Staffan Müller-Wille and Hans-Jörg Rheinberger, eds. Heredity Produced: At the Crossroads of Biology, Politics, and Culture, 1500– 1870. Cambridge, MA and London: MIT Press, 2007. x + 496 pp., illus., index. ISBN-13: 978-0-262-13476-7. $52.00 (cloth). James Schwartz. In Pursuit of the Gene: From Darwin to DNA. Cambridge, MA: Harvard University Press, 2008. xiii + 370 pp., illus., index. ISBN-13: 978-0-674-02670-4. $29.95 (cloth). Historical Studies in the Natural Sciences, 39(4):482– 490, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kolb:1993:MWA

[KH93] Adrienne Kolb and Lillian Hoddeson. The mirage of the ‘world accelerator for world peace’ and the origins of the SSC, 1953–1983. Historical Studies in the Physical and Biological Sciences, 24(1):101–124, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kim:2002:CBI

[Kim02] Dong-Won Kim. The conflict between the image and role of physics in South Korea. Historical Studies in the Physical and Biological Sciences, 33(1): 107–129, September 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kim:2006:YNT

[Kim06] Dong-Won Kim. Yoshio Nishina and two cyclotrons. Historical Studies in the Physical and Biological Sciences, 36(2):243– 273, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kim:2014:BRM

[Kim14] Mi Gyung Kim. Book review: Material enlightenments: Bruno Belhoste. Paris savant: Parcours et rencontres au temps des Lumières. Paris: Armand Colin, 2011. 311 pp., illus., index. ISBN: 978-2-200-25563-3. 25.40 (paper); Emma C. Spary. Eating the Enlightenment: Food and the Sciences in Paris, 1670–1760. Chicago: University of Chicago Press, 2012. x + 366 pp., illus., index. ISBN: 978-0-226-76886-1. $45.00 (hardcover).; Sean Takats. The Expert Cook in Enlightenment France. Baltimore, MD: Johns Hopkins University Press, 2011. x + 203 pp., illus., index. ISBN: 978-1-4214-0283-3. $60.00 (hardcover). Historical Studies in the Natural Sciences, 44(4):424–433, September 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.4.424.

Kingston:2011:BRD

[Kin11] Ralph Kingston. Book review: Do ancient artifacts have politics? The Zodiac of Paris: How an Improbable Controversy over an Ancient Egyptian Artifact Provoked a Modern Debate between Religion and Science by Jed Z. Buchwald and Diane Greco Josefowicz. Historical Studies in the Natural Sciences, 41(4):457–469, November 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.
41.4.457.

Kinukawa:2013:LVC

[Kin13] Tomomi Kinukawa. Learned vs.

commercial?: The commodification of nature in early modern natural history specimen exchanges in England, Germany, and The Netherlands. Historical Studies in the Natural Sciences, 43(5):589–618, November 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2013.
43.5.589.

Kirsch:2005:PGP

[Kir05] Scott Kirsch. Proving grounds:

Project Plowshare and the unrealized dream of nuclear earthmoving. Rutgers University Press, New Brunswick, NJ, USA, 2005. ISBN 0-8135-3666-9 (hardcover). xi + 257 pp. LCCN TA748 .K57 2005. URL http:/

/www.loc.gov/catdir/toc/
ecip057/2005002579.html.

Kirby:2011:LCH

[Kir11] David A. (David Allen) Kirby.

Lab coats in Hollywood: science, scientists, and cinema. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-01478-5 (hardcover), 0-262-51870-8 (paperback). xiv + 265 pp. LCCN PN1995.9.S265 K57 2011.

Kitcher:2012:MLS

[Kit12] Philip Kitcher. The many lessons of structure. Historical Studies in the Natural Sciences, 42(5):255–282, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.532.pdf.

Klein:1969:GC

[Kle69] Martin J. Klein. Gibbs on Clausius. Historical Studies in the Physical Sciences, 1(??): 127–149, ???? 1969. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757297.

Klein:1970:FPB

[Kle70] Martin J. Klein. The first phase of the Bohr–Einstein dialogue. Historical Studies in the Physical Sciences, 2(??):iv + 1–39, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757302.

Kleinman:2019:WEA

[Kle19] Kim Kleinman. Why Edgar Anderson visited math departments: Natural history, statistics, and applied mathematics. Historical Studies in the Natural Sciences, 49(1):41–69, February 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kleemans:2021:SGD

[Kle21] Machiel Kleemans. Secrecy and the genesis of the 1951 Dutch– Norwegian nuclear reactor. Historical Studies in the Natural Sciences, 51(1):48–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/1/48/116292/Secrecy-
and-the-Genesis-of-the-
1951-Dutch.

Klickstein:1948:TTP

[Kli48] Herbert S. Klickstein. Thomas Thomson: Pioneer historian of chemistry. Chymia: Annual Studies in the History of Chemistry, 1(??):37–53, ???? 1948. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757113.

Klickstein:1953:CCC

[Kli53] Herbert S. Klickstein. Charles Caldwell and the controversy in America over Liebig’s “Animal Chemistry”. Chymia: Annual Studies in the History of Chemistry, 4(??):129–157, ???? 1953. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757166.

Klickstein:1959:EFS

[Kli59] Herbert S. Klickstein. Edgar Fahs Smith — his contributions to the history of chemistry. Chymia: Annual Studies in the History of Chemistry, 5(??):11–30, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757173.

Kline:2013:BRR

[Kli13] Wendy Kline. Book reviews: Risk and responsibility: The brave new world of reproductive medicine. Rene Almeling. Sex Cells: The Medical Market for Eggs and Sperm. Berkeley: University of California Press, 2011. vii + 228 pp., index. ISBN 978-0520-27096-1. $24.95 (paper).; Annie Murphy Paul. Origins: How the Nine Months Before Birth Shape the Rest of Our Lives. New York: Simon and Schuster, 2010. vii + 306, index. ISBN 978-0-7432-9663-2. $15 (paper).; Heather Munro Prescott. The Morning After: A History of Emergency Contraception in the United States. New Brunswick, NJ: Rutgers University Press, 2012. vii + 153 pp., index. ISBN 978-0-8135-5163-0. $22.95 (paper).; Chikako Takeshita. The Global Biopolitics of the IUD: How Science Constructs Contraceptive Users and Women’s Bodies. Cambridge, MA: MIT Press, 2012. xi + 238, index. ISBN 978-0-262-01658-2. $30 (hardcover). Historical Studies in the Natural Sciences, 43(2): 236–242, April 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.2.236.pdf.

Kohler:1971:OGL

[Koh71] Robert E. Kohler, Jr. The origin of G. N. Lewis’s theory of the shared pair bond. Historical Studies in the Physical Sciences, 3(??):343–376, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757322.

Kohler:1974:ILO

[Koh74] Robert E. Kohler, Jr. Irving Langmuir and the “Octet” Theory of Valence. Historical Studies in the Physical Sciences, 4 (??):39–87, ???? 1974. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757327.

Kohler:1975:LLT

[Koh75] Robert E. Kohler, Jr. The Lewis–Langmuir theory of valence and the chemical community, 1920–1928. Historical Studies in the Physical Sciences, 6(??):431–468, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757346.

Kohler:1977:RSI

[Koh77] Robert E. Kohler, Jr. Rudolf Schoenheimer, isotopic tracers, and biochemistry in the 1930’s. Historical Studies in the Physical Sciences, 8(??): 257–298, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757372.

Kohler:1991:SPD

[Koh91] Robert E. Kohler. Systems of production: Drosophila, Neurospora, and biochemical genetics. Historical Studies in the Physical and Biological Sciences, 22(1):87–130, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757674.

Kohler:2008:PPC

[Koh08] Robert E. Kohler. Plants and pigeonholes: Classification as a practice in American ecology. Historical Studies in the Natural Sciences, 38(1):77–108, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kohlstedt:2010:TCS

[Koh10] Sally Gregory Kohlstedt. Teaching children science: hands-on nature study in North America, 1890–1930. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-44990-4 (hardcover), 0-226-44992-0 (e-book). xv + 363 pp. LCCN LB1585.3 .K64

2010.  URL http://site.
ebrary.com/id/10389565.

Kohler:2011:PEA

[Koh11] Robert E. Kohler. Paul Errington, Aldo Leopold, and wildlife ecology: Residential science. Historical Studies in the Natural Sciences, 41(2):216–254, Spring 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.2.216.

Koizumi:1975:EJF

[Koi75] Kenkichiro Koizumi. The emergence of Japan’s first physicists: 1868–1900. Historical Studies in the Physical Sciences, 6(??): iv + 3–108, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757339.

Koizumi:2007:TCF

[Koi07] Kenkichiro Koizumi. Technology at a crossroads: The Fifth Generation Computer Project in Japan. Historical Studies in the Physical and Biological Sciences, 37(2):355–368, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kojevnikov:1999:DAK

[Koj99a] Alexei Kojevnikov. Dialogues about knowledge and power in totalitarian political culture. Historical Studies in the Physical and Biological Sciences, 30 (1):227–248, ???? 1999. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kojevnikov:1999:FCQ

[Koj99b] Alexei Kojevnikov. Freedom, collectivism, and quasiparticles: Social metaphors in quantum physics. Historical Studies in the Physical and Biological Sciences, 29(2):295–331, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kojevnikov:2002:DBC

[Koj02] Alexei Kojevnikov. David Bohm and collective movement. Historical Studies in the Physical and Biological Sciences, 33(1): 161–192, September 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kozhevnikov:1991:PKS

[Koz91] Aleksei Kozhevnikov. Piotr Kapitza and Stalin’s government: a study in moral choice. Historical Studies in the Physical and Biological Sciences, 22 (1):131–164, ???? 1991. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757675.

Krige:1986:CCF

[KP86] John Krige and Dominique Pestre. The choice of CERN’s first large bubble chambers for the proton synchrotron (1957–1958). Historical Studies in the Physical and Biological Sciences, 16(2): 255–279, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757566.

Kaiser:2018:PGP

[KR18] David Kaiser and Dean Rickles.

The price of gravity: Private patronage and the transformation of gravitational physics after World War II. Historical Studies in the Natural Sciences, 48(3): 338–379, June 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Kragh:1979:NBS

[Kra79] Helge Kragh. Niels Bohr’s second atomic theory. Historical Studies in the Physical Sciences, 10(??):123–186, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757389.

Kragh:1982:CPT

[Kra82] Helge Kragh. Cosmo-physics in the thirties: Towards a history of Dirac cosmology. Historical Studies in the Physical Sciences, 13(1):69–108, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757506.

Kragh:1985:FSH

[Kra85] Helge Kragh. The fine structure of hydrogen and the gross structure of the physics community, 1916–26. Historical Studies in the Physical Sciences, 15 (2):67–125, ???? 1985. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757550.

Kragh:2007:CEC

[Kra07] Helge S. Kragh. Cosmology and the entropic creation argument. Historical Studies in the Physical and Biological Sciences, 37 (2):369–382, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kraft:2009:MTL

[Kra09a] Alison Kraft. Manhattan transfer: Lethal radiation, bone marrow transplantation, and the birth of stem cell biology, ca. 1942–1961. Historical Studies in the Natural Sciences, 39(2):171– 218, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kragh:2009:SAB

[Kra09b] Helge Kragh. The spectrum of the Aurora Borealis: From enigma to laboratory science. Historical Studies in the Natural Sciences, 39(4):377–417, November 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kragh:2014:NBB

[Kra14] Helge Kragh. Naming the Big Bang. Historical Studies in the Natural Sciences, 44(1):3–36, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2014.44.1.3.

Krementsov:2011:MSE

[Kre11] N. L. Krementsov. A Martian stranded on Earth: Alexander Bogdanov, blood transfusions, and proletarian science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-45412-6 (hardcover). xvi + 175 pp. LCCN QH305.2.S63 K74 2011.

Kreiner:2021:OPO

[Kre21] Jamie Kreiner. Our pigs, ourselves. Historical Studies in the Natural Sciences, 51(4):553–??, September 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/51/4/553/118594/
Our-Pigs-Ourselves.

Krige:1999:FFE

[Kri99] John Krige. The Ford Foundation, European physics, and the Cold War. Historical Studies in the Physical and Biological Sciences, 29(2):333–361, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Krige:2001:FBC

[Kri01] John Krige. Felix Bloch and the creation of a ‘scientific spirit’ at CERN. Historical Studies in the Physical and Biological Sciences, 32(1):57–69, September 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Krige:2005:PPC

[Kri05] John Krige. The politics of phosphorus-32: a Cold War fable based on fact. Historical Studies in the Physical and Biological Sciences, 36(1): 71–91, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Krige:2008:PAP

[Kri08] John Krige. The peaceful atom as political weapon: Euratom and American foreign policy in the late 1950s. Historical Studies in the Natural Sciences, 38 (1):5–44, February 2008. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Kuhn:1984:RP

[Kuh84] Thomas S. Kuhn. Revisiting Planck. Historical Studies in the Physical Sciences, 14 (2):231–252, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757534.

Kuklick:1995:MM

[Kuk95] Henrika Kuklick. Mind over matter? Historical Studies in the Physical and Biological Sciences, 25(2):361–378, ???? 1995. CODEN HSPSEW.

ISSN 0890-9997 (print), 1533-8355 (electronic).

Kuljian:2025:SSF

[Kul25]  Christa Kuljian.  She said

— feminist critiques of sociobiology begin in 1975. Historical Studies in the Natural Sciences, 55(1):53–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/53/207541/She-Said-
Feminist-Critiques-of-
Sociobiology-Begin.

Kushner:2018:QCR

[Kus18] Birdie Kushner. The question concerning reproductive technology. Historical Studies in the Natural Sciences, 48(4):526– 535, September 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Kutcher:2009:CMC

[Kut09]  Gerald Kutcher.  Contested

medicine: cancer research and the military. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-46531-4 (hardcover). x + 247 pp. LCCN RC267 .K88 2009.

Kuznetsov:1966:DBI

[Kuz66] V. I. Kuznetsov. The development of basic ideas in the field of catalysis. Chymia: Annual Studies in the History of Chemistry, 11(??):179– 204, ???? 1966. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757267.

Kuznetsov:1971:QRR

[Kuz71] Boris Kuznetsov. Quantumrelativistic retrospection and the history of classical physics: Classical rationalism and nonclassical science. Historical Studies in the Physical Sciences, 3(??):117–135, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757316.

Kuznetsov:1974:NSP

[Kuz74] Boris Kuznetsov. Nonclassical science and the philosophy of optimism. Historical Studies in the Physical Sciences, 4 (??):197–231, ???? 1974. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757331.

Kragh:1996:SST

[KW96] Helge Kragh and Stephen J. Weininger. Sooner silence than confusion: The tortuous entry of entropy into chemistry. Historical Studies in the Physical and Biological Sciences, 27 (1):91–130, ???? 1996. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kwa:1993:MG

[Kwa93] Chunglin Kwa. Modeling the grasslands. Historical Studies in the Physical and Biological Sciences, 24(1):125–155, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Kwa:2018:VGF

[Kwa18] Chunglin Kwa. The visual grasp of the fragmented landscape: Plant geographers vs. plant sociologists. Historical Studies in the Natural Sciences, 48(2):180– 222, April 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Laemmli:2022:HCM

[Lae22] Whitney E. Laemmli. How to capture movement. Historical Studies in the Natural Sciences, 52(1):132–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/1/132/120492/How-to-
Capture-Movement.

Levey:1966:CMF

[LAKAK66] Martin Levey, Noury Al-Khaledy, and Nowry Al-Khaledy. Chemistry in the medical formulary of Al-Samarqandi. Chymia: Annual Studies in the History of Chemistry, 11(??):37–44, ???? 1966. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757258.

Lalli:2013:ARA

[Lal13] Roberto Lalli. Anti-relativity in action: The scientific activity of Herbert E. Ives between 1937 and 1953. Historical Studies in the Natural Sciences, 43 (1):41–104, February 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.1.41.pdf.

Lam:2011:PFS

[Lam11] Tong Lam. A passion for facts: social surveys and the construction of the Chinese nation state, 1900–1949, volume 9 of Asia Pacific modern. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-26786-9 (hardcover). xiii + 263 pp. LCCN HN740.Z9 S6752 2011.

Lamy:2018:MAT

[Lam18] Jérôme Lamy. The measure of all things: Space oceanography and the TOPEX/POSEIDON Mission, 1980s 1990s. Historical Studies in the Natural Sciences, 48(4):403–440, September 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Lannoo:2010:LSR

[Lan10] Michael J. Lannoo. Leopold’s shack and Ricketts’s lab: the emergence of environmentalism. University of California Press, Berkeley, CA, USA, 2010. ISBN 0-520-26478-9 (hardcover). xix + 196 pp. LCCN GE195 .L37 2010.

Lane:2011:GMS

[Lan11] K. Maria D. Lane. Geographies of Mars: seeing and knowing the Red Planet. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-47078-4 (hardcover). xiii + 266 pp. LCCN QB641 .L366 2011.

Lancaster:2019:HSC

[Lan19] Cheryl Lancaster. The history of stem cell research: A philosophical review. Historical Studies in the Natural Sciences, 49(4):447–458, September 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Lans:2023:BS

[Lan23] Aja M. Lans. Bioarchaeology of the self. Historical Studies in the Natural Sciences, 53(1): 79–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/53/1/79/
195536/Bioarchaeology-of-
the-Self.

Larson:2017:NDS

[Lar17] Carolyne R. Larson. “Noble and delicate sentiments”: Museum natural scientists as an emotional community in Argentina, 1862–1920. Historical Studies in the Natural Sciences, 47(1): 42–75, February 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Laucht:2012:EGK

[Lau12] Christoph Laucht. Elemental Germans: Klaus Fuchs, Rudolf Peierls, and the making of British nuclear culture 1939–59. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire, UK, 2012. ISBN 0-230-35487-4 (hardcover). xiv + 274 pp. LCCN QC773.3.G7 L38 2012.

Launius:2014:BRA

[Lau14] Roger D. Launius. Book reviews: Alexander C. T. Geppert, Ed. Imagining Outer Space: European Astroculture in the Twentieth Century. New York: Palgrave Macmillan, 2012. xxix + 393 pp., color plates, notes, index. ISBN 978-0-230-23172-6. $105.00 (hardcover).; Matthew H. Hersch. Inventing the American Astronaut. New York: Palgrave Macmillan, 2012. x + 219 pp., illus., notes, selected biblio., index. ISBN 978-1-137-02528-9. $27.00 (paper).; David P. D. Munns. A Single Sky: How an International Community Forged the Science of Radio Astronomy. Cambridge, MA: MIT Press, 2013. x + 247 pp., illus., biblio., index. ISBN 978-0-262-01833-3. $34.00, £23.95 (hardcover).; Kendrick Oliver. To Touch the Face of God: The Sacred, the Profane, and the American Space Program, 1957–1975. Baltimore, MD: Johns Hopkins University Press, 2013. xii + 229 pp., illus., biblio. essay, index. ISBN 978-1-4214-0788-3. $35.96 (hardcover).; Phil Tiemeyer. Plane Queer: Labor, Sexuality, and AIDS in the History of Male Flight Attendants. Berkeley: University of California Press, 2013. vii + 288 pp., illus., biblio., index. ISBN 978-0-520-27477-8. $24.95 (paper).; Peter J. Westwick, Ed. Blue Sky Metropolis: The Aerospace Century in Southern California. Foreword by William Deverell. Berkeley: University of California Press; San Marino: Huntington Library, 2012. xii + 308 pp., illus., biblio., index. ISBN 978-0-87328-249-9. $44.95, £30.95 (hardcover). Historical Studies in the Natural Sciences, 44(2):187–195, April 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2014.44.2.187.

Lavine:2014:SAP

[Lav14] Matthew Lavine. The science of automatic precision: The rise and fall of spectro-chrome therapy. Historical Studies in the Natural Sciences, 44(2):140– 177, April 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2014.44.2.140.

Levey:1959:SAM

[LB59]  Martin  Levey  and

J. E. Burke. A study of Ancient Mesopotamian bronze. Chymia: Annual Studies in the History of Chemistry, 5(??):37–50, ???? 1959. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757175.

Lecuyer:2010:MMD

[LB10]  Christophe  Lécuyer

and David C. Brock. Makers of the microchip: a documentary history of Fairchild Semiconductor. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01424-6 (hardcover). xi + 312 pp. LCCN TK7871.85 .L4167 2010.

Lecuyer:1992:MSB

[Lec92] Christophe Lecuyer. The making of a science based technological university: Karl Compton, James Killian, and the reform of MIT, 1930–1957. Historical Studies in the Physical and Biological Sciences, 23(1): 153–180, ???? 1992. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757693.

Lecuyer:1995:PRI

[Léc95] Christoph Lécuyer. MIT, progressive reform, and ‘industrial service,’ 1890–1920. Historical Studies in the Physical and Biological Sciences, 26(1):35–88, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Lecuyer:2006:MSV

[Léc06] Christophe Lécuyer. Making Silicon Valley: innovation and the growth of high tech, 1930– 1970. Inside technology. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-12281-2. x + 393 pp. LCCN HC107.C22 S395

                                        2006. URL http://www.h-net.
                                        org/review/hrev-a0f4a8-aa.

Leeds:2016:DCF

[Lee16] Adam E. Leeds. Dreams in cybernetic fugue: Cold War technoscience, the intelligentsia, and the birth of Soviet mathematical economics. Historical Studies in the Natural Sciences, 46(5):633–668, November 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Lee:2018:MTJ

[Lee18] Victoria Lee. Microbial transformations: The Japanese domestication of penicillin production, 1946–1951. Historical Studies in the Natural Sciences, 48(4):441–474, September 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Leggett:2015:BRV

[Leg15] Don Leggett. Book reviews:

A Vision of Modern Science: John Tyndall and the Role of the Scientist in Victorian Culture; ThermoPoetics: Energy in Victorian Literature and Science; Pursuing Power and Light: Technology and Physics from James Watt to Albert Einstein; Communicating Physics: The Production, Circulation and Appropriation of Ganot’s Textbooks in France and England, 1851–1887. Historical Studies in the Natural Sciences, 45(3):501–510, June 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.3.501.

Leicester:1948:FWL

[Lei48] Henry M. Leicester. Factors which led Mendeleev to the Periodic Law. Chymia: Annual Studies in the History of Chemistry, 1(??):67–74, ???? 1948. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757115.

Leicester:1959:STL

[Lei59] Henry M. Leicester. The spread of the theory of Lavoisier in Russia. Chymia: Annual Studies in the History of Chemistry, 5(??):138–144, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757181.

Leicester:1961:BCA

[Lei61] Henry M. Leicester. Biochemical concepts among the Ancient Greeks. Chymia: Annual Studies in the History of Chemistry, 7(??):9–35, ???? 1961. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757202.

Lekan:2023:DEM

[Lek23] Thomas M. Lekan. “A delicate equilibrium of the most complex sort”: Competition and community in Edward F. Ricketts’s tidepool ecologies, 1923–1948. Historical Studies in the Natural Sciences, 53(2):147–??, April 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/

53/2/147/196000/A-
Delicate-Equilibrium-of-
the-Most-Complex-Sort.

Lelong:2005:IEP

[Lel05] Benoit Lelong. Ions, electrometers, and physical constants: Paul Langevin’s laboratory work on gas discharges, 1896–1903. Historical Studies in the Physical and Biological Sciences, 36(1):93–130, September 2005. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Lemay:1949:DCD

[Lem49]  Pierre Lemay.  Désormes

et Clément découvrent et expliquent la catalyse. (French) [Désormes and Clément discover and explain catalysis]. Chymia: Annual Studies in the History of Chemistry, 2(??):45–49, ???? 1949. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757132.

Lemov:2018:EHP

[Lem18] Rebecca Lemov. An episode in the history of precrime. Historical Studies in the Natural Sciences, 48(5):637–647, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Lenoir:1986:MID

[Len86] Timothy Lenoir. Models and instruments in the development of electrophysiology, 1845–1912. Historical Studies in the Physical and Biological Sciences, 17 (1):1–54, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757574.

Leopold:2009:URC

[Leo09]  Ellen Leopold.  Under the

radar: cancer and the Cold War. Critical issues in health and medicine. Rutgers University Press, New Brunswick, NJ, USA, 2009. ISBN 0-8135-4404-1. xi + 284 pp. LCCN RC276 .L46 2009. URL http://www.

loc.gov/catdir/toc/
ecip0811/2008007751.html.

Lesch:1981:CCE

[Les81] John E. Lesch. Conceptual change in an empirical science: The discovery of the first alkaloids. Historical Studies in the Physical Sciences, 11(2): 305–328, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757482.

Leslie:1987:PEG

[Les87] Stuart W. Leslie. Playing the education game to win: The military and interdisciplinary research at Stanford. Historical Studies in the Physical and Biological Sciences, 18(1): 55–88, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757596.

Leslie:1990:PLM

[Les90] Stuart W. Leslie. Profit and loss: The military and MIT in the postwar era. Historical Studies in the Physical and Biological Sciences, 21(1):59– 85, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757655.

Leslie:2001:BCS

[Les01] Stuart W. Leslie. Blue collar science: Bringing the transistor to life in the Lehigh Valley. Historical Studies in the Physical and Biological Sciences, 32(1): 71–113, September 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Lesch:2007:FMD

[Les07] John E. Lesch. The first miracle drugs: how the sulfa drugs transformed medicine. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-518775-X (cloth). x + 364 pp. LCCN RM666.S88

L47 2007. URL http://www.
loc.gov/catdir/
enhancements/fy0726/
2005036136-d.html;   http:
//www.loc.gov/catdir/toc/
ecip065/2005036136.html.

Leslie:2008:DKB

[Les08] Stuart W. Leslie. ‘A different kind of beauty’: Scientific and architectural style in I. M. Pei’s Mesa Laboratory and Louis Kahn’s Salk Institute. Historical Studies in the Natural Sciences, 38(2):173–221, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Levey:1959:RGA

[Lev59] Martin Levey. The refining of gold in Ancient Mesopotamia. Chymia: Annual Studies in the History of Chemistry, 5(??):31– 36, ???? 1959. CODEN CHYMAP. ISSN 0095-9367.

                                        URL http://www.jstor.org/
                                        stable/27757174.

Levey:1960:EMC

[Lev60a] Martin Levey. Early Muslim chemistry: Its debt to Ancient Babylonia. Chymia: Annual Studies in the History of Chemistry, 6(??):20–26, ???? 1960. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757191.

Levey:1960:GAT

[Lev60b] Martin Levey. A group of Akkadian texts on perfumery. Chymia: Annual Studies in the History of Chemistry, 6(??):11– 19, ???? 1960. CODEN CHYMAP. ISSN 0095-9367.

                                        URL http://www.jstor.org/
                                        stable/27757190.

Levey:1961:AIA

[Lev61a] Martin Levey. Alberuni and Indian alchemy. Chymia: Annual Studies in the History of Chemistry, 7(??):36–39, ???? 1961. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757203.

Levey:1961:MIL

[Lev61b] Martin Levey. The manufacture of inks, līqs, erasure fluids, and glues — a preliminary survey in Arabic chemical technology. Chymia: Annual Studies in the History of Chemistry, 7(??):57–72, ???? 1961. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757205.

Levey:1961:SDA

[Lev61c] Martin Levey. Studies in the development of atomic theory. Chymia: Annual Studies in the History of Chemistry, 7(??):40– 56, ???? 1961. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757204.

Levey:1962:AAK

[Lev62] Martin Levey. The Aqrābādhīn of al-Kindī and early Arabic chemistry. Chymia: Annual Studies in the History of Chemistry, 8(??):11–20, ???? 1962. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757214.

Levey:1964:CTC

[Lev64a] Martin Levey. Chemical technology and commercial law in early Islam. Chymia: Annual Studies in the History of Chemistry, 9(??):19–25, ???? 1964. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757228.

Levey:1964:CKA

[Lev64b] Martin Levey. Chemistry in the “Kitab Al-Sumum” (“Book of Poisons”) by “Ibn Al-Wahshiya”. Chymia: Annual Studies in the History of Chemistry, 9(??):33–45, ???? 1964. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757230.

Levey:1964:SBI

[Lev64c] Martin Levey. Some black inks in early Mediaeval Jewish literature. Chymia: Annual Studies in the History of Chemistry, 9(??):27–31, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757229.

Levey:1966:CNE

[Lev66] Martin Levey. Chemical notions of an early Ninth-Century Christian encyclopedist. Chymia: Annual Studies in the History of Chemistry, 11(??):29–36, ???? 1966. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757257.

Levey:1967:AMT

[Lev67a] Martin Levey. Arabic mineralogy of the Tenth Century. Chymia: Annual Studies in the History of Chemistry, 12(??): 15–26, ???? 1967. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757272.

Levey:1967:MAM

[Lev67b] Martin Levey. Medieval Arabic minting of gold and silver coins. Chymia: Annual Studies in the History of Chemistry, 12(??):3–14, ???? 1967. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757271.

Levitt:2000:DPA

[Lev00] Martin Levitt. The development and politicization of the American helium industry, 1917–1940. Historical Studies in the Physical and Biological Sciences, 30 (2):333–347, ???? 2000. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Levin:2013:CWU

[Lev13] Matthew Levin. Cold War university: Madison and the New Left in the sixties. Studies in American thought and culture. The University of Wisconsin Press, Madison, 2013. ISBN 0-299-29284-3 (paperback), 0-299-29283-5 (e-book). vii + 224 + 14 pp. LCCN F589.M15 L48 2013.

Lewis:2011:DFN

[Lew11]  Andrew  J.  (An-

drew John) Lewis. A democracy of facts: natural history in the early republic. Early American studies. University of Pennsylvania Press, Philadelphia, PA, USA, 2011. ISBN 0-8122-4308-0 (hardcover). 204 pp. LCCN QH21.U6 L48 2011.

Lewenstein:2012:FKF

[Lew12] Bruce V. Lewenstein. Finding Kuhn, finding myself. Historical Studies in the Natural Sciences, 42(5):538–541, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.538.pdf.

Lowe:2022:AWN

[LGSL^{+}22] James Lowe, Miguel García-Sancho, Rhodri Leng, Mark Wong, Niki Vermeulen, and Gil Viry. Across and within networks: Thickening the history of genomics. Historical Studies in the Natural Sciences, 52 (3):443–??, June 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/52/3/443/
185527/Across-and-within-
NetworksThickening-the-
History.

Lindee:2005:MTG

[Lin05] M. Susan Lindee. Moments of truth in genetic medicine. The Johns Hopkins University Press, Baltimore, MD, USA, 2005. ISBN 0-8018-8175-7 (hardcover), 0-8018-9101-9 (paperback). xii + 270 pp. LCCN RB155 .L555 2005. URL http:

//www.loc.gov/catdir/toc/
ecip054/2004028267.html.

Lindee:2009:BRT

[Lin09] Susan Lindee. Book reviews:

Transubstantiation in Science: Angela N. H. Creager, Elizabeth Lunbeck, and M. Norton Wise, eds. Science Without Laws: Model Systems, Cases, Exemplary Narratives. Durham, NC: Duke University Press, 2007. 287 pp., illus., index. ISBN 978-0-8223-4046-1. $22.95 (cloth). Jessica Riskin, ed. Genesis Redux: Essays in the History and Philosophy of Artificial Life. Chicago: University of Chicago Press, 2007. xv + 389 pp., illus., ISBN 978-0-226-72080-7. $25.00 (cloth). Historical Studies in the Natural Sciences, 39(2):248–257, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Lindee:2020:FPA

[Lin20] M. Susan Lindee. First peoples of the atomic age: Finding new kinds of data in the biobanks of the radiation effects research foundation. Historical Studies in the Natural Sciences, 50(5):554–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/5/554/114254/First-
Peoples-of-the-Atomic-
AgeFinding-New-Kinds.

Liu:2019:ACS

[Liu19] Daniel Liu. The artificial cell, the semipermeable membrane, and the life that never was, 1864–1901. Historical Studies in the Natural Sciences, 49(5):504– 555, November 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Leicester:1950:TLD

[LK50] Henry M. Leicester and Herbert S. Klickstein. Tenney Lombard Davis and the history of chemistry. Chymia: Annual Studies in the History of Chemistry, 3(??):1–16, ???? 1950. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757143.

Lui:2024:GRH

[LK24] Mary Ting Yi Lui and Theodore Kim. Global routes and hidden labor in the American Mathematical Society’s Cold War Chinese mathematics translation program. Historical Studies in the Natural Sciences, 54(3):291– ??, June 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/54/3/291/201969/
                                        Global-Routes-and-Hidden-
                                        Labor-in-the-American.

Lykknes:2005:EGD

[LKB05] Annette Lykknes, Lise Kvittingen, and Anne Kristine Børresen. Ellen Gleditsch: Duty and responsibility in a research and teaching career, 1916–1946. Historical Studies in the Physical and Biological Sciences, 36(1): 131–188, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Lowe:2022:BPG

[LLV^{+}22] James Lowe, Rhodri Leng, Gil Viry, Mark Wong, Niki Vermeulen, and Miguel García-Sancho. The bricolage of pig genomics. Historical Studies in the Natural Sciences, 52 (3):401–??, June 2022. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/3/401/185530/The-
Bricolage-of-Pig-Genomics.

Lemay:1948:PLD

[LO48] Pierre Lemay and Ralph E. Oesper. Pierre Louis Dulong, his life and work. Chymia: Annual Studies in the History of Chemistry, 1(??):171–190, ???? 1948. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757122.

Locher:2008:STE

[Loc08] Fabien Locher. Le savant et la tempête: étudier l’atmosphère et prévoir le temps au XIXe siècle. Collection Carnot. Presses universitaires de Rennes, Rennes, France, 2008. ISBN 2-7535-0696-5. xvi + 221 pp. LCCN QC857.F7 L63 2008.

Long:2008:BRP

[Lon08] Pamela O. Long. Book reviews:

Plants and Animals in History:

The Study of Nature in Renaissance and Early Modern Europe: Brian W. Ogilvie. The Science of Describing: Natural History in Renaissance Europe. Chicago: University of Chicago Press, 2006. xvi + 385 pp., illus. ISBN 0-226-62087-5. $45 (hardcover), $27 (paper). Harold J. Cook. Matters of Exchange: Commerce, Medicine, and Science in the Dutch Golden Age. New Haven: Yale University Press, 2007. xiv + 562 pp. 60 black and white illus. ISBN 978-0-300-11796-7. $35 (hardcover). Alix Cooper. Inventing the Indigenous: Local Knowledge and Natural History in Early Modern Europe. Cambridge: Cambridge University Press, 2007. xi + 218 pp., illus. ISBN 978-0-521-87087-0. $80 (hardcover). Antonio Barrera-Osorio. Experiencing Nature: The Spanish American Empire and the Early Scientific Revolution. Austin: University of Texas Press, 2006. xi + 211 pp., illus. ISBN 0-292-70981-1. $45 (hardcover). Historical Studies in the Natural Sciences, 38(2):313–323, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Long:2011:APR

[Lon11] Pamela O. Long. Artisan/ practitioners and the rise of the new sciences, 1400–1600. The OSU Press Horning visiting scholars publication series. Oregon State University Press, Corvallis, OR, USA, 2011. ISBN 0-87071-609-3 (paperback), 0-87071-647-6 (e-book). xii + 196 pp. LCCN Q124.97 .L66 2011.

URL http://www.netread.
com/jcusers/1309/2473069/
image/lgcover.3175254.jpg.

Longino:2012:LTS

[Lon12] Helen E. Longino. Lessons from teaching the structure of scientific revolutions. Historical Studies in the Natural Sciences, 42(5):542–544, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.542.pdf.

Lowood:1986:SBa

[Low86a] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 16(2):371–377, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757570.

Lowood:1986:SBb

[Low86b] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 17(1):183–190, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757579.

Lowood:1987:SB

[Low87] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 17(2):369–381, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757590.

Lowood:1988:SBb

[Low88a] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 18(2):403–414, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757608.

Lowood:1988:SBa

[Low88b] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 19(1):199–210, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757622.

Lowood:1989:SBb

[Low89a] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 19(2):403–413, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757629.

Lowood:1989:SBa

[Low89b] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 20(1):183–195, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757639.

Lowood:1990:SBb

[Low90a] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 20(2):423–436, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757649.

Lowood:1990:SBa

[Low90b] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 21(1):197–210, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757659.

Lowy:1990:VMD

[Löw90c] Ilana Löwy. Variances in meaning in discovery accounts: The case of contemporary biology. Historical Studies in the Physical and Biological Sciences, 21 (1):87–121, ???? 1990. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757656.

Lowood:1991:SB

[Low91] Henry Lowood. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 21(2):399–407, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757667.

Low:1999:SCS

[Low99] Morris Low. Science and civil society in Japan: Physicists as public men and policymakers. Historical Studies in the Physical and Biological Sciences, 30 (1):193–225, ???? 1999. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Low:2006:APP

[Low06] Morris F. Low. Accelerators and politics in postwar Japan. Historical Studies in the Physical and Biological Sciences, 36 (2):275–296, March 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Lowengard:2023:DAB

[Low23] Sarah Lowengard. On the disappearance of the animal body: Animal fat, tallow, candles, soap, and chemistry before 1830. Historical Studies in the Natural Sciences, 53(3):242– ??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/53/3/242/196657/
                                        On-the-Disappearance-of-
                                        the-Animal-BodyAnimal-Fat.

Lazaroff-Puck:2024:ELT

[LP24]  Cameron  Lazaroff-Puck.

Empire-laden theory: The technological and colonial roots of Maxwell’s theories of electromagnetism. Historical Studies in the Natural Sciences, 54(1): 42–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/54/1/42/
200202/Empire-Laden-
TheoryThe-Technological-
and-Colonial.

Lecuyer:2013:LMI

[LU13]  Christophe  Lécuyer

and Takahiro Ueyama. The logics of materials innovation: The case of gallium nitride and blue light emitting diodes. Historical Studies in the Natural Sciences, 43(3):243–280, June 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.3.243.pdf.

Lucier:2008:SSC

[Luc08] Paul Lucier. Scientists and swindlers: consulting on coal and oil in America, 1820–1890. Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-9003-9 (hardcover). xiii + 426 pp. LCCN Q127.U6 L83

2008. URL http://www.loc.
gov/catdir/toc/ecip0811/
2008006551.html.

Lukyanov:1964:FCL

[Luk64] P. M. Luk’yanov. The first chemical laboratories in Russia. Chymia: Annual Studies in the History of Chemistry, 9 (??):59–69, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757232.

Luk:2024:CFJ

[Luk24] Christine Y. L. Luk. The Chinese freshwater jellyfish unbound: Evolution, nomenclature, and bioinvasion of Craspedacusta sowerbii, 1880–1941. Historical Studies in the Natural Sciences, 54(4):493– ??, September 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

                                        https://online.ucpress.
                                        edu/hsns/article/54/4/493/
                                        203069/The-Chinese-
                                        Freshwater-Jellyfish-
                                        UnboundEvolution.

Leng:2022:SSW

[LVGS^{+}22] Rhodri Leng, Gil Viry, Miguel García-Sancho, James Lowe, Mark Wong, and Niki Vermeulen. The sequences and the sequencers: What can a mixedmethods approach reveal about the history of genomics? Historical Studies in the Natural Sciences, 52(3):277–??, June 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        52/3/277/185529/The-
                                        Sequences-and-the-
                                        SequencersWhat-Can-a-
                                        Mixed.

Lynch:2012:KSS

[Lyn12] Michael Lynch. Kuhn’s structure and science studies: a comment on commentaries. Historical Studies in the Natural Sciences, 42(5):545–550, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.545.pdf.

MacKinnon:1977:HMR

[Mac77] Edward MacKinnon. Heisenberg, models, and the rise of matrix mechanics. Historical Studies in the Physical Sciences, 8 (??):137–188, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757370.

Macuglia:2024:BBS

[Mac24] Daniele Macuglia. Blending borders and sparking change: Sidney Yip, hybridity, and the rise of molecular simulations in Cold War materials science. Historical Studies in the Natural Sciences, 54(5):569–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/54/5/569/203888/
Blending-Borders-and-
Sparking-ChangeSidney-Yip.

Madison:2017:PI

[Mad17] Paige Madison. The paleontological imagination. Historical Studies in the Natural Sciences, 47(2):255–262, April 2017. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Maltese:2000:LER

[Mal00]  Guilio  Mal-

tese. The late entrance of Relativity into Italian scientific community (1906–1930). Historical Studies in the Physical and Biological Sciences, 31(1):125–173, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Martini:2012:AOH

[Mar12] Edwin A. Martini. Agent Orange: history, science, and the politics of uncertainty. Culture, politics, and the cold war. University of Massachusetts Press, Amherst, MA, USA, 2012. ISBN 1-55849-975-X (paperback), 1-55849-974-1 (library hardcover). xvi + 302 pp. LCCN DS559.8.C5 M37 2012.

Martin:2015:FDP

[Mar15] Joseph D. Martin. Fundamental disputations: The philosophical debates that governed American physics, 1939–1993. Historical Studies in the Natural Sciences, 45(5):703–757, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2015.
                                        45.5.703.

Martin:2016:SJR

[Mar16] Jennifer A. Martin. Seeing jaws: The role of shark science in ocean conservation. Historical Studies in the Natural Sciences, 46(1):67–100, February 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Martin:2021:SAI

[Mar21] Joseph D. Martin. Science in the age of invincible surmise: Nuclear optimism and the Michigan memorial–Phoenix project. Historical Studies in the Natural Sciences, 51(2):179–??, April 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/2/179/116758/Science-
in-the-Age-of-Invincible-
SurmiseNuclear.

Marino:2022:FCW

[Mar22] Julia Marino. Fighting the Cold War and the “Market War” through critical technologies, 1979–1992. Historical Studies in the Natural Sciences, 52(4):485–??, September 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/4/485/191418/Fighting-
the-Cold-War-and-the-
Market-War-through.

Martin:2024:IFC

[Mar24a] Joseph D. Martin. Intelligence falsely so-called. Historical Studies in the Natural Sciences, 54(5):636–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/5/636/203881/
                                        Intelligence-Falsely-So-
                                        Called.

Martin:2024:WSH

[Mar24b] Joseph D. Martin. Why study the history of the natural sciences? Historical Studies in the Natural Sciences, 54(1):1–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/54/1/1/200200/Why-
                                        Study-the-History-of-the-
                                        Natural-Sciences.

Masutti:2006:FCC

[Mas06] Christophe Masutti. Frederic Clements, climatology, and conservation in the 1930s. Historical Studies in the Physical and Biological Sciences, 37(1): 27–48, September 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Mas:2022:HE

[Mas22] Catherine Mas. How not to be an expert. Historical Studies in the Natural Sciences, 52 (1):136–??, February 2022. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        52/1/136/120494/How-Not-
                                        to-Be-an-Expert.

Mateos:2015:BRC

[Mat15]  Gisela Mateos.  Book re-

view: Closing and opening doors: Spanish history of science and science historiographies in Spanish: Knowledge production, circulation, and asymmetries. Alexis De Grieff. A las puertas del universo derrotado. Bogotá: Editorial Universidad Nacional de Colombia, 2012. 182 pp., illus., index. ISBN 978-958-761-142-7. $15.46 (hardcover); Néstor Herrán and Xavier Roqué, eds. La física en la dictadura: Físicos, cultura y poder en España 1939– 1975. Barcelona: Universitat Autònoma de Barcelona, 2012. 310 pp., index. ISBN 978-84-490-3238-7. $20.75 (paper).; Stefan Pohl Valero. Energía y cultura: historia de la termodinámica en la España de la segunda mitad del siglo XX. Bogotá: Editorial Pontificia Universidad Javeriana, Editorial Universidad del Rosario, 2011. 321 pp., index. ISBN 978-958-716-498-5. $23 (paper). Historical Studies in the Natural Sciences, 45(1):198–205, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.1.198.

Mauskopf:2012:TKC

[Mau12] Seymour H. Mauskopf. Thomas S. Kuhn and the chemical revolution. Historical Studies in the Natural Sciences, 42(5):551– 556, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.551.pdf.

Mayor:2005:FLF

[May05] Adrienne Mayor. Fossil legends of the first Americans. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11345-9 (cloth). xxxix + 446 pp. LCCN E58 .M36

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0704/2004053234-d.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0704/2004053234-t.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0737/2004053234-b.html.

Mody:2013:MSN

[MC13] Cyrus C. M. Mody and Hyungsub Choi. From materials science to nanotechnology: Interdisciplinary center programs at Cornell University, 1960– 2000. Historical Studies in the Natural Sciences, 43(2): 121–161, April 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.2.121.pdf.

McCray:2013:VEH

[McC3 ] Patrick (W. Patrick) McCray. The Visioneers (eGalley): How a Group of Elite Scientists Pursued Space Colonies, Nanotechnologies, and a Limitless Future. Princeton University Press, Princeton, NJ, USA, 2013 (??). ISBN 1-4008-9963-X. 341 pp. LCCN Q125 .M417

2012. URL http://public.
eblib.com/EBLPublic/
PublicView.do?ptiID=
979998.

McConnell:1964:CMH

[McC64] Virginia F. McConnell. Clara de Milt, historian of science. Chymia: Annual Studies in the History of Chemistry, 9(??): 201–215, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757241.

McCormmach:1970:ELE

[McC70] Russell McCormmach. Einstein, Lorentz, and the electron theory. Historical Studies in the Physical Sciences, 2 (??):41–87, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757304.

McCray:2004:PVC

[McC04] W. Patrick McCray. Project Vista, Caltech, and the dilemmas of Lee DuBridge. Historical Studies in the Physical and Biological Sciences, 34(2): 339–370, March 2004. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

McCray:2008:KWS

[McC08] Patrick (W. Patrick) McCray. Keep watching the skies!: the story of Operation Moonwatch and the dawn of the space age. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-12854-5. xiii + 308 pp. LCCN QB33.U6 M33

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0835/2007941755-b.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0835/2007941755-d.html;
                                         http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0835/2007941755-t.html.

McCormmach:2015:HSP

[McC15] Russell McCormmach, editor. Historical Studies in the Physical Sciences, volume 7. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 1-4008-7018-6. 528 pp. LCCN Q125; Q125 .H578 1976eb.

McEvoy:2007:MPH

[McE07] John G. McEvoy. Modernism, postmodernism and the historiography of science. Historical Studies in the Physical and Biological Sciences, 37(2):383– 408, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

McGovern:2018:TEH

[McG18] Michael F. McGovern. Toward an epistemic history of political things: Cartography, science, and states. Historical Studies in the Natural Sciences, 48(3):380– 389, June 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

McManus:2022:SIC

[McM22] Alison Lynn McManus. Science, interrupted: Censorship and the problem of credit allocation in the American Advisory Committee on Scientific Publications, 1940–46. Historical Studies in the Natural Sciences, 52(1):80–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/1/80/120500/Science-
InterruptedCensorship-and-
the-Problem-of.

Moore:2023:I

[MCR23] Taylor M. Moore, Henry M.

Cowles, and Chitra Ramalingam. Introduction. Historical Studies in the Natural Sciences, 53(1):71–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/1/71/195531/
Introduction.

Myers:2003:WIL

[MD03]  Robert  A.  Myers

and Richard W. Dixon. Who invented the laser: An analysis of the early patents. Historical Studies in the Physical and Biological Sciences, 34(1): 115–149, September 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Meister:2023:COF

[Mei23] Anna-Maria Meister. Coded objects: The forms of protoalgorithmic thinking. Historical Studies in the Natural Sciences, 53(5):518–??, November 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        53/5/518/198194/Coded-
                                        ObjectsThe-Forms-of-Proto-
                                        Algorithmic.

Melhado:1980:MDI

[Mel80] Evan M. Melhado. Mitscherlich’s discovery of isomorphism. Historical Studies in the Physical Sciences, 11(1):87–123, ???? 1980. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757472.

Melhado:1983:OPC

[Mel83] Evan M. Melhado. Oxygen, phlogiston, and caloric: The case of Guyton. Historical Studies in the Physical Sciences, 13 (2):311–334, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757518.

Meli:1993:ERF

[Mel93] Domenico Bertoloni Meli. The emergence of reference frames and the transformation of mechanics in the enlightenment.

Historical Studies in the Physical and Biological Sciences, 23 (2):301–335, ???? 1993. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757701.

Menzel:2021:WRL

[Men21] Julia Harriet Menzel. Wilsonian renormalization in the 1970s: Labor markets, geopolitics, and the rise of a new theory. Historical Studies in the Natural Sciences, 51(5):605–??, November 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/5/605/119067/Wilsonian-
Renormalization-in-the-
1970sLabor.

Miles:1949:SKD

[Mil49] Wyndham Miles. Sir Kenelm Digby, alchemist, scholar, courtier, and man of adventure. Chymia: Annual Studies in the History of Chemistry, 2(??):119–128, ???? 1949. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757138.

Miles:1950:EAC

[Mil50] Wyndham Miles. Early American chemical societies: 1. the 1789 Chemical Society of Philadelphia 2. the Chemical Society of Philadelphia. Chymia: Annual Studies in the History of Chemistry, 3(??):95–113, ???? 1950.

CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757148.

Miles:1953:BRC

[Mil53] Wyndham Miles. Benjamin Rush, chemist. Chymia: Annual Studies in the History of Chemistry, 4(??):iv + 37–77, ???? 1953. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757159.

Miles:1959:CCS

[Mil59]  Wyndham D. Miles.  The

Columbian Chemical Society. Chymia: Annual Studies in the History of Chemistry, 5(??): 145–154, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757182.

Miles:1961:SHD

[Mil61]  Wyndham D. Miles.  “Sir

Humphrey Davie, the prince of agricultural chemists”. Chymia: Annual Studies in the History of Chemistry, 7(??):126–134, ???? 1961. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757209.

Miles:1966:JCB

[Mil66] Wyndham D. Miles. With James Curtis Booth in Europe, 1834. Chymia: Annual Studies in the History of Chemistry, 11 (??):139–149, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757264.

Mills:1990:UMC

[Mil90] Eric L. Mills. Useful in many capacities. an early career in American physical oceanography. Historical Studies in the Physical and Biological Sciences, 20(2): 265–311, ???? 1990. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757645.

Miller:1999:XLZ

[Mil99] H. Lyman Miller. Xu Liangying and He Zuoxiu: Divergent responses to physics and politics in the post-Mao period. Historical Studies in the Physical and Biological Sciences, 30 (1):89–114, ???? 1999. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Milam:2010:EWF

[Mil10a] Erika Lorraine Milam. The equally wonderful field: Ernst Mayr and organismic biology. Historical Studies in the Natural Sciences, 40(3):279–317, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Milam:2010:LFG

[Mil10b] Erika Lorraine Milam. Looking for a few good males: female choice in evolutionary biology. Animals, history, culture. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9419-0 (hardcover). 236 pp. LCCN QL761 .M48 2010.

Milam:2020:ISI

[Mil20] Erika Lorraine Milam. Introduction to special issue: Temporal horizons. Historical Studies in the Natural Sciences, 50 (1–2):1–4, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Miri:2021:FVB

[Mir21] Johnny Miri. The fall of Vannevar Bush: The forgotten war for control of science policy in postwar America. Historical Studies in the Natural Sciences, 51(4):507–??, September 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        51/4/507/118593/The-Fall-
                                        of-Vannevar-BushThe-
                                        Forgotten-War-for.

Mitchell:2021:BRB

[Mit21] M. X. Mitchell. Book review: Border Crossings. Historical Studies in the Natural Sciences, 51(3):414–??, June 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        51/3/414/117358/Border-
                                        Crossings.

Mitchell:2023:SMS

[Mit23] Elise A. Mitchell. On slavery, medicine, speculation, and the archive. Historical Studies in the Natural Sciences, 53 (1):82–??, February 2023. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/1/82/195537/On-Slavery-
Medicine-Speculation-and-
the-Archive.

Martin:2015:BCM

[MJ15] Joseph D. Martin and Michel Janssen. Beyond the crystal maze: Twentieth-century physics from the vantage point of solid state physics. Historical Studies in the Natural Sciences, 45(5):631–640, November 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.5.631.

Menzel:2020:WCW

[MK20] Julia Harriet Menzel and David Kaiser. Weimar, Cold War, and historical explanation: Rereading Forman. Historical Studies in the Natural Sciences, 50(1–2):31–40, April 2020. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

McEvoy:1975:GNP

[MM75] J. G. McEvoy and J. E.

McGuire. God and nature: Priestley’s way of rational dissent. Historical Studies in the Physical Sciences, 6(??): 325–404, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757344.

Martin:2021:SII

[MMMSD21] Joseph D. Martin, Gisela Mateos, David P. D. Munns, and Edna Suárez-Díaz. Special issue introduction: Historical peculiarity and the Order of the Phoenix. Historical Studies in the Natural Sciences, 51(2):169– ??, April 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/51/2/169/116760/
                                        Special-Issue-
                                        IntroductionHistorical-
                                        Peculiarity.

Martin-Nielsen:2018:CCW

[MN18] Janet Martin-Nielsen. Computing the climate: When models became political. Historical Studies in the Natural Sciences, 48(2):223–245, April 2018. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Mutis:2009:FNM

[MOBG09] José Celestino Mutis, Pedro Ortiz Valdivieso, J. E. (Jaime E.) Bernal Villegas, and Alberto Gómez Gutiérrez. Filosofía natural mutisiana: edición facsimilar y estudio de un manuscrito inédito de José Celestino Mutis y Bosio, correspondiente a su tercera Oración inaugural pronunciada en el Colegio Mayor de Nuestra Señora del Rosario, Santafé, octubre de 1764. Colección Mutisiana. Pontificia Universidad Javeriana, Bogotá, 2009. ISBN 958-716-263-3. 249 pp. LCCN BD581 .M88 2009.

Mody:2008:BRH

[Mod08] Cyrus C. M. Mody. Book reviews: How I Learned to Stop Worrying and Love the Bomb, the Nuclear Reactor, the Computer, Ham Radio, and Recombinant DNA: Atsushi Akera. Calculating a Natural World: Scientists, Engineers, and Computers during the Rise of U.S. Cold War Research. Cambridge, MA: MIT Press, 2007. ix + 427 pp., illus., index. ISBN 978-0262-01231-7. $40 (hardcover). Mark D. Bowles. Science in Flux: NASA’s Nuclear Program at Plum Brook Station, 1955–2005. Washington, DC: National Aeronautics and Space Administration, 2006. xxix + 335 pp., illus., index. NASA SP-2006-4317. $0 (digital) or $49 (hardcover). Kristen Haring. Ham Radio’s Technical Culture. Cambridge, MA: MIT Press, 2007. xvii + 220 pp., illus., index. ISBN 978-0262-08355-3. $27.95 (hardcover). Scott Kirsch. Proving Grounds: Project Plowshare and the Unrealized Dream of Nuclear Earthmoving. New Brunswick: Rutgers University Press, 2005. xi + 257 pp., illus., index. ISBN 978-08135-3666-8. $39.95 (cloth). Eric J. Vettel. Biotech: The Countercultural Origins of an Industry. Philadelphia: University of Pennsylvania Press, 2006. xv + 273 pp., illus., index. ISBN 978-08122-3947-8. $39.95 (hardcover). Historical Studies in the Natural Sciences, 38(3):451– 461, August 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Mody:2012:ICP

[Mod12] Cyrus C. M. Mody. Instrumental community: probe microscopy and the path to nanotechnology. Inside technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-13494-2, 0-262-29818-X (e-book). xiii + 260 pp. LCCN T174.7 .M63

                                        2011. URL http://mitpress-
                                        ebooks.mit.edu/product/
                                        instrumental-community;
                                        http://muse.jhu.edu/books/
                                        9780262298186/.

Mody:2020:HSW

[Mod20] Cyrus C. M. Mody. Historical studies in which sciences? The revolving door of engineering and technology. Historical Studies in the Natural Sciences, 50(1–2):41–49, April 2020. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Monaldi:2008:IOD

[Mon08] Daniela Monaldi. The indirect observation of the decay of mesotrons: Italian experiments on cosmic radiation, 1937–1943. Historical Studies in the Natural Sciences, 38(3):353–404, August 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Moore:2016:RWB

[Moo16] Kelly Moore. Review: The Wild Borderlands of Science and Technology. Historical Studies in the Natural Sciences, 46(1):101–109, February 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Moore:2023:CA

[Moo23] Taylor M. Moore. The cool air.

Historical Studies in the Natural Sciences, 53(1):86–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/1/86/195539/The-Cool-
Air.

Moore:2024:MHS

[Moo24] Samuel A. Moore. “Machinery hurtful to [scientific] commonality”: Automation and scientific publication. Historical Studies in the Natural Sciences, 54(5):621–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/5/621/203885/Machinery-
Hurtful-to-Scientific-
Commonality.

Morrell:1971:ISB

[Mor71] J. B. Morrell. Individualism and the structure of British science in 1830. Historical Studies in the Physical Sciences, 3 (??):183–204, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757318.

Morrell:1992:RPC

[Mor92]  Jack Morrell.  Research in

physics at the Clarendon Laboratory, Oxford, 1919–1939. Historical Studies in the Physical and Biological Sciences, 22(2): 263–307, ???? 1992. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757683.

Moran:2005:DKA

[Mor05] Bruce T. Moran. Distilling knowledge: alchemy, chemistry, and the scientific revolution. New histories of science, technology, and medicine. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01495-2. 210 pp. LCCN QD15 .M67 2005.

Morus:2009:BRS

[Mor09] Iwan Rhys Morus. Book review: Sensational, Spectacular Science: Myles W. Jackson. Harmonious Triads: Physicists, Musicians, and Instrument Makers in Nineteenth-Century Germany. Cambridge, MA: MIT Press, 2006. x + 395 pp., illus., index. ISBN 978-0-262-10116-5. $40.00 (cloth). Chris Otter. The Victorian Eye. Chicago: University of Chicago Press, 2008. x + 382 pp., illus., index. ISBN 978-0-226-64076-1. $65.00 (cloth). Jonathan Smith. Charles Darwin and Victorian Visual Culture. Cambridge: Cambridge University Press, 2006. xxiii + 349 pp., illus., index. ISBN 978-0-521-85690-4. $116.00 (cloth). Historical Studies in the Natural Sciences, 39(3):356–366, August 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Moreira:2021:BRA

[Mor21a] Tiago Moreira. Book review:

Aging: Remaking science, technology, and medicine. Historical Studies in the Natural Sciences, 51(2):274–??, April 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/2/274/116759/
AgingRemaking-Science-
Technology-and-Medicine.

Morus:2021:BRS

[Mor21b] Iwan Rhys Morus. Book review: Social Scientists. Historical Studies in the Natural Sciences, 51(5):634–??, November 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/5/634/119065/Social-
Scientists.

Mateos:2021:DPL

[MSD21] Gisela Mateos and Edna Suárez-Díaz. “The door to the promised land of atomic peace and plenty”: Mexican students and the Phoenix Memorial Project. Historical Studies in the Natural Sciences, 51 (2):209–??, April 2021. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/2/209/116756/The-Door-
to-the-Promised-Land-of-
Atomic-Peace-and.

Mukerji:2009:IET

[Muk09] Chandra Mukerji. Impossible engineering: technology and territoriality on the Canal du Midi. Princeton studies in cultural sociology. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14032-4 (hardcover). xix + 304 pp. LCCN HE445.M52 M85 2009.

Mukharji:2020:BHI

[Muk20] Projit Bihari Mukharji. Bloodworlds: A hematology of the 1952 Indo–Australian Genetical Survey of the Chenchus. Historical Studies in the Natural Sciences, 50(5):525–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/5/525/114253/
BloodworldsA-Hematology-
of-the-1952-Indo.

Mulcahy:2008:HSB

[Mul08]  Matthew Mulcahy.  Hurri-

canes and society in the British Greater Caribbean, 1624–1783.

Early America: history, context, culture. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-9079-9. ix + 257 pp. LCCN GF526. URL http://

www.gbv.de/dms/bowker/toc/
9780801882234.

Mullaney:2012:CTN

[Mul12] Thomas Shawn Mullaney. Coming to terms with the nation: ethnic classification in modern China, volume 18 of Asia - local studies/global themes. University of California Press, Berkeley, CA, USA, 2012. ISBN 0-520-27274-9 (paperback). xxiv + 232 pp. LCCN DS730 .M85 2012.

Mundy:1967:ADS

[Mun67]  B. W. Mundy.  Avogadro

on the degree of submolecularity of molecules. Chymia: Annual Studies in the History of Chemistry, 12(??):151– 155, ???? 1967. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757278.

Munns:1997:LAR

[Mun97] David Munns. Linear accelerators, radio astronomy, and Australia’s search for international prestige, 1944–1948. Historical Studies in the Physical and Biological Sciences, 27(2):299–317, ???? 1997. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Munns:2003:IWB

[Mun03] David Munns. If we build it, who will come? Radio astronomy and the limitations of “national” laboratories in Cold War America. Historical Studies in the Physical and Biological Sciences, 34(1):95–113, September 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Munz:2011:MGC

[Mun11] Tania Munz. “My Goose Child Martina”. Historical Studies in the Natural Sciences, 41(4): 405–446, November 2011. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2011.
                                        41.4.405.

Mundey:2012:CNW

[Mun12] Lisa M. Mundey. The civilianization of a nuclear weapon effects test: Operation AR-GUS. Historical Studies in the Natural Sciences, 42(4):283– 321, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2012.
                                        42.4.283.

Munns:2013:SSH

[Mun13] David P. D. Munns. A single sky: how an international community forged the science of radio astronomy. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01833-0 (hardcover). xi + 247 pp. LCCN QB475.A25 M86 2013.

Munns:2021:TSP

[Mun21] David P. D. Munns. Teaching in a swimming pool: The Ford nuclear reactor and the training of the atomic age. Historical Studies in the Natural Sciences, 51(2):232–??, April 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/2/232/116755/Teaching-
in-a-Swimming-PoolThe-
Ford-Nuclear.

Muller-Wille:2018:MUP

[MW18] Staffan Müller-Wille. Making and unmaking populations. Historical Studies in the Natural Sciences, 48(5):604–615, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Muller-Wille:2007:HPC

[MWR07] Staffan Müller-Wille and Hans-Jörg Rheinberger, editors. Heredity produced: at the crossroads of biology, politics, and culture, 1500– 1870. Transformations. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-13476-4. x + 496 + 2 pp. LCCN QH431 .H434 2007. URL http://www.

loc.gov/catdir/toc/
ecip0617/2006023090.html.

Navarro:2016:EWO

[Nav16] Jaume Navarro. Ether and wireless: An old medium into new media. Historical Studies in the Natural Sciences, 46(4):460– 489, September 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Ndiaye:2007:NBD

[Ndi07] Pap Ndiaye. Nylon and bombs: DuPont and the march of modern America. Studies in industry and society. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8444-6 (hardcover). 289 + 8 pp. LCCN HD9651.9.D8 N3513 2007. URL http://www.

loc.gov/catdir/
enhancements/fy0703/
2006004143-b.html;
http://www.loc.gov/catdir/
enhancements/fy0703/
2006004143-d.html;   http:
//www.loc.gov/catdir/toc/
ecip068/2006004143.html.

Nebeker:1994:ESH

[Neb94] Frederik Nebeker. Experimental style in high-energy physics: The discovery of the upsilon particle. Historical Studies in the Physical and Biological Sciences, 25(1):137–164, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Needell:1983:NRF

[Nee83] Allan A. Needell. Nuclear reactors and the founding of Brookhaven National Laboratory. Historical Studies in the Physical Sciences, 14(1): 93–122, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757526.

Needell:1987:PSA

[Nee87] Allan A. Needell. Preparing for the space age: Universitybased research, 1946–1957. Historical Studies in the Physical and Biological Sciences, 18(1): 89–109, ???? 1987. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757597.

Nepomucene:1960:RMC

[Nep60] St. John Nepomucene. Rydberg: The man and the constant. Chymia: Annual Studies in the History of Chemistry, 6 (??):127–145, ???? 1960. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757195.

Neswald:2013:SIC

[Nes13] Elizabeth Neswald. Strategies of international communitybuilding in early twentiethcentury metabolism research: The foreign laboratory visits of Francis Gano Benedict. Historical Studies in the Natural Sciences, 43 (1):1–40, February 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.1.1.pdf.

Neufeld:2014:FMP

[Neu14] Michael J. Neufeld. First mission to Pluto: Policy, politics, science, and technology in the origins of new horizons, 1989– 2003. Historical Studies in the Natural Sciences, 44(3):234– 276, June 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

                                        //www.jstor.org/stable/10.
                                        1525/hsns.2014.44.3.234.

Neville:1962:OMF

[Nev62] Roy G. Neville. “Observations sur la Mine de Fer de Bagory” (1767), an unpublished manuscript by P.-J. Macquer. Chymia: Annual Studies in the History of Chemistry, 8(??):89– 96, ???? 1962. CODEN CHYMAP. ISSN 0095-9367.

                                        URL http://www.jstor.org/
                                        stable/27757219.

Neville:1965:CGT

[Nev65] Roy G. Neville. Christophle Glaser and the “Traité de la Chymie,” 1663. Chymia: Annual Studies in the History of Chemistry, 10(??):25–52, ???? 1965. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757246.

Nieto-Galan:2012:RNH

[NG12] Agustí Nieto-Galan. A republican natural history in Spain around 1900: Odóe Buen (1863– 1945) and his audiences. Historical Studies in the Natural Sciences, 42(3):159–189, June 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2012.42.3.159.

Ng:2019:BSR

[Ng19] Queenie Ng. Biographies in shaping and reshaping science. Historical Studies in the Natural Sciences, 49(3):338–349, June 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Nichols:2024:HPS

[Nic24] Tiffany Nichols. Hidden in plain sight: Discerning signal from noise in the expanded laboratory environment. Historical Studies in the Natural Sciences, 54(3):335–??, June 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/3/335/201971/Hidden-in-
Plain-SightDiscerning-
Signal-from-Noise.

Nielsen:1989:AKL

[Nie89] Keld Nielsen. Another kind of light: The work of T. J. Seebeck and his collaboration with Goethe. Part I. Historical Studies in the Physical and Biological Sciences, 20(1):107– 178, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757637.

Nielsen:1991:AKL

[Nie91] Keld Nielsen. Another kind of light: The work of T. J. Seebeck and his collaboration with Goethe, part 2. Historical Studies in the Physical and Biological Sciences, 21(2):317– 397, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757666.

Nisio:1976:MTH

[Nis76] Sigeko Nisio. In memoriam: Tetu Hirosige (1928–1975). Historical Studies in the Physical Sciences, 7(??):488–489, ???? 1976. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757362.

Nofre:2021:PEP

[Nof21] David Nofre. The politics of early programming languages: IBM and the Algol Project. Historical Studies in the Natural Sciences, 51(3):379–??, June 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        51/3/379/117357/The-
                                        Politics-of-Early-
                                        Programming-LanguagesIBM-
                                        and.

Norton:1984:HEF

[Nor84] John Norton. How Einstein found his field equations: 1912– 1915. Historical Studies in the Physical Sciences, 14(2):

253–316, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757535. Reprinted in

[HS89, 101–159].

November:2012:BCD

[Nov12]  Joseph  Adam

November. Biomedical computing: digitizing life in the United States, volume 130 (series 2) of Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0468-0 (hardcover). xvi + 344 pp. LCCN R858 .N68 2012.

URL http://catalogue.bnf.
fr/ark:/12148/cb427367061.

Novoa:2015:BRI

[Nov15] Adriana Novoa. Book reviews:

The intertwined paths of science, slavery, and race: Judith Carney and Richard Nicholas Rosomoff. In the Shadow of Slavery: Africa’s Botanical Legacy in the Atlantic World. Berkeley: University of California, 2009. 280 pp., illus. ISBN: 0-520-25750-2. ISBN: 978-0-520-25750-4. $46.04 (cloth), $22.69 (paper); Andrew Curran. The Anatomy of Blackness: Science & Slavery in an Age of Enlightenment. Baltimore: Johns Hopkins University Press, 2011. 310 pp., illus. ISBN: 1-4214-0965-8. ISBN: 978-1-4214-0965-8. $65.33 (cloth), $26.96 (paper); Ann Fabian. The Skull Collectors: Race, Science, And America’s Unburied. Chicago: The University of Chicago Press, 2010. 270 pp., illus., ISBN: 978-0-226-23348-2. ISBN: 0-226-23348-0. $19.99 (cloth). Historical Studies in the Natural Sciences, 45(2):348– 356, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

                                        //www.jstor.org/stable/10.
                                        1525/hsns.2015.45.2.348.

Nestor:2013:ASP

[NR13] Herran Néstor and Xavier Roqué. An autarkic science: Physics, culture, and power in Franco’s Spain. Historical Studies in the Natural Sciences, 43 (2):202–235, April 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.2.202.pdf.

Nierenberg:2010:ECC

[NTT10] Nicolas Nierenberg, Walter R. Tschinkel, and Victoria J. Tschinkel. Early climate change consensus at the National Academy: The origins and making of changing climate. Historical Studies in the Natural Sciences, 40(3):318–349, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Nummedal:2007:AAH

[Num07] Tara E. Nummedal. Alchemy and authority in the Holy Roman Empire. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-60856-5 (cloth). xvii + 260 pp. LCCN QD13 .N86 2007. URL http:

//www.loc.gov/catdir/
enhancements/fy0827/
2007002521-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007002521-d.html;   http:
//www.loc.gov/catdir/toc/
ecip079/2007002521.html.

Nuriddin:2021:BPH

[Nur21] Ayah Nuriddin. Black public health. Historical Studies in the Natural Sciences, 51(1):151–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/51/1/151/116287/
Black-Public-Health.

Nyhart:2017:ISI

[NV17] Lynn K. Nyhart and Florence Vienne. Introduction to special issue: Revolutionary politics and biological organization in Nineteenth-Century France and Germany. Historical Studies in the Natural Sciences, 47(5):589– 601, November 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Nadim:2024:LNT

[NVHB24] Tahani Nadim, Mareike Vennen, Ina Heumann, and Filippo Bertoni. Logistical natures: Trade, traffics, and transformations in natural history collecting. Historical Studies in the Natural Sciences, 54(2):125– ??, April 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/54/2/125/200666/
                                        Logistical-NaturesTrade-
                                        Traffics-and.

Nye:1974:GLB

[Nye74] Mary Jo Nye. Gustave LeBon’s black light: a study in physics and philosophy in France at the turn of the century. Historical Studies in the Physical Sciences, 4(??):163–195, ???? 1974. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757330.

Nye:1980:REH

[Nye80] Mary Jo Nye. N-rays: An episode in the history and psychology of science. Historical Studies in the Physical Sciences, 11(1):125–156, ???? 1980. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757473.

Nye:1983:RRS

[Nye83] Mary Jo Nye. Review: Recent sources and problems in the history of French science. Historical Studies in the Physical Sciences, 13(2):401–415, ???? 1983. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757521.

nye:2007:HSS

[nye07] Mary Jo nye. Historical sources of science-as-social-practice: Michael Polanyi’s Berlin. Historical Studies in the Physical and Biological Sciences, 37(2): 409–434, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Nye:2009:BRS

[Nye09] Mary Jo Nye. Book reviews:

Scientific Families: Biographies and ‘Labographies’ in the History of Science: Deborah R. Coen. Vienna in the Age of Uncertainty: Science, Liberalism, and Private Life. Chicago: University of Chicago Press, 2007. xi + 380 pp., illus., index. ISBN 978-0-226111-72-8. $45 (cloth). Michael Hoskin. The Herschels of Hanover. Cambridge: Science History Publications, 2007. x + 182 pp., illus., index. ISBN 978-0-905193-07-6. $52 (cloth). Laura Otis. Müller’s Lab. Oxford: Oxford University Press, 2007. xix + 316 pp., illus., index. ISBN 978-0-195306-97-2. $55 (cloth). Historical Studies in the Natural Sciences, 39(1): 104–114, February 2009. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Nye:2011:MPH

[Nye11] Mary Jo Nye. Michael Polanyi and his generation: origins of the social construction of science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-61063-2 (hardcover). xxi + 405 pp. LCCN B945.P584 N94 2011.

Nye:2012:TKC

[Nye12] Mary Jo Nye. Thomas Kuhn, case histories, and revolutions. Historical Studies in the Natural Sciences, 42(5):557–561, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.557.pdf.

Nye:2017:A

[Nye17] Mary Jo Nye. Afterword. Historical Studies in the Natural Sciences, 47(3):443–449, June 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Nyhart:2009:MNR

[Nyh09] Lynn K. Nyhart. Modern nature: the rise of the biological perspective in Germany. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-61089-6 (cloth). xiv + 423 pp. LCCN QH540.83.G3 N94

                                        2009. URL http://www.loc.
                                        gov/catdir/toc/ecip0822/
                                        2008026875.html.

Nyhart:2017:POC

[Nyh17] Lynn K. Nyhart. The political organism: Carl Vogt on animals and states in the 1840s and ’50s. Historical Studies in the Natural Sciences, 47(5):602– 628, November 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

OConnor:2007:ESF

[O’C07] Ralph O’Connor. The Earth on show: fossils and the poetics of popular science, 1802– 1856. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-61668-1. xiii + 541 + 6 pp. LCCN QE13.G7 O35

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0732/2007028061-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0827/2007028061-b.html;
 http://www.loc.gov/
catdir/toc/ecip0722/
2007028061.html.

Oreskes:2010:MDH

[OC10] Naomi Oreskes and Erik M.

Conway. Merchants of doubt: how a handful of scientists obscured the truth on issues from tobacco smoke to global warming. Bloomsbury Press, New York, NY, USA, 2010. ISBN 1-59691-610-9. 355 pp. LCCN Q147 .O74 2010.

Oreskes:2008:CLD

[OCS08] Naomi Oreskes, Erik M. Conway, and Matthew Shindell. From Chicken Little to Dr. Pangloss: William Nierenberg, global warming, and the social deconstruction of scientific knowledge. Historical Studies in the Natural Sciences, 38(1):109– 152, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Onaga:2023:MAM

[OD23]  Lisa  On-

aga and Laurence Douny. Making animal materials in time. Historical Studies in the Natural Sciences, 53(3):197–??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        53/3/197/196654/Making-
                                        Animal-Materials-in-Time.

Ogilvie:2006:SDN

[Ogi06] Brian W. Ogilvie. The science of describing: natural history in Renaissance Europe. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-62087-5 (cloth). xvi + 385 pp. LCCN Q127.E8 O45 2006. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2005019254-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2005019254-d.html;   http:
                                        //www.loc.gov/catdir/toc/
                                        ecip0515/2005019254.html.

Oesper:1950:HSC

[OL50] Ralph E. Oesper and Pierre Lemay. Henri Sainte-Claire Deville, 1818–1881. Chymia: Annual Studies in the History of Chemistry, 3(??):205–221, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757153.

Olivotto:2009:GSC

[Oli09] Cristina Olivotto. The G-Stack Collaboration (1954): An experiment of transition. Historical Studies in the Natural Sciences, 39(1):63–103, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Oliver:2013:TFG

[Oli13] Kendrick Oliver. To touch the face of God: the sacred, the profane and the American space program, 1957–1975. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0788-4 (hardcover), 1-4214-0834-1 (electronic). xiii + 229 pp. LCCN TL789.8.U5 O39 2013.

Oliver:2019:LST

[Oli19] Kendrick Oliver. “The lucky start toward today’s cosmology”? Serendipity, the “Big Bang” theory, and the science of radio noise in Cold War America. Historical Studies in the Natural Sciences, 49(2):151– 193, April 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Olson:2008:SSN

[Ols08] Richard Olson. Science and scientism in Nineteenth-Century Europe. University of Illinois Press, Urbana, IL, USA, 2008. ISBN 0-252-03188-1 (cloth), 0-252-07433-5 (paperback). 349 pp. LCCN Q175.52.E85 O47

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2007005146-d.html;

 http://www.loc.gov/
catdir/toc/ecip0710/
2007005146.html.

Ottaviani:2013:F

[OM13] Jim Ottaviani and Leland Myrick. Feynman. First Second, New York, NY, USA, 2013. ISBN 1-59643-827-4 (paperback). 266 pp. LCCN ????

Onion:2015:GHB

[Oni15] Rebecca Onion. Graphic histories: Book reviews: Eugene Byrne and Simon Gurr. Darwin: A Graphic Biography. Washington, DC: Smithsonian Books, 2013. 96 pp., illus. ISBN 978-1-58834-352-9. $9.95 (paper). Apostolos Doxiadis and Christos H. Papadimitriou. Logicomix: An Epic Search for Truth. New York: Bloomsbury, 2009. 347 pp., illus. ISBN 978-1-59691-452-0. $22.95 (paper). Jonathan Fetter-Vorm. Trinity: A Graphic History of the First Atomic Bomb. New York: Hill and Wang, 2012. 154 pp., illus. ISBN 978-0-8090-9355-7. $14.95 (paper). Jim Ottaviani and Leyland Myrick. Feynman. New York: First Second, 2011. 272 pp., illus. ISBN 978-1-59643-827-9. $19.99 (paper). Lauren Redniss. Radioactive: Marie & Pierre Curie: A Tale of Love and Fallout. New York: Harper Collins, 2010. 208 pp., illus. ISBN 978-0-06135-132-7. $29.99 (hardcover). Historical Studies in the Natural Sciences, 45(4): 621–629, September 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.4.621.

Oreskes:1988:RCD

[Ore88] Naomi Oreskes. The rejection of continental drift. Historical Studies in the Physical and Biological Sciences, 18(2):311– 348, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757605.

Oreskes:2000:LTM

[Ore00] Naomi Oreskes. Laissez-tomber:

Military patronage and women’s work in mid-20th-Century oceanography. Historical Studies in the Physical and Biological Sciences, 30(2):373–392, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Orlando:1998:PJP

[Orl98] Lucida Orlando. Physics in the 1930s: Jewish physicists’ contributions to the realization of the ‘new tasks’ of physics in Italy. Historical Studies in the Physical and Biological Sciences, 29 (1):141–182, ???? 1998. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Orlando:2004:SSB

[Orl04] Lucia Orlando. The SIRIO satellite, 1968–1977: Between scientific engagement and managerial inexperience. Historical Studies in the Physical and Biological Sciences, 34(2):371– 398, March 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Otis:2007:ML

[Oti07] Laura Otis. Müller’s lab. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-530697-X. xix + 316 pp. LCCN QM16.M85

O85 2007. URL http://www.
loc.gov/catdir/
enhancements/fy0737/
2006026796-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0619/2006026796.html.

Otter:2008:VEP

[Ott08] Chris Otter. The Victorian eye: a political history of light and vision in Britain, 1800– 1910. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-64076-0 (cloth), 0-226-64077-9 (paperback). x + 382 pp. LCCN TH7703 .O88

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0805/2007043184-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0827/2007043184-b.html;
 http://www.loc.gov/
catdir/toc/ecip082/
2007043184.html.

Poster:2011:UTI

[PA11] Mark Poster and Nancy Ann Roth. Into the Universe of Technical Images. Electronic Mediations. University of Minnesota Press, Minneapolis, MN, USA, 2011. ISBN 0-8166-7696-8, 0-8166-7020-X. 225 pp. LCCN TR183.F5813 2011.

URL http://www.SLQ.eblib.
com.au/patron/FullRecord.
aspx?p=673646.

Palter:1989:RSI

[Pal89] Robert Palter. Review: Some impressions of recent work on Eighteenth-Century science. Historical Studies in the Physical and Biological Sciences, 19 (2):349–401, ???? 1989. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757628.

Pancaldi:1990:ELV

[Pan90a] Giuliano Pancaldi. Electricity and life. Volta’s path to the battery. Historical Studies in the Physical and Biological Sciences, 21(1):123–160, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757657.

Pang:1990:EBR

[Pan90b] Alex Soojung-Kim Pang. Edward Bowles and radio engineering at MIT, 1920–1940. Historical Studies in the Physical and Biological Sciences, 20(2): 313–337, ???? 1990. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757646.

Pang:1993:SB

[Pan93] Alex Soojung-Kim Pang. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 24(1):189–??, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Pang:1994:SBa

[Pan94a] Alex Soojung-Kim Pang. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 24(2):407–??, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Pang:1994:SBb

[Pan94b] Alex Soojung-Kim Pang. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 25(1):177–??, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Pang:1995:SB

[Pan95] Alex Soojung-Kim Pang. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 25(2):379–??, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Pang:1997:VRP

[Pan97] Alex Soojung-Kim Pang. Visual representation and postconstructivist history of science. Historical Studies in the Physical and Biological Sciences, 28 (1):139–171, ???? 1997. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Panofsky:2014:MSC

[Pan14] Aaron Panofsky. Misbehaving science: controversy and the development of behavior genetics. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-05831-X (hardcover), 0-226-05845-X (paperback), 0-226-05859-X (e-book). xi + 321 pp. LCCN QH457 .P36 2014.

Pancaldi:2020:RSL

[Pan20] Giuliano Pancaldi. Reframing the sciences of the long Eighteenth Century. Historical Studies in the Natural Sciences, 50 (1–2):58–66, April 2020. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Pannhorst:2024:BVI

[Pan24] Kerstin Pannhorst. Becoming visible: Insects and infrastructures in Colonial Taiwan. Historical Studies in the Natural Sciences, 54(2):187–??, April 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/2/187/200667/Becoming-
VisibleInsects-and-
Infrastructures-in.

Partington:1948:CSC

[Par48] J. R. Partington. The concepts of substance and chemical element. Chymia: Annual Studies in the History of Chemistry, 1 (??):109–121, ???? 1948. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757118.

Partington:1959:BAT

[Par59] J. R. Partington. Berthollet and the Antiphlogistic Theory. Chymia: Annual Studies in the History of Chemistry, 5 (??):130–137, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757180.

Partington:1960:JBL

[Par60] J. R. Partington. Joseph Black’s “Lectures on the Elements of Chemistry”. Chymia: Annual Studies in the History of Chemistry, 6(??):27–67, ???? 1960. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757192.

Paris:2001:LRF

[Par01] Elizabeth Paris. Lords of the ring: The fight to build the first U.S. electron–positron collider. Historical Studies in the Physical and Biological Sciences, 31 (2):355–380, March 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Parthasarathy:2007:BGM

[Par07] Shobita Parthasarathy. Building genetic medicine: breast cancer, technology, and the comparative politics of health care. Inside technology. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-16242-3 (hardcover). x + 271 pp. LCCN RC268.44.B73

P37 2007. URL http://www.
loc.gov/catdir/toc/
ecip0619/2006027956.html.

Park:2009:BAM

[Par09] Buhm Soon Park. Between accuracy and manageability: Computational imperatives in quantum chemistry. Historical Studies in the Natural Sciences, 39(1):32–62, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Park:2010:LAC

[Par10] Hyung Wook Park. Longevity, aging, and caloric restriction: Clive Maine McCay and the construction of a multidisciplinary research program. Historical Studies in the Natural Sciences, 40(1):79–124, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Parmar:2012:FAC

[Par12]  Inderjeet Parmar.  Founda-

tions of the American century: the Ford, Carnegie, and Rockefeller Foundations in the rise of American power. Columbia University Press, New York, NY, USA, 2012. ISBN 0-231-14628-0 (hardcover), 0-231-51793-9 (ebook). ix + 356 pp. LCCN HV97.F62 P37 2012.

Parthasarathy:2022:HET

[Par22] Shobita Parthasarathy. How to be an epistemic trespasser. Historical Studies in the Natural Sciences, 52(1):140–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        52/1/140/120493/How-to-Be-
                                        an-Epistemic-Trespasser.

Paul:2011:OHN

[Pau11] Annie Murphy Paul. Origins: how the nine months before birth shape the rest of our lives. Free Press, New York, NY, USA, 2011. ISBN 0-7432-9663-X. xii + 306 pp. LCCN RG560 .P38 2011.

Pearl:2021:M

[Pea21] Sharrona Pearl. The mask. Historical Studies in the Natural Sciences, 51(1):155–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        51/1/155/116286/The-Mask.

Perillan:2018:QNP

[Per18] José G. Perillán. Quantum narratives and the power of rhetorical omission: An early history of the pilot wave interpretation of quantum theory. Historical Studies in the Natural Sciences, 48(1):24–55, February 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Perez:2025:CLS

[Per25] Myrna Perez. The colonial life of sociobiology. Historical Studies in the Natural Sciences, 55(1): 74–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/55/1/74/
207544/The-Colonial-Life-
of-Sociobiology.

Petit:2016:APC

[Pet16] Axel Petit. Associating physics and chemistry to dissociate molecules: The history of the Clausius–Williamson Hypothesis. Historical Studies in the Natural Sciences, 46(3):360– 391, June 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Phillips:1966:LKC

[Phi66] J. P. Phillips. Liebig and Kolbe, critical editors. Chymia: Annual Studies in the History of Chemistry, 11(??):89–97, ???? 1966. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757261.

Phillips:2010:BRR

[Phi10]  Denise Phillips.  Book re-

views: Reconsidering the Sonderweg of German Science: Biology and Culture in the Nineteenth Century: Sander Gliboff. H. G. Bronn, Ernst Haeckel, and the Origins of German Darwinism: A Study in Translation and Transformation. Cambridge: MIT Press, 2008. xii + 259 pp., index. ISBN: 978-0-262-07293-9. $35.00 (cloth). Jonathan Harwood. Technology’s Dilemma: Agricultural Colleges between Science and Practice in Germany, 1860– 1934. Frankfurt a. M.: Peter Lang, 2005. 288 pp., illus., index. ISBN: 978-3-039-10299-0. $68.95 (paper). Lynn K. Nyhart. Modern Nature: The Rise of the Biological Perspective in Germany. Chicago: University of Chicago Press, 2009. xiv + 423 pp., illus., index. ISBN: 978-0-226-61089-4. $45.00 (cloth). Richard G. Olson. Science and Scientism in Nineteenth-Century Europe. Urbana: University of Illinois Press, 2008. 349 pp., index. ISBN: 978-0-252-07433-2. $27 (paper). Robert J. Richards. The Tragic Sense of Life: Ernst Haeckel and the Struggle over Evolutionary Thought. Chicago: University of Chicago Press, 2008. xx + 551 pp., illus., index. ISBN: 978-0-226-71216-1. $25.00 (paper). Nicolaas A. Rupke. Alexander von Humboldt: A Metabiography. Chicago: University of Chicago Press, 2008. 316 pp., illus., index. ISBN: 978-0-226-73149-0. $21.00 (paper). Historical Studies in the Natural Sciences, 40(1):136–147, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Phillips:2021:C

[Phi21] Christopher J. Phillips. The curve. Historical Studies in the Natural Sciences, 51(1): 159–??, February 2021. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
51/1/159/116288/The-Curve.

Picciotto:2014:BRN

[Pic14] Joanna Picciotto. Book review:

The Natural Philosophy of Margaret Cavendish: Reason and Fancy during the Scientific Revolution Fictions of the Cosmos: Science and Literature in the Seventeenth Century. Historical Studies in the Natural Sciences, 44(2):178–186, April 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2014.
44.2.178.

PohlValero:2011:ECH

[Poh11] Stefan Pohl Valero. Energía y cultura: historia de la termodinámica en la España de la segunda mitad del siglo XIX. Opera eximia. Editorial Pontificia Universidad Javeriana, Bogotá, D.C., Colombia, 2011. ISBN 958-716-498-9. 321 pp. LCCN QC311 .P74 2011.

Porter:1981:SSG

[Por81] Theodore M. Porter. A statistical survey of gases: Maxwell’s social physics. Historical Studies in the Physical Sciences, 12 (1):77–116, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757490.

Portuondo:2013:BRM

[Por13] María M. Portuondo. Book reviews: Mutis: Plants, ants, and visualizing nature: Daniela Bleichmar. Visible Empire: Botanical Expeditions and Visual Culture in the Hispanic Enlightenment. Chicago: University of Chicago Press, 2012. xii + 286 pp., illus., index. ISBN: 978-0-2260-5853-5. $55.00 (hardcover); J. E. Bernal and Alberto Gómez Gutiérrez. A impulsos de una rara resolución: El viaje de José Celestino Mutis al Nuevo Reino de Granada, 1760– 1763. Bogotá: Pontificia Universidad Javeriana, 2010. 333 pp., illus. ISBN: 978-9-5871-6326-1. $70.00 (hardcover); J. E. Bernal, Alberto Gómez Gutiérrez, and José Celestino Mutis. Medicina científica Mutisiana: Estudio elaborado a partir de un documento médico inédito de José Celestino Mutis y Bosio, escrito en Madrid en 1759 y conservado hoy en el Archivo Histórico Javeriano de Bogotá. Bogotá: Pontificia Universidad Javeriana, 2008. 144 pp., illus., facs. ISBN: 978-9-5871-6135-9. $50.00 (paper); Alberto Gómez Gutiérrez, Jorge Tomás Uribe Angel, Pedro Ortiz Valdivieso, and J. E. Bernal. Academia Mutisiana: Documentos preneogranadinos de José Celestino Mutis y la promoción de sociedades científicas en la Nueva Granada. Bogotá: Pontificia Universidad Javieriana and Academia Nacional de Medicina, 2011. 236 pp., illus., facs. ISBN: 978-9-5871-6490-9. $76.00 (paper); José Celestino Mutis, Pedro Ortiz Valdivieso, J. E. Bernal, and Alberto Gómez Gutiérrez. Filosofía natural Mutisiana: Edición facsimilar y estudio de un manuscrito inédito de José Celestino Mutis y Bosio, correspondiente a su tercera oración inaugural pronunciada en el Colegio Mayor de Nuestra Señora del Rosario, Santa Fé, Octubre de 1764. Bogotá: Pontificia Universidad Javeriana, 2009. 249 pp., illus,. facs. ISBN: 9-789-5871-6263-9. $70.00 (paper); Edward O. Wilson and José María Gómez Durán. Kingdom of Ants: José Celestino Mutis and the Dawn of Natural History in the New World. Baltimore, MD: Johns Hopkins University Press, 2010. 120 pp., illus., index. ISBN: 9-780-8018-9785-6. $25.95 (hardcover). Historical Studies in the Natural Sciences, 43(3): 381–389, June 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.3.381.pdf.

Porter:2018:OND

[Por18] Theodore M. Porter. Observations, narrative, and data in Nineteenth-Century asylum medicine. Historical Studies in the Natural Sciences, 48(5):594– 603, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Potters:2023:HSP

[Pot23] Jan Potters. Hazy spots on photographic plates: On the measurement of the velocitydependency of the electron’s mass. Historical Studies in the Natural Sciences, 53(1):1– ??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

                                        https://online.ucpress.
                                        edu/hsns/article/53/1/1/
                                        195540/Hazy-Spots-on-
                                        Photographic-PlatesOn-the.

Prandtl:1949:CLB

[Pra49] Wilhelm Prandtl. Das chemische Laboratorium der Bayerischen Akademie der Wissenschaften in München. (German) [The Chemical Laboratory of the Bayern Academy of Science in Munich]. Chymia: Annual Studies in the History of Chemistry, 2(??):81–97, ???? 1949. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757136.

Prandtl:1953:VMP

[Pra53] Wilhelm Prandtl. Zur Vorgeschichte des Meissner Porzellans. (German) [On the prehistory of Meissen porcelain]. Chymia: Annual Studies in the History of Chemistry, 4(??):115–127, ???? 1953. CODEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757165.

Prescott:2011:MAH

[Pre11] Heather Munro Prescott. The morning after: a history of emergency contraception in the United States. Critical issues in health and medicine. Rutgers University Press, New Brunswick, NJ, USA, 2011. ISBN 0-8135-5162-5 (hardcover), 0-8135-5163-3 (paperback). xi + 163 pp. LCCN RG137.5 .P74 2011.

Priestley:2012:MRN

[Pri12] Rebecca Priestley. Mad on radium: New Zealand in the atomic age. Auckland University Press, Auckland, NZ, 2012. ISBN 1-86940-727-X (paperback). xii + 284 pp. LCCN TK9122.5 .P75 2012.

Priest:2014:HPG

[Pri14] Tyler Priest. Hubbert’s Peak:

The great debate over the end of oil. Historical Studies in the Natural Sciences, 44(1):37–79, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2014.44.1.37.

Prelat:1950:QME

[PVP50] Carlos E. Prelat, Alberto G.

Velarde, and Carlos E. Prélat. La química en los “éléments de Chimie” de Orfila. (Spanish)

[Chemistry and the Orfila’s Elements of Chemistry]. Chymia: Annual Studies in the History of Chemistry, 3(??):77–93, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757147.

Pyenson:1974:MCC

[Pye74] Lewis Pyenson. In memoriam: Charles A. Culotta (1938– 1974). Historical Studies in the Physical Sciences, 4(??): 234–235, ???? 1974. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757332.

Pyenson:1976:EES

[Pye76] Lewis Pyenson. Einstein’s early scientific collaboration. Historical Studies in the Physical Sciences, 7(??):iv + 83–123, ???? 1976. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757352.

Pyenson:1982:AEE

[Pye82] Lewis Pyenson. Audacious enterprise: The Einsteins and electrotechnology in late Nineteenth-Century Munich. Historical Studies in the Physical Sciences, 12(2): 373–392, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757500.

Pyenson:1990:HMG

[Pye90a] Lewis Pyenson. Habits of mind:

Geophysics at Shanghai and Algiers, 1920–1940. Historical Studies in the Physical and Biological Sciences, 21(1):161– 196, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757658.

Pyenson:1990:RBM

[Pye90b] Lewis Pyenson. Review: Over the bounding main. Historical Studies in the Physical and Biological Sciences, 20(2):407– 422, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757648.

Pyenson:2007:EF

[Pye07] Lewis Pyenson. Editor’s foreword. Historical Studies in the Physical and Biological Sciences, 37(2):189–204, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Pyenson:2011:EIN

[Pye11] Lewis Pyenson. The enlightened image of nature in the Dutch East Indies: Consequences of postmodernist doctrine for broad structures and intimate life. Historical Studies in the Natural Sciences, 41 (1):1–40, Winter 2011. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2011.
                                        41.1.1.

Pyenson:2013:EPQ

[Pye13] Lewis Pyenson. The Einstein– Picasso question: Neo-idealist abstraction in the decorative arts and manufactures. Historical Studies in the Natural Sciences, 43 (3):281–333, June 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2013.43.3.281.pdf.

Quedec:1988:WMS

[Qué88] Pierre Quédec. Weiss’ magneton: The sin of pride or a venial mistake? Historical Studies in the Physical and Biological Sciences, 18(2):349– 375, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757606.

Quinn:1982:RFE

[Qui82] Arthur Quinn. Repulsive force in England, 1706–1744. Historical Studies in the Physical Sciences, 13(1):109–128, ???? 1982. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

                                        jstor.org/stable/27757507.

Radick:2012:BRE

[Rad12] Gregory Radick. Book review: The exemplary Kuhnian:

Gould’s structure revisited: The Structure of Evolutionary Theory by Stephen Jay Gould. Historical Studies in the Natural Sciences, 42(2):143–157, April 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2012.42.2.143.

Rader:2013:BRS

[Rad13] Karen A. Rader. Book review: Searching for Women Scientists and Finding STS, Margaret Rossiter. Women Scientists in America: Struggles and Strategies to 1940. Baltimore: Johns Hopkins University Press, 1984. xviii + 439 pp., illus., index. ISBN 9-780-8018-2509-5. $36.00 (paper); Margaret Rossiter. Women Scientists in America: Before Affirmative Action, 1940–1972. Baltimore: Johns Hopkins University Press, 1998. xviii + 584 pp., illus., index. ISBN 9-780-8018-5711-9. $36.00 (paper); Margaret Rossiter. Women Scientists in America: Forging a New World since 1972. Baltimore: Johns Hopkins University Press, 2012. xx + 488 pp., illus., index. ISBN 9-781-4214-0363-2. $45.00 (paper). Historical Studies in the Natural Sciences, 43 (3):371–380, June 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2013.43.3.371.pdf.

Rainger:2000:SCN

[Rai00] Ronald Rainger. Science at the crossroads: The Navy, Bikini Atoll, and American oceanography in the 1940s. Historical Studies in the Physical and Biological Sciences, 30(2):349–371, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Rainger:2010:BRN

[Rai10] Ronald Rainger. Book review: New Bones to Pick: Stanley Hedeen. Big Bone Lick: The Cradle of American Paleontology. Lexington: University of Kentucky Press, 2008. xviii + 182 pp., illus., index. ISBN 978-0-8131-2485-8. $24.95 (cloth). Adrienne Mayor. Fossil Legends of the First Americans. Princeton: Princeton University Press, 2005. xxxix + 442 pp., illus., index. ISBN 978-0-691-11345-6. $27.95 (paper). Pascal Richet. A Natural History of Time. Chicago: University of Chicago Press, 2007. xiv + 471 pp., illus., index. ISBN 978-0-226-71287-1. $29 (cloth). Martin J. S. Rudwick. Worlds Before Adam: The Reconstruction of Geohistory in the Age of Reform. Trans. John Venerella. Chicago: University of Chicago Press, 2008. xxii + 614 pp., illus., index. ISBN 978-0-226-73128-5. $49 (cloth). Ralph O’Connor. The Earth on Show: Fossils and the Poetics of Popular Science, 1802–1856. Chicago: University of Chicago Press, 2008. xiii + 542 pp., illus., index. ISBN 978-0-2266-1668-1. $45 (cloth). Historical Studies in the Natural Sciences, 40(2):268–278, May 2010. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Ramberg:1995:AMC

[Ram95] Peter J. Ramberg. Arthur Michael’s critique of stereochemistry, 1887–1899. Historical Studies in the Physical and Biological Sciences, 26 (1):89–138, ???? 1995. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Rankin:2022:HEE

[Ran22] Alisha Rankin. How to “Be Expert” in Early Modern Europe. Historical Studies in the Natural Sciences, 52(1):143–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/52/1/143/120496/
How-to-Be-Expert-in-Early-
Modern-Europe.

Rappaport:1960:GFR

[Rap60]  Rhoda  Rappaport.

G.-F. Rouelle: An Eighteenth-Century chemist and teacher. Chymia: Annual Studies in the History of Chemistry, 6(??):68– 101, ???? 1960. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757193.

Rappaport:1961:RSP

[Rap61] Rhoda Rappaport. Rouelle and Stahl — the Phlogistic Revolution in France. Chymia: Annual Studies in the History of Chemistry, 7(??):73–102, ???? 1961. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757206.

Rasmussen:1993:FAR

[Ras93] Nicolas Rasmussen. Freund’s adjuvant and the realization of questions in postwar immunology. Historical Studies in the Physical and Biological Sciences, 23(2):337–366, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757702.

Ray:1959:CK

[Rây59] Priyadaranjan Rây. Chemistry in Kautilya. Chymia: Annual Studies in the History of Chemistry, 5(??):64–72, ???? 1959. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757177.

Read:1948:SAS

[Rea48] John Read. Scottish alchemy in the Seventeenth Century. Chymia: Annual Studies in the History of Chemistry, 1(??): 139–151, ???? 1948. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757120.

Redniss:2010:RMP

[Red10] Lauren Redniss. Radioactive:

Marie and Pierre Curie, a tale of love and fallout. !t Books/Harper Collins, New York, NY, USA, 2010. ISBN 0-06-135132-6. 205 pp. LCCN QD22.C8 R395

2010. URL http://www.loc.
gov/catdir/enhancements/
fy1111/2011281920-b.html;
 http://www.loc.gov/
catdir/enhancements/
fy1111/2011281920-d.html.

Rees:2009:ICP

[Ree09] Amanda Rees. The infanticide controversy: primatology and the art of field science. University of Chicago Press, Chicago, IL, USA, 2009. ISBN 0-226-70711-3 (hardcover). 288 pp. LCCN QL762.5 .R44 2009.

Reilly:1949:IAC

[Rei49] Desmond Reilly. An Irish– American chemist, William James Mac-Neven, 1763–1841. Chymia: Annual Studies in the History of Chemistry, 2(??):17–26, ???? 1949. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757129.

Reilly:1953:CMS

[Rei53] Desmond Reilly. Contributions of Maxwell Simpson (1815– 1902) to aliphatic chemical synthesis. Chymia: Annual Studies in the History of Chemistry, 4 (??):159–170, ???? 1953. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757167.

Reingold:1987:VBN

[Rei87] Nathan Reingold. Vannevar Bush’s new deal for research: Or the triumph of the old order. Historical Studies in the Physical and Biological Sciences, 17 (2):299–344, ???? 1987. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757585.

Reingold:1995:CFU

[Rei95] Nathan Reingold. Choosing the future: The U.S. research community, 1944–1946. Historical Studies in the Physical and Biological Sciences, 25(2):301–328, ???? 1995. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Reidy:2017:MRO

[Rei17a] Michael S. Reidy. The most recent orogeny: Verticality and why mountains matter. Historical Studies in the Natural Sciences, 47(4):578–587, September 2017. ISSN 1939-1811 (print), 1939-182X (electronic).

Reinhardt:2017:OMD

[Rei17b]  Carsten Reinhardt.  “This

other method”: The dynamics of NMR in biochemistry and molecular biology. Historical Studies in the Natural Sciences, 47(3):389–422, June 2017. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Rentetzi:2009:TMG

[Ren09] Maria Rentetzi. Trafficking materials and gendered experimental practices: radium research in early 20th Century Vienna. Columbia University Press, New York, NY, USA, 2009. ISBN 0-231-13558-0, 0-231-50959-6. xxiii + 279 pp. LCCN QC795.34 .R46

2008. URL http://www.loc.
gov/catdir/toc/ecip0828/
2008040581.html.

Reti:1965:HOH

[Ret65] Ladislao Reti. How old is hydrochloric acid? Chymia: Annual Studies in the History of Chemistry, 10(??):11–23, ???? 1965. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757245.

Reynolds:2021:BRR

[Rey21] Andrew S. Reynolds. Book review: Re-Envisioning the History of Cellular and Molecular Biology. Historical Studies in the Natural Sciences, 51(5):642– ??, November 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/51/5/642/
119066/Re-Envisioning-the-
History-of-Cellular-and.

Reynolds:2022:HCH

[Rey22] Melissa Reynolds. How to cure a horse, or, the experience of knowledge and the knowledge of experience. Historical Studies in the Natural Sciences, 52(4):547–??, September 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/4/547/191416/How-to-
Cure-a-Horse-or-the-
Experience-of-Knowledge.

Raman:1969:WWI

[RF69] V. V. Raman and Paul Forman. Why was it Schrödinger who developed de Broglie’s ideas? Historical Studies in the Physical Sciences, 1(??):291–314, ???? 1969. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757299.

Roll-Hansen:2000:ACB

[RH00] Nils Roll-Hansen. The application of complementarity to biology: From Niels Bohr to Max Delbrück. Historical Studies in the Physical and Biological Sciences, 30(2):417–442, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Reid-Henry:2010:CCR

[RH10] Simon Reid-Henry. The Cuban cure: reason and resistance in global science. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-70917-5 (hardcover). xii + 200 pp. LCCN TP248.195.C9 R44 2010.

Rheinboldt:1950:BVA

[Rhe50]  Heinrich Rheinboldt.  Bun-

sens Vorlesung über allgemeine Experimentalchemie. (German) [Bunsen’s lectures on general experimental chemistry]. Chymia: Annual Studies in the History of Chemistry, 3(??):223–241, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757154.

Rheinberger:2010:ECT

[Rhe10a] Hans-Jörg Rheinberger. An epistemology of the concrete: twentieth-century histories of life. Experimental futures: technological lives, scientific arts, anthropological voices. Duke University Press, Durham, NC, USA, 2010. ISBN 0-8223-4560-9 (hardcover), 0-8223-4575-7 (paperback). xix + 330 pp. LCCN QH331 .R4513 2010. Foreword by Tim Lenoir.

Rheinberger:2010:HEE

[Rhe10b] Hans-Jörg Rheinberger. On historicizing epistemology: an essay. Cultural memory in the present. Stanford University Press, Stanford, CA, USA, 2010. ISBN 0-8047-6288-0 (hardcover), 0-8047-6289-9 (paperback). 114 pp. LCCN Q175.32.K45 R48513 2010. Translated by David Fernbach.

Richet:2007:NHT

[Ric07] Pascal Richet. A natural history of time. University of Chicago Press, Chicago, IL, USA, 2007.

ISBN 0-226-71287-7 (cloth). xiv + 471 pp. LCCN QE508 .R5413

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0707/2006033992-b.html;
 http://www.loc.gov/
catdir/enhancements/
fy0707/2006033992-d.html;
 http://www.loc.gov/
catdir/toc/ecip073/
2006033992.html.

Richards:2008:TSL

[Ric08] Robert J. (Robert John) Richards. The tragic sense of life: Ernst Haeckel and the struggle over evolutionary thought. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-71214-1 (hardcover), 0-226-71216-8 (paperback). xx + 551 + 8 pp. LCCN M09.E10966; M09.E02488; QH31.H2 R53 2008; QH31.H2 RIC 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0827/
2007039155-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007039155-d.html;   http:
//www.loc.gov/catdir/toc/
fy0804/2007039155.html.

Richmond:2012:BRS

[Ric12] Jesse Richmond. Book review: Still figuring out Nature’s economy: Nature’s Ghosts: Confronting Extinction from the Age of Jefferson to the Age of Ecology by Mark V. Barrow, Jr.; Evolutionary Restraints: The Contentious History of Group Selection by Mark E. Borrello; Looking for a Few Good Males: Female Choice in Evolutionary Biology by Erika Lorraine Milam; The Infanticide Controversy: Primatology and the Art of Field Science by Amanda Rees. Historical Studies in the Natural Sciences, 42(1):62–70, February 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.
1525/hsns.2012.42.1.62. See

[BJ10, Bar09, Bor10, Mil10b, Ree09].

Richardson:2017:NCN

[Ric17a] Alan W. Richardson. “Neither a confession nor an accusation”: Michael Polanyi, Hans Reichenbach, and philosophical modernity after World War One. Historical Studies in the Natural Sciences, 47(3):423–442, June 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Richmond:2017:WPS

[Ric17b] Marsha L. Richmond. Women as public scientists in the Atomic Age: Rachel Carson, Charlotte Auerbach, and genetics. Historical Studies in the Natural Sciences, 47(3):349– 388, June 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Rider:1984:AOE

[Rid84] Robin E. Rider. Alarm and opportunity: Emigration of mathematicians and physicists to Britain and the United States, 1933–1945. Historical Studies in the Physical Sciences, 15(1): 107–176, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757544.

Rider:1991:SB

[Rid91] Robin E. Rider. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 22(1):193–205, ???? 1991. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757677.

Rider:1992:SB

[Rid92] Robin Rider. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 22(2):399–404, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757687.

Rider:1993:SB

[Rid93] Robin Rider. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 23(2):367–374, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757703.

Rigden:1983:MBE

[Rig83] John S. Rigden. Molecular beam experiments on the hydrogens during the 1930s. Historical Studies in the Physical Sciences, 13(2):335–373, ???? 1983. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757519.

Riordan:2001:TTC

[Rio01] Michael Riordan. A tale of two cultures: Building the Superconducting Super Collider, 1988–1993. Historical Studies in the Physical and Biological Sciences, 32(1):125–144, September 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Riskin:1998:PRL

[Ris98] Jessica Riskin. Poor Richard’s Leyden Jar: Electricity and economy in Franklinist France. Historical Studies in the Physical and Biological Sciences, 28 (2):301–336, ???? 1998. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Riskin:2007:GRE

[Ris07] Jessica Riskin, editor. Genesis redux: essays in the history and philosophy of artificial life. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-72080-2 (cloth), 0-226-72081-0 (paperback). xv + 389 pp. LCCN BD418.8 .G46

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0707/2007003437-b.html;
http://www.loc.gov/
catdir/enhancements/

fy0707/2007003437-d.html;
 http://www.loc.gov/
catdir/toc/ecip079/
2007003437.html.

Riskin:2010:BRN

[Ris10] Jessica Riskin. Book reviews: Newton and Monotheism: Peter Dear. The Intelligibility of Nature: How Science Makes Sense of the World. Chicago: University of Chicago Press, 2007. xii + 242 pp., illus., index. ISBN 978-0-226-13949-4. $17.00 (paper). Stephen Gaukroger. The Emergence of a Scientific Culture: Science and the Shaping of Modernity 1210–1685. Oxford: Clarendon, 2006. ix + 563 pp., illus., index. ISBN 978-0-199-55001-2. $39.95 (paper). Peter Harrison. The Fall of Man and the Foundations of Science. Cambridge: Cambridge University Press, 2007. xi + 300 pp., index. ISBN 978-0-521-87559-2. $43.00 (paper). George Saliba. Islamic Science and the Making of the European Renaissance. Cambridge, MA: MIT Press, 2007. xi + 315 pp., illus., index. ISBN 978-0-262-19557-7. $43.00 (hardcover). Historical Studies in the Natural Sciences, 40(3): 399–408, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

RibeiroDeAndrade:2006:QBS

[RM06] Ana M. Ribeiro De Andrade and R. P. A. Muniz. The quest for the Brazilian synchrocyclotron. Historical Studies in the Physical and Biological Sciences, 36(2):311–327, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Rispoli:2020:SDA

[RO20] Giulia Rispoli and Doubravka Olšáková. Science and diplomacy around the Earth: From the Man and Biosphere Programme to the International Geosphere– Biosphere Programme. Historical Studies in the Natural Sciences, 50(4):456–??, September 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/4/456/111899/Science-
and-Diplomacy-around-the-
EarthFrom-the-Man.

Robinson:1959:TF

[Rob59] Trevor Robinson. Theophrastos on fire. Chymia: Annual Studies in the History of Chemistry, 5(??):51–63, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757176.

Robinson:1960:MT

[Rob60]  Trevor Robinson.  Michael

Tswett. Chymia: Annual Studies in the History of Chemistry, 6(??):146–161, ???? 1960. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757196.

Roberts:1976:ERC

[Rob76] Gerrylynn K. Roberts. The establishment of the Royal College of Chemistry: An investigation of the social context of early-Victorian chemistry. Historical Studies in the Physical Sciences, 7(??):437–485, ???? 1976. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757361.

Roberts:1980:LEC

[Rob80] Gerrylynn K. Roberts. The liberallyeducated chemist: Chemistry in the Cambridge Natural Sciences Tripos, 1851–1914. Historical Studies in the Physical Sciences, 11(1):157–183, ???? 1980. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757474.

Robotti:1983:SPH

[Rob83] Nadia Robotti. The spectrum of ζ puppis and the historical evolution of empirical data. Historical Studies in the Physical Sciences, 14(1):123–145, ???? 1983. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757527.

Robinson:2020:ETC

[Rob20]  Sam  Robin-

son. Early Twentieth-Century ocean science diplomacy: Competition and cooperation among North Sea nations. Historical Studies in the Natural Sciences, 50(4):384–??, September 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/4/384/111895/Early-
Twentieth-Century-Ocean-
Science.

Robertson:2024:DFR

[Rob24] Jennifer Robertson. Doing fieldwork in robotland: Combining Longue Durée with rapid ethnography. Historical Studies in the Natural Sciences, 54(5):630–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/5/630/203886/Doing-
Fieldwork-in-
RobotlandCombining-Longue-
Duree.

Rocke:1978:AEE

[Roc78] Alan J. Rocke. Atoms and equivalents: The early development of the chemical atomic theory. Historical Studies in the Physical Sciences, 9(??): 225–263, ???? 1978. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757379.

Rocke:2017:IHR

[Roc17] Alan J. Rocke. The “indifferent hypothesis” redux: The dilemmas of Pierre Duhem. Historical Studies in the Natural Sciences, 47(3):268–292, June 2017. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Rogaski:2012:ASB

[Rog12] Ruth Rogaski. Addicted to science: Book reviews: Susan Greenhalgh, Just One Child: Science and Policy in Deng’s China. Tong Lam, A Passion for Facts: Social Surveys and the Construction of the Chinese Nation-State, 1900–1949. Berkeley and Los Angeles: University of California Press, 2011. 280 pp. ISBN 978-0-520-26786-2. $60.00 (hardcover). Thomas Mullaney, Coming to Terms with the Nation: Ethnic Classification in Modern China. Berkeley and Los Angeles: University of California Press, 2011. 256 pp. ISBN 978-0-520-27274-3. $29.95 (paper). Sigrid Schmalzer, The People’s Peking Man: Popular Science and Human Identity in Twentieth-Century China. Chicago: University of Chicago Press, 2008. 368 pp. ISBN 978-0-226-73860-4. $26.00 (paper). Historical Studies in the Natural Sciences, 42(5):581– 589, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.581.pdf.

Rossiter:1982:WSA

[Ros82] Margaret W. Rossiter. Women scientists in America: struggles and strategies to 1940. The Johns Hopkins University Press, Baltimore, MD, USA, 1982. ISBN 0-8018-2443-5 (hardcover), 0-8018-2509-1 (paperback). xviii + 439 pp. LCCN Q130 .R683 1982.

Rossiter:1992:WSA

[Ros95] Margaret W. Rossiter. Women scientists in America. The Johns Hopkins University Press, Baltimore, MD, USA, 1992– 1995. ISBN 0-8018-4893-8 (hardcover), 0-8018-5711-2 (paperback). xviii + 584 + 16 pp. LCCN Q130 .R683 1992.

Rossiter:2012:STT

[Ros12a] Margaret W. Rossiter. A sophomore tackles Thomas Kuhn’s structure of scientific revolutions (1964). Historical Studies in the Natural Sciences, 42(5): 562–563, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.562.pdf.

Rossiter:2012:WSA

[Ros12b] Margaret W. Rossiter. Women scientists in America: forging a new world since 1972. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0233-5 (hardcover), 1-4214-0363-3 (paperback). xx + 426 pp. LCCN Q130 .R6835 2012.

Rosenberg:2018:DW

[Ros18] Daniel Rosenberg. Data as word. Historical Studies in the Natural Sciences, 48(5):557– 567, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Ross:2022:IFP

[Ros22] Andrew J. Ross. An icy feud in planetary science: Carl Sagan, Edward Teller, and the ideological roots of the nuclear winter debates, 1980–1984. Historical Studies in the Natural Sciences, 52(2):190–??, April 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/2/190/144553/An-Icy-
Feud-in-Planetary-
ScienceCarl-Sagan-Edward.

Rothschild:2013:EAA

[Rot13] Rachel Rothschild. Environmental awareness in the atomic age: Radioecologists and nuclear technology. Historical Studies in the Natural Sciences, 43(4):492–530, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2013.
43.4.492.

Rudwick:2008:WBA

[Rud08] M. J. S. Rudwick. Worlds before Adam: the reconstruction of geohistory in the age of reform. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-73128-6 (cloth). xxii + 614 pp. LCCN QE13.E85 R83

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2007041676-b.html;

http://www.loc.gov/
catdir/enhancements/
fy0827/2007041676-d.html;
http://www.loc.gov/
catdir/toc/ecip082/
2007041676.html.

Rueger:1992:ATI

[Rue92] Alexander Rueger. Attitudes towards infinities: Responses to anomalies in quantum electrodynamics, 1927–1947. Historical Studies in the Physical and Biological Sciences, 22(2):309– 337, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757684.

Ruger:1988:ACE

[Rüg88] Alexander Rüger. Atomism from cosmology: Erwin Schrödinger’s work on wave mechanics and space–time structure. Historical Studies in the Physical and Biological Sciences, 18(2):377– 401, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757607.

Rupke:2008:AHM

[Rup08] Nicolaas A. Rupke. Alexander von Humboldt: a metabiography. University of Chicago Press, Chicago, IL, USA, University of Chicago Press paperback edition, 2008. ISBN 0-226-73149-9 (paperback). 316 pp. LCCN Q143.H9 R87 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0828/
2008001122-b.html;
http://www.loc.gov/catdir/
enhancements/fy0828/
2008001122-d.html;   http:
//www.loc.gov/catdir/toc/
ecip088/2008001122.html.

Russo:1981:FRB

[Rus81]  Arturo Russo.  Fundamen-

tal research at Bell Laboratories: The discovery of electron diffraction. Historical Studies in the Physical Sciences, 12(1): 117–160, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757491.

Russo:1986:SII

[Rus86] Arturo Russo. Science and industry in Italy between the two World Wars. Historical Studies in the Physical and Biological Sciences, 16(2):281–320, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757567.

Russell:2011:EHU

[Rus11a] Edmund Russell. Evolutionary history: uniting history and biology to understand life on Earth. Studies in environment and history. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-76211-1 (hardback), 0-521-74509-8 (paperback). xxi + 216 pp. LCCN QH366.2 .R88 2011; QH366.2 .R88X 2011 (LC).

Russo:2011:EPM

[Rus11b] Arturo Russo. Europe’s path to Mars: The European Space Agency’s Mars Express Mission. Historical Studies in the Natural Sciences, 41(2):123–178, Spring 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.2.123.

Ruswick:2013:AWP

[Rus13] Brent Ruswick. Almost worthy: the poor, paupers, and the science of charity in America, 1877–1917. Philanthropic and nonprofit studies. Indiana University Press, Bloomington, IN, USA, 2013. ISBN 0-253-00634-1 (hardcover), 0-253-00638-4 (ebook). xi + 267 pp. LCCN HV91 .R87 2013.

Ruskin:2015:BNH

[Rus15] Steve Ruskin. The business of natural history: Charles Aiken, Colorado ornithology, and the role of the professional collector. Historical Studies in the Natural Sciences, 45(3):357–396, June 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.3.357.

Ruse:2016:RGM

[Rus16] Michael Ruse. Review: The Greatest Man Since the Deluge? Historical Studies in the Natural Sciences, 46(4):528– 537, September 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Sakurai:2024:LBE

[Sak24] Ayako Sakurai. Listing butterflies: Economic and epistemic logistics of commodification, 1880s–1910s. Historical Studies in the Natural Sciences, 54(2):216–??, April 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/2/216/200669/Listing-
ButterfliesEconomic-and-
Epistemic.

Saliba:2007:ISM

[Sal07] George Saliba. Islamic science and the making of the European Renaissance. Transformations. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-19557-7 (hardcover). xi + 315 pp. LCCN Q127.I742 S35

2007. URL http://www.loc.
gov/catdir/toc/ecip0617/
2006023618.html.

Samudzi:2023:HD

[Sam23] Zoé Samudzi. Haunted by denial. Historical Studies in the Natural Sciences, 53(1):94–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/53/1/94/195530/
Haunted-by-Denial.

Santesmases:2005:SOM

[San05]  María  Jesús

Santesmases. Severo Ochoa: de músculos a proteínas. (Spanish) [Severo Ochoa: from muscles to proteins]. Memorias y biografías. Síntesis, Madrid, Spain, 2005. ISBN 84-9756-311-5. 398 pp. LCCN QP511.8.O24 S26 2005.

Santesmases:2010:BRL

[San10] María Jesús Santesmases. Book reviews: Life and Death in the Atomic Era: Jacob D. Hamblin. Poison in the Well: Radioactive Waste in the Oceans at the Dawn of the Nuclear Age. New Brunswick: Rutgers University Press, 2008. x + 311 pp., index. ISBN: 978-0-813-54220-1. $49.95 (cloth). Néstor Herran. Aguas, semillas y radiaciones: El Laboratorio de Radiactividad de la Universidad de Madrid. Madrid: Consejo Superior de Investigaciones Científicas, 2008. 271 pp., illus., index. ISBN: 978-8-400-08692-3. $29.95 (cloth). Gerald Kutcher. Contested Medicine: Cancer Research and the Military. Chicago: University of Chicago Press, 2009. x + 247 pp., index. ISBN: 978-0-226-46531-9. $35.00 (cloth). Ellen Leopold. Under the Radar: Cancer and the Cold War. New Brunswick, NJ: Rutgers University Press, 2009. xii + 284 pp., illus., index. ISBN: 978-0-813-54404-5. $35.00 (cloth). Maria Rentetzi. Trafficking Materials and Gendered Experimental Practices: Radium Research in Early 20th Century Vienna. New York: Columbia University Press, 2008. 316 pp. ISBN: 978-0-231-13558-0. $60.00 (cloth). Online edition: http://www.gutenberge.org/rentetzi/ includes images, maps, and index. Historical Studies in the Natural Sciences, 40(3):409–418, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Santesmases:2015:HCC

[San15] María Jesús Santesmases. Human chromosomes and cancer: Tumors and the geographies of cytogenetic practices, 1951– 1956. Historical Studies in the Natural Sciences, 45(1):85–114, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

                                        //www.jstor.org/stable/10.
                                        1525/hsns.2015.45.1.85.

Sarton:1950:BBS

[Sar50] George Sarton. Boyle and Bayle the sceptical chemist and the sceptical historian. Chymia: Annual Studies in the History of Chemistry, 3(??):155–189, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757151.

Saraiva:2010:FLG

[Sar10a] Tiago Saraiva. Fascist labscapes: Geneticists, wheat, and the landscapes of fascism in Italy and Portugal. Historical Studies in the Natural Sciences, 40(4):457–498, Fall 2010. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2010.
40.4.457.

Sarasohn:2010:NPM

[Sar10b] Lisa T. Sarasohn. The natural philosophy of Margaret Cavendish: reason and fancy during the scientific revolution, volume 2 (128th series) of The Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9443-3 (hardcover). xi + 251 pp. LCCN Q130 .S266 2010.

Sepkoski:2025:SPO

[SB25] David Sepkoski and Mark Borrello. Sociobiology and the politics of objectivity. Historical Studies in the Natural Sciences, 55(1):44–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/44/207535/
Sociobiology-and-the-
Politics-of-Objectivity.

Schofield:1959:JWI

[Sch59] Robert E. Schofield. Josiah Wedgwood, industrial chemist. Chymia: Annual Studies in the History of Chemistry, 5(??): 180–192, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757184.

Schofield:1964:SMW

[Sch64]  Robert E. Schofield.  Still

more on the water controversy. Chymia: Annual Studies in the History of Chemistry, 9(??):71– 76, ???? 1964. CODEN CHYMAP. ISSN 0095-9367.

                                        URL http://www.jstor.org/
                                        stable/27757233.

Schelar:1966:TTL

[Sch66] Virginia M. Schelar. Thermochemistry and the Third Law of Thermodynamics. Chymia: Annual Studies in the History of Chemistry, 11(??):99–124, ???? 1966. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757262.

Schufle:1967:TBE

[Sch67] J. A. Schufle. Torbern Bergman, Earth scientist. Chymia: Annual Studies in the History of Chemistry, 12(??):58–97, ???? 1967. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757274.

Schubring:1981:MTT

[Sch81] Gert Schubring. Mathematics and teacher training: Plans for a Polytechnic in Berlin. Historical Studies in the Physical Sciences, 12(1):161–194, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757492.

Schweber:1986:ETR

[Sch86] S. S. Schweber. The empiricist temper regnant: Theoretical physics in the United States 1920–1950. Historical Studies in the Physical and Biological Sciences, 17(1):55–98, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757575.

Schopman:1988:ISS

[Sch88] Joop Schopman. Industrious science: Semiconductor research at the N. V. Philips’ Gloeilampenfabrieken, 1930–1957. Historical Studies in the Physical and Biological Sciences, 19(1): 137–172, ???? 1988. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757619.

Schweber:1990:YJC

[Sch90] S. S. Schweber. The young John Clarke Slater and the development of quantum chemistry. Historical Studies in the Physical and Biological Sciences, 20 (2):339–406, ???? 1990. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757647.

Scheideler:2002:SMA

[Sch02] Britta Scheideler. The scientist as moral authority: Albert Einstein between elitism and democracy, 1914–1933. Historical Studies in the Physical and Biological Sciences, 32(2): 319–346, March 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Schirrmacher:2003:ETP

[Sch03] Arne Schirrmacher. Experimenting theory: The proofs of Kirchhoff’s radiation law before and after Planck. Historical Studies in the Physical and Biological Sciences, 33(2): 299–335, March 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Schickore:2007:MEH

[Sch07] Jutta Schickore. The microscope and the eye: a history of reflections, 1740–1870. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-73784-5 (cloth). ix + 317 pp. LCCN QH204 .S34 2007. URL http:

//www.loc.gov/catdir/
enhancements/fy0732/
2007006102-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007006102-b.html;   http:
//www.loc.gov/catdir/toc/
ecip0710/2007006102.html.

Schmalzer:2008:PPM

[Sch08a] Sigrid Schmalzer. The people’s Peking man: popular science and human identity in twentieth-century China. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-73859-0 (hardcover), 0-226-73860-4 (paperback). xix + 346 pp. LCCN GN284.7 .S36

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0829/2007046296-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0829/2007046296-t.html.

Schwartz:2008:PGD

[Sch08b] James Schwartz. In pursuit of the gene: from Darwin to DNA. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02670-5. xiii + 370 pp. LCCN QH428 .S24

2008. URL http://www.loc.
gov/catdir/toc/fy0804/
2007034830.html.

Scheiding:2009:PKO

[Sch09] Tom Scheiding. Paying for knowledge one page at a time: The author fee in physics in Twentieth-Century America. Historical Studies in the Natural Sciences, 39(2):219–247, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Schmalzer:2012:PSU

[Sch12a] Sigrid Schmalzer. Popular science, a useful and productive category after all: Book reviews: Peter J. Bowler, Science for All: The Popularization of Science in Early Twentieth-Century Britain. Chicago: University of Chicago Press, 2009. ix + 339 pp., illus., app., bibl., index. ISBN 978-0-226-06863-3. $45.00 (cloth). Sally Gregory Kohlstedt, Teaching Children Science: Hands-On Nature Study in North America, 1890–1930. Chicago: University of Chicago Press,, 2010. 363 pp., illus., apps., bibl., index. ISBN 978-0-226-44990-6. $45.00 (cloth). Nikolai Krementsov, A Martian Stranded on Earth: Alexander Bogdanov, Blood Transfusions, and Proletarian Science. Chicago: University of Chicago Press,, 2011. xvi + 192 pp., illus., bibl., index. ISBN 978-0-226-45412-2. $35.00 (cloth). Asif A. Siddiqi, The Red Rockets’ Glare: Spaceflight and the Soviet Imagination, 1857– 1957. New York: Cambridge University Press, 2010. xiv + 402 pp., illus., tables, index. ISBN 978-0-521-89760-0. $85.00 (cloth). Historical Studies in the Natural Sciences, 42(5):590– 600, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.590.pdf.

Schweber:2012:NFM

[Sch12b] S. S. (Silvan S.) Schweber. Nuclear forces: the making of the physicist Hans Bethe. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06587-5 (hardcover). viii + 579 + 12 pp. LCCN QC774.B4 S39 2012. £25.95.

Schickore:2020:OEP

[Sch20] Jutta Schickore. Obsolete, extraneous, or pertinent? Philosophy of Science and Historical Studies in the Physical (Natural) Sciences. Historical Studies in the Natural Sciences, 50(1– 2):17–24, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Scott:1950:KFM

[Sco50] John Mark Scott. Karl Friedrich Mohr, 1806–1879: Father of volumetric analysis. Chymia: Annual Studies in the History of Chemistry, 3(??):191–203, ???? 1950. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757152.

Suarez-Diaz:2008:BRY

[SD08] Edna Suarez-Diaz. Book reviews: Yes! There’s Still a Lot to Learn about the History of Molecular Biology: Frederic Lawrence Holmes. Reconceiving the Gene: Seymour Benzer’s Adventures in Phage Genetics. Ed. William C. Summers. New Haven, CT: Yale University Press, 2006. ix + 334 pp., illus., index. ISBN 978-03001-1078-4. $55 (hardcover). María Jesús Santesmases. Severo Ochoa: De músculos a proteínas. Madrid: Editorial Síntesis/FECYT, 2005. 398 pp., illus., index. ISBN 978-84-9756-311-6. EUR17.79 (paperback). Bruno J. Strasser. La fabrique d’une nouvelle science: La biologie moléculaire à l’âge atomique (1945–1964). Florence: Leo O. Olschki Editore, 2006. xxxiii + 450 pp., illus., index. ISBN 978-88-222-5496-2. $74.50 (paperback). Historical Studies in the Natural Sciences, 38(3): 463–473, August 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Shin:2025:PNC

[SD25] Jieun Shin and David H. De-Vorkin. A planetarium for the nation’s capital. Historical Studies in the Natural Sciences, 55(1):1–??, February 2025. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        55/1/1/207545/A-
                                        Planetarium-for-the-
                                        Nation-s-Capital.

Sear:2025:LWS

[Sea25] Rebecca Sear. The legacy of Wilson’s sociobiology for the human evolutionary behavioral sciences, fifty years on. Historical Studies in the Natural Sciences, 55(1):77–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        55/1/77/207534/The-Legacy-
                                        of-Wilson-s-Sociobiology-
                                        for-the-Human.

Secord:2014:VSB

[Sec14] James A. Secord. Visions of science: books and readers at the dawn of the Victorian age.

University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-20328-X (hardcover). xiii + 306 pp. LCCN Q127.G4 S44 2014.

Seidel:1983:ASP

[Sei83] Robert W. Seidel. Accelerating science: The postwar transformation of the Lawrence Radiation Laboratory. Historical Studies in the Physical Sciences, 13(2):375–400, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757520.

Seidel:1986:BRN

[Sei86a] Robert W. Seidel. Book review:

Nuclear physics under Rutherford at Cambridge: Rutherford: Simple Genius by David Wilson; Cockcroft and the Atom by T. E. Allibone and Guy Hartcup; Oliphant by Stewart Cockburn and David Ellyard; Cambridge Physics in the Thirties by John Hendry. Historical Studies in the Physical and Biological Sciences, 17(1): 175–181, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757578.

Seidel:1986:HBS

[Sei86b] Robert W. Seidel. A home for big science: The Atomic Energy Commission’s laboratory system. Historical Studies in the Physical and Biological Sciences, 16(1):135–175, ????

1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757560.

Seidel:1987:GFH

[Sei87] Robert W. Seidel. From glow to flow: a history of military laser research and development. Historical Studies in the Physical and Biological Sciences, 18 (1):111–147, ???? 1987. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757598.

Seidel:2001:NLA

[Sei01] Robert W. Seidel. The national laboratories of the Atomic Energy Commission in the early Cold War. Historical Studies in the Physical and Biological Sciences, 32(1):145–162, September 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Seidel:2008:FFD

[Sei08] Robert W. Seidel. From factory to farm: Dissemination of computing in high-energy physics. Historical Studies in the Natural Sciences, 38(4):479– 507, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Sementsov:1966:EEV

[Sem66] A. Sementsov. Egor Egorovich Vagner and his role in terpene chemistry. Chymia: Annual Studies in the History of Chemistry, 11 (??):151–155, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757265.

Semendeferi:2008:LNC

[Sem08] Ioanna Semendeferi. Legitimating a nuclear critic: John Gofman, radiation safety, and cancer risks. Historical Studies in the Natural Sciences, 38(2):259– 301, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Semerdjian:2023:AW

[Sem23] Elyse Semerdjian. Archival wounds. Historical Studies in the Natural Sciences, 53(1):97– ??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/53/1/97/
195532/Archival-Wounds.

Seow:2021:BRT

[Seo21] Victor Seow. Book review:

Time/Travel. Historical Studies in the Natural Sciences, 51 (3):420–??, June 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL

https://online.ucpress.
edu/hsns/article/51/3/420/
117355/Time-Travel.

Sepkoski:2018:DTS

[Sep18] David Sepkoski. Data in time:

Statistics, natural history, and the visualization of temporal data. Historical Studies in the Natural Sciences, 48(5):581– 593, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Serwer:1977:UZP

[Ser77] Daniel Serwer. Unmechanischer Zwang: Pauli, Heisenberg, and the rejection of the mechanical atom, 1923–1925. Historical Studies in the Physical Sciences, 8(??):189–256, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757371.

Servos:1983:EBF

[Ser83] John W. Servos. To explore the borderland: The foundation of the Geophysical Laboratory of the Carnegie Institution of Washington. Historical Studies in the Physical Sciences, 14 (1):147–185, ???? 1983. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757528.

Seth:2004:QTE

[Set04] Suman Seth. Quantum theory and the electromagnetic world-view. Historical Studies in the Physical and Biological Sciences, 35(1):67–93, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See comment [Kat05] and response [Set05].

Seth:2005:RSK

[Set05] Suman Seth. Response to Shaul Katzir: “On the electromagnetic world-view”. Historical Studies in the Physical and Biological Sciences, 36(1): 193–196, September 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). See [Set04, Kat05].

Seth:2010:CQA

[Set10] Suman Seth. Crafting the quantum: Arnold Sommerfeld and the practice of theory, 1890– 1926. Transformations. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01373-8 (hardcover). viii + 378 pp. LCCN QC16.S76 S48 2010.

Seth:2011:HPA

[Set11] Suman Seth. The history of physics after the cultural turn. Historical Studies in the Natural Sciences, 41(1):112–122, Winter 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2011.41.1.112.

Sethi:2012:IEI

[Set12] Megan Barnhart Sethi. Information, education, and indoctrination: The Federation of American Scientists and public communication strategies in the Atomic Age. Historical Studies in the Natural Sciences, 42 (1):1–29, February 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2012.
42.1.1.

Seth:2016:DEL

[Set16]  Suman Seth.  Darwin and

the ethnologists: Liberal racialism and the geological analogy. Historical Studies in the Natural Sciences, 46(4):490–527, September 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Seth:2020:RME

[Set20] Suman Seth. Russell McCormmach and the empathetic worldview. Historical Studies in the Natural Sciences, 50(1–2):5– 10, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Silva:2013:CPQ

[SF13] Indianara Silva and Olival Freire, Jr. The concept of the photon in question: The controversy surrounding the HBT effect circa 1956–1958. Historical Studies in the Natural Sciences, 43(4):453–491, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.

2013.43.4.453. See corrections [SF15].

Silva:2015:CCP

[SF15] Indianara Silva and Olival Freire, Jr. Corrections: “The Concept of the Photon in Question: The Controversy Surrounding the HBT Effect circa 1956–1958, HSNS 43, no. 4 (2013). Historical Studies in the Natural Sciences, 45(3):511, June 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.

45.3.511. See [SF13].

Simoes:1999:QCQ

[SG99] Ana Simões and Kostas Gavroglu. Quantum chemistry qua applied mathematics. The contributions of Charles Alfred Coulson (1910–1974). Historical Studies in the Physical and Biological Sciences, 29(2):363–406, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Stuhrmann:2025:NES

[SG25] Cora Stuhrmann and Nayanika Ghosh. No end in sight: The sociobiology debate at fifty. Historical Studies in the Natural Sciences, 55(1):37–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/37/207538/No-End-in-
SightThe-Sociobiology-
Debate-at-Fifty.

Shell:2012:HSC

[She12] Hanna Rose Shell. Hide and seek: camouflage, photography, and the media of reconnaissance. Zone Books, New York, NY, USA, 2012. ISBN 1-935408-22-4. 239 + 16 pp. LCCN TR148 .S54 2012.

Sheldon:2016:RSS

[She16] Myrna Perez Sheldon. Review: The Study of Science and Religion as Social Justice. Historical Studies in the Natural Sciences, 46(4):538–547, September 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Shinn:1979:FSF

[Shi79] Terry Shinn. The French Science Faculty System, 1808– 1914: Institutional change and research potential in mathematics and the physical sciences. Historical Studies in the Physical Sciences, 10(??): 271–332, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757392.

Shinn:1986:RFS

[Shi86] Terry Shinn. Review: Failure or success? interpretations of 20th Century French physics. Historical Studies in the Physical and Biological Sciences, 16 (2):353–369, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757569.

Shinn:1987:RFS

[Shi87] Terry Shinn. Review: Failure or success? interpretations of 20th Century French physics. Historical Studies in the Physical and Biological Sciences, 17 (2):361–??, ???? 1987. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757588.

Shinn:1993:BGE

[Shi93] Terry Shinn. Bellevue grand électroaimant, 1900–1940: Birth of a research-technology community. Historical Studies in the Physical and Biological Sciences, 24(1):157–187, ???? 1993. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Shields:2016:MPC

[Shi16]  Brit  Shields.

Mathematics, peace, and the Cold War: Scientific diplomacy and Richard Courant’s scientific identity. Historical Studies in the Natural Sciences, 46(5):556– 591, November 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Shindell:2019:MCE

[Shi19] Matthew Shindell. Making a community of experts: The rise of consensus-based assessments for policy in Cold War America. Historical Studies in the Natural Sciences, 49(1):70–103, February 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Shih:2023:TSA

[Shi23] Ashanti Shih. Talking story with the archives. Historical Studies in the Natural Sciences, 53(1):102–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/1/102/195538/Talking-
Story-with-the-Archives.

Siddiqi:2010:RRG

[Sid10] Asif A. Siddiqi. The Red rockets’ glare: spaceflight and the Soviet imagination, 1857–1957. Cambridge centennial of flight. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-89760-2 (hardcover). 1 v. (xiii + 402) pp. LCCN TL789.8.S65 S489 2010.

Siegfried:1959:CPH

[Sie59] Robert Siegfried. The chemical philosophy of Humphry Davy. Chymia: Annual Studies in the History of Chemistry, 5(??): 193–201, ???? 1959. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757185.

Siegfried:1964:PCH

[Sie64] Robert Siegfried. The phlogistic conjectures of Humphry Davy. Chymia: Annual Studies in the History of Chemistry, 9 (??):117–124, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757236.

Siegel:1978:CER

[Sie78] Daniel M. Siegel. Classicalelectromagnetic and relativistic approaches to the problem of nonintegral atomic masses. Historical Studies in the Physical Sciences, 9(??):323–360, ???? 1978. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757382.

Siegel:1986:ODC

[Sie86] Daniel Siegel. The origin of the displacement current. Historical Studies in the Physical and Biological Sciences, 17(1): 99–146, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757576.

Sigurdsson:1992:RLD

[Sig92]  Skuli Sigurdsson.  17,000

reprints later: Description and analysis of the Vito Volterra Reprint Collection. Historical Studies in the Physical and Biological Sciences, 22(2):391– 397, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757686.

Silliman:1974:FEP

[Sil74] Robert H. Silliman. Fresnel and the emergence of physics as a discipline. Historical Studies in the Physical Sciences, 4 (??):137–162, ???? 1974. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757329.

Simon:2011:CPP

[Sim11] Josep Simon. Communicating physics: the production, circulation and appropriation of Ganot’s textbooks in France and England, 1851–1887. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-130-1 (hardcover). xi + 302 pp. LCCN ????

Simon:2016:WDG

[Sim16] Josep Simon. Writing the discipline: Ganot’s textbook science and the “invention” of physics. Historical Studies in the Natural Sciences, 46(3):392–427, June 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Simpson:2019:BCP

[Sim19] Emily Simpson. Building canals from Panama to Mars: Technological progress and scientific ideals. Historical Studies in the Natural Sciences, 49(4):420– 446, September 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Sivin:1962:WLC

[Siv62] Nathan Sivin. William Lewis (1708–1781) as a chemist. Chymia: Annual Studies in the History of Chemistry, 8(??):63–88, ???? 1962. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757218.

Slater:2000:IAS

[Sla00] Leo B. Slater. Industry and academy: The synthesis of steroids. Historical Studies in the Physical and Biological Sciences, 30(2):443–480, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Slaney:2012:ERB

[Sla12a] Patrick David Slaney. Eugene Rabinowitch, the Bulletin of the Atomic Scientists, and the nature of scientific internationalism in the early Cold War. Historical Studies in the Natural Sciences, 42(2):114–142, April 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2012.42.2.114.

Slayton:2012:DAL

[Sla12b] Rebecca Slayton. From a “dead albatross” to Lincoln labs: Applied research and the making of a normal Cold War university. Historical Studies in the Natural Sciences, 42(4):255– 282, September 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2012.
42.4.255.

Sleigh:2025:HAM

[Sle25] Charlotte Sleigh. How ants made sociobiology: Cybernetics and the organicist tradition. Historical Studies in the Natural Sciences, 55(1):71–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/71/207540/How-Ants-

                                        Made-
                                        SociobiologyCybernetics-
                                        and-the.

Sloan:2014:MCR

[Slo14] Phillip R. Sloan. Molecularizing Chicago — 1945–1965: The rise, fall, and rebirth of the University of Chicago Biophysics Program. Historical Studies in the Natural Sciences, 44(4): 364–412, September 2014. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.4.364.

Smail:2008:DHB

[Sma08] Daniel Lord Smail. On deep history and the brain. An Ahmanson foundation book in the humanities. University of California Press, Berkeley, CA, USA, 2008. ISBN 0-520-25289-6 (hardcover), 1-4337-0883-3, 0-520-25812-6 (paperback). xiv + 271 pp. LCCN D16.9 .S62 2008.

                                        URL  http://catdir.loc.
                                        gov/catdir/enhancements/
                                        fy0808/2007011729-b.html;
                                         http://catdir.loc.gov/
                                        catdir/enhancements/
                                        fy0808/2007011729-d.html;
                                         http://catdir.loc.gov/
                                        catdir/toc/ecip0713/
                                        2007011729.html.

Smeaton:1966:MCM

[Sme66] W. A. Smeaton. Macquer on the composition of metals and the artificial production of gold and silver. Chymia: Annual Studies in the History of Chemistry, 11(??):81–88, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757260.

Smith:2001:I

[Smi01] Robert W. Smith. Introduction.

Historical Studies in the Physical and Biological Sciences, 32 (1):3–9, September 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Smith:2006:CDV

[Smi06] Jonathan Smith. Charles Darwin and Victorian visual culture. Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85690-6 (hardcover). xxiii + 349 pp. LCCN QH46.5 .S65

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0703/2006283483-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0703/2006283483-t.html;
 http://www.loc.gov/
catdir/enhancements/
fy0733/2006283483-b.html.

Smocovitis:2009:PDB

[Smo09] Vassiliki Betty Smocovitis. The ‘plant Drosophila’: E. B. Babcock, the genus Crepis, and the evolution of a genetics research program at Berkeley, 1915–1947. Historical Studies in the Natural Sciences, 39(3): 300–355, August 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Smocovitis:2012:DKH

[Smo12] Vassiliki Betty Smocovitis. Dr. Kuhn, or how I learned to stop worrying and started loving the structure. Historical Studies in the Natural Sciences, 42(5): 564–569, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.564.pdf.

Soderqvist:1994:ABP

[Sod94] Thomas Soderqvist. The architecture of a biographical pathway. Historical Studies in the Physical and Biological Sciences, 25(1):165–175, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Solovev:1961:NMS

[Sol61] Yuriĭ Ivanovich Solov’ev. New materials for the scientific biography of J. J. Berzelius — the scientific relations of Berzelius with Russian scholars, from unpublished letters. Chymia: Annual Studies in the History of Chemistry, 7(??):109–125, ???? 1961. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

                                        jstor.org/stable/27757208.

Solovev:1965:ULH

[Sol65] Yurii Ivanovich Solov’ev. Unpublished letters from H. Sainte-Claire Deville to B. S.

Yakobi. Chymia: Annual Studies in the History of Chemistry, 10(??):111–120, ???? 1965. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757248.

Somsen:2020:PRH

[Som20] Geert Somsen. The philosopher and the rooster: Henri Bergson’s French diplomatic missions, 1914–1925. Historical Studies in the Natural Sciences, 50(4):364–??, September 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/4/364/111894/The-
Philosopher-and-the-
RoosterHenri-Bergson-s.

Sonnedecker:1953:SBC

[Son53] Glenn Sonnedecker. The scientific background of chemistry teachers in representative pharmacy schools of the United States during the 19th Century. Chymia: Annual Studies in the History of Chemistry, 4 (??):171–200, ???? 1953. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757168.

Sowards:2017:CSS

[Sow17] Adam M. Sowards. Claiming spaces for science: Scientific exploration and the Canadian Arctic Expedition of 1913–1918. Historical Studies in the Natural Sciences, 47(2):164–199, April 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Spary:2012:EEF

[Spa12] E. C. (Emma C.) Spary. Eating the Enlightenment: food and the sciences in Paris. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-76886-4 (hardcover), 0-226-76888-0 (ebook). xi + 366 pp. LCCN GT2853.F8 S737 2012.

Spier:2011:BHF

[Spi11] Fred Spier. Big history and the future of humanity. Wiley-Blackwell, Chichester, UK, 2011. ISBN 1-4443-3421-2 (hardcover), 1-4443-3943-5 (paperback). xv + 272 pp. LCCN CB19 .S679 2011.

Sanchez-Ron:2002:IRS

[SR02a] José M. Sánchez-Ron. International relations in Spanish physics from 1900 to the Cold War. Historical Studies in the Physical and Biological Sciences, 33(1):3–31, September 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Shama:2002:AIR

[SR02b] Gilbert Shama and Jonathan Reinarz. Allied intelligence reports on wartime German penicillin research and production. Historical Studies in the Physical and Biological Sciences, 32 (2):347–367, March 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Siegmund-Schultze:2018:AMV

[SS18]  Reinhard  Siegmund-

Schultze. Applied mathematics versus fluid dynamics: The catalytic role of Richard von Mises (1883–1953). Historical Studies in the Natural Sciences, 48(4):475–525, September 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Skulberg:2023:BHI

[SS23] Emilie Skulberg and Martin Sparre. A black hole in ink: Jean-Pierre Luminet and “Realistic” black hole imaging. Historical Studies in the Natural Sciences, 53(4):389–??, September 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/53/4/389/197392/A-
Black-Hole-in-InkJean-
Pierre-Luminet-and.

Santesmases:2015:CBE

[SSD15] María Jesús Santesmases and Edna Suárez-Díaz. A cell-based epistemology: Human genetics in the era of biomedicine. Historical Studies in the Natural Sciences, 45(1):1–13, February 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2015.45.1.1.

Sviedrys:1970:RPS

[ST70] Romualdas Sviedrys and Arnold Thackray. The rise of physical science at Victorian Cambridge with commentary and with reply. Historical Studies in the Physical Sciences, 2 (??):127–151, ???? 1970. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757306.

Sepkoski:2018:ISC

[ST18] David Sepkoski and Marco Tamborini. “An image of science”: Cameralism, statistics, and the visual language of natural history in the Nineteenth Century. Historical Studies in the Natural Sciences, 48(1):56– 109, February 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Shindell:2024:MHS

[ST24] Matthew Shindell and Samantha M. Thompson. Museums and the history of science. Historical Studies in the Natural Sciences, 54(1):112–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/54/1/112/200193/
Museums-and-the-History-
of-Science.

Stanley:2008:BRE

[Sta08] Matthew Stanley. Book reviews: Einstein: Essence or Explanation? Einstein: His Life and Universe, by Walter Isaacson. New York: Simon & Schuster, 2007. xxii + 675 pp., illus., index. $32 (cloth). Einstein: A Biography, by Jürgen Neffe. Trans. Shelley Frisch. New York: Farrar, Straus, and Giroux, 2007. x + 461 pp., illus., index. $30 (cloth). Einstein on Politics. David E. Rowe and Robert Schulmann, ed. Princeton: Princeton University Press, 2007. xxxiv + 523 pp., illus., index. $29.95 (cloth). Historical Studies in the Natural Sciences, 38(1):153–161, February 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Staley:2011:BRD

[Sta11]  Richard Staley.  Book re-

view: Discontinuous memory and the making of quantum mechanics: Quantum Theory at the Crossroads: Reconsidering the 1927 Solvay Conference by Guido Bacciagaluppi and Antony Valentini; Beyond Uncertainty: Heisenberg, Quantum Physics, and the Bomb by David C. Cassidy; Crafting the Quantum: Arnold Sommerfeld and the Practice of Theory, 1890– 1926 by Suman Seth. Historical Studies in the Natural Sciences, 41(4):447–456, November 2011. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2011.

41.4.447. See [BV09, Cas09b, Set10].

Stark:2012:HKB

[Sta12]  Laura Stark.  The house

that Kuhn built: Teaching Fleck’s genesis and development through structure. Historical Studies in the Natural Sciences, 42(5):570–575, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.570.pdf.

Staley:2016:RTH

[Sta16] Richard Staley. Review: Telling Histories: Working Worlds and the Sciences of the Twentieth Century. Historical Studies in the Natural Sciences, 46(1):110– 118, February 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Stark:2019:BCP

[Sta19] James F. Stark. Biopolitics and the collective predicament. Historical Studies in the Natural Sciences, 49(5):566–572, November 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Stewart:1988:RTC

[Ste88] Larry Stewart. Review: Texts and contextualists: The hunting of Newtonianism. Historical Studies in the Physical and Biological Sciences, 19(1):193– 197, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757621.

Stevens:2003:FPJ

[Ste03] Hallam Stevens. Fundamental physics and its justifications, 1945–1993. Historical Studies in the Physical and Biological Sciences, 34(1):151–197, September 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Stevens:2016:RSD

[Ste16] Hallam Stevens. Review: Seeing Data. Historical Studies in the Natural Sciences, 46(2):252– 259, April 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Stradins:1964:WTG

[Str64] Janis Stradins. The work of Theodore Grotthus and the invention of the Davy Safety Lamp. Chymia: Annual Studies in the History of Chemistry, 9(??):125–145, ???? 1964. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757237.

Strasser:2006:FNS

[Str06] Bruno J. Strasser. La fabrique d’une nouvelle science: la biologie moléculaire à l’âge atomique (1945–1964). Leo S. Olschki, Firenze, Italy, 2006. ISBN 88-222-5496-1. xxxiii + 450 pp. LCCN QH506 .S8255 2006.

Strasser:2008:BRM

[Str08] Bruno J. Strasser. Book reviews: Magic Bullets and Wonder Pills: Making Drugs and Diseases in the Twentieth Century: Robert Bud. Penicillin: Triumph and Tragedy. Oxford: Oxford University Press, 2007.

344 pp., illus. ISBN 978-0-19-925406-4. $46 (hardcover). Jeremy A. Greene. Prescribing by Numbers: Drugs and the Definition of Disease. Baltimore: Johns Hopkins University Press, 2007. 336 pp., illus. ISBN 978-0-8018-8477-1. $42.70 (hardcover). John E. Lesch. The First Miracle Drugs: How the Sulfa Drugs Transformed Medicine. Oxford: Oxford University Press, 2007. 376 pp., illus. ISBN 978-0-19-518775-5. $59.50 (hardcover). Elizabeth Siegel Watkins. The Estrogen Elixir: A History of Hormone Replacement Therapy in America. Baltimore: Johns Hopkins University Press, 2007. 368 pp. ISBN 978-0-8018-8602-7. $37.10 (hardcover). Historical Studies in the Natural Sciences, 38(2): 303–312, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Strasser:2010:LMC

[Str10] Bruno J. Strasser. Laboratories, museums, and the comparative perspective: Alan A. Boyden’s quest for objectivity in serological taxonomy, 1924–1962. Historical Studies in the Natural Sciences, 40(2):149–182, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Stuewer:1975:GLD

[Stu75] Roger H. Stuewer. G. N. Lewis on detailed balancing, the symmetry of time, and the nature of light. Historical Studies in the Physical Sciences, 6 (??):469–511, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757347.

Stuewer:1986:RSM

[Stu86] Roger H. Stuewer. Rutherford’s satellite model of the nucleus. Historical Studies in the Physical and Biological Sciences, 16 (2):321–352, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.
jstor.org/stable/27757568.

Stuhrmann:2025:EPF

[Stu25] Cora Stuhrmann. The ecopolitical future of sociobiology. Historical Studies in the Natural Sciences, 55(1):65–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/65/207537/The-Eco-
Political-Future-of-
Sociobiology.

Sunderland:2012:CBR

[Sun12]  Mary  E.

Sunderland. Collections-based research at Berkeley’s Museum of Vertebrate Zoology. Historical Studies in the Natural Sciences, 42(2):83–113, April 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/

                                        stable/10.1525/hsns.2012.
                                        42.2.83.

Sur:2002:SSJ

[Sur02] Abha Sur. Scientism and social justice: Meghnad Saha’s critique of the state of science in India. Historical Studies in the Physical and Biological Sciences, 33(1):87–105, September 2002. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Surita:2022:PPM

[Sur22] Gina Surita. The power of phosphate: Making and breaking bonds across the Atlantic, 1927– 1946. Historical Studies in the Natural Sciences, 52(1):1–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/52/1/1/120501/The-
                                        Power-of-PhosphateMaking-
                                        and-Breaking-Bonds.

Sutton:1981:PSE

[Sut81] Geoffrey Sutton. The politics of science in early Napoleonic France: The case of the Voltaic pile. Historical Studies in the Physical Sciences, 11(2): 329–366, ???? 1981. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757483.

Sviedrys:1976:RPL

[Svi76]  Romualdas Sviedrys.  The

rise of physics laboratories in Britain. Historical Studies in the Physical Sciences, 7(??): 405–436, ???? 1976. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757360.

Saraiva:2010:AAG

[SW10] Tiago Saraiva and M. Norton Wise. Autarky/autarchy: Genetics, food production, and the building of fascism. Historical Studies in the Natural Sciences, 40(4):419–428, Fall 2010. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2010.
40.4.419.

Swanner:2017:ISC

[Swa17] Leandra Swanner. Instruments of science or conquest? Neocolonialism and modern American astronomy. Historical Studies in the Natural Sciences, 47(3):293– 319, June 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Szczygielska:2023:RTI

[Szc23] Marianna Szczygielska. Reading teeth: Ivory as an artifact of classed whiteness. Historical Studies in the Natural Sciences, 53(3):308–??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
53/3/308/196660/Reading-
TeethIvory-as-an-Artifact-
of-Classed.

Szczygielska:2024:WEW

[Szc24] Marianna Szczygielska. War on extinction: Wildlife as statecraft in interwar Poland. Historical Studies in the Natural Sciences, 54(2):244–??, April 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/2/244/200668/War-on-
                                        ExtinctionWildlife-as-
                                        Statecraft-in.

Szymborski:1984:PIC

[Szy84] Krzysztof Szymborski. The physics of imperfect crystals — a social history. Historical Studies in the Physical Sciences, 14 (2):317–355, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757536.

Tabernero:2022:CKS

[Tab22] Carlos Tabernero. The case of the killer she-wolf: Media, the natural sciences, and the construction of the environment in late Franco’s Spain. Historical Studies in the Natural Sciences, 52(4):524–??, September 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        52/4/524/191417/The-Case-
                                        of-the-Killer-She-
                                        WolfMedia-the-Natural.

Tai:2017:LRM

[Tai17] Chaokang Tai. Left radicalism and the Milky Way: Connecting the scientific and socialist virtues of Anton Pannekoek. Historical Studies in the Natural Sciences, 47(2):200–254, April 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Takats:2011:ECE

[Tak11] Sean Takats. The expert cook in Enlightenment France, volume 2 of The Johns Hopkins University studies in historical and political science, 129th series. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0283-1 (hardcover). x + 203 pp. LCCN TX719 .T253 2011.

Takeshita:2012:GBI

[Tak12] Chikako Takeshita. The global biopolitics of the IUD: how science constructs contraceptive users and women’s bodies. Inside technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01658-3. xvi + 238 pp. LCCN RG137.3 .T35 2012.

Taltavull:2016:TKA

[Tal16] Marta Jordi Taltavull. Transmitting knowledge across divides: Optical dispersion from classical to quantum physics. Historical Studies in the Natural Sciences, 46(3):313–359, June 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Tamborini:2018:EHN

[Tam18] Marco Tamborini. Expanding the history of natural history. Historical Studies in the Natural Sciences, 48(3):390–401, June 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Turchetti:2020:IJN

[TAR^{+}20] Simone Turchetti, Matthew Adamson, Giulia Rispoli, Doubravka Olšáková, and Sam Robinson. Introduction: Just Needham to Nixon? On writing the history of “science–diplomacy”. Historical Studies in the Natural Sciences, 50(4):323–??, September 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        50/4/323/111896/
                                        IntroductionJust-Needham-
                                        to-Nixon-On-Writing-the.

Tatarchenko:2016:CSM

[Tat16] Ksenia Tatarchenko. Calculating a showcase: Mikhail Lavrentiev, the politics of expertise, and the international life of the Siberian science-city. Historical Studies in the Natural Sciences, 46(5):592–632, November 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Tatarchenko:2019:TAC

[Tat19] Ksenia Tatarchenko. Thinking algorithmically: From Cold War computer science to the socialist information culture. Historical Studies in the Natural Sciences, 49(2):194–225, April 2019. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Taylor:1948:AMM

[Tay48] F. Sherwood Taylor. The argument of Morien and Merlin: An English alchemical poem. Chymia: Annual Studies in the History of Chemistry, 1(??):23– 35, ???? 1948. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757112.

Tavares:2020:TSI

[TBV20] Heráclio D. Tavares, Alexandre Bagdonas, and Antonio A. P. Videira. Transnationalism as scientific identity: Gleb Wataghin and Brazilian physics, 1934–1949. Historical Studies in the Natural Sciences, 50(3):248– 301, June 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/50/3/248/110325/
Transnationalism-as-
Scientific-IdentityGleb.

Tchorbadjiev:1967:PRF

[Tch67] S. Tchorbadjiev. P. N. Raikov — founder of organic chemistry in Bulgaria. Chymia: Annual Studies in the History of Chemistry, 12(??):171– 181, ???? 1967. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757280.

Teasley:2023:SSP

[Tea23] Sarah Teasley. Sticky solutions: The persistence of animal glues in laboratories and workshops in Twentieth-Century Japan. Historical Studies in the Natural Sciences, 53(3):278–??, June 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        53/3/278/196659/Sticky-
                                        SolutionsThe-Persistence-
                                        of-Animal-Glues-in.

Teissier:2010:SSC

[Tei10]  Pierre Teissier.  Solid-state

chemistry in France: Structures and dynamics of a scientific community since World War II. Historical Studies in the Natural Sciences, 40(2):225– 258, May 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

TenHagen:2019:HFS

[Ten19] Sjang L. Ten Hagen. How “facts” shaped modern disciplines: The fluid concept of fact and the common origins of German physics and historiography. Historical Studies in the Natural Sciences, 49(3):300–337, June 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

TenHagen:2022:HTN

[Ten22] Sjang L. Ten Hagen. History as a tool for natural science: How Ernst Mach applied historical methods to physics. Historical Studies in the Natural Sciences, 52(1):40–??, February 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
52/1/40/120502/History-as-
a-Tool-for-Natural-
ScienceHow-Ernst.

Terrall:2014:CNA

[Ter14] Mary Terrall. Catching nature in the act: Réaumur and the practice of natural history in the eighteenth century. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-08860-X (hardcover), 0-226-08874-X (e-book). viii + 275 pp. LCCN QH15 .T47 2014.

Thomas:1984:JSF

[Tho84] Jerry Thomas. John Stuart Foster, McGill University, and the renaissance of nuclear physics in Montreal, 1935–1950. Historical Studies in the Physical Sciences, 14(2):357–377, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757537.

Thomas:2012:SDI

[Tho12] William Thomas. Strategies of detection: Interpretive methods in experimental particle physics, 1930–1950. Historical Studies in the Natural Sciences, 42(5): 389–431, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/

                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.389.pdf.

Thomas:2017:BSN

[Tho17]  Marion Thomas.  Biologi-

cal and social new orders: Charles Robin (1821–1885) and Comteanism in the Third Republic. Historical Studies in the Natural Sciences, 47(5):653– 677, November 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Thompson:2022:GHN

[Tho22] Peter Thompson. From gas hysteria to nuclear fear: a historical synthesis of chemical and atomic weapons. Historical Studies in the Natural Sciences, 52(2):223–??, April 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        52/2/223/144554/From-Gas-
                                        Hysteria-to-Nuclear-FearA-
                                        Historical.

Tiemeyer:2013:PQL

[Tie13] Philip James Tiemeyer. Plane queer: labor, sexuality, and AIDS in the history of male flight attendants. University of California Press, Berkeley, CA, USA, 2013. ISBN 0-520-27476-8 (hardcover), 0-520-27477-6 (paperback), 0-520-95530-7 (ebook). xi + 288 pp. LCCN HD6073.A432 T54 2013.

Tilley:2011:ALL

[Til11] Helen Tilley. Africa as a living laboratory: empire, development, and the problem of scientific knowledge, 1870–1950. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-80346-5 (hardcover), 0-226-80347-3 (paperback). xiv + 496 pp. LCCN Q180.A5 T55 2011.

Tinn:2024:EAE

[Tin24] Honghong Tinn. East Asia and ethics in technoscience. Historical Studies in the Natural Sciences, 54(1):118–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/118/200201/East-Asia-
and-Ethics-in-
Technoscience.

Tatarchenko:2016:MSP

[TP16] Ksenia Tatarchenko and Christopher J. Phillips. Mathematical superpowers: The politics of universality in a divided world. Historical Studies in the Natural Sciences, 46(5):549–555, November 2016. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Trenn:1975:RRA

[Tre75] Thaddeus J. Trenn. Rutherford and recoil atoms: The metamorphosis and success of a once stillborn theory. Historical Studies in the Physical Sciences, 6(??):513–547, ???? 1975. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757348.

Turner:1971:GPR

[Tur71] R. Steven Turner. The growth of professorial research in Prussia, 1818 to 1848 — causes and context. Historical Studies in the Physical Sciences, 3 (??):137–182, ???? 1971. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757317.

Turner:1982:JLV

[Tur82] R. Steven Turner. Justus Liebig versus Prussian chemistry: Reflections on early institutebuilding in Germany. Historical Studies in the Physical Sciences, 13(1):129–162, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757508.

Turner:2005:AFP

[Tur05] R. Steven Turner. After the famine: Plant pathology, Phytophthora infestans, and the late blight of potatoes, 1845–1960. Historical Studies in the Physical and Biological Sciences, 35 (2):341–370, March 2005. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Turchetti:2006:IBN

[Tur06] Simone Turchetti. The invisible businessman: Nuclear physics, patenting practices, and trading activities in the 1930s. Historical Studies in the Physical and Biological Sciences, 37(1): 153–172, September 2006. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Turner:2008:PAL

[Tur08] R. Steven Turner. Potato agriculture, late blight science, and the molecularization of plant pathology. Historical Studies in the Natural Sciences, 38(2):223– 257, May 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Turner:2013:DSM

[Tur13] Fred Turner. The democratic surround: multimedia and American liberalism from World War II to the psychedelic sixties. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-81746-6 (hardcover), 0-226-06414-X (e-book). 365 pp. LCCN P95.82.U6 T87 2013.

Turchetti:2020:SDO

[Tur20] Simone Turchetti. The (science diplomacy) origins of the Cold War. Historical Studies in the Natural Sciences, 50(4):411–??, September 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/50/4/411/111898/
The-Science-Diplomacy-
Origins-of-the-Cold-War.

Turnbull:2024:REU

[Tur24] Thomas M. Turnbull. Redefining efficiency: US physicists and the 1970s energy crisis. Historical Studies in the Natural Sciences, 54(3):365–??, June 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/3/365/201972/
                                        Redefining-EfficiencyUS-
                                        Physicists-and-the-1970s.

Tyrrell:2015:BRT

[Tyr15] Brian Tyrrell. Bred for the race: Thoroughbred breeding and racial science in the United States, 1900–1940. Historical Studies in the Natural Sciences, 45(4):549–576, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2015.
                                        45.4.549.

Tyrrell:2018:PA

[Tyr18] Brian Tyrrell. Political animals. Historical Studies in the Natural Sciences, 48(4):536– 547, September 2018. ISSN 1939-1811 (print), 1939-182X (electronic).

Ungureanu:2019:CEC

[Ung19] James C. Ungureanu. Creation, evolution, and the continuing conflict. Historical Studies in the Natural Sciences, 49(5):573– 584, November 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Urdang:1948:ECP

[Urd48] George Urdang. The early chemical and pharmaceutical history of calomel. Chymia: Annual Studies in the History of Chemistry, 1(??):93–108, ???? 1948. CODEN CHYMAP. ISSN 0095-9367. URL http://www.

jstor.org/stable/27757117.

VanKlooster:1949:BLI

[Van49] H. S. Van Klooster. The beginnings of laboratory instruction in chemistry in the U.S.A. Chymia: Annual Studies in the History of Chemistry, 2(??):1– 15, ???? 1949. CODEN CHYMAP. ISSN 0095-9367.

URL http://www.jstor.org/
stable/27757128.

VanDongen:2007:ERA

[Van07a] Jeroen Van Dongen. Emil Rupp, Albert Einstein, and the canal ray experiments on waveparticle duality: Scientific fraud and theoretical bias. Historical Studies in the Physical and Biological Sciences, 37(S1):73– 120, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.73.pdf.

VanDongen:2007:IER

[Van07b] Jeroen Van Dongen. The interpretation of the Einstein– Rupp experiments and their influence on the history of quantum mechanics. Historical Studies in the Physical and Biological Sciences, 37(S1):121– 131, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.121.pdf.

Vargha:2021:V

[Var21] Dóra Vargha. The vaccine. Historical Studies in the Natural Sciences, 51(1):162–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/51/1/162/116289/
The-Vaccine.

Vaziri:2023:TA

[Vaz23] Parisa Vaziri. Tracing absence. Historical Studies in the Natural Sciences, 53(1):106–??, February 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/53/1/106/195529/
Tracing-Absence.

Vance:2010:GMC

[VD10] Tiffany C. Vance and Ronald E. Doel. Graphical methods and Cold War scientific practice: The Stommel Diagram’s intriguing journey from the physical to the biological environmental sciences. Historical Studies in the Natural Sciences, 40(1):1–47, February 2010. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Vedanayagam:2019:BCT

[Ved19]  Joseph  Satish

Vedanayagam. Beyond the conflict thesis: Studying science and religion as public policy. Historical Studies in the Natural Sciences, 49(4):459–469, September 2019. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Vedanayagam:2020:BRE

[Ved20] Joseph Satish Vedanayagam.

Book review: Exploring the many meanings of purpose and dialogue in religion and science [Yves Gingras. Science and Religion: An Impossible Dialogue. Cambridge: Polity Press, 2017. 272 pp., illus., index. ISBN 978-1-509-51893-7. $26. 95 (paper). Edward J. Larson and Michael Ruse. On Faith and Science. New Haven, CT: Yale University Press, 2017. 312 pp., index. ISBN 978-0-300-21617-2. $30. 00 (cloth). Michael Ruse. On Purpose. Princeton, NJ: Princeton University Press, 2018. 320 pp., index. ISBN 978-0-691-17246-0. $27. 95 (cloth)]. Historical Studies in the Natural Sciences, 50(3):311–322, June 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
50/3/311/110326/Exploring-
the-Many-Meanings-of-
Purpose-and.

Velocci:2022:WTB

[Vel22]  Beans Velocci.  Wrenching

torque: On being professionally nonbinary. Historical Studies in the Natural Sciences, 52(3):476– ??, June 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

                                        //online.ucpress.edu/hsns/
                                        article/52/3/476/185526/
                                        Wrenching-TorqueOn-Being-
                                        Professionally-Nonbinary.

Vennen:2024:TTZ

[Ven24] Mareike Vennen. Tortoise traffic: Zoo animal logistics in Twentieth-Century Berlin. Historical Studies in the Natural Sciences, 54(2):268–??, April 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.
                                        ucpress.edu/hsns/article/
                                        54/2/268/200665/Tortoise-
                                        TrafficZoo-Animal-
                                        Logistics-in-Twentieth.

Vettel:2004:PNS

[Vet04] Eric J. Vettel. The protean nature of Stanford University’s biological sciences, 1946–1972. Historical Studies in the Physical and Biological Sciences, 35(1):95–113, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Vettel:2006:BCO

[Vet06] Eric James Vettel. Biotech: the countercultural origins of an industry. Politics and culture in modern America. University of Pennsylvania Press, Philadelphia, PA, USA, 2006.

ISBN 0-8122-3947-4, 0-8122-2051-X (paperback), 0-8122-0362-3 (e-book). xv + 273 pp. LCCN HD9999.B442 V48 2006.

Vienne:2017:WCC

[Vie17] Florence Vienne. Worlds conflicting: The cell theories of François-Vincent Raspail and Theodor Schwann. Historical Studies in the Natural Sciences, 47(5):629–652, November 2017. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

VonRauchhaupt:2001:CCU

[Von01] Ulf Von Rauchhaupt. Colorful clouds and unruly rockets: Early research programs at the Max Planck Institute for Extraterrestrial Physics. Historical Studies in the Physical and Biological Sciences, 32(1):115– 124, September 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

VonOertzen:2018:DSV

[Von18] Christine Von Oertzen. Datafication and spatial visualization in Nineteenth-Century census statistics. Historical Studies in the Natural Sciences, 48(5):568– 580, November 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Voskuhl:2020:IH

[Vos20] Adelheid Voskuhl. Industrialization as a historical episode. Historical Studies in the Natural Sciences, 50(1–2):50–57, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

vanSpronsen:1967:HPL

[vS67] Jan W. van Spronsen. The history and prehistory of the law of Dulong and Petit as applied to the determination of atomic weights. Chymia: Annual Studies in the History of Chemistry, 12(??):157–169, ???? 1967. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757279.

vanSpronsen:1966:OHY

[vSvS66] J. W. van Spronsen and J. W. van Spronson. One hundred years of the “Law of Octaves”: When the Italian Cannizzaro was fighting for atomic weights in Karlsruhe, Newlands fought for the liberation of Italy. Chymia: Annual Studies in the History of Chemistry, 11(??): 125–137, ???? 1966. CO-DEN CHYMAP. ISSN 0095-

9367. URL http://www.jstor.
org/stable/27757263.

Walker:1992:PPW

[Wal92] Mark Walker. Physics and propaganda: Werner Heisenberg’s Foreign Lectures under National Socialism. Historical Studies in the Physical and Biological Sciences, 22(2):339– 389, ???? 1992. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/27757685.

Walker:1994:SNS

[Wal94a] Mark Walker. Science, National Socialism, and the longue durée. Historical Studies in the Physical and Biological Sciences, 24 (2):395–402, ???? 1994. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Walters:1994:PS

[Wal94b] Alice Walters. Public science.

Historical Studies in the Physical and Biological Sciences, 24 (2):403–??, ???? 1994. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Walker:2000:TTC

[Wal00] Deborah Walker. From Tallahassee to Timbuktu: Cold War efforts to measure intercontinental distances. Historical Studies in the Physical and Biological Sciences, 30(2):393–415, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Walker:2015:BRE

[Wal15a] Mark Walker. Book review: Émigré Biographies: Anja Skaar Jacobsen, Léon Rosenfeld: Physics, Philosophy, and Politics in the Twentieth Century. Hackinsack: World Scientific, 2012. xii + 354 pp., illus., index. ISBN 978-981-4307-81-9. $87.00 (hardcover).; Christoph Laucht, Elemental Germans: Klaus Fuchs, Rudolf Peierls and the Making of British Nuclear Culture 1939–1959. London: Palgrave Macmillan, 2012. xiv + 274 pp., index. ISBN 978-0-230-35487-6. $75.00 (hardcover).; Mary Jo Nye, Michael Polanyi and His Generation: Origins of the Social Construction of Science. Chicago: University of Chicago Press, 2011. xi + 405 pp., illus., index. ISBN 978-0-226-61063-4. $50.00 (hardcover).; Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Cambridge, MA: Harvard University Press, 2012. viii + 579 pp., illus., index. ISBN 978-0-674-06587-1. $35.00 (hardcover). Historical Studies in the Natural Sciences, 45(1): 206–216, February 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.1.206.

Wallace:2015:CSP

[Wal15b] Matthew L. Wallace. A climate for science policy: Constructing postwar atmospheric research in Canada, 1945–1970. Historical Studies in the Natural Sciences, 45(4):577–609, September 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.4.577.

Walker:2021:V

[Wal21] Oriana Walker. The ventilator. Historical Studies in the Natural Sciences, 51(1):165–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/51/1/165/116290/
The-Ventilator.

Wang:1995:LPL

[Wan95a] Jessica Wang. Liberals, the progressive left, and the political economy of postwar American science: The National Science Foundation debate revisited. Historical Studies in the Physical and Biological Sciences, 26 (1):139–166, ???? 1995. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Wang:1995:PBS

[Wan95b] Zuoyue Wang. The politics of big science in the Cold War: PSAC and the funding of SLAC. Historical Studies in the Physical and Biological Sciences, 25 (2):329–356, ???? 1995. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Wang:1999:MSP

[Wan99a] Jessica Wang. Merton’s shadow:

Perspectives on science and democracy since 1940. Historical Studies in the Physical and Biological Sciences, 30(1): 279–306, ???? 1999. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Wang:1999:UCS

[Wan99b] Zuoyue Wang. U.S.–China scientific exchange: a case study of state-sponsored scientific internationalism during the Cold War and beyond. Historical Studies in the Physical and Biological Sciences, 30(1):249–277, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Wang:2008:SSP

[Wan08] Zuoyue Wang. In Sputnik’s shadow: the President’s Science Advisory Committee and Cold War America. Rutgers University Press, New Brunswick, NJ, USA, 2008. ISBN 0-8135-4331-2, (paperback), 0-8135-4332-0 (paperback). xix + 454 pp. LCCN Q127.U6 W365

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip0726/
                                        2007035777.html.

Wang:2012:PES

[Wan12] Jessica Wang. Physics, emotion, and the scientific self: Merle Tuve’s Cold War. Historical Studies in the Natural Sciences, 42(5):341–388, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsns.2012.42.5.341.pdf.

Wang:2020:MPC

[Wan20] Zuoyue Wang. From modern physics to the Cold War and beyond. Historical Studies in the Natural Sciences, 50(1–2):

25–30, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Wan:2024:TAH

[Wan24] Shu Wan. TikTok Asian history of science. Historical Studies in the Natural Sciences, 54 (1):115–??, February 2024. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
54/1/115/200194/TikTok-
Asian-History-of-Science.

Watkins:2007:EEH

[Wat07] Elizabeth Siegel Watkins. The estrogen elixir: a history of hormone replacement therapy in America. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8602-3. ix + 351 pp. LCCN RM295 .W38 2007.

Weart:1997:GWC

[Wea97] Spencer R. Weart. Global warming, Cold War, and the evolution of research plans. Historical Studies in the Physical and Biological Sciences, 27(2): 319–356, ???? 1997. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Weart:2007:MKM

[Wea07] Spencer R. Weart. Money for Keeling: Monitoring CO_{2} levels. Historical Studies in the Physical and Biological Sciences, 37 (2):435–452, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Weisberg:2022:PPB

[Wei22]  Sarah Weisberg.  Pipeline,

pathway, burrow: Reworking science’s metaphorical terrain. Historical Studies in the Natural Sciences, 52(5):629–??, November 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/52/5/629/194654/
Pipeline-Pathway-
BurrowReworking-Science-s.

Weidman:2025:WSF

[Wei25] Nadine Weidman. Wilson, sociobiology, and feminism. Historical Studies in the Natural Sciences, 55(1):57–??, February 2025. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.
ucpress.edu/hsns/article/
55/1/57/207542/Wilson-
Sociobiology-and-Feminism.

Wellerstein:2012:BRN

[Wel12] Alex Wellerstein. Book review: Nuclear others: Being Nuclear: Africans and the Global Uranium Trade by Gabrielle Hecht. Historical Studies in the Natural Sciences, 42(3):235–243, June 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/
hsns.2012.42.3.235.

Werner:1997:LAW

[Wer97] Petra Werner. Learning from an adversary? Warburg against Wieland. Historical Studies in the Physical and Biological Sciences, 28(1):173–196, ???? 1997. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:1996:ASC

[Wes96a] Peter J. Westwick. ’abraded from several corners:’ medical physics and biophysics at Berkeley. Historical Studies in the Physical and Biological Sciences, 27(1):131–162, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:1996:SB

[Wes96b] Peter J. Westwick. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 27(1):177–??, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:1997:SBa

[Wes97a] Peter J. Westwick. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 27(2):357–??, ???? 1997. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:1997:SBb

[Wes97b] Peter J. Westwick. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 28(1):197–??, ???? 1997. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:1998:SBa

[Wes98a] Peter J. Westwick. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 28(2):353–??, ???? 1998. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:1998:SBb

[Wes98b] Peter J. Westwick. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 29(1):183–??, ???? 1998. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:2000:SB

[Wes00] Peter J. Westwick. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 30(2):481–??, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westfall:2001:CTS

[Wes01a] Catherine Westfall. Collaborating together: The stories of TPC, UA1, CDF, and CLAS. Historical Studies in the Physical and Biological Sciences, 32 (1):163–178, September 2001. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westfall:2001:F

[Wes01b] Catherine Westfall. Foreword.

Historical Studies in the Physical and Biological Sciences, 32 (1):1, September 2001. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westfall:2002:TTM

[Wes02] Catherine Westfall. A tale of two more laboratories: Readying for research at Fermilab and Jefferson Laboratory. Historical Studies in the Physical and Biological Sciences, 32(2): 369–407, March 2002. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Westwick:2007:BJA

[Wes07] Peter J. Westwick. Into the black: JPL and the American space program, 1976–2004. Yale University Press, New Haven, CT, USA, 2007. ISBN 0-300-11075-8. xviii + 392 pp. LCCN TL862.J48 W47 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0643/
2006013349-b.html;
http://www.loc.gov/catdir/
enhancements/fy0643/
2006013349-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0612/2006013349.html.

Westfall:2008:ISI

[Wes08a] Catherine Westfall. Introduction to the special issue: Surviving the squeeze: National Laboratories in the 1970s and 1980s. Historical Studies in the Natural Sciences, 38(4):475– 478, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Westfall:2008:RFL

[Wes08b] Catherine Westfall. Retooling for the future: Launching the Advanced Light Source at Lawrence’s Laboratory, 1980– 1986. Historical Studies in the Natural Sciences, 38(4):569– 609, November 2008. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Westfall:2010:STT

[Wes10] Catherine Westfall. Surviving to tell the tale: Argonne’s Intense Pulsed Neutron Source from an ecosystem perspective. Historical Studies in the Natural Sciences, 40(3):350–398, August 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Westwick:2012:BSM

[Wes12] Peter J. Westwick, editor. Blue sky metropolis: the aerospace century in Southern California, volume 4 of Western histories. Published for Huntington-USC Institute on California and the West by University of California Press and Huntington Library, San Marino, California, Berkeley, CA, USA, 2012. ISBN 0-87328-249-3 (hardcover). xii + 308 pp. LCCN HD9711.5.U63 C23 2012.

Westfall:2020:RC

[Wes20] Catherine Westfall. Reimagining 20th-Century physics. Historical Studies in the Natural Sciences, 50(1–2):209–216, April 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Wilson:2010:KAJ

[WG10]  Edward  O.

Wilson and José María Gómez Durán. Kingdom of ants: José Celestino Mutis and the dawn of natural history in the New World. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9785-8, 0-8018-9973-7 (e-book). viii + 95 pp. LCCN QK31.M8 W55 2010eb.

Wheaton:1978:PLP

[Whe78] Bruce R. Wheaton. Philipp Lenard and the photoelectric effect, 1889–1911. Historical Studies in the Physical Sciences, 9(??):299–322, ???? 1978. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757381.

Wheaton:1981:IXR

[Whe81] Bruce R. Wheaton. Impulse Xrays and radiant intensity: The double edge of analogy. Historical Studies in the Physical Sciences, 11(2):367–390, ???? 1981. CODEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic). URL http://www.

jstor.org/stable/27757484.

Wheaton:1982:BR

[Whe82] Bruce R. Wheaton. Book reviews. Historical Studies in the Physical Sciences, 13(1): 163–??, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757509.

Wilson:1982:EAM

[Wil82] David B. Wilson. Experimentalists among the mathematicians: Physics in the Cambridge Natural Sciences Tripos, 1851–1900. Historical Studies in the Physical Sciences, 12(2): 325–371, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757499.

Wilson:1983:RSG

[Wil83]  David Wilson.  Rutherford,

simple genius. Hodder and Stoughton, London, UK, 1983. ISBN 0-340-23805-4. 638 + 8 pp. LCCN QC16.R8 W54 1983. US$14.95. URL http:/

                                        /catalog.hathitrust.org/
                                        api/volumes/oclc/10560773.
                                        html.

Williams:1999:FLB

[Wil99] James H. Williams. Fang Lizhi’s big bang: a physicist and the state in China. Historical Studies in the Physical and Biological Sciences, 30(1):49–87, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Williams:2013:THE

[Wil13] Rosalind H. Williams. The triumph of human empire: Verne, Morris, and Stevenson at the end of the world. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-89955-1 (hardcover), 0-226-89958-6 (ebook). xi + 416 pp. LCCN PN3435 .W495 2013.

Wilson:2015:C

[Wil15] Benjamin Wilson. The consultants: Nonlinear optics and the social world of Cold War science. Historical Studies in the Natural Sciences, 45(5):758– 804, November 2015. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2015.
45.5.758.

Williams:2024:SFG

[Wil24] Damien P. Williams. Scholars are failing the GPT review process. Historical Studies in the Natural Sciences, 54(5):625–??, November 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/54/5/625/203882/
Scholars-are-Failing-the-
GPT-Review-Process.

Winderlich:1949:BHC

[Win49] Rudolf Winderlich. Brenngläser als Hilfsmittel chemischen Forschens. (German) [????]. Chymia: Annual Studies in the History of Chemistry, 2(??):37–43, ???? 1949. CODEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757131.

Wise:1979:WTM

[Wis79] M. Norton Wise. William Thomson’s mathematical route to energy conservation: a case study of the role of mathematics in concept formation. Historical Studies in the Physical Sciences, 10(??):49–83, ???? 1979. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757387.

Wise:1982:MLB

[Wis82] M. Norton Wise. The Maxwell literature and British dynamical theory. Historical Studies in the Physical Sciences, 13 (1):175–205, ???? 1982. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27757511.

Wise:1987:RWD

[Wis87]  M. Norton Wise.  Review:

What did 19th Century British physics owe to Cambridge? Historical Studies in the Physical and Biological Sciences, 17(2): 363–368, ???? 1987. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/27757589.

Wise:2012:C

[Wis12a] M. Norton Wise. Crisis. Historical Studies in the Natural Sciences, 42(5):576–580, November 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.576.pdf.

Wisnioski:2012:ECC

[Wis12b] Matthew H. Wisnioski. Engineers for change: competing visions of technology in 1960s America. Engineering studies series. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01826-8 (hardcover). xvii + 286 pp. LCCN T14.5 .W5664 2012.

Wittje:2004:PAT

[Wit04] Roland Wittje. A proton accelerator in Trondheim in the 1930s. Historical Studies in the Physical and Biological Sciences, 35(1):115–152, September 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Witkowski:2021:BRN

[Wit21] Nicholas Witkowski. Book review: Nature and Modernity in an East Asian Key: An Appreciation of Two Recent Challenges to Eurocentric Scholarship on the Genre of “Natural History”. Historical Studies in the Natural Sciences, 51(2):269– ??, April 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/51/2/269/116757/

                                        Nature-and-Modernity-in-
                                        an-East-Asian-KeyAn.

Werner:1967:CTA

[WK67] Alfred Werner and George B. Kauffman. Contributions to the theory of affinity and valence. Chymia: Annual Studies in the History of Chemistry, 12 (??):189–216, ???? 1967. CO-DEN CHYMAP. ISSN 0095-

                                        9367. URL http://www.jstor.
                                        org/stable/27757282.

Wolff:2003:PKG

[Wol03] Stefan L. Wolff. Physicists in the ‘Krieg der Geister’: Wilhelm Wien’s ‘Proclamation’. Historical Studies in the Physical and Biological Sciences, 33 (2):337–368, March 2003. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Wolfe:2010:WDI

[Wol10] Audra J. Wolfe. What does it mean to go public? The American response to Lysenkoism, reconsidered. Historical Studies in the Natural Sciences, 40(1): 48–78, February 2010. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Wolfe:2013:BRG

[Wol13a] Audra J. Wolfe. Book reviews: Giving philanthropy a new history. Nick Cullather. The Hungry World: America’s Cold War Battle Against Poverty in Asia. Cambridge, MA: Harvard University Press, 2010. xii + 348 pp., illus., index. ISBN 978-0-674-05078-5. $36.50 (hardcover); ISBN 978-0-674-72581-2; $21.95 (paper).; Inderjeet Parmar. Foundations of the American Century: The Ford, Carnegie, and Rockefeller Foundations in the Rise of American Power. New York: Columbia University Press, 2012. xii + 356 pp., index. ISBN 978-0-231-14628-9. $40.00 (hardcover).; Brent Ruswick. Almost Worthy: The Poor, Paupers, and the Science of Charity in America, 1877–1917. Bloomington: Indiana University Press, 2013. xiv + 270 pp., illus., index. ISBN 978-0-253-00634-9. $37.00 (hardcover).; Olivier Zunz. Philanthropy in America: A History. Princeton, NJ: Princeton University Press, 2012. xii + 384 pp. ISBN 978-0-691-12836-8. $29.95 (hardcover). Historical Studies in the Natural Sciences, 43(5):619–630, November 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2013.
43.5.619.

Wolfe:2013:CSS

[Wol13b] Audra J. Wolfe. Competing with the Soviets: science, technology, and the state in Cold War America. Johns Hopkins introductory series in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0769-8 (hardcover), 1-4214-0771-X (paperback). viii + 166 pp. LCCN Q127.U6 W65 2013.

Wright:2014:ABI

[Wri14] Aaron Sidney Wright. The advantages of bringing infinity to a finite place: Penrose diagrams as objects of intuition. Historical Studies in the Natural Sciences, 44(2):99–139, April 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1525/hsns.2014.
                                        44.2.99.

Wright:2018:GSN

[Wri18] Will Wright. Geological semantics and the naming of the Anthropocene. Historical Studies in the Natural Sciences, 48(1): 110–122, February 2018. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Wise:1986:MWI

[WS86] M. Norton Wise and Crosbie Smith. Measurement, work and industry in Lord Kelvin’s Britain. Historical Studies in the Physical and Biological Sciences, 17(1):147–173, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

                                        jstor.org/stable/27757577.

Wu:2020:BRK

[Wu20] Harry Yi-Jui Wu. Book review: Korean medicine: A timehonored brand that transformed itself. John P. Dimoia, Reconstructing Bodies: Biomedicine, Health and Nation-Building in South Korea since 1945. Stanford, CA: Stanford University Press, 2013. 296 pp., illus. ISBN 978-0-8047-8411-5 (cloth). Theodore Jun Yoo, It’s Madness: The Politics of Mental Health in Colonial Korea. Oakland: University of California Press, 2016. 248 pp., illus. ISBN: 978-0-520-28930-7 (cloth). Soyoung Suh, Naming the Local: Medicine, Language, and Identity in Korea since the Fifteenth Century. Cambridge, MA: Harvard University Press, 2017. xv + 229 pp., illus. ISBN: 978-0-674-97696-2 (cloth). Historical Studies in the Natural Sciences, 50(5):633–??, November 2020. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://

online.ucpress.edu/hsns/
article/50/5/633/114258/
Korean-MedicineA-Time-
Honored-Brand-that.

Yagi:1984:CMM

[Yag84] Eri Yagi. Clausius’s mathematical method and the mechanical theory of heat. Historical Studies in the Physical Sciences, 15 (1):177–195, ???? 1984. CO-DEN HSPSAS. ISSN 0073-2672 (print), 2327-7394 (electronic).

URL http://www.jstor.org/
stable/27757545.

Yeang:2003:SLD

[Yea03] Chen-Pang Yeang. The study of long-distance radio-wave propagation, 1900–1919. Historical Studies in the Physical and Biological Sciences, 33(2):369– 403, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Yeang:2012:MOI

[Yea12] Chen-Pang Yeang. From mechanical objectivity to instrumentalizing theory: Inventing radio ionospheric sounders. Historical Studies in the Natural Sciences, 42(3):190–234, June 2012. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

                                        jstor.org/stable/10.1525/
                                        hsns.2012.42.3.190.

Yi:2007:CRD

[Yi07]  Doogab Yi.  The coming

of reversibility: The discovery of DNA repair between the atomic age and the information age. Historical Studies in the Physical and Biological Sciences, 37(S1):35–72, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

                                        URL http://www.jstor.org/
                                        stable/pdfplus/10.1525/
                                        hsps.2007.37.S.35.pdf.

Yi:2024:DSR

[Yi24] Jongsik Christian Yi. Decolonization and self-reflection: Teaching the Korean history of science in South Korea. Historical Studies in the Natural Sciences, 54(1):121–??, February 2024. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

                                        tronic). URL https://online.

ucpress.edu/hsns/article/
54/1/121/200191/
Decolonization-and-Self-
ReflectionTeaching-the.

Young:2006:SBb

[You06a] Stephanie Young. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 37(1):173–183, September 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Young:2006:SBa

[You06b] Stephanie C. Young. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 36(2):397–402, March 2006. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Young:2007:SBa

[You07a] Stephanie Young. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 37(2):479–492, March 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Young:2007:SBb

[You07b] Stephanie C. Young. Selected bibliography. Historical Studies in the Physical and Biological Sciences, 37(S1):133– 139, September 2007. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsps.2007.37.S.133.pdf.

Yu:2023:ABC

[Yu23] Gloria B. Yu. Addiction beyond a cure. Historical Studies in the Natural Sciences, 53(2):189– ??, April 2023. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https:

//online.ucpress.edu/hsns/
article/53/2/189/195999/
Addiction-Beyond-a-Cure.

Zakariya:2012:MKW

[Zak12] Nasser Zakariya. Making knowledge whole: Genres of synthesis and grammars of ignorance. Historical Studies in the Natural Sciences, 42(5):432– 475, November 2012. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/pdfplus/10.1525/
hsns.2012.42.5.432.pdf.

Zakariya:2013:BRH

[Zak13] Nasser Zakariya. Book review: Is history still a fraud? Cynthia Stokes Brown. Big History: From the Big Bang to the Present. New York and London: New Press, 2007. xvi + 288 pp., illus., bibl., index. ISBN 978-1-595-58196-9. $25.95 (hardcover).; David Christian. Maps of Time: An Introduction to Big History. 2nd edn. Berkeley: University of California Press, 2011. xxvii + 642 pp., illus., bibl., index. ISBN 978-0-520-27144-9. $26.95 (paper).; Edmund Russell. Evolutionary History: Uniting History and Biology to Understand Life on Earth. Cambridge: Cambridge University Press, 2011. xxi + 216 pp., illus., bibl., index. ISBN 978-0-521-74509-3. $23.99 (paper).; Daniel Lord Smail. On Deep History and the Brain. Berkeley: University of California Press, 2011. xiv + 271 pp., bibl., index. ISBN 978-0-520-25812-9. $21.95 (paper).; Frederick Spier. Big History and the Future of Humanity. Chichester, U.K. and Malden, MA: Wiley-Blackwell, 2010. xv + 272 pp., illus., bibl., index. ISBN 978-1-4443-3943-7. $130.95 (hardcover). Historical Studies in the Natural Sciences, 43(5):631– 641, November 2013. CO-DEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

URL http://www.jstor.org/
stable/10.1525/hsns.2013.
43.5.631.

Zulueta:2009:MMG

[Zul09] Benjamin C. Zulueta. Master of the master gland: Choh Hao Li, the University of California, and science, migration, and race. Historical Studies in the Natural Sciences, 39(2):129– 170, May 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Zunz:2012:PAH

[Zun12] Olivier Zunz. Philanthropy in America: a history. Politics and society in twentieth-century America. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-12836-7 (hardcover). x + 381 pp. LCCN HV91 .Z86 2012.